A Complete Bibliography of Computers and Security, 1990–1999

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
      WWW URL: http://www.math.utah.edu/~beebe/

23 September 2019 Version 1.00

Title word cross-reference

#38 [Ano92-90].

$18 [Ano92-78]. $24.95 [Zaj94b]. $29.95 [Hig92a]. $2m [McC98-48]. $32.95 [Lat93]. $4 [Mey97-56]. $44.95 [Zaj94a]. $495 [Zaj94c]. $640 [McC98-43]. $90 [McC98-53]. ^{T M} [Fol97]. E [Hin99c]. ϵ [Ano92-90].

0 [Hig92a, Lat93, Zaj94a, Zaj94b]. 0-07-707805-5 [Zaj94b]. 0-471-54777-8 [Lat93]. 0-471-59756-2 [Zaj94a]. 0-939195-72-2 [Hig92a].

1 [Ano92-45, Ano92-76, Ano92-81, Col99a, Zaj94c]. 1-881585-01-8 [Zaj94c].

1. [Ano92-89]. 1.5 [Mey96-49]. 10 [Ano92-84]. 10126 [JvdV96]. 11 [Ano91-39, Ano92a, Hig92b, Hig92c, Hig93c, Hig93d, Hig93e, Hig93f, Hig93g, Hig93h, Hig94a, Hig95-30, Hig95-31, Hig96a, Hig97a, Hig97h]. 11.1 [Ano91-40]. 111 [Ano92-92]. 117 [Ano97-74]. 11rch [Ano96-59]. 12 [Ano97v, Col95-135]. 12th [Kow91c]. 13th [HP90a]. 14. [Ano92-66]. 16 [Ano92s, Ano92-46, Ano92-95, Ano97-85]. 16th [Hig93a]. 17 [Ano92-88]. 17th [Hei94a, Hig95-48]. 190 [Ano97-84]. 1989 [SB90]. 1990 [Ano90-38, Ano90-42, Ano90-45, Ano90-40, Ano90-41, Ano90-39, Ano90-44, Ano90-43, Woo90a]. 1990s [Ano90-47, Zaj90a]. 1990s-what [Zaj90a]. 1991 [Ano91-54, Ano91-50, Ano91-53, Ano91-48, Ano91-49, Ano91-47, Ano91-52, Ano91-51]. 1992 [Ano92y, Ano92-68, Ano92-75, Ano92-71, Ano92-70, Ano92-69, Ano92-74, Ano92-73, Ano92-72, Ano92-86, Ano92-96]. 1993 [Ano93-49, Ano93-52, Ano93-46, Ano93-48, Ano93-45, Ano93-47, Ano93-51, Ano93-50, Mey93]. 1994 [Ano94-50, Ano94-44, Ano94-43]. 1997 [Ano97u, Ano97-60]. 19th [Hig97b]. 1st [Hig95-48].

2 [Ano92-65, Ano92-90, Cog95, Col99b, Hig92a, Zaj94a]. 2.0 [Mey96-71]. 20 [Ano92-48]. 2000 [Han99-58]. 21 [Kov99a]. 21-st [Nos99, Kov99a]. 21st [Ano96-84]. 22 [Ano92-85]. 232 [Smu90]. 24 [Ano92-64]. 2nd [Zaj94b].

3 [Ano92-29, Ano92-47]. 3-89209-000-9 [SB90]. 30 [Ano97-47]. 33 [Ano92-79]. 34.10 [MNP96]. 35 [Ano92-83]. 37 [Ano92-55].

4 [Ano92w, Ano92-91]. 4-step [Mey97-42]. 4.0 [Mey99-45]. 40 [Ano97-55]. 44 [Ano97-35]. 45M [Han99-84].

5 [Zaj94b]. 52 [Ano97-73]. 54 [Ano97-76]. 56-bit [Mey99b].

6 [Ano92h, Ano92-87]. 60. [Ano92j].

70 [Ano97-72]. 71 [Ano97-27]. 77 [Ano97-36].

8 [Lat93, Zaj94c]. 8. [Ano92-78]. 80 [Ano97c]. ’89 [EM90].

’90 [Ano90-48, Ano91q]. ’90s [Lat93]. ’91 [Ano92u]. ’92 [BN92, She91, Ano92-94]. ’93 [Hig93i]. 94966-1219 [Zaj94c]. 95 [Cur95]. ’96 [Ano96y]. ’97 [Ano97q]. ’98 [Hin98h]. ’99 [Ano99r, Ano99s, Hin99f].

= [Ano96-75].

A. [Ano96-30]. ABCs [Kow91a]. Abilene [Ano96-50, Kru96]. abound [Col95-61]. Abrams [Hig95b, Hig95z]. abroad [Nic99-59]. Abstract [Ano90a, Col96a]. Abstracts [Ano90b, Ano90c, Ano90d, Ano90e, Ano90f, Ano90g, Ano90h, Ano91a, Ano91b, Ano91c, Ano91d, Ano91e, Ano91f, Ano91g, Ano91h, Ano92b, Ano92c, Ano92d, Ano92e, Ano92f, Ano92g, Ano93a, Ano93b, Ano93c, Ano93d, Ano93e, Ano93f, Ano93g, Ano94a, Ano99a, Col94a, Col94b, Col94c, Col94d, Col94e, Col94f, Col94g, Mes93, Ano99s]. abuse [DFIR99, Han99z, Mul97c]. abuses [Beq98b]. abusive [Han99-99]. academia [McC98a]. Academy [Ano96l]. accelerators [Mey98m]. acceptability [DBHM95]. Accepted [Ano96-50, Kru96]. Access [Ano96a, CSW94, GS91, Mey96a, Ano96j, Ano96k, Ano96-106, Ano96-108, Ano97t, Ano97p, Ano97-75, Ano99b, BE95, BS94, Bor96, Col95-132, Col95-130, Col95-161, DF91, GKKT92, Han95a, Han99-36, HMCC95, HTY96, JCW91, Kim95, Klo90b, LE97, Mad95, Mey96n, Mey96y, Mey96-55, Mey96-53, Mey97-78, Mey98-60, O’S95, Oba93, OA90, Par90, Par95a, SHW94, She95c, Ste91, ZH90]. Accessibility [PK91]. accessible [Nic99-33]. accessing [Col95z]. accidentally [Han99-52]. accords [McC97o]. account [Ano97-56]. accountability [Oli95b]. accountants [Col95-42]. Accounting [Ano92-83, Hug95a]. accreditation [Har93, Hig95o]. accused [McC97l]. accuses [Han99-36]. ACF2/MVS [Cro98]. achieve [Woo91b]. Achieving [Hop94, Smi95b, Smi96, Ano96-111]. Achilles [McC98b]. across [Col95-120, Mey99r]. Act [Ano92-27, Pou99a]. action [Ano92z, Col95-67, Col95-95, Faw92a, Pou98b, Pou99a]. activation [LE98]. Active [Mey99c, AM95]. ActiveX [Mey96z, Mey96-77, Hop97]. acts [McC98-38]. adapting [Ano97-83]. adaptive [BH93]. add [Ano92w]. Added [Sch93, Mit95]. Added-on [Sch93]. Adding [Col96b]. address [Han99p]. addresses [Col95-141]. addresses’net [Mey96-91]. Addressing [Pep97]. administration [Ano99q, Ass92, Ano92-64]. Adoption [McC98c, Lil94]. Adrenaline [Han98o]. ADS [Ano96d]. ADSL [Ano99b]. Advanced [Osb97a]. advantage [Woo91b]. adversary [Men91b]. Aegean [Ano96-66]. AES [Han99x]. afar [Mey98-43]. affected [CM92]. Affeldt [Hig95-48]. afford [Mey99e]. Africa [Ano96k]. Afrikaans [Ano96k]. After [Ano92h, Ano92t, DS99, McC98-78, Mey98j, Nor93]. AG [Hig95k, Mey97-79]. again [Col95-103, Col95-146]. against [Ano97-47, Coh92b, Col95c, Han99n, Har92, Mey96-50, Mey98-49, Moe92]. age [Ano96-86, Lin94b, Mey97-51]. Agency [Hig95-34, Col95-152, Hig95-48]. agents [Ano97-75, Bai97, Mey98-57, SSP90, vS98]. aggregation [Han99-85]. agreement [Mey99t, Ano96i]. ahead [Mey97-62]. AI [Gal95]. Aid [Ano96w, Ano97z, Col95-149, PvS91]. aids [Men90b]. aim [Col95-90, Mey98-74]. aims [Col95-80, Mey98-80]. ain’t [McC97v, Hig95f]. air [Hin99b]. Airports [McC98d]. Airtight [McC98e]. AL [Hig95u]. Alamos [Nic99b]. Alan [Ano96c, Ano96-50, Ano96-110, Ano97c, Hig95-35]. Alarm [Col95b, Han99-68]. alarms [Nic99-68]. Albanese [Ano96-34]. Albert [Hig95h]. Alert [Ano92-86, Col95-121, Mey97-40]. alerts [Col95-44]. Alexander [Hig95k]. Alexandris [Ano96i]. Algorithm [DF91, BH92, BH93, DFHR90, Mey99b, PBGV90, Ram90]. Algorithm-sequenced [DF91]. algorithms [GDNC94, RR91, Zaj94a]. Alien [Hig95s]. alive [Han99-71]. alliance [Ano99-46, Han99-42, Han99-75, Nic99a]. alliances [Mey99u]. Almos [Hig95c]. almost [Col96f]. alone [Woo97c]. along [Col95-158]. Already [Han99-58]. alternative [Nic99c]. America [Col95c, Han99-80]. American [Ano92-79, Woo90a]. among [Rat99b]. Analysing [Som94]. Analysis [CPS^{+}98, Hig95-29, Hig95-34, Hig95-45, Mey97c, Ano96-98, Ano97k, Ano97x, BE90, Bas91, Bol91, Cie97, Coh91a, ELB93, Hig95i, Hig95v, KJ95, Kow91b, Kow91c, KS97, LE98, LS93, Lun92a, Mer90a, Mey96g, Mey98-71, ZHM90, dRE96, vdME98]. analyzing [Ano96-115]. Anderson [Ano96-101, Hig95u]. Andres [Ano96-34]. Andrew [Ano92h, Ano96-33]. Annapolis [Hig95-47]. announcement [Ano90m]. announces [Han99m]. Annual [Pou98a, Ano90m, Mey99a]. anonymity [Oli95b]. Anonymizer [Han99d]. anonymous [HS95a]. ANSA [Mey98-96]. ante [Mey98t]. antennas [Muf90]. Anthes [Ano92h, Ano92-64, Col95-135]. Anti [Col95e, McC98f, Mey97d, Mey98b, Nic99a, Ano92-65, Ano92-96, Col95-57, Hig95f, Mey98l, Mey98-104, Mey99g, PK98, Zaj92b, Zaj94d]. anti-hacker [Mey98l]. anti-intrusion [Hig95f]. Anti-piracy [Nic99a]. anti-viral [Zaj92b, Zaj94d]. Anti-virus [Col95e, McC98f, Mey97d, Mey98b, Ano92-65, Col95-57, Mey98-104, Mey99g, PK98]. anti-viruses [Ano92-96]. Anticipating [Kru95a, Kru95c]. antics [Mey98-29]. Antivirus [Mey96b, Ano92w, Col95-143, McC98-44, Mey96-101]. any [McC98-27, Pea91]. anymore [Col95-64]. anything [Col95-64]. AOL [McC98g, Mey97-31]. apathy [Col95f]. APIs [CGO93]. app [Mey96d]. app-level [Mey96d]. Appeals [Han99e]. Appeals-court [Han99e]. Appendices [Mul97g]. applets [Han99-56, McC97k, Ano97-73]. appliance [Mey97-87]. applicability [SPP98]. Application [CGO93, CPS^{+}98, Han99f, Hig95u, Par90, Ano96-98, Ano97p, Hig95-27, PG94, VSU93]. Applications [HS95a, Hig95-47, Ano90m, Kap98, Lem96, Pre99, SGI^{+}99, Hig95t]. Applied [Ano96-51, Zaj94a]. apply [Mey98-57]. Approach [Ano96b, DKGS94, Hig95c, BE94, CJC98, Far91, Han99-83, Hig95k, Hig97c, Hit95, JJ96, Lil94, Mey96g, Øln93, SG94, TCWH95, vSvdM94, Hig95n, Hig95p]. Approaches [Hig95o, McN95b]. approval [Mey97-60]. approved [Mey98-92]. approves [Ano92-67]. Approximating [Ano96-66]. approximation [KP99]. April [Ano90-38, Ano91-54, Ano92-68, Ano94-50]. Arca [Hig95f]. Archer [Hig95t]. architecture [Ano96-103, Ano99-48, Cur95, Hig95e, McC98-62, Ram90, SS93, She96, Hig95g]. architectures [Hig95j, LHB90, Hig95h]. Area [Hig95x, McC97a, Hig95p, Lat92]. areas [Mey96a]. aren’t [Mey98-27, Nic99-39]. arithmetic [BH92, BH93, CIRM95]. Arlington [Hig95s, Hig95-46]. arm [Mey97d]. armour [Col96u, Mey97-37]. arms [Mey98i, Nic99q]. arrives [Nic99z]. art [Par97a, Pre99]. Arthur [Ano92-79]. article [Col96a]. Articles [Ano92e, Ano90i, Ano90b, Ano90c, Ano90d, Ano90a, Ano90e, Ano90f, Ano90g, Ano90h, Ano90-49, Ano90-50, Ano91a, Ano91b, Ano91c, Ano91d, Ano91e, Ano91f, Ano91g, Ano91h, Ano91-55, Ano91-56, Ano92b, Ano92c, Ano92d, Ano92f, Ano92g, Ano92-82, Ano93a, Ano93b, Ano93c, Ano93d, Ano93e, Ano93f, Ano93g, Ano93-54, Ano94a, Ano99a, Col94a, Col94b, Col94c, Col94d, Col94e, Col94f, Col94g, Mes93]. Artificial [Hig95a]. artists [Mey98-93]. Ascom [Col95-31]. Asia [Col95-136]. Asian [Eva92]. Ask [Ano96c, Han99-86]. ASP [Han99f]. aspects [Ano96-36, Col95-123, Gil92, Lat92, Ste91]. assailed [Mey96-93]. assails [Col95-43]. assault [Ano99-45]. assessing [Ano96-101, Bol95]. assessment [Ano96-39, Far91, JATC93, Rat99a, Svi94]. assets [McC98p, Mey96-89]. assignment [SHW94]. assistance [SB90]. associate [Han99m]. associated [Tay95]. Associates [Hig92a, McC98-43]. associative [BPH^{+}97]. Assurance [Hig95x, Ano96-111]. Asymmetric [Lau92]. AT&T [Hig95m]. Athens [Ano96l, Ano96-111]. ATM [Col95-40, Col96d, Han95c, SG98]. Attack [Mey96c, Ano96d, Ano97a, BH93, Col95-49, Han99-37, Hig91a, McC98-34, Mey97h, Mey97-27, Mey98-100, Mey99c, Nic99t, Nic99-70]. attacker [McC98-54, Mey99-49]. Attacks [Mey98c, Ano97-85, Bon96, Coh96a, Coh97a, CPS^{+}98, Coh99, Col95-160, DS99, Han99-38, Hig95-32, HDW97, McC98-75, MS90, Mey97u, Mey98-49, Mey99-56, Nic99l]. attempt [Klo90a]. attempts [Han99-61]. attend [Car90a, Han99-49]. attends [Col95-45]. attitudes [DFIR99]. attractive [McC98-41]. attributes [Pre91a]. Aubrey [Ano96-75]. Aubrey-Jones [Ano96-75]. auction [Han99s]. Audit [Ano96-27, Nic99b, Poo97, Rus94, Ano96w, Ano96-77, Ano97-69, Cro95a, Cro95b, Jon97a, Jon97b, Kru95b, Lun92a, Men90b, Woo97e]. Audit-by-receiver [Rus94]. Auditing [Hug95a, Osb97b, Osb98, PH90, PdL91, Ano96-65, Col95-106, Mer90b, Mit95, Osb97a]. auditor [Ano96c]. auditors [CM92, Oli95a]. audits [FW95, McC98-72]. August [Ano90-42, Ano91-50, Ano93-49, Ano96v, Ano96-84]. Aussie [Nic99c]. Australia [Hig95p]. Austria [Ano96-51]. authentic [DFHR90]. authenticated [Ran93]. Authenticating [TCHS96]. Authentication [GBD93, LCC94, McC98h, ABD93, Ano96-30, Ano96x, Ano97-59, CL94, CH94, CHT97, Col95-32, HZS94, HLY92, Kan94, Kim95, Lau92, Lea95, Lyn91, Mey97i, Mey97-87, Mey98-39, Mey99-30, Rei96, SJJF96, YS99, YA96]. Author [Ano90i, Ano91i, Ano93h, Ano94b, Ano95a, Ano96e, Ano97b, Ano98a, Ano99c, Col95-50, Han99-53, Hig95i, KS97]. Authorities [Han99-84, Ano97n, Han99-65, Mey98-83, Par97d]. authorities/trusted [Ano97n]. authorization [Ano97x, LK92, Rus93, Rus94]. authors [Ano91-28, Ano91-29, Ano91-30, Ano92-38, Ano92-39, Ano92-40, Ano92-41, Ano92-42, Ano92-43, Ano93q, Ano93-27, Ano93-28, Ano93-29, Ano93-30, Ano93-31, Ano93-32, Ano94t, Ano94u, Ano94v, Ano94w, Ano94x, Ano94y, Ano94z, Ano94-27, Sol92, SW93]. Authorship [KS97, Hig95i]. Automated [Cro95a, Lun92a, HJS^{+}93]. automatic [LR92, LV94]. automation [Hig92r]. Aventail [Mey98d]. Averting [Mey97e]. avoid [Faw92b, Mey96-90]. awaiting [Col95-45]. Award [Mey93]. Awareness [Hig95s, Hig95-46, DFIR99, Klo90d, Woo97c]. away [Col95-138, Mey99-48]. axe [Nic99r]. Axent [Mey98e]. axiomatic [Ran92].

B [Ano96-107]. B1 [Paa91]. Back [Ano99d, Han99g, Ame97, Sol92]. Backdoor [Mey97f, Nic99-53]. BackOrifice [Mey99-31]. backs [Ano99-67, Han99n, Mey97-34, Mey98-98]. backup [Ano97w]. bad [Han98j, Hig95z, Hin98g, Sch95b, Zaj92a]. bag [Mey98-103]. baggage [vSG98]. Bailey [Ano96-32]. bait [Mey96-95]. bake [Han99x]. bake-off [Han99x]. Balancing [Beq98a]. ballot [JL96, NSS91]. balloting [JJ96]. ban [Han99e]. bandits [Col95m]. Bang [Bac97a]. Bank [Mey96d, Ano97-56, Mey96-78]. Banking [McC98i, Ano92-45, Han99q, Osb95]. Banks [Col95g, Mey97g, McC97a, Moe92]. Baranca [Hig95c]. Barbara [Hig95m]. bark [Mey99-28]. Barriers [Jar99]. Bartlett [Hig95v]. Base [Ano96-78, Ano96-85, Mey97-61, SPP98, Ste91]. Based [CPS^{+}98, Hig95m, Ano96x, Ano97-70, Ano97-72, CL94, CH94, CHT97, Coh91b, Col95-32, Hig95h, HTY96, Kap98, Laf91, Man96, Mey99-44, MD94, Par95a, PMRZ99, PRZ99, Rei96, Ric99, San92, SS93, Sch98, SJJF96]. Baseline [Zaj94c, Gri97]. baselines [Ano97k]. basic [Col95-109]. Baskerville [Ano96-115]. Battening [Col96c]. battle [McC98-77, Mey97-81, Mey98i]. battles [Col95v, McC97m, Mey96f]. Bauknecht [Ano96-30]. BBS [Col95-65]. BC [Ano99-59]. BCM [Nos99]. Be [McC98j, Nic99d, Col95o, Col95-91, Col95-65, Eva90, Han99h, Han99g, Han99-100, Hig95y, Hig97y, McC97k, McC98r, McC98w, Mey96-59, Mey97-81, Mey99-33, Mey99-55, Nic99-33, Nic99-67, Nic99-68, Thi91, Zaj90b]. bearer [Nic99d]. beat [Ano99-46, Col95m]. beater [Col95-148]. Beavercreek [Hig95v]. because [Nic99-52]. Beckman [Mey93]. becoming [Mey99-49]. beef [Mey98-52, Mey99-38]. beefing [Mey97-48]. beefs [Mey98j]. been [Han99-49]. before [Ano97-79, Mey99-49, Nic99o]. begin [Men91c]. begins [Han99-75]. behind [Ano96-118, McC98u, Mey96-78]. beholder [Nic99-58]. Being [Han99h, Han99-79, McC97l, Mey97l, Mey98-37]. Beker [Ano96-86]. Belfired [Ano96d]. Belief [Hig95b]. Bell [Hig95m, Hin99b, LH94]. bell-shaped [Hin99b]. belongs [Mey98b]. benchmarking [Lil94]. benefit [Ano99o]. Benelux [Ano92k]. Bennett [Ano92-79]. Benson [Hig95t, Hig95-27]. Bequai [Ano96v, Ano96-84]. Berkeley [Ano96-34]. Bernard [Ano92-83]. Berne [Ano96-34]. Bertino [Hig95n]. best [Ass92, Mey98-46]. bets [Mey98-32]. Better [Mey99d, Han99-39, JvdV96, Ken93, McC98-36, Mey97-64, Mey98f, MG92, Han99i]. Between [Ano97c, CJB94, Hop97]. Beware [Ano96f, Mey96e, Nic99-45]. Beyond [Hic95, LS93, Lyn91, Par97a]. bibliography [Mul97g]. Big [Col95h, Nic99e, Ano99x, Han99b, Mey97h]. biggest [Ano92-88]. Bill [Ano99-67, Hig95o, Nic99s, Nic99r, Mey96-88, Nic99-65]. billing [Han99-84]. billion [Woo90a]. binary [CSW94]. Binder [Ano96-101]. Biology [Mey96f]. biometric [DBHM95, Mey98-65]. Biometrics [Kim95, She92a, Bov95, Han99y]. bit [Mey99b]. bite [Mey98-54, Mey99-28]. Bits [Hig95-37, Hig96j, Ano93-53, Ano97-71, Col95-132, Hig90a, Hig90b, Hig90c, Hig90d, Hig90e, Hig90f, Hig90g, Hig90h, Hig91f, Hig91g, Hig91h, Hig91i, Hig91j, Hig91k, Hig92j, Hig92k, Hig92l, Hig92m, Hig92n, Hig92o, Hig92p, Hig92q, Hig93j, Hig93k, Hig93l, Hig93m, Hig93n, Hig93o, Hig93p, Hig94b, Hig94c, Hig94d, Hig94e, Hig94f, Hig94g, Hig94h, Hig94i, Hig95-36, Hig95-38, Hig95-39, Hig95-40, Hig95-41, Hig95-42, Hig95-43, Hig96h, Hig96i, Hig96k, Hig96l, Hig96m, Hig96n, Hig97d, Hig97p, Hig97o]. bitten [Han99r]. Bl [Hig95m]. blacklist [Faw92b]. blackmail [Ano92-89, Mey97-29]. Blaming [Ano96g, Spa97a]. blanket [Mey96-75]. blasé [Mey96p]. blast [Mey98v]. blasts [Mey99-36]. Blaustein [Ano96-107]. Blind [Ano96h, Ano96-40, vSN92]. block [ABD93, McC97e, MV91, PC99]. blocks [Han99-69, Mey96-48, Mey98-63]. Bloombecker [Woo90a]. Bloomberg [McC98-57]. blows [Han99-82]. blunder [Ano99-62]. Board [Ano91-41, Ano92-62, Ano95-29, Ano95-30, Ano95-32, Ano95-35, Ano96-69, Ano96-71, Ano97-48, Ano97-50, Ano97-51, Ano97-52, Ano98-28, Ano98-29, Ano98-30, Ano99-39, Ano99-40, Ano90-31, Ano90-32, Ano90-33, Ano90-37, Ano90-34, Ano90-35, Ano90-36, Ano90-49, Ano91-42, Ano91-43, Ano91-44, Ano91-45, Ano91-46, Ano92-56, Ano92-57, Ano92-58, Ano92-59, Ano92-60, Ano92-61, Ano92-82, Ano93-41, Ano93-42, Ano93-43, Ano93-44, Ano93-55, Ano93-56, Ano93-57, Ano94-36, Ano94-37, Ano94-38, Ano94-39, Ano94-40, Ano94-41, Ano94-42, Ano94-51, Ano95-28, Ano95-31, Ano95-33, Ano95-34, Ano96-67, Ano96-68, Ano96-70, Ano96-72, Ano96-73, Ano96-74, Ano97-49, Ano97-53, Ano97-54, Ano98w, Ano98x, Ano98y, Ano98z, Ano98-27, Ano99-37, Ano99-38, Ano99-41, Ano99-42, Col95-34, Ano90o, Ano91r, Ano92x, Ano93r, Ano94k, Ano95j, Ano96-37, Ano97y, Ano98l, Ano99v]. Bob [Ano92-66, Ano96-87, Ano96-109]. body [Mad98]. bombard [Nic99t]. Bombed [Ano92i]. bombing [Ano92t, Ano99-27, Han99-37]. bombs [Col95-147]. bonds [Ano99e]. Book [Ano92j, Ano96i, Ano96j, Ano96k, Ano96l, Hig92a, Hig95h, Hig95f, Hig95i, Hig95j, Hig95s, Hig95k, Hig95l, Hig95c, Hig95m, Hig95o, Hig95n, Hig95g, Hig95d, Hig95q, Hig95e, Hig95p, Hig95r, Hig95t, Hig95u, Hig95v, Hig95w, Hig95x, Lat93, SB90, Zaj94a, Zaj94b, Zaj94c, Col95-50]. booming [Mey98-88]. boost [Col95-60, Mey97-28]. boosts [McC98-30]. bootstrapping [Coh91c, DFHR90]. Booz [Hig95s]. borders [Mol97]. Borough [Col95-71]. Borrett [Hig95-35]. Boston [Ano92s]. both [LK92]. Bottom [Ano92-85]. Boucher [Hig95-44]. bought [Hig97v]. Boulahia [Hig95w]. Boulahia-Cuppens [Hig95w]. braces [Nic99l]. Brand [Ano92-96]. Brassard [Ano92-79]. brave [Car90a]. Bray [Ano97u]. Breach [Gil92, Col95w]. breadth [DD94]. break [Col95c, Col95-96, McC98-81, McL92, Mey98a, Mey99p]. break-ins [Col95c, Mey99p]. breakers [McC97c]. Breaking [Gor98]. brick [Col95d]. Bridging [Mey97i]. brief [MNP96]. Brigham [Hig95l]. bring [Ano97-85, Col95-40]. brings [Col95-134]. Britain [Ano96-110, Lai97]. British [Han99j]. broadcast [CH94, CHT97]. broadcasting [Lia94]. broader [Col96p]. Broderick [Ano96-76]. broken [Mey99b]. brother [Ano99x, Mey97-30, Nic99e]. Brown [Ano92-66]. Browser [Mey97j, Ano97-56, Hig96d, Mey98q]. browsers [Mey97-72]. Bruce [Zaj94a]. BS7799 [Pou99b]. BSA [Ano92k, Col95i, Col95j, Nic99f]. BSykB [Col95-48]. BT [Ano96-116]. Buck [Woo90a]. budget [Col95-81, McC98-32]. Bug [Ano99e, Nic99g, Nic99h, Mey98-47, Nic99-35]. bugged [Col95-81]. Bugs [Mey97k, Mey97j]. Build [Col95k, Han99-92]. Building [Fol97, OvS92, Han99-96, Men92b]. builds [Col95-98]. built [Mey96-102]. built-in [Mey96-102]. Bull [Ano96-64, Ano96-110]. bullet [Cor95b]. Bulletin [Ano92-29, Ano92-47, Ano92-90, Ano92-96, Col95-34]. Bulletproof [McC97b, Mey98g]. bump [Hin99b]. Bundled [Mey98h]. burden [Pea91]. Bureau [Han99i]. burglar [Han99-68]. Burgoon [Hig95o]. Burmester [Ano96i]. Burned [McC98k, Mey96-90]. Burning [Woo91a]. Business [Ano96m, Ano96-115, Ano97-60, Han99i, Mey98-88, RH91, She95b, SS95, Ada96, Ano96-36, Ano96-81, Bac97a, Buc95, Buc96, Bul99, CM93, Col95h, Col95-68, Col95-127, Dev99, HS95b, Kov97d, Kov98a, McC98-28, Men94, Mey96g, Mey96-43, Mey96-86, Mey97i, MM96, Mul95, Par97b, Par97c, Woo90a]. business-to-business [Mey97i]. Businesses [Ano99f, Ano97-85, CM92, Mey97-29]. bust [Han99-84, McC98-53]. busy [Mey97o]. buy [McC98-43, Mey98-56, Moe92, Pla93, Mey98-100]. buy-in [Pla93]. buys [Mey96-71, Mey98-104]. Bytes [Hig95-37, Hig96j, Ano93-53, Ano97-71, Hig90a, Hig90b, Hig90c, Hig90d, Hig90e, Hig90f, Hig90g, Hig90h, Hig91f, Hig91g, Hig91h, Hig91i, Hig91j, Hig91k, Hig92j, Hig92k, Hig92l, Hig92m, Hig92n, Hig92o, Hig92p, Hig92q, Hig93j, Hig93k, Hig93l, Hig93m, Hig93n, Hig93o, Hig93p, Hig94b, Hig94c, Hig94d, Hig94e, Hig94f, Hig94g, Hig94h, Hig94i, Hig95-36, Hig95-38, Hig95-39, Hig95-40, Hig95-41, Hig95-42, Hig95-43, Hig96h, Hig96i, Hig96k, Hig96l, Hig96m, Hig96n, Hig97d, Hig97p, Hig97o].

C [Ano96-30, Ano97-72, Hig95l, Hig95-45, Zaj94a]. C.Y.A. [Kin92b]. C2 [Hig95h]. CA [Hig92a, Hig95f, Hig95u, Hig95-32, Zaj94c, Ano96-34, Cro98, Hig95a, Hig95r, Hig95t, Hig95-44]. CA-ACF2 [Cro98]. CA-ACF2/MVS [Cro98]. CAACF2 [Cro97]. CAACF2/MVS [Cro97]. cables [Smu90]. cache [Col95-83]. cage [Mey98-68]. Calendar [Ano90j, Ano90k, Ano90l, Ano91j, Ano91k, Ano91l, Ano91m, Ano91n, Ano91o, Ano91p, Ano92l, Ano92m, Ano92n, Ano92o, Ano92p, Ano92q, Ano92r, Ano93i, Ano93j, Ano93k, Ano93l, Ano93m, Ano93n, Ano93o, Ano93p, Ano94c, Ano94d, Ano94e, Ano94f, Ano94g, Ano94h, Ano94i, Ano94j, Ano95b, Ano95c, Ano95d, Ano95e, Ano95f, Ano95g, Ano95h, Ano95i, Ano96n, Ano96o, Ano96p, Ano97d, Ano97e, Ano97f, Ano97g, Ano98b, Ano98c, Ano98d, Ano98e, Ano98f, Ano98g, Ano98h, Ano98i, Ano99g, Ano99h, Ano99i, Ano99j, Ano99k, Ano99l, Ano99m, Ano99n]. Calender [Ano96q, Ano96r, Ano96s, Ano96t, Ano96u, Ano97h, Ano97i, Ano97j]. California [Hig95r, Hig95t, Hig95a]. Call [Mey98i, Han98n, Mey96x, Han99-36]. called [Col95-99, McC98-34]. Calling [Col95l, McC98l]. calls [Ano99q, Han99-46]. Cambridge [Col95-96]. Can [Ano97k, Hig95y, McC98m, Mey99e, SW93, Thi91, Zaj90b, Ban95, Bou97, Col95-91, Col96o, Eva90, For98b, Han99h, Hin99d, Mey96y, Mey97-28, Mey97u, Mey97-55, Mey97-81, Mey98-94, Mey98-82, Mey98-75, Mey99s, Moe92, Nic99-68]. Canada [Ano99p, Ano99o, Nic99i, Nic99-60]. Canadian [Han99-83, Hig95-29]. capabilities [Col96i]. capability [SS93]. capability-based [SS93]. capitalize [Klo90c]. capsizing [Ano96-102]. capture [Han99-48]. Card [Nic99j, Col95-123, Han99-84, Mey99a, SJJF96]. cards [Bov95, Col95-31, YS99]. care [Mey97-32]. career [Kov97a]. careful [McC98j]. Carl [Ano96-29, Ano96-49]. Carlyn [Ano92-66]. Carnegie [Hig95-32, Hig95-33]. carriers [Ano92-88, Col95m]. Carroll [Ano92-85]. Case [Ano92s, Har92, Ano97-81, Col95-28, Col95-78, GTKK91, LS93, McC98-59, Men92b]. cases [McC98a]. Cashless [Ano96v]. cast [McC98a]. casts [McC98-83]. cat [Mey98-103]. catch [Ano92-78, Mey99-50, Nor92b]. catches [Mey98-72]. catching [Han99y]. categories [Fin98b]. caught [Col95-133, Mey98a]. cause [CPS^{+}98, Han99-56, Mey96-37]. causes [Col95-85, Col95-86]. causing [Han99-63]. caution [Han99w]. CD [Ano92-67, Col95-96, New95]. CD-ROM [Ano92-67, New95]. CE [Mey99-51]. cease [BdAR99, McC98l]. cease-fire [McC98l]. Cedex [Hig95w]. Cellular [Col95m, Col95r, Mey96-76]. Center [Hei93c, Hig95-48, McC98-48, McC98-54]. centers [Han99f, Sch91]. central [McC98h]. centralized [Mey99-53]. Centre [Col95o, Hig95-27, Ano92-80]. centric [Ano96a]. Century [Kov99a, Ano96-84, Nos99]. CERT [Hig95w]. Certifiable [McC98n]. Certificate [Hug99, Kap98, McC98-38]. Certificates [Mey97l, Ano97-37, McC98-68, McC98-73, Mey97-71, Mey98-85, Ano96j]. Certification [Hig95o, Ano97n, HH99, Par97d]. certifies [Mey99-29]. chain [Fin96, Fin97a, Gor98]. chaining [MV91]. challenge [McC97c, Smi95a]. challenges [Ano96-84, Hic95]. Chang [Ano97-72]. change [Hin99b, PH90]. changed [Ano92-86]. changes [Mey97-30, Mey98-69]. changing [Ano97-83, Tay95]. channel [TCHS96]. Chao [Hig95m]. Chao-Yeuh [Hig95m]. chaos [Ano99-43, Col95-85, Col95-86, Han99-49]. Chapin [Hig95q]. charges [Mou98]. Charles [Zaj94c, Ano92-79]. cheaper [McC97s]. check [Mey98z]. checking [BK95, McC97k, MS90]. checklist [MM96]. checksum [PBGV90]. cheese [Ano97-60, Col96v]. Cheltenham [Hig95-35]. Chem [Nic99k]. Chen [CHT97]. Chen-Hwang [CHT97]. cherished [Mey97m]. Chernobyl [Mey99-37]. Cheyenne [Mey96h]. Chicago [BN92, Men92d]. child [Col95-83, Han99p]. China [Col95n, Han99k, Han99l, Han99c, Hei96a, Mey96-34, Nic99-31]. Chinese [Col95-151, Fol97, Han99-37, HTY96, San92]. chinks [Col96u]. chip [Ano96-28, Col95-98, Mey99-52]. chip-level [Mey99-52]. choices [OA90, Spa92]. chosen [BH93]. Chow [Ano96j]. Chris [Ano96-27]. Chrissikopoulos [Ano96i]. Christmas [Ano99-69]. Christopher [Ano97-76]. Church [Hig95-45]. CIBC [Mey96-45]. CICS [Ban91]. CID [CPS^{+}98]. CIMA [DMC92]. CIPFA [Ano96w]. Cipher [DG98, MV91, PC99]. CipherNET [Mey99-29]. ciphers [ABD93, Col95-36, Pre99, Hig95l]. circle [Col95-32]. circumvents [Nic99g]. Cisco [Han99m, Han99-95, McC98o]. CISS [MH92]. CISSP [Han98n]. cited [Nic99-48]. Citicorp [McC98-59]. citizens [Ano99p]. City [Ano92t, Ano99-45, Nic99l]. civil [Col95b]. civilian [Woo90b]. claims [Col95-76]. Clark [Hig95m]. Clark-Wilson [Hig95m]. Classification [BE95, Wis95, Coh97a, Coh97b, DMGK95, EHT96]. clean [McC98-38]. Cleansing [Mey97n]. Clearly [Ano96-28]. clears [Han99-74]. Client [Cor95a, Ano99-64, Lem96, Mey97-71, Mey98d, Nic99-67, Sym94]. client/server [Lem96, Sym94]. Clifford [Zaj90c]. Clinton [Ano96-62, Ano99q, Beq98d, Han99n, Han99o, Mey97-34]. clock [LC94]. clog [Mey97u]. cloning [Col95r]. closer [Mey97-88, VE98]. closes [Mey97f]. Closing [Ano96x, Har92]. Club [Han99-49]. CMA [Mey96-64]. Coast [Mey98j]. COBIT [Poo97]. Code [Ano97l, McC97c, Col95-48, Col96p, Han99-56, Han99-96, Hic95, LLO95, LS93, Mey97s, Mey98-101, SW93, SG94, Zaj94a, vdME98]. codes [Ano97-70, FGAR93, McC98y]. coding [BH92, BH93, CIRM95]. Cognitive [BPH^{+}97, ZH90]. Cohen [LT94]. cold [BGK92, Col95-90]. Colin [Ano96-108]. Collecting [vSG98]. College [Ano96d, Hig95u, Hig95-50]. collision [ABD93, JL96]. collision-free [JL96]. Color [Hig95x]. Columbia [Hig95x]. Combating [Mey99f, Han99p]. combats [Mey97-94]. come [Col95-39, Han99-87]. coming [Col95q]. comm [Mou98]. commandments [Lau98]. comment [Law92]. Comments [LT94]. Commerce [Han99-81, Hin98e, Ano96-86, Dos97, Han99-57, Jar99, Mey96-54, Mey98-90, Nic99s, Nic99r, Nic99-38, Rat99b, Woo97a, Woo97b, vS98]. Commercial [Har93, Mey96i, Pre99, Wat95]. Commission [Ano96-27, Beq98d, Pou98b]. commitment [McC98-45]. Committee [Ano91-32, Ano91-33, Ano91-35, Ano93-36, Ano93-38, Ano93-39, Ano95t, Ano95w, Ano95x, Ano96-58, Ano97-39, Ano97-45, Ano98t, Ano98u, Ano99-32, Ano99-33, Ano99-35, Ano90x, Ano90y, Ano90z, Ano90-27, Ano90-28, Ano90-29, Ano90-30, Ano91-31, Ano91-34, Ano91-36, Ano91-37, Ano91-38, Ano92-49, Ano92-50, Ano92-51, Ano92-52, Ano92-53, Ano92-54, Ano93-33, Ano93-34, Ano93-35, Ano93-37, Ano93-40, Ano94-28, Ano94-29, Ano94-30, Ano94-31, Ano94-32, Ano94-33, Ano94-34, Ano94-35, Ano95u, Ano95v, Ano95y, Ano95z, Ano95-27, Ano96-52, Ano96-53, Ano96-54, Ano96-55, Ano96-56, Ano96-57, Ano96-59, Ano97-40, Ano97-41, Ano97-42, Ano97-43, Ano97-44, Ano98o, Ano98p, Ano98q, Ano98r, Ano98s, Ano98v, Ano99-29, Ano99-30, Ano99-31, Ano99-34, Ano99-36, Mey97-34]. common [Ano97m, Har95]. Commonsense [Zaj94b]. communication [Ano96-30, Ano97-83, CKM91, Men90a, Men90c, Ram90, Ran92, Ran93]. Communications [Ano96-51, Hei94b, Hig95s, Ano96-97, Beq99c, Men90d, Mey97-68, MM96, Pou97, Sto95]. communities [McC98-35]. community [Lin97]. Companies [Mey98k, Mey98l, Mey99g, Han99-87, Lic98, McC98-63, Nic99-64]. company [Ano96-64, Ano96-87, Fin96, Mey97-75, Por96]. Compaq [Hig97v, Mey99h]. comparative [ELB93, Kan94]. comparison [Ano97n, Hop97]. competitive [Woo91b]. compile [Ano97o]. compile-time [Ano97o]. complementary [LE97]. Complex [Hig95-34, Col95-131, Mey96q]. complexity [Men90a]. compliance [Ano97-78, Pou99a, Tan95]. comply [Ano99f]. Component [Hig95j]. components [Mey96-77]. comprehensive [Ano97p, SSP90]. compression [BH92, BH93]. compromise [Han99d]. compromised [Han99s]. COMPSEC [Ano96y, EM90, Ano90-48, Ano91q, Ano92u, Ano97q, Ano98j, Ano98k, Ano99r, Ano99s, Hin98h, Hin99f]. compulsory [Dev99]. computation [VSU93]. Computer [AD90, Ano92-27, Ano92-29, Ano92a, Ano92-55, Ano92-64, Ano92-76, Ano92-81, Ano92-84, Ano92-90, Ano92-89, Ano96j, Ano96k, Ano96z, Ano96-30, Ano96-34, Ano96-80, Ano96-59, Ano97r, BE90, Buz99, CM92, Col95o, DFIR99, Eva91, Fåk91, FW99, Ha92, Han99-49, Hei94a, Hei96a, Hig95c, Hig95t, Hig95u, Hig95-27, Hig95y, Hig95-48, Hil91, HS94b, Ken93, Kow91c, Mey96k, Mey96l, Mey97o, Mul97a, OA90, Rob92, Tab90, Wol95, Zaj90b, Zaj92a, AM95, Ano90m, Ano90-46, Ano96d, Ano96-31, Ano96-66, Ano97a, Ano97s, BD95, Coh91a, Coh92b, Coh92a, Col95-67, Col95-71, Col95-78, Col95-85, Col95-86, Col95-129, Col95-156, DMC92, DMGK95, GDNC94, Han99-44, Hig93a, Hig97e, Hig97f, Hig97g, Hig97n, JJ96, JW90, JATC93, KJ95, Kov95, Kow91a, Kow91b, Lat90, Mad92b, Mag94, McC97h, McC98-29, MS90, Men91c]. computer [Men92a, Men92c, Mey96j, Mey97n, Mey98-30, Mey99f, Moe92, MvS94, Mul97c, Nic99i, Nic99y, NSS91, Oba93, PK91, RR91, She90, SG94, Tay95, Woo90a, Woo91a, WFZ92, Zaj90d, Zaj94b, vSvdM94, Ano92-86, Ano92-88]. Computer- [Hig95c]. computer-generated [Kov95]. Computer-Industry [Hig95-27]. Computer-related [Mul97a]. computerized [JL96, PvS91]. Computers [Ano93q, Hig91a, Hig95h, Ano92-98, Col95-46, Nic99-39, Nic99-50]. Computerworld [Ano92w, Ano92h, Ano92-46, Ano92-95, Ano97c, Ano97v, Ano97-35, Ano97-36, Ano97-73, Ano97-85, Col95-135, Ano92-48]. Computing [Ano92-78, Ano92-87, Ano97-55, Ano97-27, Ano97-47, Ano97-76, Sch91, AJ95b, All91, Ano92-92, Ano96a, Col95-117, Hig97r, Kan94, McC98x, VSU93, Ano92-65, Ano92-91]. Con [Ano97-35]. concept [HTY96]. concepts [AJ95c, Ano96-49, Ano97-78, Col95-109, SSP90, She90]. conceptual [Fin98a]. concern [Han99-80]. Concerns [McC97d, Men92a, Nic99m, Ano92-46, Beq98a, Beq99b, Col96e, Cro97, Cro98, Han99-63, Mey96-62, Mey96-91, Mey97j, Nel99, Nic99-51, Sch99a]. concludes [Han99z]. Conclusions [Mul97f]. conditions [BS94]. Conference [Ano90m, Ano90n, Ano92u, Car90a, EM90, Han99p, Hei94a, Hei95b, Hig93b, Hig97b, Ano96-60, CH94, CHT97, Hig93i, Hig93a, Hig97i, Kow91c]. Conferences [Hig95-48]. conferencing [Haw99]. Confessing [Mey97p]. Confessions [Jac90]. confidence [Col95l]. confidential [CKM91]. confidentiality [Ano92-87, TMO96, Wis90]. conflicts [Cie97]. confused [LL98]. confusion [McC98-73]. Congressional [Han99q]. connected [Gor97]. connections [Col95-72, Sch97a, Woo94]. Conputer [Hig95-48]. conscious [Mey97v]. consequences [Coh99]. considerations [Pre96]. consistent [Hop94]. consolidations [Mey98-58]. constitute [Woo97c]. constraints [LK92]. Constructing [Ano97s, Mey96m]. constructs [Ano97-78]. consultant [Ano96c, Jac90]. consulting [Mey98e]. consumer [Mey98-58]. consumer-friendly [Mey98-58]. consumers [Mad92a]. containment [AD90]. contains [Col95-29]. contemporary [Osb95]. content [Han99u]. contents [Hig95-48]. Contingency [Hig95z, Hin98a, Ano96-85, CM92, Klo90d, Nos99]. continue [Nic99w]. continuity [Ano96m, Dev99, Men94, She95b, SS95]. contract [Mey96-34]. Contracting [AJ97]. contracts [She97b, She97c]. contrary [Koe93]. Control [Ano92-55, Ano92-84, Ano96-104, Men91a, Wis90, Ano92-83, Ano96a, Ano96j, Ano97p, Ano97-75, BE95, BS94, Bor96, CSW94, Col95-130, Col95-161, DF91, GS91, Hin99a, HMCC95, HTY96, JCW91, Kim95, Klo90b, LE97, McC98-77, Mey96y, OA90, Par90, Par95a, SHW94, She90, Ste91, Woo97e, ZH90, Ano96k]. controlled [Hig95-27]. Controlling [Ano97t, Hol95a, Hol95c, Hol95b, Hol95d, Mey96n]. Controls [Kru95a, Kru95c, CL96, Cur95, Han99o, Han99-88, Hop94, Lyn91, Osb97a, Rat97, Woo97a]. conventional [BPH^{+}97, JJ96]. converge [Mey98-61]. convicted [Col95-83]. cookie [Mey97-91]. Cooperating [WP96]. coordinated [Coh96a]. cop [Col95-148]. Cope [Mey96o]. Copenhagen [Ano96-115]. Cops [Col95p]. Copy [Col95q]. copycat [Nic99-41]. copyright [Col95-150, Col95-157, Mey99r]. copyrightable [Col95-156]. Cordant [Hig95h, Hig95j, Hig95g]. core [Nic99n]. Cornelius [Hig95h]. corners [Mey98-44]. cornerstone [Dun95]. Corp [Hig95k, Hig95v, Hig95z]. Corporate [Ano96-27, Hin98b, Ano92-28, Ano96-99, Ano96-108, Boy97, Han99z, Han99-68, Han99-80, HS95b, Lin96, Pre91b]. Corporates [Mey96p]. Corporation [Ano96-107, Hig95b, Hig95o, Hig95-47, Hig95-45, Ano96-109, Fra96]. corporations [HS95b]. corral [Mey96-83]. correction [FGAR93]. correctly [Mey98-37]. Correctness [Hig95b, AZ95]. corruption [LT94, VPC93]. Corum [Ano96-102]. Cost [Hig95c, Zaj92b, Zaj94d, Coh91a, EOO95, Pea91, Woo90a]. Cost-effective [Hig95c]. Cost-effectiveness [Zaj92b]. Costly [Col95r]. could [Ano97-56, Han99d, Mey97-80, Mey99-33]. counsel [Col95-66]. counter [Gal95]. counterespionage [Han99-89]. counterfeit [Nic99f]. counterfeits [Eva92]. Countering [Ano96-28]. countermeasures [LE98, MH97]. countries [Nic99p]. course [Col95t]. court [Han99e]. Courtney [Col95-111]. cover [Moe92]. Covering [McC98p]. Crabb [Hig95-45]. crack [Ano92-77, Col95-48, Col96p, Man96]. cracked [Han99-70, Mey97z]. cracker [Nic99-48]. crackers [Han99-79]. Cracking [Col95s, Mey96q, Mey98-101]. cracks [Ano97c, Col95i, Dun95, Mey96t, Nic99h]. Crash [Col95t]. crashes [Ano99-48]. create [Mey97-73]. creates [Han99-80]. Creating [Kap98, Pre91b, Wil97]. credibility [Car90c]. credit [Han99-84, Mey97-35]. Cresson [Zaj94c]. Crime [Hig95y, Kov99b, War92, Ano97u, Ano99-67, BD95, Beq98a, Beq98c, Beq99a, Col99a, Col99b, Col95b, Col95w, Col95-67, Col95-129, DFIR99, Gor95d, McC97j, McC97y, Mey98-54, Moe92, Mul97b, She90, Som92, Tab90]. crimes [Beq98f, Woo90a, vSN92]. criminal [Mey98k]. criminalize [Möh95]. Criminological [She90]. crimoid [Nor92b]. Criteria [Hig95-29, Fåk91, Rih91, Str94, Ano97m]. Critical [Beq98d, Hig95-28, McC98f]. criticise [Mey98w]. criticized [Col95-76, Mey97-49]. Crocker [Ano96-65]. cross [Col95u, Col95-133, JATC93]. cross-cultural [JATC93]. Cry [McC98q]. Crypt [Mey97q, Ano96-114]. Cryptanalysis [PBGV90, Yen96, Col95-36]. Cryptic [Col95u]. Crypto [Ano96-29, Col95v, Mey95, Mey99i, Mey99-34, Ano96-38, Ano99t, Col95-131, McC97c, McC97e, Mey97f, Mey97-60, Mey97b, Mey98-74]. cryptochip [Nic99-61]. Cryptographic [Ano96-30, CGO93, CL96, Mey98m, BD92, GKM^{+}97, Han99-27, HL90, Hol95b, Hol95d, HMCC95, Lia94, McN95b, PMRZ99, PBGV90, Pre96, SHW94]. Cryptography [Ano97u, McN95a, Mey98n, Ano92-79, CDR99, Car90b, Han99t, McC98-30, Mey97-46, Mey97-53, Mey99-36, SPP98, Zaj94a]. Cryptologia [Ano97-72]. cryptosystem [LB91]. cryptosystems [Ano97-70]. CSI [Han99a]. CSI/FBI [Han99a]. CSIS [Col95w]. Cui [Hig95t]. cultural [JATC93]. culture [McC98-61, Wil97]. Cuppens [Hig95w]. cure [Mey96h]. Curing [Mey96r]. Current [Hig95a, Hig97b, Hug95b, PA95, Ano92-63, Bla93, Bov95]. Curry [Hig95v]. Curtis [Ano96-99]. curve [CDR99, Han99t, Hin99b]. custodians [Ano96-63]. cut [Ano99e, McC98-42, Nic99m]. cuts [Col95-132]. cutting [Mey98-44]. Cyber [Beq99a, Boy97, Hin98c, Ano99-67, Beq98c, Coh99, FW99, Han99n, McC98l, McC98-54, Nic99l, Nic99q, Nic99-62, McC98r]. cyber-attacker [McC98-54]. Cyber-crime [Beq99a, Ano99-67, Beq98c]. Cyber-IDs [McC98r]. cyber-plan [Han99n]. cyberattack [Han99-87]. Cyberattacks [Han99a, Han99-65]. cybercash [Ano97v]. Cyberclash [Ano97v]. Cybercop [Mey96s, Mey99j]. cybercops [Col95-135]. cybercrime [Ano96v, Han99-86]. cybercrooks [Col95-47]. cyberlaw [Nic99-36]. Cybernetic [Kow91b, Kow91c]. cyberpolice [Han99j]. cyberpunk [Lin94b]. cybersecurity [Mey96t]. cybershield [Mey96-99]. cybersmut [Col95-139]. Cyberspace [Col95x, Col95y, Col95-50, Col95-91, Mey98-93, Beq98a, Beq98b, Beq99c, Col95-150, Han99-71]. Cybersquatting [Han99r]. cybervirgins [Mey97-31]. cyberwalls [Han99-62]. cyberwar [Han99h, Han99c, Mey97-63]. CyberWarfare [CPS^{+}98]. Cylink [Col96d]. cyphers [RR91]. Cypherwiz [Mey96t].

D [Ano96i, Hig95z, SB90]. D-5000 [SB90]. Dad [Han99-36]. daemon [Coh96b]. Daily [Ano97-60]. Daly [Ano92w, Ano92-46, Ano92-95]. damage [Hin99d]. Dan [Ano97-56]. Danger [Ano96-31, Ano99-65]. dangerous [Ano92-47]. dangerously [Col95-69]. Daniel [Hig95u, Hig95-27]. Darlene [Hig95-48]. Darzentas [Ano96-66]. Data [Ano97w, BH92, Fra95, GTKK91, Hig95y, Hig97c, LR93, Men92e, Mey96u, Mey98-89, Nic99n, Pou98a, Pou98e, Pou99a, Rih94, RR91, She95a, VE98, All91, Ano92-80, Ano92-91, Ano97-55, Ano99f, Ano99-48, Car91, Col95-55, Col95-122, Col95-118, Col96w, Fin97b, Gil92, Han99-88, Han99-91, Han95c, HL90, MPRZ98, McC98-57, McC98-77, Men90a, Men90c, Men90d, Men94, Mey96-36, Mey96-40, Mey97-75, Mey97-71, Mey98f, Mey98-56, Mey98-83, Mey99w, Mey99-54, Nic99w]. data-protection [Mey99-54]. Database [HZS94, CRR92, GS91, HS94a, Laf90, Lun92b, OvS92, SJ91, SU93, Thu91, TR92, Thu94, Hig95w]. databases [Ano97o, Col95-109, HS95a, Hig95n, HDW97, LHB90, Wis90]. datagram [Ran93]. DATAKONTEXT [SB90]. DATAKONTEXT-VERLAG [SB90]. Datamation [Ano97-74, Ano97-86]. datasafe [Col95-70]. dataway [Col95p]. date [Ano96h]. David [Ano92-84, Ano96-75, Hig95x, Ano92-76, Ano92-89]. Davis [Hig95r, Hig95t, Hig95a]. day [Ano99-69]. DBMS [Laf91, Wis95]. DC [Hig95-28]. DE-light [Col96e]. dead [Col95-147, Mey97-74]. Deadly [Ano96-31]. deal [Ano92-48]. Dealing [Ano92-27, Ano96-32, Men91c]. dear [Ano92-91]. debate [McC98s, Mey98o]. debates [Han99-98]. debiting [Col95-28]. Debra [Hig95u]. DEC [Ano92-48]. December [Ano97v, Ano97-47, Ano90-45, Ano91-53, Ano92-75, Ano93-52, Ano97-79]. decentralization [Sto95]. deception [Coh98]. decision [Fin97b, Fin98b]. decision-making [Fin98b]. declared [Col95-155]. declines [Col95-136]. DECnet [Cly93]. Decoding [Mey96v]. defaces [Mey98-28]. DefCon [Col95-45]. defence [Ano96-61, Mey98-49]. defences [Coh97b, Coh99, McC97a, Mey97-48]. defenders [McC98-39]. defending [Cor97]. Defense [Coh92b, Hig95n, Han99-85]. Defense-in-depth [Coh92b]. defenses [Coh91a]. defensive [Mey96-103]. Deficiencies [Hit95]. Defining [Mol97]. definition [BE90, Coh92a, For98b]. delayed [McC98-51]. Demand [SG98, Ano99e, PK91]. democracy [Nic99-31]. demons [McC97g]. demonstrate [Ano92-45]. Denial [Mey96w, Sch99b]. Denmark [Ano96-115]. Denning [Hig95-28]. Dennis [Hig95-46]. dentify [Boo92]. Department [Ano96d, Ano96j, Ano96k, Ano96-30, Ano96-66, Ano96-80, Han99-85, Hig95n, McC98-67, Ano92-27, Ano92-90]. deploying [SGI^{+}99]. Deployment [Nic99o]. Dept [Hig95t]. depth [Coh92b]. Derivatives [Bol95]. Deriving [Ano97x]. describing [Col95-29]. Design [Ano96d, DMGK95, Hig95r, Hig95w, MD94, Ste91, Ano96-49, Mor96, Pre96, TOFS97, Win90, Woo90d, Woo93b, dK95]. designed [HS95b, Mey96-70, Mey97a]. desktops [Klo90a]. Desperate [Mey96x]. details [Hoa97]. Detect [Hig95u, MS90]. Detecting [Mey97r, HJS^{+}93, LT94, VPC93]. Detection [Hig95a, Hig95r, MH97, Ano96d, Ano96-66, Ano97a, Ano97-69, Coh97c, DKGS94, Eva91, Han99-60, Han99-90, LV92, Lun92a, Lun93, Mer90a, Mer90b, Mey98c, Mey98-37, Ric99, SD95, WP96, Hig95d, Hig95e]. detectors [Han99-62]. determine [CJB94]. Determining [GKKT92]. Deterred [Hig95y]. deterrence [Mul97b, Mul97d]. deterrent [MH97]. develop [Lem96]. Developers [Mey96y, Mey96z, Mey96-96]. Developing [Kel95, Nic99p, Rob99, Ano96-103, She96]. Development [Øln94, Ano96-101, Ano97-46, Kov97a, Mey98-69, Mit95]. Developments [Rob92, PA95, Pou98c]. device [LC94]. devices [Col95-147, Han99y, Mey98-65]. devil [Col95-111]. Dial [Col95z, Han95a, Mul95, Woo94]. Dial-up [Han95a, Woo94]. Didio [Ano92-92]. differences [Nic99n]. difficult [Han99-89, Mey96m]. difficult-to-guess [Mey96m]. Digital [Ano96-108, Bla93, Col95-27, Gan94, Han99-40, Hig95k, Hig95v, Law93a, Mey96-27, Mey97-76, Mey98o, Ano97-37, Ano97-72, CDR99, Col95-133, DFHR90, Hol95a, Hol95c, McC98-38, McC98-68, McC98-73, Mey96-89, Mey98-45, Mey98-35, Mey98-85, Pou98c]. digitalme [Han99-63]. diminished [Nic99-68]. dips [Nic99-60]. Direct [Col95-28]. direction [Nic99s]. Directions [Hig95a]. Directorate [Hig95-46]. directory [Hig96c, Hig97u, TO91]. DiRK [Ano96-34]. Dirty [Ano92v, McC98g, Mey96w]. disagree [Nic99x]. Disaster [Ano96-33, Nor93, Pea91, Ano92-48, Ano92-81, Ano96-85, Col95-145, Hil95, Men94, Mey96o, Mey97e, Tay95]. disasters [Ano97r, Hin98b, Hin98d]. disclose [Bou97, McC98w]. disconnect [Col95-91]. Discretionary [BS94]. discussion [Ano90-46, Ano90-47, Laf90]. discussion-into [Ano90-47]. disk [Fer92, Mey98-54]. disks [Col95-83]. displacement [Mul97b]. dispute [Nic99n, Nic99w]. Distributed [Ano96b, HS94a, Hig95q, Hig95t, SHW94, She92b, WP96, Ano96l, Ano97-77, Coh96a, Col90, Col95-55, Don93, GBD93, GS91, Gri97, Kru91, Mey96-82, Par90, SS93, Sym94, Thu91, TR92, VB91, Ano96-34]. distribution [Ano96-30, Ano96-29, Ano96-34, BB98, Hin99b, Mey95, TO91]. Do [Ano97-55, Bul99, Car90b, Mey97s, Mey98p, Ban95, CW94, For98b, Han98n, Han99-36, Hig95z, Mey97-93, Mey98-42, Nic99-70, Woo97c]. Do-it-yourself [Car90b]. Doc [Ano96h, Ano96-77]. doctor [Mey97o]. Document [Col95-29, Hig95s, LL98, Mey98-63]. documenting [Ano97z]. documents [EHT96, Rus93, Rus94]. dodge [Mey99-39]. Does [Ano96-35, Mey97-91]. Doing [Ano96-36, JvdV96]. dollars [Woo90a]. domain [Han99r, Mey97-80]. Don [Hig95n, Woo90a]. Donaldson [Hig95h]. Don’t [Ano99t, Mey98q, Han99-94]. door [Mey98-66, Nic99-30]. Doors [Mey99k, Mey96-78]. Dorothy [Hig95-28]. DOS-based [Coh91b]. DOS/Windows [Hig95h]. DOS/Windows-based [Hig95h]. Doug [Ano97-79]. Douglas [Hig95x]. down [Ano92-95, Ano97-86, Ano97-85, Col95i, Col96c, Hin99e, Mey97-84, Mey98d, Mey98-43, Mey99-39, Nic99-36]. downloaded [McC97k]. downsized [Ano92-86]. downsizing [Nor93]. DP [Pea91]. Dr [Hig95s, McC98-43, SB90]. Dr. [Ano96-76, Ano96-75]. draws [Han99-90]. dream [Col96r]. drive [Col95-41]. drivers [Hin98e]. drives [Mey99u]. Driving [Mey97t, Woo90b]. drop [Ano96-62, Col95-147, Mey97-74]. drug [Ano99-59, Ano92-64]. dry [BN92]. DSA [MNP96]. DSS [ZHM90]. Dublin [Ano96d]. Duchy [Nic99q]. due [Ano99-27]. Duress [Woo93a]. during [Jon97a, Jon97b]. Dutch [CvdRvZvD98, McC97e, SL96]. duties [Mey97-59]. duty [McC98t].

Dynamic [Lea95, HMCC95, JCW91, LE98, Lia94, MD94].

E-business [McC98-28]. E-comm [Mou98]. E-Commerce [Hin98e, Jar99, Mey98-90, Nic99s, Nic99r, Nic99-38, Han99-57, Mey96-54]. E-Mail [Mey96-49, Ano99x, Ano99u, Col95-30, Col96f, Hin99d, McC97t, McC98e, Mey97-38, Mey97u, Mey97-94, Mey98-94, Mey98r, Mey98-87, Nel99, Woo97d, Col95-100, Han99z, Mey96-71, Mey97-67, Mey97-92, Ano92-85]. E-signatures [Mey97v]. E. [Hig95-32]. each [Han99-36, Han99-47]. Earls [Ano97c]. Early [Hig95r]. ears [Nic99z]. ease [McC97d]. easier [McC97s]. easily [Nic99-33]. East [Eva92, Gli90]. easy [Mey96-71, Mey97-57, Mey97-65, Pou98d, ZH90, Zaj94c]. eavesdropping [Ano99p, Mad92a, Mey98-35]. eavesdropping-should [Mad92a]. Ebay [Nic99-48, Han99s]. eclipse [Ano99-43]. ECMA [Hig95k]. Economical [Col95-31]. Economist [Ano92-94]. EDI [Col95-116, Dos95, Men92f, Rat97, Rat98, Rat99a]. EDIFACT [Di 93, Øln93]. edifice [Beq98f]. EDIMED [Øln93]. edition [Zaj94b]. editor [And92, Ano90p, Ano90q, Ano90r, Ano90s, Ano90t, Ano90u, Ano90v, Ano90w, Ano91t, Ano91u, Ano91v, Ano91w, Ano91x, Ano91y, Ano91z, Ano91-27, Ano92-30, Ano92-31, Ano92-32, Ano92-33, Ano92-34, Ano92-35, Ano92-36, Ano92-37, Ano93s, Ano93t, Ano93u, Ano93v, Ano93w, Ano93x, Ano93y, Ano93z, Ano94l, Ano94m, Ano94n, Ano94o, Ano94p, Ano94q, Ano94r, Ano94s, Ano95l, Ano95m, Ano95n, Ano95o, Ano95p, Ano95r, Ano95s, Ano95-36, Ano96-41, Ano96-42, Ano96-43, Ano96-44, Ano96-45, Ano96-46, Ano96-47, Ano96-48, Ano97-28, Ano97-29, Ano97-30, Ano97-31, Ano97-32, Ano97-33, Ano99z, Ant92, Bow92, Coh92c, Cou94, CR93, Dav93, Han98a, Han98c, Han98d, Han98e, Han98f, Han98g, Han98h, Han99-29, Han99-30, Han99-31, Han99-32, Han99-33, Han99-34, Han99-35, Lin94a, Mey97-95, Mur91, Rob93, Zaj92c, Ano91s, Ano95q, Han98b]. Editorial [Ano90o, Ano91r, Ano92x, Ano93r, Ano94k, Ano95j, Ano96-37, Ano97y, Ano98l, Ano99v]. EDP [Ano96c]. EDS [Han99-67]. education [Ano96-80, Han99h, Mey99g]. educational [Ano97t]. Edward [Ano96z]. Effect [LV94, CPS^{+}98]. effective [Hig95c, Mey96g, Poo97]. effectiveness [PK98, Zaj92b, Zaj94d]. efficiency [Fin97a]. Efficient [BB98, Col95-32, Han99t, MPRZ98, PA93, Ano96i]. effort [Woo97c]. efforts [Mey98c]. Ekert [Ano92-79]. elaborate [McC97a]. elections [JL96, NSS91]. electromagnetic [Smu90]. Electronic [Ano92y, Car91, Col95-33, Dos97, Han99-81, Hig95s, Kov98a, Mol97, MM96, Nic99t, Rob91, vS98, Ano96-86, Ano97-77, Ano97-81, Col95-47, Col95-87, Col95-147, Hig95-33, McC97a, McC97r, McC98m, Mey96-58, Mey97-69, Mor99, Pou97, Rat99b, Rus93, Rus94, Woo97a, Woo97b]. Electronic-Internet [Kov98a]. electronic-mail [McC97r, Mey97-69]. electronically [McC98p]. ElGamal [CL94]. Eliminating [SJ92]. Elisa [Hig95n]. Elite [Ano96-38, Mey99-41]. elliptic [CDR99]. Elliptical [Han99t]. elsewhere [Col95-136]. embassy [Han99-37]. Embedded [Kem97a, Kem97b, TOFS97]. emergence [Mor99]. Emergency [Hig95s, Mey96j, RH91]. emerging [HS95b]. Employee [Beq98b, Beq99c]. employer [Nic99-37]. Employers [Beq99b]. enable [LE98]. enabled [Gor95d, Kru95a, Kru95c]. enables [Ano99b]. Enabling [SG98, Ano97-75, BP90]. encipherment [JvdV96]. encrypt [Col95-144]. encrypting [LL98]. Encryption [Bul99, Col96g, FGAR93, Hic97a, Hic97b, McC97f, McC98s, Men92e, Mey98s, Nic99u, She95a, Ano92-67, Ano96-79, Ano96-62, Ano97-86, Col96d, Col96h, Doc97, GDNC94, Han99e, Han99-62, Hig91b, Hig91d, Hig95-28, Hig97c, Jar99, MPRZ98, McC98a, McC98-51, McC98-69, McC98-77, Mey96v, Mey96-61, Mey96-71, Mey96-66, Mey96-93, Mey97-34, Mey97-43, Mey97-49, Mey97-51, Mey97-65, Mey98o, Mey98u, Mey98-98, Mey98-78, Mey99o, Mey99t, Mey99-46, Nic99-69, Nic99z, PRZ99, Pou98c, RR91]. encryption-technology [Mey97-49]. end [War92]. endangered [Col95-99]. endangering [Rot92]. Enemy [Ano92z, Ano99w, Bac97b, Hig96b, Mey96-50, Mey99v]. Enforcement [Ano92-64, Hig95m, San92]. Enforcing [BdAR99, Bor96]. engage [Han99c]. engineer [Col95s]. Engineering [Ano96j, Hig95k, Gor95c, Hig95-47]. enhance [HS95b]. enhanced [Ano97-69, Col95-134, PG94]. Enhancement [Col95-162]. enough [Ano96-110, Ano97-74, Col96m, Han99c, McC98-46, Mey96-46, Mey98-40, Mey98-79]. ensure [Mey98-31]. Ensuring [Tan95]. enter [Mey98o]. enterprise [Ano96-103, Ano97-83, Col95-117, Hig95g, McC98o, Mey98-76, Mey99d, Mey99-44, Sha94, She96]. enters [Mad98, McC98-59]. Entropy [DG98]. entrust [Mey98-75]. environment [Ano92-86, Col95-124, Don93, Gor97, Kov97d, PdL91, Pre91a, Rus93, Sch97b, vSG99]. environments [Kan94, Mey96-43, Sta93]. Ephraim [Ano92-96]. epidemic [Ano96f, Han99z]. epidemics [Nic99-49]. Epstein [Hig95h, Hig95j, Hig95g]. Equipment [Hig95k, Hig95v]. era [Bol90, Fra95, Mey97-53]. error [FGAR93, WB93]. ESA [Paa91]. eSafe [Mey98t]. escalate [Col95-154]. Escrow [Nic99v, Han99-81, Hig95-28, Nic99c]. escrowing [McN95b]. espionage [Nic99v, Som94]. Essential [Woo97a, Pea91]. establishes [Mey97-47]. Establishing [Kov96, Kov98b, Mil96]. eternal [CL96]. Ethical [Bol90, Bol91]. EU [Han99u, Han99-40, Han99-98, Nic99w]. EU-US [Nic99w]. Eugene [Hig95i, Hig95d, Hig95q, Hig95e, Hig95-32]. Europe [Ano97u, Han99-98, Hom90, Mey99r, Nic99x]. European [Hig95-27, CRR92, Han99v, Han99-77, Mey99l, Pou97, Pou98b, Rob92, She91]. Europeans [Mey98u]. Evaluating [Hig91b]. Evaluation [Cur95, Hig95-29, Ano92-63, Ano97m, CRR92, Fåk91, GGG^{+}94, Hig95-35, Llo90, Rat97, Rih91, Rob92, Str94]. evaluations [Hig95k]. Evaluting [Hig95l]. Evans [Ano92-76, Ano92-89]. event [Han99-85]. ever [Thi91]. everywhere [Bac97b]. Evidence [Ano92-27, Ano96z, Kov95]. evil [Col95-112, Han99j]. evolution [Coh93, Mul97a]. evolving [Ano96-116, Win90]. ex [Ano99p]. ex-spy [Ano99p]. examining [Men92e]. example [vdME98]. examples [Woo93b]. Excel [Mey96-52, Mey99s]. exchange [BP90, DFHR90, Mey98-81]. excuse [Ano99d]. executing [Ano96-107]. Executive [ILM94]. Exercising [RH91]. Exhibitors [Ano95k, Ano98m, Ano90-48, Ano96y, Ano97q, Ano99r]. existence [LT94, VPC93]. existent [Nic99-64]. existing [VE98]. Exorcising [McC97g]. expands [Mey98y]. expansion [JvdV96, RR91]. expensive [Col95-34]. Experience [Jon97a, Beq99a, Jon97b, Tay95]. Expert [Col96h, Ano99t, Col95-125, KJ95]. Experts [Han99w, Mey98v, Mey98w]. explained [Sch98]. Explanation [Law93a]. exploitation [Mey97-82]. Explorer [Mey97-39]. exponential [Hin99c]. expor [Han99e]. Export [Pre96, Mey96-93, Mey98s, Mey99-39]. Exporting [Hig95-29]. exposed [McC98v]. exposes [Mey97-27, Nic99n]. exposure [Han99-85]. Extended [PG94]. Extending [Lyn91, VC96]. extends [Mey97-79]. extensions [WFZ92]. external [FW95]. extinguishes [Col95-93]. extortion [Boy97]. ey [Hig95h]. eye [Ano96-40, Boo92, McC98-40, Nic99-58]. eye-dentify [Boo92]. eyes [Mey96-70, Nic99-27].

F [Hig95w]. F.R.G [SB90]. FAA [McC98-55]. face [Ano92-66, Dav96, McC97j, Mey96-96, Mey98m, Mey98k, Nic99p]. Faced [Col95-35]. faces [Ano99-44, Col95z, Nic99r]. facilities [Han95a]. facility [Pre96]. Facing [Smi95a, Beq99c]. FACM [Mey97-96]. Factors [Ano96-39, Hig95-28]. facts [HP90a]. fail [Ano92-78, Ano99f, Nic99-63]. failing [Mey99-43]. Failings [Beq98f, Beq98d, Mey98v]. failsafe [Mey98-27]. failure [Ano92-27]. Fairfax [Hig95n]. Fairwall [HH99]. fall [McC98g, Mey96-94]. Falls [Hig95-45, Mey98-80]. false [Ano92a, Mey98x, Nic99-68]. Falun [Han99l]. famous [Hig97f]. fancies [HP90a]. FAQ [Mey97-27]. Farmer [Ano96h, Ano96-77]. fast [Col95-36, Han99-73, PBGV90, Ano92-28, Faw92a, Mey97w]. faults [McC98-55]. FBI [Han99a, Mey97p, Nic99y, Rot92]. Fear [Mey96-28, Mey97g]. fears [Ano99e, Col95-93, Mey99y]. feature [Pea91]. features [Ano92w, Col95-134, RR91]. February [Ano97c, Ano97-55, Ano97-27, Ano97-35, Ano97-36, Ano97-73, Ano97-76, Ano90-40, Ano91-48, Ano93-46, Ano94-44, Ano97-56]. fed [Mey98-98]. federal [Col95-43, Col95-152, Mey98-92]. federated [Thu94]. feds [Han99-86, Ano96-40, Col95-37, Col95-38, Han99-42]. Feeling [Mey96-29]. felony [Col95z]. few [Col95-46, Col96u]. Fiber [CKM91]. Fiber-optic [CKM91]. FICS [Mey97-96]. field [Pou98c]. Fifteen [Woo90b]. Fifty [Woo94]. fight [Ano99-49, Hor92]. fighters [Kov99a]. Fighting [McC97h, Col95w, Mey97o]. fights [Col95-67, Mey96-31]. file [FGAR93, MS90, Mey99o, TMO96]. files [Ha92]. filing [Col95-44]. filter [LLO95]. final [Mey98-71, MH97]. Finalists [Han99x]. finally [Han99-50, Mey98-78]. Financial [Hig95c, Gal95]. find [Col95-89, Han99-89]. Finding [Som92]. findings [She90]. finds [Ano97-60, Col95w, Col96l, Han99q, Mey98-96, Mey98-70]. fine [Ano96k]. fine-grained [Ano96k]. finger [Mey98-65]. fingered [Nic99b]. Fingerprint [Han99y, Mey99m, Col95-148]. fingerprints [Mey99h]. fingers [Mey98-42]. fingertips [McC98r]. Finjan [McC98t]. fire [Col95-39, Col95-133, Han99-90, Hil91, McC98l, Men92b, Mey97-36, Mey97-90, Nic99a]. Firestorm [Mey96-30]. Firev [Hig95o]. Firewall [Col96i, Col96j, Col96k, Mey96-31, Mey98y, Mey98z, Ano96-76, Ano97-27, Col95d, Col95-56, Col95-90, Col95-138, Col96x, McC98k, McC98u, McC98-37, McC98-52, Mey96y, Mey96-85, Mey96-73, Mey97-33, Mey98p, Rob99, Sch95a]. Firewalls [Col95-39, Mey96-32, Mey96-33, Mey98-27, Mey98-91, Ano97-47, Col95k, Han99-62, McC98-46, Mey96b, Mey96-30, Mey96-50, Mey96-90, Mey97-62, Nac97, Nac98, She95c]. Firm [Mey96-34]. Firms [Ano99x, McC98v, McC98w, Ano92t, Ano92-81, Ano99-65, Han99b, McC98x, Nic99-63]. firmware [McC98-52]. First [Ano96-78, Ano96-85, Han99z, Mey98-92, Pou97, Han99-93, McC97l, Mey97-56, Mey97-60]. fit [Som92]. Five [Mey97x]. fix [Ano99-28]. fixed [BH92]. fixed-model [BH92]. fixes [Col95-94]. FL [Ano96j]. Flak [Mey99n]. flank [Mey99-44]. flaw [Ano99-28, Col95-94, Col95-126, McC98d, Mor96]. flawed [Mey97z, Nic99s]. flaws [Ano99-59, Col96h, Han99q, LR92, LV94, Mey97-39, Mey99-51]. fledgling [Col95-149]. flies [Mey99n]. flood [BN92, Men92d, Mey97-52]. Florida [Ano96j]. flow [Ano97o, Gor97, VI97]. flying [Mey98-97]. focus [Ano99y, Col95g]. fodder [Col95-89]. folk [Par97a]. force [Han99-60]. forced [McC98w]. forces [Woo90b]. Forensic [McC98x, Ano96z]. forensics [SW93]. Formal [Hig95t, Ano97z, BE94, Coh92a, Han99h]. Former [McC97i]. forms [Ano97-57, MV91]. formulas [Sch98]. Fort [Hig95-34, Ano97-36]. Fortifyingyour [Ano97-27]. Forum [Col95-40]. forward [Dos97, Han99-87]. found [Col95-73, Col95-126, Mey98k, Nic99-35]. foundation [Mey97-59]. Four [Mey97y]. Framework [SG98, Ano96-99, Ano96-111, Ano97m, Bol91, ELB93, Fin98a, Har95, Kan94, Mey98-69, Mor99, Pfl91]. France [Ano99y, Han99-27, Hig95w]. Frangos [Ano96-111]. Frank [Ano96-87, Ano96-109, Hig95a]. Fraud [Ano92-29, Ano92-90, Hor92, Mer90a, Mey98-93, Ano96-104, Ano96-116, Ano99w, Col95r, Col95-28, Col95-41, Col95-71, Col95-148, DMC92, Eva91, Gal95, Han99-57, Han99-66, Hoa97, McC98x, McC98-59, Mer90b, Mey96-86, Mey97-56, Mey97-82, Mey99a, Mul95, Mul97a, MH97]. fraud-beater [Col95-148]. fray [Mey99-27]. Frederick [Hig95h]. Frederiksberg [Ano96-115]. Free [Han99-28, Mey99o, Han99-93, JL96, Mey96-51]. Free-mail [Han99-28]. freebie [Mey98-51]. Freedom [Kov99a, Woo91a]. French [Hig95k, Ano92-29, McC97e]. Frequency [Mul97c, Mey98-95]. Fresh [Col95-41]. fret [Col95-62]. Friday [HP90a]. friendly [Mey98-58]. Frizzell [Hig95s]. front [Han99-41, Mey99-42]. frontiers [War92]. Ft [Hig95n]. FTC [Han99-36]. Fujitsu [Hig95k]. full [Col96f]. fun [Wol95]. function [Ass92, Osb97b, Osb98]. functions [BD92, Man96]. Fundamental [Ano96-49, SG98, Men91a]. funding [Ano92-78]. funds [Nic99m]. furious [Han99-44]. Further [Pou98c]. Future [Han95b, Hig95a, Mey97z, Ano97-34, Bir97, Hug95b, Kru95a, Kru95c, Mey96w, Mey98m]. futures [She92a]. fuzzy [dRE96]. FX [Ano96-28].

G [Ano96h, Ano96-77]. Gabillon [Hig95w]. gain [Mey96-102, Mey97-60, Mey98-95]. gained [Jon97a, Jon97b]. Gainesville [Ano96j]. gains [Col95-153, Col96g]. Gaithersburg [Hig95-49]. Gale [Hig95c]. game [Dav90b]. games [Ano92-93, Ano97-35]. Gang [Eva92, McC98f, Smi92]. Gannon [Ano92-29]. gap [Mey96-77, Mey97i]. Gaping [Col95-42]. Garfinkel [Hig92a]. Gary [Hig95h, Hig95j, Ano92h, Ano92-64, Ano96-118, Col95-135]. Gaudin [Ano97-35]. Gene [Ano96g, Ano96-112, Hig92a]. general [JL96]. Generalized [MH92, PA93]. Generally [Ano96-50, Kru96]. Generally-Accepted [Ano96-50, Kru96]. generated [Kov95]. Generating [PMRZ99]. generation [HL90, Lia94, McC98b, vS96]. Generators [DG98, LV94]. geometric [CJC98, TCWH95]. George [Hig95n, Hig95-45]. Georgetown [Hig95-28]. Georgiadis [Ano96l]. German [Faw92b, Han99-70, Mey98-28]. Germany [Gli90, Hig95-27]. Get [Nic99-28, Col95-106, Han99j, McC97t, McC98-44, Mey96-55, Mey96-81, Mey97-88, Mey98-90, Mey98-94]. gets [Col96i, Col96q, Han99h, Han99i, Han99r, Han99s, Han99-50, McC97s, McC98-36, Mey97d, Mey97-56, Mey99p, Mey99-34, Nic99-59]. Getting [Han99-42, Mey96-35, Mey96-36, Pla93, Han99-55, Lai97, Mey96-90, Mey98r]. giant [McC98-64, Mey98-76]. Gill [Hig95v]. Gilles [Ano92-79]. Gilooly [Ano92-66]. Gilooly/Bob [Ano92-66]. Girard [Ano97-85]. girds [Col95c]. give [Han99-93, Mey97-28]. glass [Col96f]. Glenn [Hig95-27]. Glenwood [Hig95-27]. glides [Han99k]. Global [Kap98, McN95a, Nic99-40]. globally [Gor97]. Glos [Hig95-35]. Gmbh [SB90]. go [Ano99t, Han99-67, Hin99b, Hin99e]. goals [Law92]. goes [Bac97a, Han98i, McC97x, Mey98-86, Mey99-39]. going [Han99-45]. gone [Mul97b, Mul97c, Mul97d, Mul97e, Mul97f, Mul97g]. Gong [Han99l]. Good [Han98j, McC98z, Mey97-64, Col95-112, Han99j, Hig95z, Mey98-88]. good-guy [Mey98-88]. gory [Hoa97]. GOST [MNP96]. Gotcha [Ano97-35]. governance [Ano96-27, Hin98b, Lin96]. Government [Ano92-87, Ano99-27, Han99-37, Mad92a, Mad92b, Nic99-29, Han99-90, Han99-82, Han99-87]. government-owned [Han99-90]. Government-sponsored [Mad92b]. grades [Mey97-30]. grained [Ano96k]. Gray [Hig95g]. Graz [Ano96-51]. Great [BN92, Men92d, Col95k, Han99-28]. greater [Mey99-31]. Greece [Ano96i, Ano96l, Ano96l, Ano96-66, Ano96-111]. Greek [GTKK91]. Greensboro [Hig95m]. Gregory [Hig95t]. Gritzalis [Ano96l]. grocery [McC98-48]. Grossman [Hig95h, Hig95j, Hig95g]. Group [Col95-43, AAL^{+}93, Ano96-30, Ass92, Col95-67, Han99t, Han99-70, Mey99-36, Ano96-102]. Groups [Küh99, BN92, Mey98-52]. groupware [Woo93b]. Grover [Hig95s]. growing [Han99-73, Han99-80, Mey98-97]. grows [McC98-49, Nic99-60]. growth [Ano97u, Col95-141, Hin98e, Hin99c, Jar99].

GSSP [Ano96-50, Kru96, Ozi98]. GTE [Col96d, McC98-27]. Guard [Mey98j, Mey99-44]. guards [Col95-46]. guess [Mey96m]. guessing [BPH^{+}97]. guidance [Poo97]. Guide [Ano91-28, Ano91-29, Ano91-30, Ano92-38, Ano92-39, Ano92-40, Ano92-41, Ano92-42, Ano92-43, Ano93q, Ano93-27, Ano93-28, Ano93-29, Ano93-30, Ano93-31, Ano93-32, Ano94t, Ano94u, Ano94v, Ano94w, Ano94x, Ano94y, Ano94z, Ano94-27, Ano92-92, Beq98c, Zaj94b]. guidelines [Han99-91]. guilty [Mey98k]. guy [Mey98-88]. guys [Sch95b].

H [Ano92h, Ano92-64, Ano92-79, Ano96j, Ano96k, Ano97-72, Hig95i, Hig95e]. H. [Ano96-51]. H5a4BRTES7S [Col95a]. hack [Ano97-85, Han99-38, Mey98j]. hacked [Ano99-27, Nic99-29]. Hacker [Col95-44, Col95-45, Col95-46, Col95-47, Mey96-37, Mey97-27, Mey97-28, Mey98-28, Mey98-29, Mey99p, Mey99q, Ano92-77, Ano99b, Col95c, Col95-49, Col95-78, Col95-105, Han99-83, Lin97, McC98g, McC98-59, Mey97-91, Mey98l, Mey98-49, Mey99y, Mey99-31, Mey99-49, Mey99-56, Mey97-30]. Hacker-trackers [Col95-47]. Hackers [Ano92-45, Ano92-44, Col95-48, Col95-50, Han99-37, Kov99a, Mey97-29, Mey97-31, Mey98-94, Nic99-30, Ano99-45, Col95-53, Col95-80, Col95-135, Col95-149, Col95-155, Col96k, Han99w, Han99-28, Han99-48, McC98v, Mey98a, Mey96k, Mey97-48, Mey97-57, Mey98-88, Mey98-97, Mey98-102, Mul97b, Mul97c, Mul97d, Mul97e, Mul97f, Mul97g, Nic99c, Nic99y, Men91b]. Hacking [Möh95, Nic99-31, Ano99d, FW99, Mey98-103]. hackles [McC98-39]. Hacks [Mey98-95, Han99-78, McC98-39, Nic99-48]. Hains [Ano92-84]. Hal [Hig95h]. half [Col96f, Woo90a]. Halme [Hig95f]. halting [Col95-80]. Hamilton [Hig95s]. Hamlyn [Ano96-101]. Handle [Mey97-32]. hands [Col95-138]. hangs [Ano99-64]. happen [Hig95z]. harassment [Mou98]. hard [Mey97-65, Nic99-34]. Hardening [Han99-38]. harder [Man96]. Hardwall [Rob99]. hardware [RR91]. Hardwick [Col95-58]. Hardy [Ano96-118]. Harmonized [Mey99r, Rih91]. harmonizing [Ano97z]. Harold [Mey97-96]. Hartmut [SB90]. Haruki [Hig95k]. Hash [PRZ99, BD92, PRK98]. Hash-based [PRZ99]. Hassler [Ano96-51]. hatches [Col96c]. haul [Men90a]. haven [Gli90]. havoc [McC98q]. Hazards [Mey96-38, Ano96-115, EOO95]. head [Mey96-45]. headache [Mey96r]. heads [Mey99-35]. health [Ano97-77, Hin99d]. healthcare [Gri97]. healthchecks [Ano96-98]. heartburn [Col95-105]. heat [Col96k, Mey98-91]. heats [McC98s]. heaven [Col95-46]. Heckman [Hig95t]. hedges [Mey98-32]. Heel [McC98b]. Heffernan [Ano92j]. Hellenic [Ano96l]. Help [Col95-51, Ass92, Han99-86, Mey96-90, Mey99-44]. helps [McC98x]. Henry [Ano96-86]. here [Ano90-46, Ano92-97, Col95-133, Mey97-76]. heroes [Ano92-44]. Herrigel [Hig95k]. heterogeneous [GS91, Kan94]. hiccup [McC98-55]. hidden [McC97j]. hierarchical [Ano96-51]. hierarchy [CJB94, HMCC95, SHW94]. Higgins [Hig95l]. High [Ano92-46, Beq98d, Mag94, McC97j, Ano96-98, BN92, Coh91c, Col95w, Col95-31, Col96l, Han99-82, Llo90, McC97i, McC98x, Men92b, Mey98-80, Ran93, She95a, TMO96]. high-integrity [Coh91c]. High-level [Mag94]. high-security [Col95-31]. high-speed [TMO96]. High-tech [Ano92-46, Beq98d, McC97j, Col95w, McC97i, McC98x]. higher [AM95]. Highland [Mey97-96]. highlights [Ano92-46, Mey96c]. highway [Col95-152, Mey97t]. hijack [McC98-78]. Hiles [Ano96-33]. Hill [Hig95u, Zaj94b, McC97f]. Historical [Hig97d]. history [Hig97e, Hig97f, Hig97g, Hin98h, Mey97-33]. hit [Mey97-31, Nic99-34]. Hitchings [Ano92-83]. hits [Han99-68, Mey96-52, Nic99-50]. Hobby [Ano92a, Ano92-81]. hold [Mey96-100, Nic99-54, Zaj90a]. Holding [Ano97-36]. Holdings [Ano96-86]. holds [Han99u]. hole [Mey99h, Mey99s, Nic99-45]. Holes [Col95-52, Col95-42, Col95-61, McC98-31, McC98-50, Mey98-73, Rei96]. holistic [Ano97-83]. Holloran [Hig95-48]. Holm [Hig95-34]. holograms [Eva92]. Homewood [Woo90a]. Homeworking [Pou98d]. hooking [Mey96-38]. Hopkinson [Ano92-55]. horror [Ano96-76]. hospitals [Ha92]. host [Ano96a, Ano96x, Mey99-30, Rei96]. host-based [Ano96x, Rei96]. host-centric [Ano96a]. hosting [Han99f, Mey98-33]. Hot [Col95-53, Hin98d, Hin99b, McC97f]. Hotmail [Ano99-28]. House [Col95-157, Mey96-99, Mey97-34]. Howard [Hig95-34]. HP [CRR92, Col96l, Mey96-97]. HP-UX [CRR92]. HTML [Mey99s]. Hugo [CM92]. Human [WB93, Fåk91]. hunt [Col95-47]. hurdle [Han99-74]. hurdles [Nic99o]. Hurford [Ano96-27]. Hurricane [CM92, Ano92h]. hurt [Han99-94]. Hutchison [Ano96-30]. Hutte [SB90]. Hwang [CHT97]. hype [McC98-63].

I-team [Mey96-41]. I-way [Kov99b]. I. [Ano96d]. I/O [RR91]. IBM [Ano92-48, Ano96-29, Ano96-49, Ban91, Col95-56, McC98-28, Mey98-32, Mey98-33]. IBM/DEC [Ano92-48]. icebergs [Hin98d]. ID [Han99-70, Mey97-76, TO91]. ID-related [TO91]. ideas [Ano92-78]. identification [Bon98, CN98]. identify [Ass92]. Identifying [Hig95i, KS97]. Identity [CH94, Kan94, Mey97-35, CHT97, Fåk91, Kim95, Pre91a]. Identity-based [CH94, CHT97]. IDs [DD92, McC98r]. IE5 [Nic99g]. IETF [Han99-60]. if [Han99c]. IFIP [Mey93, Ano90x, Ano90y, Ano90z, Ano90-27, Ano90-28, Ano90-29, Ano90-30, Ano91-31, Ano91-32, Ano91-33, Ano91-34, Ano91-35, Ano91-36, Ano91-37, Ano91-38, Ano91-39, Ano91-40, Ano92-49, Ano92-50, Ano92-51, Ano92-52, Ano92-53, Ano92-54, Ano93-33, Ano93-34, Ano93-35, Ano93-36, Ano93-37, Ano93-38, Ano93-39, Ano93-40, Ano94-28, Ano94-29, Ano94-30, Ano94-31, Ano94-32, Ano94-33, Ano94-34, Ano94-35, Ano95t, Ano95u, Ano95v, Ano95w, Ano95x, Ano95y, Ano95z, Ano95-27, Ano96-60, Ano96-52, Ano96-53, Ano96-54, Ano96-55, Ano96-56, Ano96-57, Ano96-59, Ano96-58, Ano97-39, Ano97-40, Ano97-41, Ano97-42, Ano97-43, Ano97-45, Ano98o, Ano98p, Ano98q, Ano98r, Ano98s, Ano98t, Ano98u, Ano98v, Ano98n, Ano99-29, Ano99-30, Ano99-31, Ano99-32, Ano99-33, Ano99-34, Ano99-35, Ano99-36, Hig92d, Hig92e, Hig92f, Hig92g, Hig92h]. IFIP [Hig92b, Hig92i, Hig92c, Hig93i, Hig93c, Hig93d, Hig93e, Hig93f, Hig93g, Hig93h, Hig94a, Hig95-30, Hig95-31, Hig96a, Hig97i, Hig97a, Hig97h, Ano97-44]. IFIP/Sec [Hig93i]. IFIP/SEC’96 [Ano96-60]. IFIP/SEC’97 [Hig97i]. IFIP/TC [Ano91-39]. IFIP/WG [Ano91-40]. ignored [Col95-30]. II [Thu91]. II. [VC96]. IIA [DMC92]. III [Han99-70, Hig95h, Nic99-51, TR92]. illustrated [Men90b]. IMB [Col95-57]. Immigration [McC98-29]. Immunizing [Mey98-34]. impact [Han95c, Hig92r]. impacts [She91]. implementable [BR96]. implementation [Coh91b, HMCC95, Laf90, LB91, LL98, Par95b, Poo97, She95a, dV92]. implementations [PA93]. implemented [Fin97a]. Implementing [Hig95m]. implications [Lat90, Rat97, Woo95e]. Implicit [Rat99a]. importance [Ano92-55, Men94, Par97e]. important [Ano96-117]. improve [Ban91]. Improved [LE97, BE95]. improves [Klo90d]. Improving [BK95, Oba93, She90]. incidents [CvdRvZvD98]. include [MvS94]. incorporating [LK92]. increase [Col95b, Mey99a]. increases [Nic99-62]. increasing [Ano92-85, Han99-66]. index [Ano90i, Ano90-50, Ano91i, Ano91-57, Ano93h, Ano93-58, Ano94b, Ano94-52, Ano95a, Ano95-45, Ano96e, Ano96-113, Ano97b, Ano97-80, Ano98a, Ano98-40, Ano99c, Ano99-61]. Industrial [Col95-59, Som94]. Industrial-strength [Col95-59]. Industry [Ano96-62, Hig95m, Hig95-27, Ano99-46, Nic99-34, Ano92-27, Ano96w]. inexpensive [Nic99z]. infect [Ano92-98]. infected [Col95f]. infection [Mey99l]. infections [Mey99c]. inference [HDW97]. infestation [Han99-58]. influences [Rat98]. info [Ano97-58, Col95-37]. info-system [Ano97-58]. Information [Ano96j, Ano96-51, Ano96-63, Ano96-64, Ano96-110, Ass92, Coh97a, Coh97b, Cor97, Eva96, Fin97a, Gor97, Han99-39, Han99-40, Han99-41, Hig95c, Hig95m, Hig95u, Hig95-27, Hig97b, Kov97c, Kov97e, Kov98b, Mey96-43, Por96, Sta93, TOFS97, vS96, AJ95a, Ano91-40, Ano96-86, Ano96-97, Ano96-115, Ano97m, Ano97p, Ano97l, Ano97o, Ano97x, Ano97z, Ano97-46, Ano97-57, Ano97-83, BE94, Bol90, Bol91, Car90c, CPS^{+}98, Coh98, Col95-152, Dun95, EOO95, Fin96, Fin97b, Fin98a, Fin98b, Fra95, GGG^{+}94, GTKK91, GKKT92, Gri97, Han99-43, Han99-68, Han99-80, Har93, HLY92, Hei95a, Hig93q, Hig95-47, Hit95, ILM94, Kru95b, Lia94, Lin95, Lin96, McN95a, Men90b, Men91a, Mer90b, Mey96g, Mey97t, Par97b, Par97c, Pep97, She91, Smu90, TO91, Woo90b, Woo90c, Woo90d, Woo91b, Woo93a, WB93, Woo93b, WS95, Woo95a, Zaj94b, Ano92-45]. Information [Zaj94c]. information-intensive [Car90c]. Information-Security [Hig95c]. Informationstechnik [SB90]. infosec [Woo95c, Woo95d, Par97a]. InfraGard [Han99-42]. Infrastructure [Ano96-82, Beq98d, Han99-90, McN95a, Woo95e]. initial [Ano96-29, Mey95, Pou99a]. initialization [Ano96-29, Mey95, Mey97b]. initiative [Mey97-83]. inmates [Boo92]. Inner [Mey96-44]. innocent [Col95-71]. Inquirer [Ano97-56]. inquiry [McC98-57]. ins [Col95c, McC98-81, Mey99p]. insecure [Men90d, TCHS96]. insecurity [CIRM95, Mey97-82]. insiders [Nic99b]. installation [Pea91]. Institute [Ano96-34, Ano96-51, Ano96-80, Col95-162, Hig95m, Hig95-49]. institution [Ano97t]. insurance [Col95-92, Mey98-100]. insurer [Mey98l, Mey98-46]. integrated [Kan94, Rus93]. Integrating [Ano97-46]. Integration [Hig95o, Mey98-60]. Integrity [ILM94, SJ91, SU93, AAL^{+}93, Coh91c, Col95-122, CW94, HLY92, Hig95n, Llo90, MS90, Men90b, O’S91]. intellectual [Ano99-49]. Intelligence [Hig95a]. Intelligent [Ano96-65, Bai97, Ano97-75, vS98]. intensive [Car90c]. Intention [Ano96-66]. intentions [Ano96-66]. inter [Hug99, Smi95b, TMO96]. inter-operability [Hug99]. inter-operable [Smi95b]. inter-PC [TMO96]. intercepting [VE98]. interception [Mey97-57]. interchange [Car91, Rih94]. interconnected [Ano96l]. interconnection [Ram90]. interest [Hig91e, Hig96e, Hig97l, Hig97m, McC98-35, Mey98-48]. interesting [For98a]. Interfaces [CGO93]. interior [Col95n]. Internal [Ano97-47, FW95, Han99-43, Han99l, Mey97-75, Mey99-46]. International [Ano90-31, Ano90-32, Ano90-33, Ano90-37, Ano90-34, Ano90-35, Ano90-36, Ano90-48, Ano90-49, Ano91q, Ano91-42, Ano91-43, Ano91-44, Ano91-45, Ano91-46, Ano92-56, Ano92-57, Ano92-58, Ano92-59, Ano92-60, Ano92-61, Ano92-82, Ano93-41, Ano93-42, Ano93-43, Ano93-44, Ano93-55, Ano93-56, Ano93-57, Ano94-36, Ano94-37, Ano94-38, Ano94-39, Ano94-40, Ano94-41, Ano94-42, Ano94-51, Ano95-28, Ano95-31, Ano95-33, Ano95-34, Ano96-34, Ano96-67, Ano96-68, Ano96-70, Ano96-71, Ano96-72, Ano96-73, Ano96-74, Ano97-49, Ano97-51, Ano97-52, Ano97-53, Ano97-54, Ano98w, Ano98x, Ano98y, Ano98z, Ano98-27, Ano98-30, Ano99-37, Ano99-38, Ano99-41, Ano99-42, Hig95u, Hig95-32, Hig95-47, Hig95-44, Mey99t, Ano92u, Vin95, Ano91-41, Ano92-62, Ano95-29, Ano95-30, Ano95-32, Ano95-35, Ano96-69, Ano96-87, Ano96-109, Ano97-48, Ano97-50, Ano98j, Ano98k, Ano98-28]. International [Ano98-29, Ano99-39, Ano99-40, Han99-57]. Internet [Ano96-36, Ano96-76, Ano96-75, Ano96-100, Ano97t, Ano97-35, Ano99y, AJ97, Beq99b, Bla96, CvdRvZvD98, Col95-33, Col95-38, Col95-60, Col95-61, Col95-59, Col95-62, Col95-56, Col95-68, Col95-69, Col95-76, Col95-79, Col95-73, Col95-84, Col95-141, Col95-116, Col95-119, Col96k, Col96n, Col96m, Cri97, Doc97, Dos95, Gor95a, Han95b, Han99p, Han99k, Han99l, Han99v, Han99z, Han99-44, Han99-61, Han99-84, Han99-82, Hig95-32, Hig96c, Hig96f, Kov98a, Kov99b, Lic98, Mad95, McC97d, McC97k, McC97l, McC97o, McC98i, McC98-30, McC98-32, McC98-31, McC98-33, McC98-34, McC98-62, Mey96i, Mey96-48, Mey96-45, Mey96-46, Mey96-47, Mey96-53, Mey96-54, Mey97-37, Mey97-39, Mey97-38, Mey98n, Mey98-31, Mey98-64, Mey99-50, Nel99, Nic99t, Nic99-32, Nic99-33, Nic99-55, Oli95a, Pou98b, She95c, VE98, Woo97a, Woo97d, Woo97b, vSG98]. Internet/web [Hig96c]. Internet2 [Han99-45]. Internetwork [Ano96j]. Interoperability [Mey99u]. interoperable [Smi96]. Interpol [Han99-46]. InterScan [Mey96-48, Mey96-49]. Interview [Zaj90c]. intranet [Ano96-87, Mey96-31, Mey97-71, Mey98p, vSG99, Ano96-77, Ano96-78, Mey96-50]. Intranets [Col96n, Mey96-29, Mey98-43]. Intrasoft [Ano96-111]. Introduction [Hig97e, Men92c, Dev99, Mul97b]. Intruder [Mey97-40]. intruders [Can90, Col95-125, Gor95c, Har92, Mey97r]. Intrusion [Coh97c, DKGS94, Hig95a, Hig95s, Hig95d, Hig95r, LV92, Ano97-69, Han99-60, Han99-90, Hig95f, Hig95e, HJS^{+}93, Lun92a, Lun93, Mey98-37, Ric99, SD95, SD97, WP96]. intrusion-detection [SD95]. intrusions [Ano96-66]. intrusive [Ano97-59]. intuit [Col95-44]. invade [Mey99s]. Invasions [Col95-63]. invest [Mey99e]. investigation [Jon97a, Jon97b]. investigations [Beq98c]. Investor [Ano97-60, Han99-36]. Invite [Han99-83]. involved [Han99i]. IP [Ano96-106, Hig95-50, McC97b, Mey96-83, Mey96-84, Mey98g]. IPM [JW90]. IPSec [McC98-35, Mey99x]. IPv6 [Ano99-43]. IRA [Ano92t, Hig95-47]. Ireland [Ano96d]. Ironing [Mey98-36]. Irwin [Woo90a]. ISBN [Hig92a, Lat93, SB90, Zaj94a, Zaj94b, Zaj94c]. ISDN [TMO96]. ISO [JvdV96, WS95]. ISO-10126 [JvdV96]. ISP [McC98-27, Mey99p]. ISPs [Han99u, Han99-47]. Israel [Ano92-96]. Israeli [Nic99-47]. ISS [Mey97-42, Mey98-96]. ISSO [Kov98b, Kov97b, Kov97a, Kov97d]. issue [Ano96-81, Col95-65, Hei96c, Mey96-57]. issuers [Nic99j]. Issues [Lat93, Ano96-33, Ano96-109, Bai97, Bol90, Col95-141, Col96p, DKGS94, Fag93, Fra96, Men92f, Mey96-65, Mey96-78, Mil96, New95, Nic99-46, Nor92a, Osb95, Sha94, Thu91, TR92, Thu94, VSU93, Woo91a]. IT-enabled [Kru95a, Kru95c]. Italy [Hig95n]. ITSEC [Hig95-35, CRR92, Rih91, Rob92]. itself [Ano96-117, Nic99f, Nic99l]. Ivan [Hig95i]. Iyer [Ano96-104].

J [Ano96k, Ano96-66, Ano96-102, Hig95z]. J.K [Ano97-72]. Jack [Hig95-48]. jail [Mey99p]. Jajodia [Ano96-107, Hig95n]. James [Ano92-85, Ano92w, Ano92-46, Ano92-95]. Jan [Ano97-72]. Jansen [Hig95-49]. January [Ano97u, Ano97-60, Ano97-72, Ano97-74, Ano97-85, Ano97-37, Ano97-82, Ano97-86, Mey99i]. Japan [Hig95k]. Japanese [Ano99e]. Jason [Ano92a, Ano92-81]. Jasper [Hig95-50]. Java [Col96u, Hop97, McC97k, McC98-36, McC98-63, McC98-75, Mey96y, Mey96z, Mey98t, Mey98-39, Mey98-69, Mey99-44, Nic99-35]. Javascript [Col96o]. Jay. [Hig95z]. Jean [Ano92-83]. Jeanne [Hig95o]. Jeremy [Hig95a, Hig95h, Hig95j, Hig95g]. Jerry [Hig95-45]. Jim [Ano96x]. job [Col95-66]. jobs [Ano97-47]. Joel [Hig95x]. Johannesburg [Ano96k]. Johansson [Hig95-50]. John [Hig95l, Lat93, Zaj94a, Ano92-55, Ano96m, Ano96-103]. Join [Han99-49]. joins [Mey99-27]. Joint [Col95-67, DMC92]. Jones [Ano96-75, Woo90a]. Jones-Irwin [Woo90a]. Jose [Hig95f]. Joseph [Mey97-96]. Journal [Ano92y, Ano97u]. Jr [Hig95-45]. July [Ano92-71]. June [Ano90-41, Ano91-49, Ano93-48]. just [Han99y, Han99-48, Mey96-85, Mey98-79]. Justice [Ano92-90].

Kahn [Hig95z]. Kantzavelou [Ano96d]. Kaplan [Ano96-82]. Karl [Hig95t]. Kbps [Col95-70]. Keep [Mey96-51, Ano97-82, Col95-90, McC98-40]. keeper [Mey97q]. Keeping [Ano96-83, Col95-68]. keeps [Can90, Mey96-73]. Kelem [Hig95u]. Kent [Hig95q]. Kerberos [Col96p]. Kevin [Ano96-98, Han99-50]. Key [Cla97, Cro97, Cro98, DG98, TO91, Ano96i, Ano96j, Ano96-30, Ano96-34, Ano97u, Ano97-70, BB98, CH94, CHT97, DFHR90, GKM^{+}97, Han99-46, HL90, Hig95-28, JCW91, Lia94, LR92, LV94, MPRZ98, McN95b, Mey96-92, Mey96-97, Mey97-85, Mey98w, Mey98-73, Mos93, Nic99c, Nic99v, PC99, Ram90, ZH90]. key-lock-pair [JCW91]. keys [Ano96-29, Ano96-51, CSW94, Hol95b, Hol95d, LCC94, Mey95, Mey96-100, Mey97m, Mey98-45, PMRZ99, SHW94]. Kicking [McC98-37]. kill [Ano96-31]. killers [Ano97-84]. Kim [Ano97-85]. kingdom [Mey97m, Jon97a]. Kirk [Ano97-79]. Kista [Ano96-80]. Kiswanath [Hig95r]. kit [McC98-36, Mey98-69]. Know [Men91b, Mey99v, Ano97p, Ano97-55, Bou97, Han99-94, Mey97s]. knowledge [Bas91, CN98, GKG92, Ste91]. Koln [SB90]. Kosovar [Han99d]. Kowng [Ano97-72]. Kristian [Mey93]. Krsul [Hig95i]. Krull [Ano96c, Ano96-50]. Kudos [Sch99a]. Kui [Hig95r]. Kumar [Hig95d, Hig95e].

L [Ano96-107]. L. [Ano96-80]. labeling [PG94]. labels [Paa91]. Laboratories [Hig95m]. Labour [McC98-38, Mey97-43]. labs [Han99-89]. lack [Hig95-50]. lacking [Nic99s]. Lafayette [Hig95i, Hig95d, Hig95q, Hig95e]. lag [Col95-129]. lagging [McC98c]. Laird [Ano96-110]. LAN [Ano92s, Ano92-92, Ano97-37, Ano97-79, Ano97-82, Ban95, Col95-120, CKM91, Dav92, McC98-56, Mey96-39, Mey98-40]. LAN-to-LAN [Mey98-40]. LAN/WAN [CKM91]. Land [For99b, Mey98-41]. Landoll [Hig95x]. language [Mag94]. Lankewicz [Hig95u]. LANS [Col95-122, Han95a]. lap [Mey98l]. large [Han99-67]. Laroux [Mey96-52]. last [Ano96-106]. later [Mey97-68]. latest [Mey96-101]. Lattice [San92, WFZ92]. Lattice-based [San92]. launch [Mey99g, Mey99i, Nic99-51]. launches [Mey96-99]. Laura [Ano92-92]. law [Ano92-76, Ano99-49, Ano99-68, Col95-97, Han99-100, Hei94b, Mey98-92, Pou98e]. Lawrence [Hig95f]. laws [Han99r, McC97e, Mey99-36, Mey99-39]. lawsuits [Col95x]. Lax [Ano92-64, Nic99-52]. layer [Ban95]. Laying [Nic99-36]. leader [McC97o]. leadership [HS95b]. leak [McC98-47]. leaks [McC98-64]. Leap [Ano92-65, Nic99o]. learned [Jon97a]. lecturer [Ano92-89]. legacy [Ano97x]. Legal [Beq99b, Gil92, Mey97-44, Beq98a, Beq98b, Beq98f, Bor96, Cri97, Kel95, Mil96, Mor99, Rih94]. legalize [Ano99d]. Legislation [Mul97e, Nic99-37, Ano99y, Han99-61]. Legislative [Nic99-38]. legislature [Han99-74, Möh95]. legit [Ano96-83]. less [Bou97, McC98-29, Mey98-43]. lesson [Col95-34]. lessons [Jon97a, Men92d]. Let [Mey98-42, Ano97-56]. lets [Mey98-103]. Letter [And92, Ano95-36, Ant92, Bow92, Coh92c, Cou94, Dav93, Lin94a, Mur91, Rob93, Zaj92c]. Letters [CR93]. level [AM95, Ano96-98, Laf91, Mag94, Mey96d, Mey99-52, Ran93, vSG99]. levels [Han99z]. Levitt [Hig95t]. liable [Han99u, Par97d]. liberty [War92]. libraries [MH92]. licences [Ano99-44]. licked [Col95-89]. lifeguard [McC98t]. light [Mey96-33, Col96e]. lightens [Han99-27]. like [Ano97-76, Mey97-50]. limit [Col95-37]. Linda [Hig95u]. Line [Ano92-85, Ano92j, Col95-75, Col95-127, Mey96-103, Mey97-36, Mey98-52, Mey98-104]. linear [Ano97-70]. lines [Col95-49]. link [Mey96-67, Mey98-82]. links [McC98-79, Mey98h]. Lisa [Ano97v]. list [Ano90-48, Ano99-65, Mey97-77]. Listen [Ano96c]. Literature [Ano92e, Ano90b, Ano90c, Ano90d, Ano90a, Ano90e, Ano90f, Ano90g, Ano90h, Ano91a, Ano91b, Ano91c, Ano91d, Ano91e, Ano91f, Ano91g, Ano91h, Ano92b, Ano92c, Ano92d, Ano92f, Ano92g, Ano93a, Ano93b, Ano93c, Ano93d, Ano93e, Ano93f, Ano93g, Ano94a, Ano99a, Col94a, Col94b, Col94c, Col94d, Col94e, Col94f, Col94g, Col96a, Hig97b, Mes93]. little [For98b]. live [For98a]. Living [Col95-69, Col96n]. loathing [Mey96-28]. local [Han99-65, Hig95p, Lat92]. Lock [Mey97-45, Mey99w, Col95-51, JCW91, Klo90a, Mey96-92, Mey98f]. lockdown [Ano96-119, Mey96-98]. locked [Ano96-35]. Locking [Mey98-43]. logging [Mey98-42]. logic [dRE96]. Logicode [Col95-70]. login [GKG92]. logs [Ano96-65]. London [Ano96i, Col95-71]. long [Men90a]. longer [Pou98d]. longing [Han99-49]. Longstaff [Hig95-32]. Look [Ame97, Mey97-88, Nic99o, PK91, VE98]. looks [Mey97-38]. loopholes [Ano96x]. looser [McC98-69]. lose [Ano99-68]. losers [Col95-159]. losing [McC98-77]. loss [Ano92-91]. losses [Ano92-29, Mey97g]. Low [Mey97-46, Mey98-97, Rat99b]. Low-overhead [Mey97-46]. Ltd [Ano96-98, Hig95k]. Ludger [SB90]. Lunt [Hig95-44]. lured [McC97c]. lurk [Mey96-78].

M [Ano96i, Ano96k, Ano96-30, Hig95-50]. MA [Hig95y]. Mac [Mey96-74]. machine [McL92]. Machlis [Ano97-73]. Macro [Bon98, DS99, Ano96f, AJ97, Bon96, Mey96e, Mey96-52, Mey97h]. macros [Col95-160]. macrovirus [Han99-53]. macs [Ano92v]. made [Mer90b, Mey97-65, Nic99-67, Øln93, Zaj94c]. Madron [Lat93]. Madsen [Ano92-90]. magistrates [Col95y]. Magnetics [Ano96-32, Ano96-75, Col95-31]. magnets [Hin99b]. Mail [Mey96-49, Ano99x, Ano99u, Col95-30, Col95-100, Col96f, Col96t, Han99z, Han99-28, Hin99d, McC97r, McC97t, McC98e, Mey96-71, Mey97-38, Mey97u, Mey97-69, Mey97-67, Mey97-94, Mey97-92, Mey98-94, Mey98r, Mey98-87, Nel99, Nic99t, Nic99-37, Rob91, TO91, Woo97d, Ano92-85]. mainframe [Mey97-79, PvS91]. mainframes [AD90]. Mainland [Han99c]. mainstream [Han99k, Mey99-35]. maintained [Eva90]. Maintaining [Hig95-33, Men90b, Hig95n]. Major [McC98-39, Ano92-77, Han99-57, Woo90b]. make [Man96]. makeover [Han99-55]. maker [Nic99-59]. makes [McC98-41]. Making [Col95-72, Han99-51, Mey96-53, Mey96-54, Mey99x, Mey99y, Fin98b]. Malampy [Hig95c]. malice [Mey98-95]. Malicious [Col95-73, Han99-56, LLO95, LS93, McC98-75]. Malware [For99a]. Man [McC97l]. managed [Mey99-27]. Management [Ano91-40, Ano92j, Ano92-83, Ano96z, Hig95c, Kap98, All91, Ano97l, Ano97-57, Ass92, BE94, Beq98e, Beq99c, Fin98a, GS91, Hei93b, HS95b, JW90, KJ95, Kov97d, Kov97c, Laf90, LR92, LV94, Mey96g, Mey98-61, Mos93, MM96, PH90, Pla93, PvS91, Ram90, SJ91, Sym94, Thu91, TR92, Woo97b, vSvdM94, vS96, Ano96-104, Beq98b, Hig95c, Ano92-83]. manager [McC98-48]. Managers [Hig95-46, McC98-40, Mey96-55, Mey98-44, Han99-42, Han99-39, Han99-56, Mey98-47, MM96, WP96, Lat93]. Managing [BD95, Mey98-45, She97b, She97c, vSG99, EOO95]. manual [Ano92-63]. manual-the [Ano92-63]. Many [Nic99-39, Col95-46, McC98-72, Woo90c]. Mapping [FW95, Mey99z]. March [Ano97-84, Ano92-70, Ano93-45]. Mark [Ano96-99, Hig95t]. market [Ano92-27, Col95-57, Fin97a, Kov97b, Mey98u, Woo90b]. Marketing [Kru95b, Ano96-36]. marketplace [HS95b]. Markowitz [Fin97a]. Marks [Hig95n]. Marshall [Hig95z, Hig95b]. Martin [Zaj94b]. Maryland [Hig95-50]. mask [Mey99-56]. Mason [Hig95n]. master [Ano96-96]. MasterCard [McC98-41]. match [Mey96-85]. matches [PRK98]. Matching [Hig95d, Mey98-83]. mathematical [Hig91d, Hig97c]. Mathematics [Ano96-66]. Maxwell [Hig95h]. may [Ano99o, Col95-65, Han99-81, Han99-100, Mey99-56, Nic99-54, Ano90-39, Ano91-47, Ano92-69, Ano93-47, Ano94-43, For98a]. maybe [Han99-63]. mayhem [Nic99-43]. McAfee [Mey96h]. McCarthy [Ano97-74]. MCF [LLO95]. McGraw [Zaj94b]. McGraw-Hill [Zaj94b]. MCI [Mey99-27, Nic99-40]. MCII [Col95-74]. McLean [Ano96-107, Hig95b, Hig95s, Hig95o, Hig95z]. MD [Hig95-27, Hig95-47, Hig95-50, Hig95x, Hig95-34, Hig95-49]. Meade [Hig95n, Hig95-34]. mean [CW94]. means [BS94]. measure [Col95-143]. measures [Col96y, Gil92, Mey96x, Sch93]. Measuring [DG98, GDNC94]. mechanisms [SJ91]. Medical [Col95-75, Han99-52, Hei96b, Ano97p, GTKK91, GKKT92]. Melissa [DS99, For99b, Han99-53, Nic99-43, Nic99-41, Nic99-42]. Mellon [Hig95-32, Hig95-33]. membership [Ano92-28]. men [McC98-58]. menace [Ano99-63]. Menio [Hig95u, Hig95-44]. Menlo [Hig95-32]. Mergers [McC98-42]. mess [McC98-38]. Message [ABD93, JvdV96]. messages [Di 93]. Messiahs [Hig95q]. Metal [Han99-62]. Method [DG98, Ano96-39, Ano96-103, Col95-36, She96]. Methodologies [Hig95r]. methodology [BE90, Hit95, MvS94, Mul97b, Oba93, dK95]. methods [Col95-33, ELB93]. metrics [Kov97c]. metropolitan [Hig95p]. Meyer [Ano96-29, Ano96-49]. Miastkowski [Ano97-84]. Michael [Ano97-60]. Microcomputer [Hig91c, Klo90a]. Microsoft [Ano96-35, Ano99-44, Col95-76, Col95-77, Eva92, Han99-54, Han99-55, McC97m, Mey97-47, Mey97-83, Mey98-98, Nic99-44]. might [Han99-94, Nic99b]. migration [Woo95e]. Mike [Hig95o]. Milano [Hig95n]. Military [Mey97-48, All91, vSG99]. millennium [FW99, Hin98a, Kru95a, Kru95c, Ano99-45]. million [Ano92-78, Mey97-56, McC98-43]. Milosevic [Han99n]. MIME [McC98c, McC98-71]. mind [Ano92-93]. minimal [JvdV96]. minimum [CN98]. minimum-knowledge [CN98]. Minnesota [Col95-78]. misbehaving [Hig95f]. Misuse [Ano92-27, Hig95d, Hig95e, HJS^{+}93, Mey98-63, Hig95u]. Mitch [Ano97-85]. Mitchell [Ano92s]. MITI [Mey97-49]. Mitnick [Han99-50]. MITRE [Ano96-107, Hig95b, Hig95o, Hig95z]. mixing [MPRZ98]. MLS [Hig95o]. Mobile [Ano96-84, Han99-56, Hig95u, Mey96-56, Col95-41, Sto95]. MOBIUS [Col95-80]. MOD [Mey96-57, Ano96-61]. mode [McC98-34, Mey96-66]. Model [Hig95d, Hig95w, Ano96k, Ano97o, Ano97z, Ass92, BH92, CPS^{+}98, Col90, EOO95, Fåk91, FH93, JW90, LT94, LK92, LH94, Lin95, SS93, SD97, VPC93, WFZ92, VC96]. model-II. [VC96]. modelling [Ano96-66, dRE96]. models [Sch98, VSU93]. modem [Col95-70]. modern [Nac97, Nac98, Hig95o]. modification [MS90]. Modified [CHT97, MV91]. modulo [SPP98]. MoFAC [Ano96k]. mollify [Col96y]. Mom [Han99-36]. momentum [Col95-153, Col96g, Han99-92]. money [Buz99, Hil95]. Monica [Ano97-82]. Monitor [Ano92-45, Han99b]. Monitoring [Beq99c, Mey97-50, Ano99q]. Morrissey [Ano97-27]. Moskowitz [Ano97-55]. most [Mey99y]. Motivation [Mul97d, Mul97b]. Motorola [Col95-80, Mey99-29]. mount [McC97a]. Mountain [Hig95u]. mousetrap [MG92]. move [Col95n, Col95-110, McC97e]. moves [McC98-64]. movie [Ano92-46]. moving [Han99-88]. MPs [Nic99s]. MS [Eva92, Han99-58]. MS-DOS [Eva92]. much [Ano97-74, Col96m, Man96, Mey96-46]. Mukherjee [Hig95r]. Mulhall [Ano96-116]. Multi [Ano99-46, Laf91, MS90, Yen96, vSG99]. Multi-industry [Ano99-46]. multi-level [Laf91]. multi-user [MS90]. multi-verifier [Yen96]. Multidomain [VG94]. Multilevel [Hig95o, Thu91, TR92, VB91, All91, Ano96-107, HL90, Hig95n]. multimedia [Haw99, Kel95, Sha94]. multinational [Hop94]. multiple [Hig97t]. Mundv [Hig95o]. Munich [Hig95-27]. Murphy [Hei93a]. Music [Han99-59]. must [Hig97y, Kov97d, Nic99-33]. mutates [Nic99-41]. mutating [Mey97-73]. MVS [Cro95b, Cro97, Cro98, Paa91, Sch91, Sch93]. MVS/ESA [Paa91]. MVS/SMF [Cro95b]. my [Han98o, Mey97-85]. Myla [Hig95t]. Mystery [Col95-81]. myth [Mey96l].

N [Ano96i, Ano96l, Hig95w]. nabbed [Col95-78, Mey98-29]. NADIR [HJS^{+}93]. nail [Han99-83]. name [Nic99-31]. names [Han99r, Mey97-80]. Nancy [Hig95u]. NASA [Ano99-47, Han99-78, Nic99-52]. nasty [Mey97-52]. Natick [Hig95y]. nation [Cor97]. National [Hei94a, Hig95s, Hig95-34, Hig95-48, Hig95-49, Hig97b, Kow91c, Car90a, Hig93a, Kow91a, Kow91b, Mol97, Win90, Hig95-48]. NATO [Ano99-27]. naughty [Col95-132]. Naval [Ano96l]. NC [Hig95m]. near [Col95q, Han99-45]. NEC [Ano99-48]. necessarily [Cly93, Han99-100]. need [Ano97p, Ha92, Han99d, Har95, Mey98p, Mey98-49, Mey98-57, Mey99-30, Sto95]. need-to-know [Ano97p]. needed [Han99-87]. needs [Ano96-64, Klo90c, Nic99-67, Por96]. neglect [Mey98q]. neglected [McC98-72]. Nepean [Hig95p]. ’Net [Col95-95, Mey98f, Col95t, Col95-93, Mey96-42, Mey96-94, Mey96-97, Col95-83, Col95-82, McC98-78, Mey96-59, Mey96-60, Mey96-67, Mey97z, Mey97-29, Mey98-92, Mey98-46, Nic99-45, Nic99-44, Ano96a, Col95h, Col95-52, Col95-102, Col96b, Col96l, McC98-76, Mey96-35, Mey97g]. net-work [Ano96a]. Netbus [Mey99-31]. nets [Mey96-36]. Netscape [Col95-84, Col95-126, Mey98-99]. NetWare [Col95-143, Hig95j]. Network [Ano92-66, Ano96z, Ano96-85, Ano96-104, Ano97-55, Ano97-27, Ano97-76, McC97n, Mey98-47, Mey98-48, Mey98-100, Mey99-32, Ric99, SG98, Ano97-34, Ano97-75, Ano99q, Ano99-48, Ano99-59, Ban95, BdAR99, BM90, Col95-88, DFHR90, DMGK95, FH93, GW92, Han99-45, Han95c, Hig95g, Hig95p, Hig95v, HJS^{+}93, Jan92, Kov96, Mey96s, Mey96-51, Mey97r, Mey97-81, Mey98-37, Mey98-38, Mey99z, Mos93, Sch97b, Woo95e, Woo95b, McC98-43, Lat93]. networked [Haw99]. Networks [Mey98-49, Ano96-97, Ano97-34, Han99-83, Hig92r, JJ96, LR93, Lat92, Mey96-69, Mey96-84, Mey96-82, Mey97-52, Mey97-86, Mey98-77, Mol97, Nic99-66, NSS91, Sto95]. Neugent [Hig95o]. neural [DMGK95]. News [Hig97j, Hig97k, Ano98n, Hig92d, Hig92e, Hig92f, Hig92g, Hig92h, Hig92b, Hig92i, Hig92c, Hig93c, Hig93d, Hig93e, Hig93f, Hig93g, Hig93h, Hig94a, Hig95-30, Hig95-31, Hig96a, Hig97a, Hig97h, McC98z]. next [DS99, Han99-41, Han99-74, Kru95a, Kru95c, McC97n, McC98b]. next-generation [McC98b]. niche [Col96l]. Nicholas [Hig95r, Hig95v, Ano97u]. Nigel [Ano96-104]. night [Col95-125]. night-time [Col95-125]. nightmare [Mey96-59]. Nikitakos [Ano96l]. NM [Hig95c]. No [Ano92-55, Ano92-84, For99b, McC98-45, Mey97-55, Pou98d, Col96h, McC97v, Mey96-50, Hig96d]. Nobel [Hig95h]. Non [Ano97-59, Hig91d, Hig97c, Nic99-64]. non-existent [Nic99-64]. Non-intrusive [Ano97-59]. non-mathematical [Hig91d, Hig97c]. Norman [Ano96-65]. Notargiacomo [Ano96-107]. note [Ano91s, Coh91c, Coh96a, Coh98, LC94]. notes [Hig97k, Mey98-81, Nic99-53]. Novell [Hig95j, Hig95g, Ano92-66, Dav90a, Han99-63, Mey97-56, Mey98-51, Mil95, Nic99-46, Nic99-47]. November [Ano90-44, Ano91-52, Ano92-74, Ano93-51]. NSA [Ano92-67]. NT [Bir97, Col95-88, Col96q, Cur95, Han99-64, Han99-97, Hig97w, McC98-46, McC98-74, McC98-83, Mey96r, Mey96-102, Mey97-57, Mey98v, Mey99-40, Nic99h, Sch97b, Sch99a, Woo97e]. Ntis [Hig96e, Hig91e, Hig97l, Hig97m]. nuclear [Han99-89]. number [Col95r]. numbers [Col95z, Dav90b, Han99-73, McC98-60, PMRZ99]. NZ [Col95i].

O [RR91]. Object [Cor95b, Hig95w, Pre96, Ano96b, Ano97o, SU93, VB91]. Object-Oriented [Hig95w, Ano97o, SU93, VB91]. objects [BE95, OvS92]. observations [Ano96-112, Spa97b]. October [Ano92y, Ano92-29, Ano92a, Ano92-45, Ano92-65, Ano92-64, Ano92-66, Ano92-78, Ano92-76, Ano92-79, Ano92-86, Ano92-83, Ano92-81, Ano92-88, Ano92-87, Ano92-90, Ano92-89, Ano92-91, Ano92-96, Ano92-94, Ano90-43, Ano91-51, Ano92-73, Ano93-50]. odds [Han99-47]. OECD [McC97o]. off [Han99x, Nic99-54]. offer [Han99-54, McC98-66, Mey96-50]. offered [Mey99o]. offering [Mey98h]. offerings [Mey99-54]. offers [Col95-74, Col95-118, Mey98-99, Mey98-64, Nic99-40]. Office [Han99-58, Hig92r, VSU93]. officer [KL96]. offices [Mey96-60]. often [McC98-72]. OH [Hig95q, Hig95v]. OILZ [LS93]. OK [Nic99-59]. Old [Col95-89, Hru95]. Olivier [Ano96k]. Olson [Hig95-29]. Olsson [Hig95r, Hig95t]. On-line [Mey98-52, Col95-127]. One [Ano96-87, Far91, Han99-65, Hin99e, Ano96-81, Col95-29, Man96]. one-way [Man96]. ONERA [Hig95w]. ONERA/CERT [Hig95w]. ones [Col95-89]. Online [Ano96-106, Ano97-60, Col95c, Col95-90, Col96r, Han99-66, Hig95m, McC98-47, Mey96-62, Col95-157, Han99i, Han99q, Han99-45, Han99-94, Mey96-65]. Only [Col95-91, Hig95-47, McC97t, Mey96-70, Mey99-46, Nic99-27, Nic99z]. onto [Han99l]. Open [Ada96, Rei95a, Rei95b, Rei93, Ano96l, Bla92, Buc95, Buc96, Hig95g, Mey99k, Nic99h, Nic99-30, PA95, Pre91a, Ram90, Smi95b, Smi96]. OpenPGP [McC98-71]. operability [Hug99]. operable [Smi95b]. Operating [Coh93, Hig95h, O’S91, Ano92-84, Han99-51, Hru95]. Operational [Col95-92]. operations [Mey97b, Ano96-98]. opinion [Hei94a]. Oppliger [Ano96-34]. opportunities [Bov95]. optic [CKM91]. Optimal [Hig95c, SD95]. optimization [BE94]. Optimized [LB91]. option [Pou98d]. options [Klo90b, McC98-42]. OPUS [Spa92]. Oracle [Col95-93, All91]. ORACLE-multilevel [All91]. O’Reilly [Hig92a]. Organization [Kov98b, BdAR99, Hop94]. Organizational [Fag93, Ano97-57]. organizations [Ano96w, CvdRvZvD98, Mey99l, Rat97, Woo95a]. Oriented [Hig95w, Ano97o, Cor95b, Pre96, SU93, VB91, vSvdM94]. Orifice [Ano99d, Han99g]. OS/2 [Cog95]. OSes [Hig97t]. other [Han99-47, Hin98c, Hin98d, Kem97a, Kem97b, McC97k, Zaj90d]. our [Hin99b]. outlaws [Ano92-44]. outline [Ano97a]. outlook [BD95]. Outsourced [Han99-68, Han99-67, Nic99-40]. Outsourcing [Tan95, Ano97-78, She97b, She97c]. overestimated [Bru95]. overhead [Mey97-46]. overkill [Col95-128]. overlooked [WB93]. Overview [Bla93, Ano96z, Bov95, MNP96, Ozi98]. own [Col96n]. owned [Han99-90]. owners [Ano96-63]. ownership [Bor96]. oxymoron [Cly93].

P [Ano96l, Ano96-107, Ano92w, Ano92h, Ano92j, Ano92s, Ano92-27, Ano92-29, Ano92a, Ano92-45, Ano92-47, Ano92-46, Ano92-48, Ano92-65, Ano92-64, Ano92-78, Ano92-76, Ano92-81, Ano92-88, Ano92-87, Ano92-92, Ano92-89, Ano92-91, Ano92-95, Ano92-94, Ano97v, Ano97-37, Ano97-79, Ano97-82, Ano97-86, Ano97-85, Col95-135]. P. [Ano96-102]. PA [Hig95-32, Hig95-33, Hig95c]. Paans [Law92]. Pack [Han99-97]. package [GDNC94, Hig91b]. packages [McC98-44]. packet [VE98]. packet-intercepting [VE98]. Pact [Mey97-58, Col95-80]. Padula [LH94]. page [Han99-36, Han99-42]. pager [Mey96k]. pagers [Mey98-82]. Pages [Ano90-38, Ano90-42, Ano90-45, Ano90-40, Ano90-41, Ano90-39, Ano90-44, Ano90-43, Ano91-54, Ano91-50, Ano91-53, Ano91-48, Ano91-49, Ano91-47, Ano91-52, Ano91-51, Ano92-68, Ano92-75, Ano92-71, Ano92-70, Ano92-69, Ano92-74, Ano92-73, Ano92-72, Ano93-49, Ano93-52, Ano93-46, Ano93-48, Ano93-45, Ano93-47, Ano93-51, Ano93-50, Ano94-50, Ano94-44, Ano94-43, Ano94-45, Ano94-46, Ano94-47, Ano94-48, Ano94-49, Ano95-38, Ano95-44, Ano95-37, Ano95-39, Ano95-40, Ano95-41, Ano95-42, Ano95-43, Ano96-89, Ano96-95, Ano96-88, Ano96-90, Ano96-91, Ano96-92, Ano96-93, Ano96-94, Ano97-62, Ano97-68, Ano97-61, Ano97-63, Ano97-64, Ano97-65, Ano97-66, Ano97-67, Ano98-32, Ano98-38, Ano98-31, Ano98-33, Ano98-34, Ano98-35, Ano98-36, Ano98-37, Ano99-51, Ano99-57, Ano99-50, Ano99-52, Ano99-53, Ano99-54, Ano99-55, Ano99-56, Lat93, Mey98-85, Zaj94a, Zaj94b, Zaj94c]. pair [JCW91]. palate [Mey97n]. Panel [Ano90-46, Ano90-47, Ano96-38, Han99e]. paper [Ano99s, Hug99, LT94, Law92, Mey99i, Rob99]. papers [Ano93-55, Ano93-56, Ano93-57, Ano94-51, Ano99-58]. paradigm [Ano97-83, Han99-62, Hig95-27]. Paradigms [Rus93, Gor95d, Rus94]. paranoid [McC97t]. Parents [Han99-69]. Paring [McC97p]. Park [Ano96j, Hig95u, Hig95-32, Hig95-44, Hig95-50]. Parkhurst [Ano97-36]. Parkin [Ano96-81]. Parliament [Han99v]. Part [Mey97-59, Mul97b, Mul97c, Mul97d, Mul97e, Mul97f, Mul97g, MvS94, Nic99u, Col99a, Col99b]. partial [Ano96-51]. parties [Ano96-86, Ano97n]. Partitioning [Hig95-34]. partners [Rat99b]. partnership [Han99m]. party [Sch97a, SGI^{+}99]. Pass [SG94]. Pass-sentence [SG94]. passions [Mey96-65]. Passports [DD92]. Password [Ano96-96, Nic99-48, YS99, BK95, CL94, Col95-32, Col95-35, Col95-51, Col95-94, Col95-103, McC97p, Mey97-57, SJJF96, Spa92]. Passwords [Hig96f, Mey98-53, Mey99-33, BPH^{+}97, Han99-64, Man96, Mey96m, Nic99g, TCHS96, ZH90]. past [Mey99-33]. patched [Nic99-35]. patches [Col95-113]. Patel [Ano96d]. path [FH93]. Pathways [Ano96-108]. patrols [Han99-62]. Pattern [Hig95d]. Paul [Ano92-29, Ano96-101, Hig95-29]. pay [Mad92a]. payment [Col95-33, Mey96-58]. Payne [LT94]. payphone [Col95-31]. pays [Ano92-91]. PBX [MH97]. PC [Ano97-84, Ano96-31, Ano96-99, Ano97-76, Coh91c, Col95q, GW92, Han99-93, McC98-48, McC98-64, Mey96f, Mey96-101, Nor92a, TMO96, Tay92, Ano92-84]. PCs [Don93, Mey99s]. PDAs [Nic99-49]. peeking [Mey97-55]. pees [Mey99t]. Penetration [Har97, Jon97a, Jon97b, Ano96-32]. Pennsylvania [Hig95c]. Pentium [Han99-70, Mey99-42, Nic99-51]. People [Zaj90d, Ano96-33, Woo90c]. PeopleSoft [Col95-94]. Peppes [Ano96i]. Perceived [DBHM95]. perception [Mul97c]. Perceptions [JATC93]. Perestroika [Lat90]. perfect [vSN92]. performance [DKGS94, Ran93]. peril [Mey97-44]. perimeter [Mey96-82]. Personal [Hig95h, Hin98h, Mey96-35, Mey99-34]. perspective [Hic97a, Hic97b, Hig95-35, Lun92b, Sch95a, Vin95]. Persuasion [Hig95-46]. Pessimism [Han99-71]. Peter [Hig95v, Hig95-44, Ano96-78, Ano96-85, Ano97-27]. PGP [McC98-49, Mey97-60, Mey98-54, Mey99-34, Mor96]. phase [GKM^{+}97]. Philadelphia [Hig95c, Ano97-56]. Philip [Ano96-28]. Phillips [Hig95s]. phone [Ano92-45, Col95r, Col95-41, Hor92]. phones [Mey96-76]. Photon [Mey96-63]. phreaking [Ano99-63]. phylogenies [Ano97s]. Physical [Ban95, Ben98, Klo90b]. Picarille [Ano97v]. pickings [Mey98-102]. piece [Col95-95]. pierce [McC98m]. pileup [McC97p]. pipe [Col96r]. Piracy [Web95, Ano99-46, Col95i, Col95-104, Col95-136, Faw92a, Gli90, Mey96-34, Mey96-47, Mey97-77, Nic99a, Nic99-34, Nic99-47, Nic99-60, Wit92]. Piraeus [Ano96i, Ano96l]. pirate [Col95-96, McC97l]. pirates [Col95j, Col95n, Col95v, Col95-77, Col95-142, Col95-137, Col95-151, Faw92b, McC97m, Smi92]. pirating [Beq98e]. pitfalls [Beq99c, Rei95a, Rei95b]. pits [Mey96-88]. Pittsburgh [Hig95-32, Hig95-33]. PKI [CDR99, Mey97-61, Mey99-35]. place [Mey97-35]. plague [Col95-79]. plagues [Mey97-94]. plaintext [BH93]. Plan [Mey97-62, Ano96-117, Ano99-45, Col95-145, Han99n, Han99-90, McC98-56, Mey97f, Mey98l, Mey98s, Mey99-52, Nic99j, Pou99a, Pre91b, ZHM90]. Planned [Ano92-76]. planning [Ano96m, Ano96-85, Dev99, Hig95z, Hin98a, Klo90d, Men94, Mey96-94, Nos99, RH91, She95b, SS95]. plans [CM92, Col95-129, Col95-114, Hig95-47, McC98-78, Mey96-93, Mey97-43, dK95]. Platinum [Mey98-59]. play [Ano99x, Lin97]. plays [Jan92]. plc. [Ano96-86]. plea [Han99-50, McC98-59]. plot [McC98-78]. plug [McC98-64]. Plugging [McC98-50, Rei96]. Pohl [SB90]. Poindexter [Hig95-46]. point [McC98-69, Mey98-65]. points [Mey99h]. pol [Mey98-28]. Police [Ano92-77, Col95-96, Mey96-64, Nic99-33]. Policies [Woo97c, AM95, Kow91a, Lin95, Mey98-78, Øln94, PG94, Woo95c, Woo95d, Zaj94c]. Policy [Küh99, Ano96-38, Ano96-62, Ano96-97, Ban90, Gri97, Hic97a, Hic97b, Mey96j, Mey96v, Mey96-61, Mey97-49, Mey97-59, Mey98-46, OA90, Pou97]. Polish [Ano99-49]. politic [Mad98]. polyinstantiation [Laf91, SJ92]. Poole [Ano96w]. porn [Ano99-47, Col95-83, Mey96-59]. Pornography [Hig96g, Col95-91, Han99p]. Pornotopia [Bla96]. Portable [McC97q]. porter [Fin97a, Ano96-64]. portfolio [Fin97a]. ports [RR91]. Posch [Ano96-51]. pose [McC98-34, Mey96-62, Mey98-97]. poses [Ano99w, Mey96-45, Mey96-86]. POset [Coh91b]. Possible [Bon96, Col95w]. posted [Han99-52]. Postfach [SB90]. postponed [Mey99q]. potential [McC97k]. Power [Jan92]. Powerful [McC98-51, Nic99-50]. powers [Han99j]. pp [Ano92-55, Ano92-66, Ano92-79, Ano92-85, Ano92-83, Ano92-84, Ano92-90, Ano97c, Ano97-55, Ano97-27, Ano97-35, Ano97-36, Ano97-47, Ano97-72, Ano97-73, Ano97-74, Ano97-76, Ano97-84]. Practical [Ano96-98, Hig95o, Ano96-99, Hig95n, Hig95p, Zaj94b, Hig92a]. practice [Ano97l, Hic95, SL96]. practices [Ass92]. practitioner [Sch98]. Predicate [Laf91]. prediction [Ano96-66]. preface [Ozi98]. preface/overview [Ozi98]. Preliminary [Ano90-48, Coh97a, Coh97b, LS93]. Preparedness [Hig95s]. Preparing [Mey97-63, Tab90]. prescription [Ano99-59]. prescriptions [Sch99a]. Present [Col95-97]. presentation [McN95b]. presents [Mey99-31]. President [Beq98d]. Pressure [Col95-98, Ano99-44]. Pretty [Mey97-64]. prevent [Bon96, McC98-81, Nic99-70]. prevented [Zaj90b]. Preventing [Spa92]. prevention [AD90, Ano96-104, Ano97r, Han99g]. preventive [Gil92]. preview [Mey99-42]. prey [McC98g]. Price [Hig92a]. prime [Han98i]. prime-time [Han98i]. principle [Koe93]. Principles [Ano96-50, Hei95a, Kru96, Woo90d, Woo93b, Ano99f, SJ91]. Privacy [Ano96-100, Col95-99, Col95-158, Han99-72, Hin98e, Mey96-65, Mey97-64, Mey99-36, Nic99-51, Oli95b, Ano92-76, Ano96-80, Ano97-69, Ano99o, Ano99u, Col95-27, Col95-63, Col95-75, Col95-101, Col96f, Han99d, Han99i, Han99o, Han99-36, Han99-54, Han99-63, Han99-65, Han99-88, Han99-91, Han99-93, Hei94b, Hig95-33, Hig96g, Mad95, Mey96-88, Mey98-90, Mey98-92, Mey98-52, Mey98-79, Nic99x, Nic99w, Puj93, Rot92, Sto95, Woo91a, Ano92y]. private [Col96n, Mey96-36, Mey97-86, Mey98-77, Nic99-66]. privatization [Ano96-61]. privilege [Pre91a]. Pro [Mey99-34]. proactive [BK95]. probabilistic [GKG92, Mor96]. probabilities [PRK98]. probably [Ano96i]. Probe [Ano96-28, Han99q, Han99-44]. probes [McC98-67]. problem [Han99f, Joa97, Kem97a, Kem97b, Mey98i, Mey98-62, Mul97c, Nic99p, WB93]. problems [Ano96h, Bon98, Bon99, Cie97, Col95-44, Cor95a, Han99-56, Han99-72, Mey99m]. procedure [MPRZ98]. Procedures [Hig97n, Ano96z]. proceeding [Ano96-60]. proceedings [Hig97i]. process [Ano97x, CM93, PH90, Rat99a]. processes [SD97]. Processing [Ano96-51, Hig95m, Cro95b, Men94]. products [McC98-28, Mey99-44, Smi95b, Smi96]. Prof [Hig95p]. Prof. [Ano96g, Ano96-112]. professional [Bas91, Eva90, Han99-79, Kov97e]. Professionalization [Hig95-46]. Professor [Ano96-86, Mey97-96]. profile [Ano96l]. Profiles [Hig95u, vSvdM94]. profit [Wol95]. profitability [RR91]. program [Coh93, KS97, MS90, Mey97-42, Hig95i]. programme [Han99m, Kov96, Mey99g, Rob92]. Programming [CGO93]. Progress [Kru91]. Project [Hig95-44, Ven95, BE90, Ano92-78]. projects [Ano96-101]. Promises [Cor95a, Mey97-58]. promotes [Han99v]. proof [Col95-146, Mey97-80, Tay92]. property [Ano99-49]. proposals [Col95-43]. propose [Ano99y, Col95-152]. proposed [Col95y, Col95-87, Mey96-66]. proposes [Han99o]. proposition [Kim95]. proprietary [Ano99t]. prosecute [Han99-86]. prosecutions [Mey96-59, Mey98k]. prosper [Mey97-66]. Protect [Col95-100, Ano97-47, Col95-55, Mey96-40, Mey98-89, Mey98-80]. Protecting [Col95-101, HDW97, Mey96-67, Mol97, OvS92]. Protection [Mey96-68, Mey98-55, Mey99-37, Pou98a, Pou98e, Pou99a, Ano97-75, Ano99f, CRR92, Coh93, Coh98, Col95q, Col95-157, Eva96, Fra95, Jan92, LCC94, LK92, McC97q, McC97t, McC98-66, McC98-80, Mey96-50, Mey96-63, Mey96-101, Mey98l, Mey98-41, Mey98-53, Mey98-81, Mey99-54, Muf90, Nic99n, VC96, Beq98d, Zaj94b]. protections [Han99-87]. protects [Ano97-76, Nic99-37]. protest [Han99-37]. Protocol [McC98-62, GKG92, JL96, Ran93]. Protocols [Zaj94a, Ano96-30, Ran92, SSP90]. prototype [KJ95, Hig95-44, Hig95r]. Provable [Sch98]. provide [She95c]. Provider [Han99f]. Providers [Mey99-38]. provides [Col95-70, Mey97-61]. Providing [Col95-102, Mey96-69]. Provo [Hig95j, Hig95l, Hig95g]. Pseudonymous [Ano97-69]. Pssst [Col95-103]. Psst [Mey98-56]. Pstt [Mey96-70]. Public [Ano97-70, Mos93, Ano96j, Ano96w, DFIR99, Han99-86, LCC94, MPRZ98, Mul97c]. Public-key [Ano97-70, MPRZ98]. Publication [Gor95b]. published [Ano92-27, Col95-157]. publisher [Woo95a]. publishers [Eva96]. Puffer [Mey96-71]. Puketza [Hig95r]. punishment [Som92]. Purdue [Ano96g, Ano96-112, Hig95i, Hig95d, Hig95q, Hig95e]. purview [Mey98y]. push [Mey98-67, Nic99j]. pushes [Faw92a]. put [Mey98-87]. puts [Ano92-81]. Putting [McC98-52, Wit92, Han99-38]. puzzle [Bou97].

quadratic [Ano97-72]. quadruple [Han99-77]. quagmire [Beq98b, Beq98e, Cri97]. quality [Ano96-111]. quandary [Mey96-96]. Quantitative [YA96]. Quantum [Ano92-79]. Quarterly [Ano92-55, Ano92-84]. questions [Ano96c]. Quick [Mey98-101].

R [Ano96j, Ano96-51, Hig95f, Law92]. R. [Ano96-34, Ano96-115]. R.H [Ano96i]. R.H.-University [Ano96i]. R3 [Hig95k]. RACF [Lyn91]. RACIF [Law92]. RAD [Lem96]. radiation [Hin98g, Smu90]. Raids [Col95-104, Col95n]. Rain [Ano92-80]. raise [McC98-39, Mey97j]. raises [Col96p, Nic99-46]. rally [Ano92t]. RAMeX [KJ95]. ramifications [Mul95]. Rand [Ano96k]. Random [Ano93-53, Ano97-71, Hig90a, Hig90b, Hig90c, Hig90d, Hig90e, Hig90f, Hig90g, Hig90h, Hig91f, Hig91g, Hig91h, Hig91i, Hig91j, Hig91k, Hig92j, Hig92k, Hig92l, Hig92m, Hig92n, Hig92o, Hig92p, Hig92q, Hig93j, Hig93k, Hig93l, Hig93m, Hig93n, Hig93o, Hig93p, Hig94b, Hig94c, Hig94d, Hig94e, Hig94f, Hig94g, Hig94h, Hig94i, Hig95-36, Hig95-37, Hig95-38, Hig95-39, Hig95-40, Hig95-41, Hig95-42, Hig95-43, Hig96h, Hig96i, Hig96j, Hig96k, Hig96l, Hig96m, Hig96n, Hig97p, Hig97o, PMRZ99]. rapidly [Han99-66]. rate [Col95-129]. rates [BPH^{+}97]. ratings [Mey97-35]. rattles [Col95-151]. Ray [Ano96-82]. RCMP [McC98-53]. re [Han99-96, Rat97, Men92e]. re-building [Han99-96]. re-evaluation [Rat97]. Re-examining [Men92e]. readied [Mey99-54]. ready [Ano99-69, Nic99k, Nic99-39]. Real [Mey96-72, BdAR99, FW99, Han99y, Hei96c, Hig95-47, LE98, LT94, Mey96-36, VPC93]. real-time [LE98]. Realisation [CN98]. realistic [Ano92-46]. reality [Mey96l, Mey98z]. reap [Ano99o]. reawakening [Col95-82]. rebound [Mey98-48]. Recall [BPH^{+}97]. receiver [Rus94]. reception [Smu90]. Recession [Ano92-81]. recipe [Col95-105]. recognition [DMGK95]. Recognizing [Gor95c]. record [Mey99b]. records [Ano97-77, Col95-75, Han99-52, Hei96b]. recover [MS90]. Recovering [Kov95]. Recovery [Hei93b, Ano92-48, Ano96-33, Ano97r, Cla97, Col95-145, GKM^{+}97, Han99-46, Hil95, Men94, Mey98w, Nor93, Pea91, Tay95]. reduce [Hig97n]. Reducing [Mou98]. reduction [SPP98]. redundancy [JvdV96]. redundancy-doing [JvdV96]. Redundant [O’S95]. Refereed [Ano90-49, Ano91-55, Ano91-56, Ano92-82, Ano93-54, Ano93-55, Ano93-56, Ano93-57, Ano94-51, Ano99-58, Ano90i, Ano90-50]. referees [Ano90-31, Ano90-32, Ano90-33, Ano90-34, Ano90-35, Ano90-36, Ano90-49, Ano91-42, Ano91-43, Ano91-44, Ano91-45, Ano91-46, Ano92-56, Ano92-57, Ano92-58, Ano92-59, Ano92-60, Ano92-61, Ano92-82, Ano93-41, Ano93-42, Ano93-43, Ano93-44, Ano93-55, Ano93-56, Ano93-57, Ano94-36, Ano94-37, Ano94-38, Ano94-39, Ano94-40, Ano94-41, Ano94-42, Ano94-51, Ano95-28, Ano95-31, Ano95-33, Ano95-34, Ano96-67, Ano96-68, Ano96-70, Ano96-71, Ano96-72, Ano96-73, Ano96-74, Ano97-49, Ano97-51, Ano97-52, Ano97-53, Ano97-54, Ano98w, Ano98x, Ano98y, Ano98z, Ano98-27, Ano98-30, Ano99-37, Ano99-38, Ano99-41, Ano99-42, Ano91-41, Ano92-62, Ano95-29, Ano95-30, Ano95-32, Ano95-35, Ano96-69, Ano97-48, Ano97-50, Ano98-28, Ano98-29, Ano99-39, Ano99-40]. referres [Ano90-37]. refinement [CJC98]. reflected [Tay95]. Reflex [Ano96-32, Ano96-75]. Reg [Ano92-87]. region [Ano92k]. Registrar [Pou98a]. registration [Ano96-34]. registry [Hig97w]. regression [Mer90a]. regulated [Col95-91]. Reid [Ano96x]. related [Coh92a, Mul97a, Mul97c, TO91]. relationship [CJB94, Col95-131]. relaxes [Mey98-78]. released [Han99g]. releases [Han99-91]. relevant [She90]. reliable [Ano96-111, FGAR93]. remainder [HTY96]. remains [Col96t]. remote [Ano96-108, CL94, Han95a, Mey98-51]. remove [Fer92]. Reno [McC98-54]. repair [Yen96]. replace [Ano97k]. replicate [Ano97-79]. Report [AAL^{+}93, Han99n, McC98-55, Nic99-52, Nic99-53, Pou98a, Ano91-39, Ano92u, Car90a, DMC92, EM90, Han99z, Han99-82, Hig93b, Hig93i, Hin99f, Mey96u]. reported [Ano99-59]. reporting [Ano96-65, PvS91]. reports [Han99-87]. repossession [Col95-147]. require [Han99-81]. requirement [MvS94]. Requirements [BD92, Ano97z, Ano97-77, Cie97, Hit95, Hug95a, MvS94, Pfl91]. requires [Han99-96]. rescue [McC98-56]. Research [Hig95-27, AAL^{+}93, BD95, BdAR99, Han99-75, Lun92b, She90]. residents [Han99d]. residues [Ano97-72]. resilience [MvS94]. resolution [Han99-98]. resource [Ano96-77, BM90]. response [Boy97, Coh97c, Mey96j, RH91]. responsibility [Woo95a]. restless [Mey96d]. Reston [Hig95h, Hig95j, Hig95g]. restricted [SPP98]. restriction [Mey98s]. Results [Hig95r, Coh92a]. resumption [RH91]. retailing [Men90b]. rethinks [McC98-78]. Retinal [Boo92]. retreat [Han99-49]. retrospective [Hei93c]. return [McC97r, Mey97-69]. reuse [Hig95-45]. Reuters [McC98-57]. reveals [Mey96-64]. revenge [Col95-50]. revenue [McC98-58]. Reversible [MPRZ98]. Review [Ano92j, Ano92-85, Ano92-83, Ano92-84, Ano96i, Ano96j, Ano96k, Ano96l, Ano97-72, Ano97-73, Ano97-74, Hig92a, Hig95h, Hig95f, Hig95i, Hig95j, Hig95s, Hig95k, Hig95l, Hig95c, Hig95m, Hig95o, Hig95n, Hig95g, Hig95d, Hig95q, Hig95e, Hig95p, Hig95r, Hig95t, Hig95u, Hig95v, Hig95w, Hig95x, Lat93, SB90, Zaj94a, Zaj94b, Zaj94c, Cro97, Cro98, Ric99]. Reviewing [Ano96-101]. reviews [Col95-124]. revised [Pou99b]. revisited [Col95-109, For99a, HZS94]. Revocation [VC96]. Rich [Mey98-102]. Richard [Ano92j]. Right [Ano96-102, Col95-72, Han99-100, Mey96-85, Osb97a]. rights [Bor96, GKKT92, Han99-94, O’S95]. Rik [McC98-33]. Rindos [Col95-58]. ring [Ano92-77, Col95-96]. rings [McC98y]. RIP [Han98o]. rise [Ano92-29, Han99a, Han99-43, Han99-72, Men92b]. rises [Col95r, Tay95]. Risk [BE90, Bas91, Cie97, Hig95c, HS95b, dRE96, Ano96-39, Ano96-98, Ano97k, Ano99w, BE94, Car91, Col96t, ELB93, Far91, Hin99b, KJ95, LE98, Mey96g, Mey96-45, Mey96-79, Mey97-35, Mey99l, MM96, Rat99a, Som94, vdME98]. Risks [Han99-73, Rat99b, Ano92-85, Bol95, Fer92, Fin97b, Hil91, Kel95, Lic98, McC98w, McC98-63, Mey96c, Mey96-87, Mil96, Nel99, Rat98, Zaj92a]. risks-how [Zaj92a]. risky [Mey98w]. rivals [McC98f]. Rivka [Ano97-47]. RNS [SPP98]. RNS-modulo [SPP98]. road [Ano96-79]. roadmap [Ano96-87]. Roaming [Ano97-75]. Rob [Hig95t]. robbery [Kov99b]. Robert [Ano97-55]. Rod [Ano96-81]. Roger [Hig95j, Hig95k, Hig95g]. Rogers [Hig95-45]. Role [Hig95m, Par95a, Ano99x, Coh98, Col95-37, Col95-62, Jan92, Law93b, LV92, Lin96, MD94, vSvdM94]. role-oriented [vSvdM94]. Roles [Ano92-86, Law93b]. roll [Mey99-53]. ROM [Ano92-67, New95]. Ronald [Hig95r, Hig95t]. round [Ano92t]. roundup [Ano90n]. Router [Mey97-65]. Routers [Col95-106]. roving [Han99-62]. row [Nic99u]. Royal [Ano96-80]. RS [Smu90]. RS-232 [Smu90]. RSA [Hei95b, LB91, Mey99-39, SPP98]. RTF [Nic99-41]. RTFM [Ano96x]. Rudell [Hig95o]. rug [Mey98b]. rule [Ano99o, Ano99u]. Rules [Hig95m, Col95-150, McC98-69, Mey97x, Mey98-90, Mey98u, Mey99r]. ruling [Mey96-64]. rumbled [Col95-142]. running [Mey97w]. Rush [Han98o]. Russia [Mey98-35, MNP96, Nic99-54]. Russian [McC97m, McC98-59, Nic99-39]. Rx [McC97k].

S [Ano96h, Ano96k, Ano96l, Ano96-77, Hig95-47, McC98c, McC98-71]. S. [Ano96-107, Ano96-111]. S/MIME [McC98c, McC98-71]. sabotage [Mad92b, Nic99-62]. sabre [Col95-151]. Sachs [Hig95x]. Safe [Col95-108, Col95-107, Ano92-92, Ano97o, Ano97-82, Col95-64, Col95-46, Col95-68, Mey96-39, Mey96-53, Mey96-54, Rob91, Han99-74]. safe-guards [Col95-46]. Safeguarding [Col95-109, Mey96-89]. Safeguards [Hig95c, Ha92]. safely [Mey97t]. safer [Col95-127, Han99v]. Safety [Col95-110, McC98-60, Ano97-83, Col95-102, McC97n, McC98-64, Mey96z, Mey96-88, MvS94]. sale [Han99-64, Mey98-51]. sales [Mey96-62]. SALSA [Ano96-103, She96]. Salvage [Hig97q]. Sam [Col96w]. same [Han99-42]. San [Hig95f]. Sandeep [Hig95d, Hig95e]. Sanford [Hig95y]. sans [War92]. SAP [Ano96-104]. SATAN [Col95-111, Col95-112, Ven95]. satellites [McC98b]. Sausalito [Zaj94c]. Save [Mey97-66]. saves [Ano99-48]. say [McC98j, Nic99s]. says [Ano99p, Ano99-62, Col95-148, Han99e, McC98-77, Mey96-45, Nic99-53]. scam [Han99-84, McC98-48]. scams [McC97i]. scan [Hig97j, Hig97k, Mey98r]. scanner [Mey99j]. scanners [Boo92, Hig97t, Zaj94d]. scare [Nic99-44]. Schell [Hig95j, Hig95g]. schematic [VC96]. Scheme [Hig95-35, Ano96-51, Ano97-72, BR96, BM90, CL94, CJB94, CN98, Coh97a, Coh97b, Col95-32, Col96w, DD94, HL90, HTY96, JCW91, Lia94, LL98, Man96, SJJF96]. schemes [Ano97n, Ano97-81, CH94, CHT97, LR92, LV94, YS99]. Schneier [Zaj94a]. School [Ano96-115, Han99-62]. Schultz [Hig95-32]. Science [Ano96d, Ano96k, Ano96-30, Ano96-34, Hig95t, Hig95-47]. Sciences [Ano96j, Ano96-80]. Scientific [Ano92-79]. SCO [Col95-113]. Scrambling [Mey97-67]. Screening [Ano97-73]. Sea [Hig95-44]. search [LR92, LV94, Tab90]. seat [Mey99-42]. Seattle [Mey96-73]. Sebastopol [Hig92a]. SEC [Nic99-55, Hig93i]. SEC’96 [Ano96-60]. SEC’97 [Hig97i]. second [vS96]. secrecy [Mey99n]. Secret [Col95-114, Mey98-57, NSS91, Ano96-97, Ano96-105, Col95-108, Col96h, Col96s, DD94, JJ96, JL96, Mey98-101]. secret-sharing [DD94]. secretary [Hig95n]. secrets [CJC98, Col95w, TCWH95]. sect [Han99l]. sector [Han99-86, Nic99k]. Secure [Ano96-106, Col95-115, Col95-116, Col95-117, DFHR90, Hig95t, Hig95w, McC97r, Mey96-74, Mey97-68, Mey97-69, Mey97-70, Nic99-56, Pre91a, VI97, Wat95, Win90, All91, Ano96d, Ano96i, Ano96-30, Ano96-99, Ano96-111, Ano97a, BR96, Coh96b, Col95-33, Col95-54, Col95-74, Col95-107, Col95-152, Col96q, GW92, Hin99f, Kru91, Laf90, Laf91, Lem96, LC94, McC98i, McC98o, McC98-29, McC98-28, Mey96-29, Mey96-58, Mey96-56, Mey96-60, Mey97-54, Mey97-59, Mey97-61, Mey97-41, Mey98-40, Mey99-55, OvS92, Pou98b, Ram90, Ran92, Ran93, She95c, SJJF96, SGI^{+}99, VSU93, Wis95, Woo90d, Woo93b, Woo94, vS98]. SecureDesktop [Ano97-76]. Securely [Ada96, Ano96-107, Mey98-82, SJ92]. SECURENET [DKGS94]. secures [McC98-79, Mey98d, Mey98h]. Securicor [Col95-118]. Securing [Ano96-108, CM93, Col95-119, Hig95h, Mad95, Mey97-71, Mey97-72, Mey99-40, Sch97a, Sch97b, Ano96-87, Pou97, Hig95p]. Security [Ano92j, Ano92-29, Ano92-45, Ano92-86, Ano92-90, Ano96z, Ano96-50, Ano96-59, Ano96-99, Ano96-104, Ano96-109, Ano97-77, Ano97-78, Ano98-39, Ano99-59, Ano99-60, Ban90, Car90c, Col95o, Col95-121, Col95-123, Col95-120, Col95-131, Col95-122, Col95-124, Col95-125, Col95-126, Col95-127, Col95-128, Col95-129, Col95-130, Col95-162, Col96t, Di 93, Don93, EHT96, Fra96, GGG^{+}94, Han98i, Han98j, Han98k, Han98l, Han98m, Han99-59, Han99-75, Han99-76, Haw99, Hei94a, Hei95b, Hig95s, Hig95c, Hig95o, Hig95q, Hig95y, Hig95-34, Hig95-46, Hig95-45, Hig95-48, Hig95-49, Hig97b, Hig97r, Hug95b, Kov98b, Kow91c, Kru96, LHB90, Lat92, Lun92b, McC97s, McC98-61, McC98-63, McC98-64, McC98-65, McC98-66, Mey96-75, Mey96-76, Mey96-77, Mey96-78, Mey96-79, Mey98-58, Mey98-60, Mey98-61, Mey98-59, Mey98-62, Mey98-63, Mey99-41, Mey99-45, Mey99-42, Mey99-43, MG92, Nic99-58, Nic99-57, Osb95, Pou98e]. Security [Puj93, SG98, Sha94, Str94, Sym94, Thu94, Woo95e, AM95, AJ95a, Ano90m, Ano91-40, Ano92a, Ano92-46, Ano92-63, Ano92-55, Ano92-64, Ano92-66, Ano92-84, Ano93q, Ano96b, Ano96-36, Ano96-35, Ano96-79, Ano96-40, Ano96-81, Ano96-76, Ano96-80, Ano96-78, Ano96-65, Ano96-86, Ano96-98, Ano96-103, Ano96-119, Ano97m, Ano97l, Ano97k, Ano97z, Ano97-34, Ano97-46, Ano97-57, Ano97-58, Ano97-60, Ano97-75, Ano97-74, Ano97-83, Ano99y, Ano99-28, Ass92, BE90, Bai97, Ban95, Ban91, Ben98, Beq98a, Beq98d, BH92, BdAR99, Bir97, BK95, Bla92, BP90, BM90, Buz99, CvdRvZvD98, Car91, Car90a, Cly93, Cog95, Col90, Col95g, Col95t, Col95-31, Col95-40, Col95-43, Col95-54, Col95-60, Col95-61, Col95-59, Col95-66, Col95-70, Col95-76, Col95-30, Col95-79, Col95-82, Col95-84, Col95-88, Col95-93, Col95-94, Col95-106, Col95-141, Col95-65]. security [Col95-134, Col96b, Col96l, Col96m, Col96e, Col96r, Col96w, Col96y, Cur95, Dav92, DBHM95, Dos95, Dun95, EOO95, Fag93, Fer92, Fin96, Fin97a, Fin97b, Fin98a, Fin98b, FH93, Gil92, GTKK91, Gri97, Han95b, Han99f, Han99m, Han99-39, Han99-44, Han99-55, Han99-68, Han99-67, Han99-73, Han99-77, Han99-82, Han99-97, Han95c, Har93, HL90, HS94a, Hei94b, Hei95a, Hei96a, Hei96b, Hei96c, Hig92r, Hig93a, Hig93q, Hig95k, Hig95v, Hig95-47, Hig95-50, Hig97q, HS95b, Hin98e, Hin98f, Hit95, Hop94, Hop97, HS94b, Jac90, Jan92, KJ95, Kap98, Kin92a, Klo90a, Klo90c, Kov96, Kov97c, Kov97e, Kov98a, Kow91a, Kow91b, KL96, LR93, Lat90, Lat93, Lau98, Law92, Lil94, Lin95, Lin96, LR92, McC97d, McC97k, McC97n, McC97o, McC98d, McC98m, McC98-27, McC98-32, McC98-31, McC98-33, McC98-36, McC98-42]. security [McC98-40, McC98-71, McC98-47, McC98-50, McC98-62, McC98-67, McC98-72, McC98-78, McC98-76, Men90a, Men90c, Men91a, Men91c, Men92a, Men92c, Men92f, Mey96c, Mey96d, Mey96q, Mey96i, Mey96r, Mey96u, Mey96-44, Mey96-43, Mey96-46, Mey96-55, Mey96-57, Mey96-62, Mey96-71, Mey96-69, Mey96-72, Mey96-80, Mey96-84, Mey96-82, Mey96-91, Mey96-94, Mey96-96, Mey96-102, Mey96-97, Mey96-98, Mey97j, Mey97s, Mey97x, Mey97y, Mey97-39, Mey97-38, Mey97-42, Mey97-47, Mey97-58, Mey97-79, Mey97v, Mey97-83, Mey97-81, Mey97-89, Mey97-91, Mey98e, Mey98j, Mey98q, Mey98v, Mey98t, Mey98x, Mey98-30, Mey98-31, Mey98-96, Mey98-33, Mey98-44, Mey98-99, Mey98-51, Mey98-46, Mey98-48, Mey98-100, Mey98-54, Mey98-65, Mey98-69, Mey98-68, Mey98-72, Mey98-64, Mey98-76, Mey98-75, Mey98-87, Mey99d, Mey99h, Mey99e, Mey99m, Mey99s, Mey99u, Mey99-29, Mey99z, Mey99-27, Mey99-32, Mey99-38, Mey99-44, Mey99-47, Mey99-48, Mey99-51, Mey99-52, Mey99-53, Mil95, MD94, Mos93, MvS94, MH92, New95, Nic99j]. security [Nic99i, Nic99m, Nic99-32, Nic99-35, Nic99-44, Nic99-46, Nic99-52, Nic99-64, Nor92a, Oba93, Øln93, Øln94, Osb97b, Osb98, Paa91, Par97c, Par95b, Par97e, Pep97, Pfl91, Pla93, PA95, PK91, PvS91, Rat97, Rat98, Rei95a, Rei95b, Rei93, Rih94, Rob92, RR91, Sch91, Sch93, Sch95a, Sch99a, Sch98, She97a, SSP90, She91, She92b, She96, She97b, She97c, Smi95b, Smi95a, Smi96, SL96, Sta93, Svi94, Tan95, TOFS97, Thu91, TR92, VB91, VSU93, VG94, Vin95, WP96, Whi92, Wil97, Wol95, Woo90b, Woo90c, Woo91a, Woo91b, Woo93a, WB93, WS95, Woo95a, Woo95b, Woo97d, Woo97b, Woo97e, WFZ92, Zaj90d, Zaj94b, Zaj94c, ZHM90, dK95, vSvdM94, vS96, vSG98, vSG99, Mey96-70, Ano92-83, Hig92a, Ano96l]. secutity [Col96v]. See [Mey97-91]. seek [Han99-69]. seeks [Col95-157]. seizes [McC98-53]. selected [Han99x]. Selecting [Hig95c, Osb97a]. selection [Ano96-39]. Self [Mey97-73, OvS92]. Self-mutating [Mey97-73]. self-protecting [OvS92]. Selling [Mey96-80, Mey97x]. Senate [McC98-67]. sender [McC97r, Mey97-69]. sending [Ano96-97]. sends [Ano99-47]. sense [Ano92a, Mey98x]. sensors [Mey99m]. sentence [Han99-50, SG94]. Sentry [Col95-132]. separation [Mey97-59]. September [Ano92w, Ano92h, Ano92j, Ano92s, Ano92-47, Ano92-46, Ano92-48, Ano92-85, Ano92-92, Ano92-95, Col95-135, Ano92-72]. sequel [Col95q]. sequenced [DF91]. Serbs [Nic99t]. serious [Col95-28, McC98-44]. seriously [Hin99d]. Server [Mey97-94, Mey99-44, Mey99-45, Ano96-119, Cor95a, Lem96, Mey96-60, Mey96-98, Mey97-67, Mey98r, Mey99-40, Sym94, Woo97e]. Server-based [Mey99-44]. Servers [Mey96-81, Mey97-74, Ano97-85, Col95-56, Wat95]. Service [Han99f, Han99-97, Ano96w, Col95b, Col95-74, Han99d, Mey96w, Sch99b]. services [Col96j, GBD93, Hol95a, Hol95c, McC98-65, Mey99-53, Smi95b, Smi96, SGI^{+}99, Wis95]. Serving [McC98-68]. session [McN95b]. set [Col95o, Eva90, Hig95-27, McC98-54, Mey99i, SPP98, McC98-41, McC98-45, Mey98-32, Mey98-36]. sets [Col96x]. Setting [Hom90, Mey96-82, SD95, Mey98-59]. shadow [McC98a]. shadowed [Col96e]. shadows [Nic99-65]. shaky [McC98-83]. shall [Mey96-42]. Shamir [DD94]. Shannon [KP99]. shaped [Hin99b]. shapes [Col95-88]. shared [CJC98]. sharing [BM90, DD94, TCWH95]. Sharon [Ano97-73, Ano97-35]. sharply [Ano92-29]. Shaw [Hig95t]. Sherizen [Hig95y]. Sherwood [Ano96m, Ano96-103]. Shielding [Mey97-75]. Shifting [Gor95d, Woo95a]. shine [Mey99-43]. ship [Nic99-59]. Shoot [Mey96-83]. Shoot-out [Mey96-83]. shops [Mey99-31]. short [Mey96-94, Mey98-80]. shortages [Pep97]. shortest [FH93]. shot [Mey97d]. Should [McC97k, McC97t, Buz99, Mad92a, Woo90c]. shouldn’t [Mey98o]. show [Han99-77]. shows [Mey97-77]. Sicherheit [SB90]. sickens [Ano92-80]. side [Han99-100, Zaj90d]. Sidewinder [Mey99-45]. Sign [Col95-133, Mey97-76, Ano96-118, Col95-134, Mey96-81]. sign-on [Ano96-118, Col95-134, Mey96-81]. Signature [Bla93, Ano97-72, CL94, CN98, Hol95a, Hol95c, Rih94, Yen96]. signatures [CDR99, Col95-133, Gan94, Han99-40, Law93a, Mey96-27, Mey97v, Mey98o, Pou98c, vSN92]. significant [WB93]. signitures [DFHR90]. Signs [McC98-69, Col95-84]. silver [Cor95b]. Simon [Hig92a]. simple [Ano92-47, Man96, Øln93]. simple-the [Øln93]. simpler [Mey99k, Nic99-67]. Simplified [Col95-134, Mey97-58]. simplify [Mey98-60]. Simplifying [Ano96-29, Mey95]. Simulating [Coh99]. simulation [SD97]. Singapore [Mey97-77]. Single [Mey98-64, Ano96-118, Mey96-81]. SingTel [Han99-44]. SISL [Ano96-98]. site [Han99s, Han99-36, Hig97u, Mey97-41, Mey98-28, Nic99-29, Nic99-48]. sites [Ano99-27, Han99-37, Han99-69, Mil96]. situation [Eva91]. Six [Col95-135, Mey98-65]. Sixth [Ano90m]. sizing [Ano96-102]. skills [Pep97]. sky [Han99-82]. Slam [Mey98-66]. slams [Ano96-38]. Sloan [Ano96-98]. sloppier [HP90b]. sloppiest [HP90b]. Sloppy [HP90b]. slots [RR91]. Slovenia [HS94b]. slow [Ano99-28]. slowly [Mey98-72]. slows [Ano97-86]. smaller [Mey96-60]. Smart [Bov95, SJJF96, Ano96-110, YS99]. smartcard [Hol95a, Hol95c]. smartcards [HS95a, Lea95, Ano96-110, Lai97, McC98-70, Mey97-78, Svi94]. smarter [Col95-51, Lai97]. SMF [Cro95b]. Smith [Ano96-107, Ano97-76, Hig95c, Hig95m, Zaj94b]. Smith-Thomas [Hig95m]. smurfing [McC98-34]. SNA [PdL91]. SNA/SNI [PdL91]. snapshots [Mey99d]. Snell [Ano97-82]. SNI [PdL91]. sniffer [Hig95-32]. SNMP [McC98-82]. snooping [Nic99-37]. snoops [Mey96s, Mey97-55]. Social [Gor95c, Hig95-47, Col95s]. society [Car90c, Han99-40]. SOCKS [Mey98-67]. Software [Ano96-111, Beq98e, Col95-136, Col95-137, Klo90c, Llo90, Mey97-79, Mey97-80, Nic99-59, Nic99-60, SW93, Web95, Zaj94c, vdME98, Ano92-90, Ano99o, Ano99-64, Col95e, Col95n, Col95-57, Col95-106, Col95-126, Col95-147, Col95-143, Col95-156, Col95-157, Gli90, Han99e, HLY92, Hig95e, Hig95-45, LT94, McC98-53, McC98-80, Mey96p, Mey96h, Mey96-47, Mey96-73, Mey97d, Mey97-77, Mey98-37, Nic99-28, Nic99-34, Nic99-47, Nic99-67, PA93, PK98, RR91, She95a, VPC93, Wit92, Zaj92b]. Solar [Hin98g]. solicitors [Col95-42]. Solms [Ano96k]. Solomon [McC98-43]. solution [Ano97a, Ano97-77, Buc95, Buc96, Col95-35, HS95b, Mey96-79, MH97]. solutions [Cor95a, Ken93, Lat93, Mey96-87]. solve [Mey99m]. SOM [Ano96b]. Some [Ano96-112, CPS^{+}98, Spa97b, WFZ92, Coh92a, Dun95, Fåk91, Mey96-71, Mey98-56]. someone [Mey98-38]. Sons [Lat93, Zaj94a]. soon [McC98r]. sophisticated [McC98y]. sorry [Ano99-62]. sought [Col95-33]. source [Bas91, Han99-57, Rus93, Rus94, Zaj94a, vdME98]. South [Ano96k]. space [Ano99-47]. Spafford [Ano96g, Ano96-112, Hig92a, Hig95i, Hig95d, Hig95q, Hig95e]. spam [Mey97-52, Mey97-74, Mey98-66]. SPAN [ZHM90]. Spate [Mey98-103]. spec [Col96g, McC97v, Mey96-55]. special [Hig97g]. species [Col95-99, Hin99b]. specific [Tab90]. specification [Yen96]. specs [McC98-66]. Spectacular [Woo90a]. spectrum [Nac97, Nac98]. speed [She95a, TMO96]. spend [Buz99]. spending [Han99-77]. spie [Col95-62]. spies [Col95-90, Mey96-100]. spills [Han99l]. sponsored [Mad92b]. spoof [Mey97-80]. spoof-proof [Mey97-80]. spook [Buc95, Buc96]. spread [Nic99-49]. Spring [Hig95u]. Sprint [Hor92]. spur [McC98-39, Mey98c]. spy [Ano99p]. Spyrou [Ano96-66]. SRI [Ano96-87, Ano96-109, Hig95u, Hig95-32, Hig95-44]. SSL [Mey97-71]. st [Nos99, Kov99a]. stack [McC98-33]. staff [Woo90c]. stalemate [Mey98-74]. stalking [Mey98-38]. stamp [HTY96, Mey98-87]. Stan [Ano97-84]. stand [Mey98-91]. Standard [Bla93, Men92e, She95a, Ano97l, Col95-87, Fol97, McC97v, McC98-51, Mey98-98, Nic99d, Nic99-69]. Standards [Har95, Hig95-49, Ano92-55, Dav92, Eva90, GGG^{+}94, Han99t, Hom90, Lil94, Par90, Rei93]. Star [Col95z]. starts [Col95-130]. State [Hig95q, McC98-67, Pre99, Ano96-84, Cor97, Lin97]. State-of-the-art [Pre99]. states [Han99-57, Han99-87]. station [Ano92s]. stations [Col95-115]. Statistical [Ano92-87]. Statistics [Han99-77]. status [Ano92-63, Bla93]. stay [Ano90-46, Ano92-97, Ano97-85, BN92]. Steel [Mey98-68]. Steel-cage [Mey98-68]. stenography [Col95-140]. Step [Col95-138, Hin99f, Mey97-42, Mey98-76, Nic99i]. Stephen [Hig95c]. steps [Mey97y, Pou97]. Sterne [Hig95u, Hig95-27]. Stets [Ano97-56]. Steve [Hig95q, Ano96-32]. Still [Col96u, McC98c, Mey97l, Nic99-39]. sting [Nic99y]. stir [Mey96-65]. stock [Col95-51, Fin97a]. Stockholm [Ano96-80]. stole [Ano92-90]. Stoll [Zaj90c]. Stomping [Col95-139]. Stop [Ano97-79, McC97u, Col95-97, Wit92]. storage [FGAR93]. stories [Ano96-76]. story [McC97y]. strain [Mey96-52, Mey97-73]. Strategic [Mey96-84, Par97b, Par97c]. strategies [Ano96-104, Fer92, KP99, Laf90, Mey97-27, Mey97-88]. strategy [Ano96-103, Mey96-97, Mey99z, Par95b, She96, Win90]. Street [Ano97u]. strength [Col95-59, GDNC94, Hig95l]. strict [Mey98-90]. strike [Sol92]. Striking [Mey96-85]. strives [McC98o]. Striving [AZ95]. Strong [Col95-140, Mey99-46]. Stronghold [Mey99-47]. strongroom [Col95-118]. structured [PC99]. struggle [Col96s]. Stuart [Ano96-76]. Students [Han99-78]. Study [McC98-72, AAL^{+}93, Ano92s, BdAR99, Har92, Kan94, LS93, McC98f, McC98-77, Men92b, Mul97b]. stung [Col95-135]. stupid [Ano96c, Ano96-82]. Subject [Ano90-50, Ano91-57, Ano93-58, Ano94-52, Ano95-45, Ano96-113, Ano97-80, Ano98-40, Ano99-61]. subsidiary [McC98-57]. substitution [Col95-36]. subtransport [Ran93]. successful [Men91a]. sucker [Mey96-95]. sue [Han99-100]. sued [Mey96-47]. sufficient [Woo97c]. suggestions [Möh95]. suite [Mey99-34]. Summary [Hig95-44, ILM94]. summer [Han99-49]. Summit [Col95-141]. Sun [Mey98-69]. superhighway [Fra95]. Suppliers [Ano92-88, Col95f]. supplying [LC94]. support [CKM91, Fin97b, Hig95e, PG94, Hig95q]. suprises [For99b]. surf [Mey98-93]. surfer [Col95-83]. surfers [Nic99-45]. surfing [Ano97-82, AJ97]. SurfinGate [McC98t]. Surrey [Ano96i]. surrogates [Rih94]. survey [Ano96w, Ano97-60, CvdRvZvD98, CM92, DMC92, DFIR99, Han99a, Lun93, Mey97-77, VSU93]. surveys [Hin98f, Mul97c]. Survival [Dev99]. Survive [Ano96-33, Mey98-89]. Surviving [Sch99b]. Sushil [Hig95n]. suspects [Col95-104]. Suspicious [Mey98-70]. Suzanne [Hig95c]. SWAT [Mey99-41]. Sweden [Ano96-80]. Swedish [Col95-142]. sweep [Ano92k]. Swiss [Ano97-60, Col96v]. Swiss-cheese [Col96v]. Switzerland [Ano96-34, Hig95k, Ano96-30]. SWORD [Wis95]. Sydney [Hig95p]. Symantec [Mey98-104]. Symmetric [DG98, PC99]. symmetric-key [PC99]. Syntegra [Ano96-98]. System [Hig95s, Hig95t, Kap98, Koe93, Kru96, SD97, AJ95c, Ano96-29, Ano96-32, Ano96-50, Ano97a, Ano97p, Ano97-58, BK95, Bor96, Coh93, Coh97a, Coh97b, Col95v, Col95-44, Col95-85, Col95-86, Col95-84, Col95-148, DKGS94, GKM^{+}97, Gil92, Han99-51, Han99-90, HLY92, HJS^{+}93, Jon97a, Jon97b, KJ95, LE97, McC98d, McC98-29, MS90, Mey95, Mey97b, Mey98-34, Mit95, Nic99-52, O’S91, PRZ99, PvS91, Rei93, SSP90, Ste91, TO91, TMO96, VB91, Hig95h, Hig95m]. system-access [LE97]. Systems [Ano96-80, Ano96-64, Hig95f, Hig95q, Hig95r, Hig95u, Hig95x, Hig95y, Hig95-34, Hig97b, Kov98b, Ano96a, Ano96d, Ano96l, Ano96-40, Ano96-98, Ano96-111, Ano96-108, Ano96-115, Ano97a, Ano97m, Ano97p, Ano97x, Ano97-38, Ano97-46, Ban91, Bla92, BP90, CPS^{+}98, Col90, Col95-31, Col95-123, CKM91, DBHM95, Fin97b, GBD93, GS91, GTKK91, GKKT92, Gri97, Han99-28, Han99-73, Hei95a, Hig95-27, Hig95-28, Hig95z, Hig95-47, Hru95, HS94b, ILM94, Jon97a, Jon97b, Kem97a, Kem97b, Klo90a, Kov97c, Kov97e, Kru95b, Kru91, Laf90, Lia94, Llo90, Lun92b, MS90, Mer90b, Mey96k, Mey97u, Mey97-59, Mey97-82, Mey98-62, Mey99-52, Nic99-33, Par90, PA95, Pre91a, Ram90, Rei95a, Rei95b, SJ91, She92b, SU93, Smi95b, Smi96, Sym94, TO91, Tay95, TOFS97, Thu91, Thu94, VSU93, WP96, Win90, Woo90d, Woo93b, Woo95a, Ano96-110]. systems [TR92, Ano92-84]. systems-III [TR92]. systems-outline [Ano97a].

T [Ano96-107]. T. [Ano96-66]. table [Hig95-48]. Tabuchi [Hig95k]. tackles [Col95-77]. Tackling [Mey97-81]. Tadjer [Ano97-47]. Tailor [Mer90b]. Tailor-made [Mer90b]. Taipei [Hig95m]. Taiwan [Han99c, Hig95m]. take [Col95-27, Col95-50]. takes [Mey96s, Mey98-54, Mey99-42, Pou98b]. Taking [Col95-143, McC98-73, Nos99, Nic99q]. Tales [Ano96-114, Col95-144, Nic99-61, Gor99]. talks [Han99-95]. Tankard [Ano96-108]. tap [Ano97-56]. Tapping [Mey99-48]. Target [McC98-74, Col95-38, Col95-104, Col96d, McC98-57]. targets [Col95-57, Han99-28, McC97o, Mey98s, Nic99y]. Tarsala [Ano97-60]. Taschenlexicon [SB90]. task [Han99-60]. tax [Col95-44]. Taxman [Ano99-62]. Taxonomy [Hig95-49, Ano96-115, Ano97-81, Hig95f]. TC [Hig92b, Hig92c, Hig93c, Hig93d, Hig93e, Hig93f, Hig93g, Hig93h, Hig94a, Hig95-30, Hig95-31, Hig96a, Hig97a, Hig97h, Ano91-39]. TC11 [Ano98n, Hig92d, Hig92e, Hig92f, Hig92g, Hig92h, Hig92i]. TCP [Ano96-106, Hig95-50]. TCP/IP [Hig95-50]. team [Col95-111, Mey96j, Mey96-41]. teams [Mey99-43, RH91]. teardrop [Mey98-41]. tech [Ano92-46, Beq98d, Col95w, McC97i, McC97j, McC98x]. Technical [Ano91-32, Ano91-33, Ano91-35, Ano93-36, Ano93-38, Ano93-39, Ano95t, Ano95w, Ano95x, Ano96-58, Ano97-39, Ano97-45, Ano98s, Ano98t, Ano98u, Ano99-32, Ano99-33, Ano99-35, Ano90x, Ano90y, Ano90z, Ano90-27, Ano90-28, Ano90-29, Ano90-30, Ano91-31, Ano91-34, Ano91-36, Ano91-37, Ano91-38, Ano92-49, Ano92-50, Ano92-51, Ano92-52, Ano92-53, Ano92-54, Ano93-33, Ano93-34, Ano93-35, Ano93-37, Ano93-40, Ano94-28, Ano94-29, Ano94-30, Ano94-31, Ano94-32, Ano94-33, Ano94-34, Ano94-35, Ano95u, Ano95v, Ano95y, Ano95z, Ano95-27, Ano96-81, Ano96-52, Ano96-53, Ano96-54, Ano96-55, Ano96-56, Ano96-57, Ano96-59, Ano97-40, Ano97-41, Ano97-42, Ano97-43, Ano97-44, Ano98o, Ano98p, Ano98q, Ano98r, Ano98v, Ano99-29, Ano99-30, Ano99-31, Ano99-34, Ano99-36]. techniques [Ano96z, Cro95a, Gal95, Han99-85, Lun93, Hig95f]. Techno [Beq98f]. Techno-crimes [Beq98f]. Technologically [Gor95d]. technologies [Ano97-83, Bol91, Bov95, LE97, Ric99]. Technology [Ano96-80, Hig95-49, Mey98-71, AJ95a, Ano97m, Ano97w, Ano99-47, BE94, GGG^{+}94, Mey96g, Mey97-33, Mey97-49, MG92, Rob99]. Ted [Hig95s]. Teen [Han99-79, Col95-78]. Telcos [Ano99-63]. Telecommunications [Ano96-116, Col99a, Col99b, Hoa97, Ano96-98, Hig92r, Hig95s]. Telecoms [Mey97-82]. telematic [Ano97-46, Hor92]. telematics [TOFS97]. telemedical [SGI^{+}99]. Telephone [Gor99, Mey96-86, Nic99z]. telephones [Mey96-56]. telephony [Mul97a]. Teleworking [Mey96-87, Ano96-99]. ten [Lau98]. tense [Mey99-33]. Teresa [Hig95-44]. Termination [Ano99-64]. terminations [Woo93a]. Terrorism [Mey96-88, FW99]. TESPAR [KP99]. test [Hig95-47, Nic99-63]. testbed [Han99-45]. Testing [Ano96-117, Col95-145, Har97, Jon97a, Jon97b, Osb97a, Hig95r]. tests [Nic99-54]. text [Col95-157]. textual [LHB90]. TGB [Hig95-46]. thanks [Mey98u]. theatened [Ano92-84]. Theft [Han99-80, Ano96-28, Ano99-49, Gil92, Han99-43, McC98-57, Mey96p, Smu90]. their [Bol95, Gan94]. them [Bon96]. theorem [Bou97, HTY96]. theoretical [Col95-140, Fåk91]. theory [Ano97-72, Fin97a, Ran92]. There [Dun95, Han99-100, McC97x, Mey98-86]. thief [McC97i, McC97u, Mey99-50]. thieves [Col95-98, Han99-79, Mey97-35]. thin [Ano99-64]. things [Hig95z, Hin99b]. Think [Col95-146, Mey99-49, Col95-127, Col95-146]. thinking [AJ95a]. third [Ano96-86, Ano97n, Mey96-55, Sch97a, SGI^{+}99]. Thomas [Hig95m, Hig95-32, Lat93, Ano97-36]. Threat [Hig95s, Nic99-62, Ano92-66, Ano92-89, Ano92-95, Ano96-116, Ano99-64, Bru95, Car90a, Col95-153, Dav96, Hig97n, Mag94, McC98-34, Mey96-31, Mey96-86, Mey97-92, Mey98-97, Mey98-50, Mey99-31, Nic99-38, Nic99-65, Smu90, Zaj92a]. threats [Bov95, FW99, Hig97s, Hin98c, Koe93, Mey96-87, Mey97-75]. Three [Hig97g, Col96w, Fin98b, Han99-65]. three-tier [Col96w]. threshold [Ano96-51]. thresholds [SD95]. throughout [Hop94]. thru [Mul95]. thwart [McC98-54, Mey97-55]. Thwarting [McC98-75]. tier [Col96w]. tightrope [Ano92-81]. Tim [Ano96-64]. Time [Col95-147, Hin98h, Sch98, Ano97o, Col95-125, Han98i, HTY96, Kin92b, LE98, LT94, Mey97w, Mey99b, Mey99p, VPC93]. Time-based [Sch98]. Times [Ano92s, Ano92-92, Ano97-37, Ano97-79, Ano97-82, For98a, Mey96x].

Tips [Mey96-89]. tires [McC98-37]. TIS [Mey97-83]. Titans [Mey98-74]. title [McC98-71]. today [Bir97, Sch97b]. token [Mey98-39]. told [Ha92, McC98d]. tolls [Hin99b]. Tom [Ano96-116]. tomorrow [Hig93q]. toned [Ano92-95]. too [Mey98w]. tool [Han99g, Mey98-99, Mey98-63, Mey99-31, Hig95v]. tooled [Mey96-57]. toolkits [Ano92-47]. Tools [Ano97-82, Cro95a, Ano96z, Col95-149, Gor95b, Han99-54, Lem96, McC98-66, Mey97-28, Mey99-53, Oli95a, VE98]. Top [Ano99-65, Col95-148, Nic99-63]. topic [McC97f]. TOPM [BE94]. tops [Mey97-77]. touch [Mey97-89]. Tough [Mey96-90, Faw92a, McC98-46]. toughens [Col95j]. toy [McC98y]. trace [Col96o]. track [McC98x, SW93]. trackers [Col95-47]. Tracking [Mey97-84]. tracks [Col96o]. Tracy [Ano92s]. Trade [Ano92-27, Hei93c, McC98-76, Nic99p, vS98]. tradeshows [Han98j]. trading [Nic99-55, Rat99b]. traditional [Hit95]. traffic [Hig95v]. Traigh [Hig95s]. trail [Col95-53, Cro95b, Lun92a, Men90b]. trails [Nic99b]. training [Eva91]. Transaction [Hig95m, Muf90, Col95-74]. transactions [Ano96-107, Hig95-33, Mey97-61, Mor99]. Transarc [Mey96-91]. transfer [Pre91a, TMO96]. transform [SS93]. Traps [Rei95a, Rei95b]. treasure [Col95-137]. trend [Han99-51]. trial [Col95-45, Mey97-56, Mey99q]. Trials [McC97w]. triangle [CL96]. tribulations [McC97w]. tricks [Col95m, McC98g, Mey96w]. tries [Han99-67]. trio [Hig97f]. trip [Ano96-50, Kru96]. triple [Mey97b, Mey96-66]. triple-DES [Mey96-66]. TriStrata [Mey98-76]. trouble [McC98-33, Mey97-84]. Troxell [Hig95v]. Troy [For99a]. Trust [Hig95-29, Mey97-85, Ano97-38, BB98, Mey96-42, Mey97-93, Mey98-71, Nic99-64, Ran92, Rat98, Rat99a, Rat99b]. Trusted [AJ95b, AJ95c, All91, Hig95u, SGI^{+}99, Ano96-86, Ano97n, Han99-51, Hig95j, Hig95g, Hig95-27, LC94]. trustworthy [BB98]. truth [Ano96-118]. try [Col95m, Mey98-52]. TSS [dV92]. tuned [Mey97l]. turf [Mey98-93]. turn [Ano96-40, McC98-76, Mey96k]. turns [Col96k, Mey96-34]. TV [Ano92s]. twice [Mey99-49]. Twinkle [Ano99-66]. twists [Mey97-54]. Two [GKM^{+}97, CJB94, Col95-27, Mey98-82]. Two-phase [GKM^{+}97]. two-way [Mey98-82]. Typeface [Smi92]. typical [Coh91a]. typing [VI97].

U [McC98-76]. U-turn [McC98-76]. U.S [Hig92a, Kow91c]. U.S. [Han99-85]. U.S.A [Hig92a]. U.S.S.R. [Lat90]. UK [Hig95-35, Ano92-27, Ano96i, Ano96w, Bul99, Cor97, Han99-65, Han99-81, Han99-82, Hic97a, Hic97b, Hig95-35, Jon97b, McC97l, Mey96-86, Mey98-83, Nic99-64, Pou98a, Pou99a]. uncertain [Mey98m]. Uncle [Col96w]. uncommon [Mey97-89]. uncovered [Mey96-77]. uncovers [Nic99f]. undercover [Col95-135]. Underground [Col95-149, Ano92y]. underlying [Ano96-111]. undermine [Sch91]. understand [Kov97d]. Understanding [Ban91, Men90c, Men92f, Whi92]. Undesirable [Hig95h]. unification [She91]. Union [Pou97]. unique [Ano96-51, Han99-83]. unit [Col95-32, Han99-67, Mey98-64]. unite [Mey98-47]. United [Jon97a]. Universita [Hig95n]. University [Ano96d, Ano96i, Ano96j, Ano96k, Ano96l, Ano96-30, Ano96-34, Ano96-66, Ano96-80, Hig95i, Hig95l, Hig95c, Hig95n, Hig95d, Hig95q, Hig95e, Hig95p, Hig95r, Hig95t, Hig95-28, Hig95-32, Hig95-33, Hig95-50, Ano92-89, Ano96g, Ano96-112, Hig95a]. UNIX [CRR92, Hig92a, Ano96-32, Col95-88, Col95-117, Faw92a, Fol97, Mey96-69, She97a]. unlawful [Han99e]. unlock [Mey97-43, Mey98u]. Unlocking [Mey97-86, Mey98-77]. unnecessary [Pea91]. unobservability [BP90]. Unravelling [Col95-150]. unsaid [Ano96c]. unveiled [Col96j]. unveils [Col95-56, McC98-28, Mey97-42]. upconversion [Bon99]. Update [Bla92, Lin97, Mey97b, Web95, AJ95b, Hin99g]. upgrade [Mey97-56]. Upgrades [Hig95c]. uphill [Mey97-81]. upon [Ano97-72, CL94, SPP98]. ups [Mey98t, Mey98-33]. urged [Nic99i]. urges [Ano99t]. Ursula [SB90]. USA [Hig95j, Hig95k, Hig95g, Hig95-27, Zaj94b, Zaj94c, Hig95f, Hig95i, Hig95j, Hig95l, Hig95g, Hig95e, Hig95v, Hig95-28, Hig95z, Hig95-32, Hig95-33, Hig95-47, Hig95-45, Hig95-50]. Usage [Hug95a, BS94, Han99z, Hol95b, Hug95b, Law93a]. Use [Han99w, Han99-27, Han99-99, Hol95d, LE98, Mey99-46, PK98, Rob91, dRE96]. used [Han99-79, Mey98-37]. useful [Thi91]. User [Ano92-92, Mey97-87, Ano96-66, Ano99-44, CH94, CHT97, Col95-34, Col96o, Fåk91, HMCC95, Lau92, Lyn91, MS90, Mey97-72, Mey98-60, MD94, MvS94, PMRZ99, Woo95a]. user-based [PMRZ99]. user-role-based [MD94]. Users [Ano99-68, Mey96-94, Mey99-51, Ano92-66, Ano96-63, CJB94, Col95-62, Col95-84, Col96y, Han99-44, McC98g, Mey96-47, Mey96-90, Mey97-55]. uses [Ano92-89, Gan94]. Using [Hig95q, Lem96, She95c, Woo91b, Ano92-85, Ano96j, CRR92, Hol95a, Hol95c, LE97, McC98-34, OvS92, TO91, vSvdM94, Hig95u, Hig95v]. usurped [Nic99-66]. UT [Hig95j, Hig95l, Hig95g]. Utahn [Col95z]. utilities [Hig91c, Mey96-101]. UX [CRR92].

V [Ano96i, Ano96-51, Col95-58]. VA [Hig95j, Hig95s, Hig95n, Hig95g, Hig95z, Hig95-45, Ano96-107, Hig95b, Hig95h, Hig95s, Hig95o, Hig95-46]. vaccine [Ano92-65]. validation [LV92]. validity [LH94]. Value [BP90, Mit95, Fin97a, Par97b, PRK98, SPP98]. values [Par97c]. Vance [Ano97-74]. vandals [McC98-34]. vanish [Col95-158]. varying [Ano96l]. vault [Ano97-76]. VAX [Whi92]. VAX/VMS [Whi92]. veil [McC98m]. Veirs [Hig95h]. vendor [Sch98, Woo95a]. vendor/publisher [Woo95a]. Vendors [Mey99-52, Mey99-53, Col95g, Klo90c, Smi92]. verifiable [HS95a]. Verification [Hig95t, Col95-87, Fåk91, Rus93, Rus94]. verifier [Yen96]. Verisign [McC98-79]. VERLAG [SB90]. Vernon [Ano96w]. version [Pou99b]. versus [DD92, Hig95-35]. Veto [Nic99-65]. via [BK95, Han99-85, LC94, Mer90a, Mey96-71, Nic99-41]. viable [HS95b, Kim95]. victim [Ano96g, Col95-71, Spa97a]. video [Col95v]. vie [McC98-71]. View [Hig95u, Hig95w, Hig95-44, Ano97-83, Ano99-60, Hig91d, Hig93q, Hin99b, Sym94, Woo97b]. Views [Han98k, Ano98-39, Han98l, Han98m, Han99-76]. Vijay [Hig95p]. violations [Han99-36, McC97k]. Viral [Ano92-93, Zaj92b, Zaj92a, Zaj94d]. Virginia [Han99-78]. Virtual [Mey98-79, Ano96-106, Ano96-109, Fra96, Mey96-36, Mey97-86, Mey98-77, Nic99-46, Nic99-66, Hig95w]. virtually [Han99h]. virtuous [Ano92-94]. viru [Col95-81]. Virus [Ano92-96, Ano92-95, Ano97-84, Ano99-69, Bon99, Col95-153, Col95-154, Hig97t, McC98-80, Mey97-88, Mey98-80, Mey98-81, Tay92, AD90, Ano92-47, Ano92-44, Ano92-65, Ano92-88, Ano92-89, Ano92-94, Ano96f, Ano97s, AJ97, Bon96, Bon98, Car90a, Col95e, Col95-29, Col95-57, Col95-73, Col95-97, Col95f, Col95-146, Col95-160, Dav90a, Dav96, Han99-58, Hig97n, JW90, McC98f, Mey96e, Mey96-51, Mey96-52, Mey97d, Mey98b, Mey98-50, Mey98-104, Mey99g, Mey99l, Mey99n, Mey99-37, Nic99-41, Nic99-42, Nic99-49, Nic99-50, Nic99-63, PK98, Pre91b, Sol92, Ano92-47]. virus-free [Mey96-51]. Virus-proof [Tay92, Col95-146]. viruse [Mey97o]. Viruses [Ano92-96, Ano92-97, Ano92-98, Bru95, Mey96-95, Ano90-46, Ano96-31, Ano97-79, Coh91a, Coh92b, Col95-89, DMGK95, Hig97e, Hig97f, Hig97g, Hru95, JATC93, Ken93, Mag94, McC97h, McC97k, Mey96f, Mey96h, Mey96l, Mey96-48, Mey97-73, Mey97-92, Mey98-55, Mey99f, Mey99-41, Nic99-28, Thi91, Zaj90b]. Viruses-here [Ano92-97]. Virusnet [Ano96-75]. Viruswall [Mey96-49]. Visa [Han99-57]. visas [DD92]. vistas [Ano97-58]. visual [vdME98]. vital [Jan92]. VLANs [Nic99-66]. VMS [Can90, Whi92]. Voas [LT94]. vogue [Han99-48]. Vol [Ano92-55, Ano92-84, Ano90i, Ano90-50, Ano91i, Ano91-57, Ano93h, Ano93-58, Ano94b, Ano94-52]. volume [Ano95a, Ano95-45, Ano96e, Ano96-113, Ano97b, Ano97-80, Ano98a, Ano98-40, Ano99c, Ano99-61]. volumes [Hig91e, Hig96e, Hig97l, Hig97m]. voting [BR96]. vouching [YA96]. voyage [Hin98h]. VPN [McC98-79, Mey97-87, Nic99-67, Nic99-68]. VPNS [Mey97-54, Mey97-89, Han99-47, Han99-92, Han99-95, Mey97-70, Mey98d, Mey98-40, Mey98-82, Mey99k, Mey99-38, Nic99-40]. Vradharajan [Hig95p]. vs [FW95, Han99j, Mey96l, Mey96-88, PK91]. vulnerabilities [Ano96-112, Gor95b, Koe93, Men90c, Spa97b]. Vulnerability [Klo90d, Nic99-48, Ano96l]. vulnerable [Han99-65, Nic99-52].

W [Hig95-45, Lat93]. wages [Nic99-47]. Wake [BGK92, Ano99-63]. wakes [Nic99-42]. Walking [Mey97-90]. Wall [Ano97u, Mey97-90]. wallet [Col95-33, Nic99-46]. walls [Fol97, San92]. WAN [CKM91]. Wang [Hig95m]. WANK [LS93]. wanna [Mey98-56]. WANs [Col95-122]. Want [Han99-93, McC98-81, Mey97-91, Ano96-31, Han99-39]. War [Col95-155, Ano96-105, BGK92, FW95, Han99l, Hin98h, Nic99-47]. Warburg [Ano96h, Ano96-77]. Ward [Ano92-87]. warehousing [Fin97b]. warfare [Ano96-64, Col95-37, Cor97, Han99-41, Kov97e, Mad92b, Por96]. warn [Mey97-57]. warned [Ano99x, Mey99-51]. Warning [Mey98-83]. warns [Col95-125, McC97i, Nic99-55]. warrant [Tab90]. Wars [Hin98c, Col95-50, Col95-154, Nic99q]. was [Mey97a]. Washington [Hig95-28, McC98s]. waste [Hil95]. Watch [Mey97-92, Han99-48, Mey96-73, Mey98-84]. watchdog [Col95w]. Watching [Mey98-84, Nic99e]. water [Hin98d]. Watermark [Col95-31]. wave [Ano92-84]. Wavner [Ano97-85]. way [Dos97, Kov99b, Man96, McC98e, Mey96-76, Mey97-52, Mey97-65, Mey98-82]. Wayne [Ano92-90, Hig95-49]. ways [Woo94]. Weak [McC98-82, Spa92]. weakend [Ano92-76]. weakness [Ano92-45, McC98-51, Mey96-64]. weaknesses [Sch93]. weapon [Hig96d]. web [Col95-56, Han99u, Han99-37, Han99-52, Han99-69, Han99-99, Hig96c, Hig96f, Hig96d, Hig97s, Hig97y, Mey96-68, Mey97k, Mey97-44, Mey97-55, Mey98-29, Mil96, Ano96-119, Ano97-82, Ano97-85, Coh96b, Col96i, Col96t, Hig97u, McC98-83, Mey96c, Mey96-28, Mey96-96, Mey96-97, Mey96-98, Mey97-72, Mey97-41, Mey98-33, Mey99-54, Ano97-74]. webmaster [Han99-36]. webs [BB98]. Webscan [Col96x]. Weekly [Ano92a, Ano92-76, Ano92-81, Ano92-88, Ano92-89]. Weigh [Nic99-69]. Weiler [SB90]. well [Han99-71]. West [Hig95i, Hig95d, Hig95q, Hig95e]. Western [Hig95p]. WG [Ano91-40]. Where [Mul97b, Mul97c, Mul97d, Mul97e, Mul97f, Mul97g, Ano97-55]. whilst [AJ97]. White [Col95-157, Mey96-99, Mey98-85, Rob99, Han99w, Hug99]. white-hat [Han99w]. Who [Han98n, McC97x, Mey98-86, Cla97, Ano92j]. Whom [Mey97-93, Mey98-75]. Wichers [Hig95x]. Wide [Coh96b, Mey96-28, Han99-99]. widen [McC98-31]. widespread [Han99-99]. Wilding [Ano96z]. Wiley [Lat93, Zaj94a]. Will [Col95-158, Mey96-100, Mey99-55, Col95-97, Han99-77, Hin99b, McC98r, Zaj90a]. William [Hig95c]. Williams [Ano92-83, Ano96-101]. Wilson [Hig95m]. Win [Col95-77, Col95-137, Col96y, Han99-100]. window [Har92]. Windows [Han98o, Han99-96, Mey99-40, Bir97, Col95-89, Cur95, Han99-55, Han99-97, Hig97w, McC98-83, Mey96-101, Mey96-102, Mey98-68, Mey99-31, Nic99-30, Sch97b, Sch99a]. Windows-2000 [Han99-96]. Windows-98 [Han98o]. Windows-based [Hig95h]. Windows-NT [Han99-97]. Winkler [Hig95-47]. Winners [Col95-159]. winning [Sch95b]. Wireless [McC97y, Woo95b, Lat92]. wires [Col95-158]. Wiretaps [Han99-98]. wise [Nic99-28]. within [Ano99w, Ass92, CvdRvZvD98, Gor97, Haw99, Hig96b, Jon97a, Jon97b, Mey96-50, Mul97a]. without [Ano96-102]. woes [Beq99b, Col95-79, Han99-97]. wonder [Ame97]. Wood [Ano96-78, Ano96-85, Zaj94c]. Word [Col95-73, Col95-160, Mey96e, Mey97-73]. WordMacro [Bon99]. words [Col95u]. work [Ano96a, Ano97-37, Mey96-27, Mey99x]. workbench [KL96]. worked [Hei93a]. workers [Han99b]. working [Col95-51]. workplace [Han99-80]. works [Mey98-30]. Workstation [Col95-161, GW92, Sta93]. World [Ano92-66, Ano97-84, Col95-162, Han99-99, Hei93c, Mey96-28, Ano92-64, Han99y, Hin99f, Mey96q, Mey97-78, Mey99-55, Nic99-42, Coh96b]. World-Wide-Web [Coh96b]. WorldCom [Mey99-27, Nic99-40]. WorldSecure [Mey97-94]. Worldwide [Ano96-101]. worms [Coh92a, LS93]. worried [Mey96z]. worries [Han99-85]. would [Mey96-59]. wrapped [Ano99-69]. wrinkles [Mey98-36]. writers [Ano92-44, Col95-97]. writers-are [Ano92-44]. Writing [Woo95c, Woo95d]. WSS [Mey98-87]. WWW [SGI^{+}99].

X.25 [LR93, Mos93]. X.400 [Hug95b, Kin92a].

Y2000 [Mad98]. Y2K [Ano99-43, Ano99-65, McC98j, Mey98-62, Mey99-56, Nic99k, Nic99m, Nic99p, Nic99-39, Nic99-54, Nic99-65, Nos99]. Yazdaman [Hig95w]. Year [Gor95d, Hei96c, Hig97x, Hin99g, McC98-55, Ano96h, Ano99-62, Joa97, McC98w, Mey98i, Woo90a]. years [Mey97a]. Yeuh [Hig95m]. Yngstrom [Ano96-80]. York [Lat93]. Young [Hig95l, Han99-36]. you’re [Col95-146]. yourself [Car90b, Kov97b, Nic99-56].

Zergo [Ano96-86, Ano96-76, Ano96-118]. zero [GKG92]. Zhang [Hig95r, Hig95t]. zone [FW95]. zones [Mey97-47]. Zurich [Ano96-30].

References

Abrams:1993:RIR

[AAL^{+}93] Marshall D. Abrams, Edward G. Amoroso, Leonard J. La-Padula, Teresa F. Lunt, and James G. Williams. Report of an integrity research study group. Computers & Security, 12(7):679–689, November 1993. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

             sciencedirect.com/science/article/pii/016740489390085J.

Allinson:1993:MCB

[ABD93] Caroline Allinson, Helen Bergen, and Ed Dawson. Message collision in block ciphers with message authentication. Computers & Security, 12(8):781–787, December 1993. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

             tronic). URL https://www.sciencedirect.com/science/
             article/pii/0167404893900446.

Al-Dossary:1990:CVP

[AD90] Ghannam M. Al-Dossary. Computer virus prevention and containment on mainframes. Computers & Security, 9(2):131– 137, April 1990. CODEN CPSEDU. ISSN 0167-4048 (print),

             1872-6208 (electronic). URL https://www.sciencedirect.
             com/science/article/pii/0167404890900858.

Adams:1996:OBS

[Ada96]  Dean Adams. Open for business — securely!  Com-

puters & Security, 15(8):673–682, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404897897003.

Abrams:1995:NTA

[AJ95a] Marshall D. Abrams and Michael V. Joyce. New thinking about information technology security. Computers & Security, 14(1): 69–81, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/0167404895970278.

Abrams:1995:TCU

[AJ95b] Marshall D. Abrams and Michael V. Joyce. Trusted computing update. Computers & Security, 14(1):57–68, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/0167404895970267.

Abrams:1995:TSC

[AJ95c] Marshall D. Abrams and Michael V. Joyce. Trusted system concepts. Computers & Security, 14(1):45–56, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/0167404895970256.

Aubrey-Jones:1997:CMV

[AJ97] David Aubrey-Jones. Contracting a macro virus whilst surfing the Internet. Computers & Security, 16(6):518, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404897846743.

Allen:1991:TOM

[All91] R. J. Allen. Trusted ORACLE-multilevel secure data management for military computing. Computers & Security, 10(3):271– 275, May 1991. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/016740489190044E.

Abrams:1995:HLC

[AM95] Marshall D. Abrams and Jonathan D. Moffett. A higher level of computer security through active policies. Computers & Security, 14(2):147–157, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/016740489597048F.

Amery:1997:LBW

[Ame97] Chris Amery. Look back in wonder. Computers & Security, 16 (6):516, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404897846676.

Anderson:1992:LE

[And92] Ross Anderson. Letter to the editor. Computers & Security, 11(8):710, December 1992. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404892900915.

Anonymous:1990:ARAd

[Ano90a] Anonymous. Abstract of recent articles and literature. Computers & Security, 9(4):349–358, June 1990. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/0167404890901064.

Anonymous:1990:ARAa

[Ano90b] Anonymous. Abstracts of recent articles and literature. Computers & Security, 9(1):87–96, February 1990. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/016740489090162M.

Anonymous:1990:ARAb

[Ano90c] Anonymous. Abstracts of recent articles and literature. Computers & Security, 9(2):177–187, April 1990. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/0167404890900928.

Anonymous:1990:ARAc

[Ano90d] Anonymous. Abstracts of recent articles and literature. Computers & Security, 9(3):265–273, May 1990. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/016740489090174R.

Anonymous:1990:ARAe

[Ano90e] Anonymous. Abstracts of recent articles and literature. Computers & Security, 9(5):449–458, August 1990. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/0167404890900766.

Anonymous:1990:ARAf

[Ano90f] Anonymous. Abstracts of recent articles and literature. Computers & Security, 9(6):549–556, October 1990. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/016740489090133E.

Anonymous:1990:ARAg

[Ano90g] Anonymous. Abstracts of recent articles and literature. Computers & Security, 9(7):647–654, November 1990. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/016740489090064Z.

Anonymous:1990:ARAh

[Ano90h] Anonymous. Abstracts of recent articles and literature. Computers & Security, 9(8):739–747, December 1990. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/016740489090117C.

Anonymous:1990:AIR

[Ano90i] Anonymous. Author index for refereed articles in vol. 9. Computers & Security, 9(8):751, December 1990. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/016740489090119E.

Anonymous:1990:Ca

[Ano90j] Anonymous. Calendar. Computers & Security, 9(2):189–191, April 1990. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/0167404890900939.

Anonymous:1990:Cb

[Ano90k] Anonymous. Calendar. Computers & Security, 9(5):461–463, August 1990. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/0167404890900788.

Anonymous:1990:Cc

[Ano90l] Anonymous. Calendar. Computers & Security, 9(8):749–750, December 1990. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/016740489090118D.

Anonymous:1990:CAS

[Ano90m] Anonymous. Conference announcement: Sixth annual computer security applications conference. Computers & Security, 9(5): 464, August 1990. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/0167404890900799.

Anonymous:1990:CR

[Ano90n] Anonymous. Conference roundup. Computers & Security, 9 (6):513–516, October 1990. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/016740489090128G.

Anonymous:1990:EB

[Ano90o] Anonymous. Editorial Board. Computers & Security, 9(1):2, February 1990. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/016740489090147L.

Anonymous:1990:Ea

[Ano90p] Anonymous. From the editor. Computers & Security, 9(1):2–3, February 1990. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/016740489090148M.

Anonymous:1990:Ec

[Ano90q] Anonymous. From the editor. Computers & Security, 9(2):100– 101, April 1990. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/016740489090080D.

Anonymous:1990:Ed

[Ano90r] Anonymous. From the editor. Computers & Security, 9(3):194– 195, May 1990. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/016740489090163N.

Anonymous:1990:Ee

[Ano90s] Anonymous. From the editor. Computers & Security, 9(4):280– 281, June 1990. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/016740489090094A.

Anonymous:1990:Ef

[Ano90t] Anonymous. From the editor. Computers & Security, 9(5):364– 365, August 1990. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/0167404890900652.

Anonymous:1990:Eg

[Ano90u] Anonymous. From the editor. Computers & Security, 9(6):468– 469, October 1990. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/0167404890901219.

Anonymous:1990:Eh

[Ano90v] Anonymous. From the editor. Computers & Security, 9(7): 568–569, November 1990. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/016740489090053V.

Anonymous:1990:Ei

[Ano90w] Anonymous. From the editor. Computers & Security, 9(8): 658–659, December 1990. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404890901075.

Anonymous:1990:ITCa

[Ano90x] Anonymous. IFIP technical committee 11. Computers & Security, 9(1):85–86, February 1990. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/016740489090161L.

Anonymous:1990:ITCb

[Ano90y] Anonymous. IFIP technical committee 11. Computers & Security, 9(2):175–176, April 1990. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404890900917.

Anonymous:1990:ITCc

[Ano90z] Anonymous. IFIP technical committee 11. Computers & Security, 9(3):263–264, May 1990. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/016740489090173Q.

Anonymous:1990:ITCd

[Ano90-27] Anonymous. IFIP technical committee 11. Computers & Security, 9(4):347–348, June 1990. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404890901053.

Anonymous:1990:ITCe

[Ano90-28] Anonymous. IFIP technical committee 11. Computers & Security, 9(5):447–448, August 1990. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/0167404890900755.

Anonymous:1990:ITCf

[Ano90-29] Anonymous. IFIP technical committee 11. Computers & Security, 9(7):645–646, November 1990. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/016740489090063Y.

Anonymous:1990:ITCg

[Ano90-30] Anonymous. IFIP technical committee 11. Computers & Security, 9(8):737–738, December 1990. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/016740489090116B.

Anonymous:1990:IBRa

[Ano90-31] Anonymous. International board of referees. Computers & Security, 9(1):51, February 1990. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/016740489090156N.

Anonymous:1990:IBRb

[Ano90-32] Anonymous. International board of referees. Computers & Security, 9(2):143, April 1990. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/016740489090087A.

Anonymous:1990:IBRc

[Ano90-33] Anonymous. International board of referees. Computers & Security, 9(3):233, May 1990. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/016740489090169T.

Anonymous:1990:IBRe

[Ano90-34] Anonymous. International board of referees. Computers & Security, 9(6):517, October 1990. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/016740489090129H.

Anonymous:1990:IBRf

[Ano90-35] Anonymous. International board of referees. Computers & Security, 9(7):611, November 1990. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404890900593.

Anonymous:1990:IBRg

[Ano90-36] Anonymous. International board of referees. Computers & Security, 9(8):697, December 1990. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404890901127.

Anonymous:1990:IBRd

[Ano90-37] Anonymous. International board of referres. Computers & Security, 9(5):409, August 1990. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/016740489090071Z.

Anonymous:1990:PAa

[Ano90-38] Anonymous. Pages 100–191 (April 1990). Computers & Security, 9(2):??, April 1990. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:1990:PM

[Ano90-39] Anonymous. Pages 194–273 (May 1990). Computers & Security, 9(3):??, May 1990. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:1990:PF

[Ano90-40] Anonymous. Pages 2–96 (February 1990). Computers & Security, 9(1):??, February 1990. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:1990:PJ

[Ano90-41] Anonymous. Pages 280–358 (June 1990). Computers & Security, 9(4):??, June 1990. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:1990:PAb

[Ano90-42] Anonymous. Pages 364–464 (August 1990). Computers & Security, 9(5):??, August 1990. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:1990:PO

[Ano90-43] Anonymous. Pages 468–561 (October 1990). Computers & Security, 9(6):??, October 1990. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:1990:PN

[Ano90-44] Anonymous. Pages 568–654 (November 1990). Computers & Security, 9(7):??, November 1990. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:1990:PDb

[Ano90-45] Anonymous. Pages 658–756 (December 1990). Computers & Security, 9(8):??, December 1990. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:1990:PDC

[Ano90-46] Anonymous. Panel discussion: Are computer viruses here to stay. Computers & Security, 9(4):305–307, June 1990. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/016740489090098E.

Anonymous:1990:PDa

[Ano90-47] Anonymous. Panel discussion-into the 1990s. Computers & Security, 9(1):37–40, February 1990. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/016740489090153K.

Anonymous:1990:PLE

[Ano90-48] Anonymous. Preliminary list of exhibitors for Compsec International ’90. Computers & Security, 9(6):559–561, October 1990. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/016740489090135G.

Anonymous:1990:RAI

[Ano90-49] Anonymous. Refereed articles: International board of referees. Computers & Security, 9(4):317, June 1990. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/016740489090101X.

Anonymous:1990:SIR

[Ano90-50] Anonymous. Subject index for refereed articles in vol. 9. Computers & Security, 9(8):753–756, December 1990. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/016740489090120I.

Anonymous:1991:ARAa

[Ano91a] Anonymous. Abstracts of recent articles and literature. Computers & Security, 10(1):78–83, February 1991. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/016740489190060Q.

Anonymous:1991:ARAb

[Ano91b] Anonymous. Abstracts of recent articles and literature. Computers & Security, 10(2):163–177, April 1991. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/016740489190010B.

Anonymous:1991:ARAc

[Ano91c] Anonymous. Abstracts of recent articles and literature. Computers & Security, 10(3):277–284, May 1991. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/016740489190045F.

Anonymous:1991:ARAd

[Ano91d] Anonymous. Abstracts of recent articles and literature. Computers & Security, 10(4):373–378, June 1991. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/016740489190113R.

Anonymous:1991:ARAe

[Ano91e] Anonymous. Abstracts of recent articles and literature. Computers & Security, 10(5):469–481, August 1991. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/016740489190072L.

Anonymous:1991:ARAf

[Ano91f] Anonymous. Abstracts of recent articles and literature. Computers & Security, 10(6):569–579, October 1991. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/016740489190085R.

Anonymous:1991:ARAg

[Ano91g] Anonymous. Abstracts of recent articles and literature. Computers & Security, 10(7):675–682, November 1991. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/016740489190127Y.

Anonymous:1991:ARAh

[Ano91h] Anonymous. Abstracts of recent articles and literature. Computers & Security, 10(8):779–788, December 1991. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/016740489190097W.

Anonymous:1991:AIV

[Ano91i] Anonymous. Author index for vol. 10. Computers & Security, 10(8):791, December 1991. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/016740489190099Y.

Anonymous:1991:Ca

[Ano91j] Anonymous. Calendar. Computers & Security, 10(1):84–86, February 1991. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/016740489190061H.

Anonymous:1991:Cb

[Ano91k] Anonymous. Calendar. Computers & Security, 10(2):178– 180, April 1991. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/0167404891900112.

Anonymous:1991:Cc

[Ano91l] Anonymous. Calendar. Computers & Security, 10(3):285– 287, May 1991. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/016740489190046G.

Anonymous:1991:Cd

[Ano91m] Anonymous. Calendar. Computers & Security, 10(4):379– 380, June 1991. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/016740489190114S.

Anonymous:1991:Ce

[Ano91n] Anonymous. Calendar. Computers & Security, 10(6):580–581, October 1991. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/016740489190086S.

Anonymous:1991:Cf

[Ano91o] Anonymous. Calendar. Computers & Security, 10(7):683–684, November 1991. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/016740489190128Z.

Anonymous:1991:Cg

[Ano91p] Anonymous. Calendar. Computers & Security, 10(8):789, December 1991. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/016740489190098X.

Anonymous:1991:CI

[Ano91q] Anonymous. Compsec International 90. Computers & Security, 10(1):75–77, February 1991. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/016740489190059M.

Anonymous:1991:EB

[Ano91r] Anonymous. Editorial Board. Computers & Security, 10(1):2, February 1991. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/016740489190047H.

Anonymous:1991:EN

[Ano91s] Anonymous. Editor’s note. Computers & Security, 10(7):611, November 1991. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/016740489190118W.

Anonymous:1991:Ea

[Ano91t] Anonymous. From the editor. Computers & Security, 10(1):2– 3, February 1991. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/016740489190048I.

Anonymous:1991:Ec

[Ano91u] Anonymous. From the editor. Computers & Security, 10(2):90– 91, April 1991. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/016740489190001T.

Anonymous:1991:Ed

[Ano91v] Anonymous. From the editor. Computers & Security, 10 (3):182–183, May 1991. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/016740489190028C.

Anonymous:1991:Ee

[Ano91w] Anonymous. From the editor. Computers & Security, 10 (4):290–291, June 1991. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/016740489190102J.

Anonymous:1991:Ef

[Ano91x] Anonymous. From the editor. Computers & Security, 10 (5):382–383, August 1991. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/016740489190063J.

Anonymous:1991:Eg

[Ano91y] Anonymous. From the editor. Computers & Security, 10 (6):484–485, October 1991. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/016740489190073M.

Anonymous:1991:Eh

[Ano91z] Anonymous. From the editor. Computers & Security, 10(7): 584–585, November 1991. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/016740489190115T.

Anonymous:1991:Ei

[Ano91-27] Anonymous. From the editor. Computers & Security, 10(8): 688–689, December 1991. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/016740489190087T.

Anonymous:1991:GAa

[Ano91-28] Anonymous. Guide for authors. Computers & Security, 10(1): 87, February 1991. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/016740489190062I.

Anonymous:1991:GAb

[Ano91-29] Anonymous. Guide for authors. Computers & Security, 10 (7):685, November 1991. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404891901292.

Anonymous:1991:GAc

[Ano91-30] Anonymous. Guide for authors. Computers & Security, 10 (8):798, December 1991. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/016740489190101I.

Anonymous:1991:ITCa

[Ano91-31] Anonymous. IFIP technical committee 11. Computers & Security, 10(1):73–74, February 1991. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/016740489190058L.

Anonymous:1991:ITCb

[Ano91-32] Anonymous. IFIP technical committee 11. Computers & Security, 10(2):161–162, April 1991. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404891900093.

Anonymous:1991:ITCc

[Ano91-33] Anonymous. IFIP technical committee 11. Computers & Security, 10(3):269–270, May 1991. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/016740489190043D.

Anonymous:1991:ITCd

[Ano91-34] Anonymous. IFIP technical committee 11. Computers & Security, 10(4):371–372, June 1991. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/016740489190112Q.

Anonymous:1991:ITCe

[Ano91-35] Anonymous. IFIP technical committee 11. Computers & Security, 10(5):467–468, August 1991. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/016740489190071K.

Anonymous:1991:ITCf

[Ano91-36] Anonymous. IFIP technical committee 11. Computers & Security, 10(6):561–562, October 1991. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/016740489190082O.

Anonymous:1991:ITCg

[Ano91-37] Anonymous. IFIP technical committee 11. Computers & Security, 10(7):671–672, November 1991. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/016740489190125W.

Anonymous:1991:ITCh

[Ano91-38] Anonymous. IFIP technical committee 11. Computers & Security, 10(8):777–778, December 1991. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/016740489190096V.

Anonymous:1991:ITR

[Ano91-39] Anonymous. IFIP/TC 11 report. Computers & Security, 10 (6):563–564, October 1991. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/016740489190083P.

Anonymous:1991:IWM

[Ano91-40] Anonymous. IFIP/WG 11.1: Management of information security. Computers & Security, 10(7):673–674, November 1991. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/016740489190126X.

Anonymous:1991:IBRa

[Ano91-41]  Anonymous.  International Board of Referees.  Com-

puters & Security, 10(1):35, February 1991. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/016740489190053G.

Anonymous:1991:IBRb

[Ano91-42] Anonymous. International board of referees. Computers & Security, 10(2):115, April 1991. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/016740489190005X.

Anonymous:1991:IBRc

[Ano91-43] Anonymous. International board of referees. Computers & Security, 10(4):329, June 1991. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/016740489190108P.

Anonymous:1991:IBRd

[Ano91-44] Anonymous. International board of referees. Computers & Security, 10(5):411, August 1991. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/016740489190067N.

Anonymous:1991:IBRe

[Ano91-45] Anonymous. International board of referees. Computers & Security, 10(6):533–534, October 1991. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/016740489190078R.

Anonymous:1991:IBRf

[Ano91-46] Anonymous. International board of referees. Computers & Security, 10(8):725–726, December 1991. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/016740489190092R.

Anonymous:1991:PM

[Ano91-47] Anonymous. Pages 182–287 (May 1991). Computers & Security, 10(3):??, May 1991. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:1991:PF

[Ano91-48] Anonymous. Pages 2–87 (February 1991). Computers & Security, 10(1):??, February 1991. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:1991:PJ

[Ano91-49] Anonymous. Pages 290–380 (June 1991). Computers & Security, 10(4):??, June 1991. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:1991:PAb

[Ano91-50] Anonymous. Pages 382–481 (August 1991). Computers & Security, 10(5):??, August 1991. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:1991:PO

[Ano91-51] Anonymous. Pages 484–582 (October 1991). Computers & Security, 10(6):??, October 1991. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:1991:PN

[Ano91-52] Anonymous. Pages 584–685 (November 1991). Computers & Security, 10(7):??, November 1991. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:1991:PD

[Ano91-53] Anonymous. Pages 688–798 (December 1991). Computers & Security, 10(8):??, December 1991. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:1991:PAa

[Ano91-54] Anonymous. Pages 90–180 (April 1991). Computers & Security, 10(2):??, April 1991. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:1991:RAa

[Ano91-55] Anonymous. Refereed articles. Computers & Security, 10(3): 237, May 1991. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/016740489190039G.

Anonymous:1991:RAb

[Ano91-56] Anonymous. Refereed articles. Computers & Security, 10 (7):637, November 1991. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/016740489190121S.

Anonymous:1991:SIV

[Ano91-57] Anonymous. Subject index for vol. 10. Computers & Security, 10(8):793–797, December 1991. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/016740489190100R.

Anonymous:1992:FSS

[Ano92a] Anonymous. ‘A false sense of security’: Jason Hobby, Computer Weekly, 8 October 1992, p. 11. Computers & Security, 11(8):726, December 1992. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404892901217.

Anonymous:1992:ARAa

[Ano92b] Anonymous. Abstracts of recent articles and literature. Computers & Security, 11(1):93–103, March 1992. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/016740489290224F.

Anonymous:1992:ARAb

[Ano92c] Anonymous. Abstracts of recent articles and literature. Computers & Security, 11(2):185–192, April 1992. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/016740489290045S.

Anonymous:1992:ARAc

[Ano92d] Anonymous. Abstracts of recent articles and literature. Computers & Security, 11(3):279–287, May 1992. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/0167404892902089.

Anonymous:1992:ARAd

[Ano92e] Anonymous. Abstracts of recent articles and literature. Computers & Security, 11(4):375–383, July 1992. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/016740489290178T.

Anonymous:1992:ARAe

[Ano92f] Anonymous. Abstracts of recent articles and literature. Computers & Security, 11(6):535–542, October 1992. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/016740489290188W.

Anonymous:1992:ARAf

[Ano92g] Anonymous. Abstracts of recent articles and literature. Computers & Security, 11(7):629–636, November 1992. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/016740489290141D.

Anonymous:1992:AHA

[Ano92h] Anonymous. After Hurricane Andrew: Gary H. Anthes, Computerworld, 7 September 1992, p. 6. Computers & Security, 11(8):725, December 1992. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/0167404892901158.

Anonymous:1992:B

[Ano92i] Anonymous. Bombed out. Computers & Security, 11(5):488, September 1992. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/016740489290025M.

Anonymous:1992:BRW

[Ano92j] Anonymous. Book review: Who’s on the line?: Richard Heffernan, Security Management, September 1992, p. 59 and 60. Computers & Security, 11(8):726, December 1992. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/0167404892901239.

Anonymous:1992:BSB

[Ano92k] Anonymous. BSA sweep on Benelux region. Computers & Security, 11(5):486–487, September 1992. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/016740489290019N.

Anonymous:1992:Ca

[Ano92l] Anonymous. Calendar. Computers & Security, 11(1):104–105, March 1992. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/016740489290225G.

Anonymous:1992:Cb

[Ano92m] Anonymous. Calendar. Computers & Security, 11(3):288– 289, May 1992. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/016740489290209A.

Anonymous:1992:Cc

[Ano92n] Anonymous. Calendar. Computers & Security, 11(4):384– 385, July 1992. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/016740489290179U.

Anonymous:1992:Cd

[Ano92o] Anonymous. Calendar. Computers & Security, 11(5):493–494, September 1992. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/016740489290260X.

Anonymous:1992:Ce

[Ano92p] Anonymous. Calendar. Computers & Security, 11(6):543–544, October 1992. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/016740489290189X.

Anonymous:1992:Cf

[Ano92q] Anonymous. Calendar. Computers & Security, 11(7):637, November 1992. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/016740489290142E.

Anonymous:1992:Cg

[Ano92r] Anonymous. Calendar. Computers & Security, 11(8):729, December 1992. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/016740489290127D.

Anonymous:1992:CSB

[Ano92s] Anonymous. Case study: Boston TV station: Tracy Mitchell, LAN Times, 14 September 1992, p. 16. Computers & Security, 11(8):724, December 1992. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/0167404892901136.

Anonymous:1992:CFR

[Ano92t] Anonymous. City firms rally round after IRA bombing. Computers & Security, 11(5):485, September 1992. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/016740489290014I.

Anonymous:1992:CRC

[Ano92u] Anonymous. Conference report: Compsec international ’91. Computers & Security, 11(1):16–17, March 1992. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/016740489290214C.

Anonymous:1992:DM

[Ano92v] Anonymous. Dirty macs. Computers & Security, 11(5):485, September 1992. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/016740489290015J.

Anonymous:1992:AAF

[Ano92w] Anonymous. DOS to add antivirus features: James Daly, Computerworld, 7 September 1992, p. 4. Computers & Security, 11(8):723–724, December 1992. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404892901073.

Anonymous:1992:EB

[Ano92x] Anonymous. Editorial Board. Computers & Security, 11(1): 2, March 1992. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/0167404892902119.

Anonymous:1992:EUP

[Ano92y] Anonymous. Electronic underground: Privacy Journal, October 1992. Computers & Security, 11(8):722, December 1992. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/016740489290099D.

Anonymous:1992:EAa

[Ano92z] Anonymous. Enemy action. Computers & Security, 11(5):490, September 1992. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/016740489290245M.

Anonymous:1992:EMF

[Ano92-27] Anonymous. Evidence of market failure of UK’s computer misuse act?: Dealing with computer misuse — published by the Department of Trade and Industry, p. 24. Computers & Security, 11(8):723, December 1992. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/016740489290104Y.

Anonymous:1992:FCM

[Ano92-28] Anonymous. FAST corporate membership. Computers & Security, 11(5):487, September 1992. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/016740489290023K.

Anonymous:1992:FLR

[Ano92-29] Anonymous. French losses rise sharply: Paul Gannon, Computer Fraud and Security Bulletin, October 1992, p.

3. Computers & Security, 11(8):724, December 1992. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/0167404892901114.

Anonymous:1992:Ea

[Ano92-30] Anonymous. From the editor. Computers & Security, 11(1):2– 3, March 1992. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/016740489290212A.

Anonymous:1992:Ec

[Ano92-31] Anonymous. From the editor. Computers & Security, 11 (2):108–109, April 1992. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/016740489290034O.

Anonymous:1992:Ed

[Ano92-32] Anonymous. From the editor. Computers & Security, 11 (3):198–199, May 1992. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/016740489290196X.

Anonymous:1992:Ee

[Ano92-33] Anonymous. From the editor. Computers & Security, 11(4): 292–293, July 1992. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/016740489290168Q.

Anonymous:1992:Ef

[Ano92-34] Anonymous. From the editor. Computers & Security, 11 (5):388–389, September 1992. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404892900018.

Anonymous:1992:Eg

[Ano92-35] Anonymous. From the editor. Computers & Security, 11 (6):496–497, October 1992. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/016740489290181P.

Anonymous:1992:Eh

[Ano92-36] Anonymous. From the editor. Computers & Security, 11(7): 590–591, November 1992. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/016740489290134D.

Anonymous:1992:Ei

[Ano92-37] Anonymous. From the editor. Computers & Security, 11(8): 680–681, December 1992. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404892900856.

Anonymous:1992:GAa

[Ano92-38] Anonymous. Guide for authors. Computers & Security, 11(1): 106, March 1992. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/016740489290226H.

Anonymous:1992:GAb

[Ano92-39] Anonymous. Guide for authors. Computers & Security, 11(3): 290, May 1992. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/016740489290210I.

Anonymous:1992:GAc

[Ano92-40] Anonymous. Guide for authors. Computers & Security, 11(4): 386, July 1992. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/016740489290180Y.

Anonymous:1992:GAd

[Ano92-41] Anonymous. Guide for authors. Computers & Security, 11(6): 588, October 1992. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/016740489290195W.

Anonymous:1992:GAe

[Ano92-42] Anonymous. Guide for authors. Computers & Security, 11 (7):678, November 1992. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/016740489290148K.

Anonymous:1992:GAf

[Ano92-43] Anonymous. Guide for authors. Computers & Security, 11 (8):768, December 1992. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/016740489290133C.

Anonymous:1992:HVW

[Ano92-44] Anonymous. Hackers and virus writers-are they outlaws or heroes? Computers & Security, 11(5):489, September 1992. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/016740489290032M.

Anonymous:1992:HDP

[Ano92-45] Anonymous. Hackers demonstrate phone banking weakness: Information Security Monitor, October 1992, p 1. Computers & Security, 11(8):721, December 1992. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/0167404892900948.

Anonymous:1992:HTM

[Ano92-46] Anonymous. High-tech movie highlights realistic security concerns: James Daly, Computerworld, September 14 1992, p. 1 and 16. Computers & Security, 11(8):727, December 1992. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/016740489290126C.

Anonymous:1992:HDS

[Ano92-47] Anonymous. How dangerous are simple virus toolkits?: Virus Bulletin, September 1992, p. 3. Computers & Security, 11 (8):726, December 1992. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/016740489290120G.

Anonymous:1992:IDD

[Ano92-48] Anonymous. IBM/DEC disaster recovery deal: Computerworld, 28 September 1992, p. 20. Computers & Security, 11 (8):725, December 1992. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404892901147.

Anonymous:1992:ITCa

[Ano92-49] Anonymous. IFIP technical committee 11. Computers & Security, 11(1):91–92, March 1992. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/016740489290223E.

Anonymous:1992:ITCb

[Ano92-50] Anonymous. IFIP technical committee 11. Computers & Security, 11(4):371–373, July 1992. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/016740489290177S.

Anonymous:1992:ITCc

[Ano92-51] Anonymous. IFIP technical committee 11. Computers & Security, 11(5):481–483, September 1992. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/016740489290013H.

Anonymous:1992:ITCd

[Ano92-52] Anonymous. IFIP technical committee 11. Computers & Security, 11(6):585–587, October 1992. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/016740489290194V.

Anonymous:1992:ITCe

[Ano92-53] Anonymous. IFIP technical committee 11. Computers & Security, 11(7):675–677, November 1992. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/016740489290147J.

Anonymous:1992:ITCf

[Ano92-54] Anonymous. IFIP technical committee 11. Computers & Security, 11(8):765–767, December 1992. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/016740489290132B.

Anonymous:1992:ISS

[Ano92-55] Anonymous. The importance of security standards: John Hopkinson, Computer Control Quarterly, Vol. 7, No. 3, 1992, pp. 33–37. Computers & Security, 11(8):725–726, December 1992. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/016740489290119C.

Anonymous:1992:IBRa

[Ano92-56] Anonymous. International board of referees. Computers & Security, 11(1):39, March 1992. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/016740489290219H.

Anonymous:1992:IBRb

[Ano92-57] Anonymous. International board of referees. Computers & Security, 11(2):147, April 1992. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/016740489290041O.

Anonymous:1992:IBRc

[Ano92-58] Anonymous. International board of referees. Computers & Security, 11(3):251, May 1992. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404892902045.

Anonymous:1992:IBRd

[Ano92-59] Anonymous. International board of referees. Computers & Security, 11(4):345, July 1992. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/016740489290174P.

Anonymous:1992:IBRe

[Ano92-60] Anonymous. International board of referees. Computers & Security, 11(6):545, October 1992. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404892901903.

Anonymous:1992:IBRf

[Ano92-61] Anonymous. International board of referees. Computers & Security, 11(7):639, November 1992. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/016740489290143F.

Anonymous:1992:IBRg

[Ano92-62]  Anonymous.  International Board of Referees.  Com-

puters & Security, 11(8):731, December 1992. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/016740489290128E.

Anonymous:1992:ISE

[Ano92-63] Anonymous. IT security evaluation manual-the current status. Computers & Security, 11(5):492, September 1992. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/016740489290257R.

Anonymous:1992:LSD

[Ano92-64] Anonymous. Lax security at Drug Enforcement Administration: Gary H. Anthes, Computer world, 5 October 1992, p.

24. Computers & Security, 11(8):724, December 1992. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/0167404892901125.

Anonymous:1992:LAV

[Ano92-65] Anonymous. Leap in anti-virus vaccine: Computing, 22 October 1992, p. 2. Computers & Security, 11(8):722, December 1992. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/016740489290098C.

Anonymous:1992:NUF

[Ano92-66] Anonymous. Novell users face threat to security: Carlyn Gilooly/Bob Brown, Network World, 5 October 1992, pp. 1 and 66, 12 October 1992, pp. 13 and 14. Computers & Security, 11(8):723, December 1992. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/016740489290105Z.

Anonymous:1992:NAC

[Ano92-67] Anonymous. NSA approves CD-ROM encryption. Computers & Security, 11(5):487, September 1992. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/016740489290021I.

Anonymous:1992:PA

[Ano92-68] Anonymous. Pages 108–192 (April 1992). Computers & Security, 11(2):??, April 1992. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:1992:PMb

[Ano92-69] Anonymous. Pages 198–290 (May 1992). Computers & Security, 11(3):??, May 1992. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:1992:PMa

[Ano92-70] Anonymous. Pages 2–106 (March 1992). Computers & Security, 11(1):??, March 1992. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:1992:PJ

[Ano92-71] Anonymous. Pages 292–386 (July 1992). Computers & Security, 11(4):??, July 1992. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:1992:PS

[Ano92-72] Anonymous. Pages 388–494 (September 1992). Computers & Security, 11(5):??, September 1992. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:1992:PO

[Ano92-73] Anonymous. Pages 496–588 (October 1992). Computers & Security, 11(6):??, October 1992. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:1992:PN

[Ano92-74] Anonymous. Pages 590–678 (November 1992). Computers & Security, 11(7):??, November 1992. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:1992:PD

[Ano92-75] Anonymous. Pages 680–768 (December 1992). Computers & Security, 11(8):??, December 1992. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:1992:PPL

[Ano92-76] Anonymous. Planned privacy law weakend: David Evans, Computer Weekly, 15 October 1992, p. 1. Computers & Security, 11(8):724, December 1992. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404892901084.

Anonymous:1992:PCM

[Ano92-77] Anonymous. Police crack major hacker ring. Computers & Security, 11(5):485, September 1992. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/016740489290016K.

Anonymous:1992:PIF

[Ano92-78] Anonymous. Project ideas fail to catch $18 million funding: Computing, 15 October 1992, p. 8. Computers & Security, 11(8):724, December 1992. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404892901095.

Anonymous:1992:QCC

[Ano92-79] Anonymous. Quantum cryptography: Charles H. Bennett, Gilles Brassard and Arthur K. Ekert Scientific American, October 1992, pp. 26–33. Computers & Security, 11(8):725, December 1992. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/016740489290117A.

Anonymous:1992:RSD

[Ano92-80] Anonymous. Rain sickens data centre. Computers & Security, 11(5):488, September 1992. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/016740489290028P.

Anonymous:1992:RPF

[Ano92-81] Anonymous. Recession puts firms on disaster tightrope: Jason Hobby, Computer Weekly, 22 October 1992, p. 1. Computers & Security, 11(8):724, December 1992. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/016740489290110D.

Anonymous:1992:RAI

[Ano92-82] Anonymous. Refereed articles international board of referees. Computers & Security, 11(5):443, September 1992. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/016740489290010O.

Anonymous:1992:RMC

[Ano92-83] Anonymous. Review: Management control and security: Jean Hitchings and Bernard Williams Management Accounting, October 1992, pp. 34 and 35. Computers & Security, 11(8):725, December 1992. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/0167404892901169.

Anonymous:1992:RPS

[Ano92-84] Anonymous. Review: PC security theatened by new wave operating systems: David Hains. Computer Control Quarterly, Vol. 10, No. 2, 1992, pp. 8–10. Computers & Security, 11(8):722, December 1992. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/016740489290101V.

Anonymous:1992:RIR

[Ano92-85] Anonymous. Review: The increasing risks of using e-mail: James Carroll The Bottom Line, September 1992, pp. 21 and 22. Computers & Security, 11(8):722–723, December 1992. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/016740489290103X.

Anonymous:1992:RCD

[Ano92-86] Anonymous. Roles changed in downsized environment: Computer Security Alert, October 1992. Computers & Security, 11(8):722, December 1992. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/016740489290102W.

Anonymous:1992:SCR

[Ano92-87] Anonymous. Statistical confidentiality: Reg Ward Government Computing, October 1992, p. 6. Computers & Security, 11(8):726, December 1992. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404892901228.

Anonymous:1992:SBV

[Ano92-88] Anonymous. Suppliers are biggest virus carriers: Computer Weekly, 29 October 1992, p. 17. Computers & Security, 11 (8):721, December 1992. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/016740489290097B.

Anonymous:1992:ULU

[Ano92-89] Anonymous. University lecturer uses virus threat to blackmail: David Evans. Computer Weekly, 8 October 1992, p.

1. Computers & Security, 11(8):721, December 1992. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/0167404892900959.

Anonymous:1992:UJD

[Ano92-90] Anonymous. US Justice Department ‘stole’ software: Wayne Madsen. Computer Fraud and Security Bulletin, October 1992, pp. 1 ϵ #38; 2. Computers & Security, 11(8):722, December 1992. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/0167404892901006.

Anonymous:1992:UPD

[Ano92-91] Anonymous. US pays dear for data loss: Computing, 29 October 1992, p. 4. Computers & Security, 11(8):721, December 1992. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/016740489290096A.

Anonymous:1992:UGS

[Ano92-92] Anonymous. User’s guide to safe computing: Laura Didio, LAN Times, 14 September 1992, p. 1, 110 and 111. Computers & Security, 11(8):727, December 1992. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/016740489290125B.

Anonymous:1992:VMG

[Ano92-93] Anonymous. Viral mind games. Computers & Security, 11 (5):489, September 1992. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/016740489290030U.

Anonymous:1992:VVE

[Ano92-94] Anonymous. The virtuous virus: The Economist, 17 October 1992, p. 92. Computers & Security, 11(8):725, December 1992. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/016740489290118B.

Anonymous:1992:VTT

[Ano92-95] Anonymous. Virus threat toned down: James Daly, Computerworld September 14 1992, p. 16. Computers & Security, 11(8):723, December 1992. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404892901062.

Anonymous:1992:VAV

[Ano92-96] Anonymous. Viruses and anti-viruses in Israel: Ephraim Brand, Virus Bulletin, October 1992. Computers & Security, 11(8):726–727, December 1992. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/016740489290124A.

Anonymous:1992:VHS

[Ano92-97] Anonymous. Viruses-here to stay. Computers & Security, 11 (5):489, September 1992. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/016740489290031L.

Anonymous:1992:VIC

[Ano92-98] Anonymous. Viruses infect 37 computers. Computers & Security, 11(5):488, September 1992. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/016740489290026N.

Anonymous:1993:ARAa

[Ano93a] Anonymous. Abstracts of recent articles and literature. Computers & Security, 12(1):37–42, February 1993. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/016740489390009T.

Anonymous:1993:ARAb

[Ano93b] Anonymous. Abstracts of recent articles and literature. Computers & Security, 12(2):143–149, March 1993. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/016740489390096N.

Anonymous:1993:ARAc

[Ano93c] Anonymous. Abstracts of recent articles and literature. Computers & Security, 12(3):263–269, May 1993. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/016740489390114K.

Anonymous:1993:ARAd

[Ano93d] Anonymous. Abstracts of recent articles and literature. Computers & Security, 12(5):459–464, August 1993. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/016740489390065D.

Anonymous:1993:ARAe

[Ano93e] Anonymous. Abstracts of recent articles and literature. Computers & Security, 12(6):556–560, October 1993. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/0167404893900516.

Anonymous:1993:ARAf

[Ano93f] Anonymous. Abstracts of recent articles and literature. Computers & Security, 12(7):647–652, November 1993. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/016740489390080O.

Anonymous:1993:ARAg

[Ano93g] Anonymous. Abstracts of recent articles and literature. Computers & Security, 12(8):740–744, December 1993. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/0167404893900376.

Anonymous:1993:AIV

[Ano93h] Anonymous. Author index for vol. 11. Computers & Security, 12(2):199, March 1993. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/016740489390102B.

Anonymous:1993:Ca

[Ano93i] Anonymous. Calendar. Computers & Security, 12(1):45–47, February 1993. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/0167404893900103.

Anonymous:1993:Cb

[Ano93j] Anonymous. Calendar. Computers & Security, 12(2):151–153, March 1993. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/016740489390097O.

Anonymous:1993:Cc

[Ano93k] Anonymous. Calendar. Computers & Security, 12(3):270– 272, May 1993. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/016740489390115L.

Anonymous:1993:Cd

[Ano93l] Anonymous. Calendar. Computers & Security, 12(4):371– 372, June 1993. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/016740489390024Y.

Anonymous:1993:Ce

[Ano93m] Anonymous. Calendar. Computers & Security, 12(5):465–466, August 1993. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/016740489390066E.

Anonymous:1993:Cf

[Ano93n] Anonymous. Calendar. Computers & Security, 12(6):561–562, October 1993. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/0167404893900527.

Anonymous:1993:Cg

[Ano93o] Anonymous. Calendar. Computers & Security, 12(7):653–654, November 1993. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/016740489390081F.

Anonymous:1993:Ch

[Ano93p] Anonymous. Calendar. Computers & Security, 12(8):745–746, December 1993. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/0167404893900387.

Anonymous:1993:CSG

[Ano93q] Anonymous. Computers and security: Guide for authors. Computers & Security, 12(1):102, February 1993. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/016740489390016X.

Anonymous:1993:EB

[Ano93r] Anonymous. Editorial Board. Computers & Security, 12(1):2, February 1993. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/016740489390001L.

Anonymous:1993:Ea

[Ano93s] Anonymous. From the editor. Computers & Security, 12(1):2– 3, February 1993. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/016740489390002M.

Anonymous:1993:Ec

[Ano93t] Anonymous. From the editor. Computers & Security, 12 (2):104–105, March 1993. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/016740489390088M.

Anonymous:1993:Ed

[Ano93u] Anonymous. From the editor. Computers & Security, 12 (3):214–215, May 1993. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/016740489390106F.

Anonymous:1993:Ee

[Ano93v] Anonymous. From the editor. Computers & Security, 12 (4):320–321, June 1993. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/016740489390017Y.

Anonymous:1993:Ef

[Ano93w] Anonymous. From the editor. Computers & Security, 12 (5):422–423, August 1993. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/016740489390059E.

Anonymous:1993:Eg

[Ano93x] Anonymous. From the editor. Computers & Security, 12 (6):516–517, October 1993. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404893900457.

Anonymous:1993:Eh

[Ano93y] Anonymous. From the editor. Computers & Security, 12(7): 610–611, November 1993. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/016740489390073E.

Anonymous:1993:Ei

[Ano93z] Anonymous. From the editor. Computers & Security, 12(8): 696–697, December 1993. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/016740489390031Y.

Anonymous:1993:GAa

[Ano93-27] Anonymous. Guide for authors. Computers & Security, 12(2): 211, March 1993. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/016740489390105E.

Anonymous:1993:GAb

[Ano93-28] Anonymous. Guide for authors. Computers & Security, 12(3): 318, May 1993. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/016740489390121K.

Anonymous:1993:GAc

[Ano93-29] Anonymous. Guide for authors. Computers & Security, 12(4): 419, June 1993. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/0167404893900309.

Anonymous:1993:GAd

[Ano93-30] Anonymous. Guide for authors. Computers & Security, 12(5): 514, August 1993. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/016740489390072D.

Anonymous:1993:GAe

[Ano93-31] Anonymous. Guide for authors. Computers & Security, 12(6): 608, October 1993. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/016740489390058D.

Anonymous:1993:GAf

[Ano93-32] Anonymous. Guide for authors. Computers & Security, 12 (7):694, November 1993. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/016740489390087L.

Anonymous:1993:ITCa

[Ano93-33] Anonymous. IFIP technical committee 11. Computers & Security, 12(1):99–101, February 1993. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/016740489390015W.

Anonymous:1993:ITCb

[Ano93-34] Anonymous. IFIP technical committee 11. Computers & Security, 12(2):207–209, March 1993. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/016740489390104D.

Anonymous:1993:ITCc

[Ano93-35] Anonymous. IFIP technical committee 11. Computers & Security, 12(3):315–317, May 1993. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/016740489390120T.

Anonymous:1993:ITCd

[Ano93-36] Anonymous. IFIP technical committee 11. Computers & Security, 12(4):373–375, June 1993. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/016740489390025Z.

Anonymous:1993:ITCe

[Ano93-37] Anonymous. IFIP technical committee 11. Computers & Security, 12(5):511–513, August 1993. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/016740489390071C.

Anonymous:1993:ITCf

[Ano93-38] Anonymous. IFIP technical committee 11. Computers & Security, 12(6):605–607, October 1993. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/016740489390057C.

Anonymous:1993:ITCg

[Ano93-39] Anonymous. IFIP technical committee 11. Computers & Security, 12(7):691–693, November 1993. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/016740489390086K.

Anonymous:1993:ITCh

[Ano93-40] Anonymous. IFIP technical committee 11. Computers & Security, 12(8):747–749, December 1993. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/0167404893900398.

Anonymous:1993:IBRa

[Ano93-41] Anonymous. International board of referees. Computers & Security, 12(2):155, March 1993. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/016740489390098P.

Anonymous:1993:IBRb

[Ano93-42] Anonymous. International board of referees. Computers & Security, 12(5):467, August 1993. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/016740489390067F.

Anonymous:1993:IBRc

[Ano93-43] Anonymous. International board of referees. Computers & Security, 12(7):655, November 1993. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/016740489390082G.

Anonymous:1993:IBRd

[Ano93-44] Anonymous. International board of referees. Computers & Security, 12(8):751, December 1993. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/016740489390040C.

Anonymous:1993:PMa

[Ano93-45] Anonymous. Pages 104–211 (March 1993). Computers & Security, 12(2):??, March 1993. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:1993:PF

[Ano93-46] Anonymous. Pages 2–102 (February 1993). Computers & Security, 12(1):??, February 1993. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:1993:PMb

[Ano93-47] Anonymous. Pages 214–318 (May 1993). Computers & Security, 12(3):??, May 1993. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:1993:PJ

[Ano93-48] Anonymous. Pages 320–419 (June 1993). Computers & Security, 12(4):??, June 1993. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:1993:PA

[Ano93-49] Anonymous. Pages 422–514 (August 1993). Computers & Security, 12(5):??, August 1993. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:1993:PO

[Ano93-50] Anonymous. Pages 516–608 (October 1993). Computers & Security, 12(6):??, October 1993. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:1993:PN

[Ano93-51] Anonymous. Pages 610–694 (November 1993). Computers & Security, 12(7):??, November 1993. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:1993:PD

[Ano93-52] Anonymous. Pages 696–787 (December 1993). Computers & Security, 12(8):??, December 1993. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:1993:RBB

[Ano93-53] Anonymous. Random bits and bytes. Computers & Security, 12(4):322–333, June 1993. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/016740489390018Z.

Anonymous:1993:RA

[Ano93-54] Anonymous. Refereed articles. Computers & Security, 12(1): 49, February 1993. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/016740489390011S.

Anonymous:1993:RPIa

[Ano93-55] Anonymous. Refereed papers: International board of referees. Computers & Security, 12(3):273, May 1993. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/016740489390116M.

Anonymous:1993:RPIb

[Ano93-56] Anonymous. Refereed papers: International board of referees. Computers & Security, 12(4):377, June 1993. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/0167404893900262.

Anonymous:1993:RPIc

[Ano93-57] Anonymous. Refereed papers: International board of referees. Computers & Security, 12(6):563, October 1993. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/0167404893900538.

Anonymous:1993:SIV

[Ano93-58] Anonymous. Subject index for vol. 11. Computers & Security, 12(2):201–205, March 1993. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/016740489390103C.

Anonymous:1994:ARA

[Ano94a] Anonymous. Abstracts of recent articles and literature. Computers & Security, 13(1):39–43, February 1994. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/0167404894900906.

Anonymous:1994:AIV

[Ano94b] Anonymous. Author index for vol. 12. Computers & Security, 13(1):85, February 1994. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/016740489490099X.

Anonymous:1994:Ca

[Ano94c] Anonymous. Calendar. Computers & Security, 13(1):45–46, February 1994. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/0167404894900914.

Anonymous:1994:Cb

[Ano94d] Anonymous. Calendar. Computers & Security, 13(2):133– 134, April 1994. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/0167404894900612.

Anonymous:1994:Cc

[Ano94e] Anonymous. Calendar. Computers & Security, 13(3):226– 227, May 1994. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/0167404894900760.

Anonymous:1994:Cd

[Ano94f] Anonymous. Calendar. Computers & Security, 13(4):312– 313, ???? 1994. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/0167404894900213.

Anonymous:1994:Ce

[Ano94g] Anonymous. Calendar. Computers & Security, 13(5):407– 408, ???? 1994. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/0167404894900345.

Anonymous:1994:Cf

[Ano94h] Anonymous. Calendar. Computers & Security, 13(6):490– 491, ???? 1994. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/0167404891901362.

Anonymous:1994:Cg

[Ano94i] Anonymous. Calendar. Computers & Security, 13(7):570, ???? 1994. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/016740489490006X.

Anonymous:1994:Ch

[Ano94j] Anonymous. Calendar. Computers & Security, 13(8):657, ???? 1994. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/0167404894900469.

Anonymous:1994:EB

[Ano94k] Anonymous. Editorial Board. Computers & Security, 13(1):2, February 1994. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/0167404894900833.

Anonymous:1994:Ea

[Ano94l] Anonymous. From the editor. Computers & Security, 13(1):2– 3, February 1994. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/0167404894900841.

Anonymous:1994:Ec

[Ano94m] Anonymous. From the editor. Computers & Security, 13(2):96– 97, April 1994. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/016740489490054X.

Anonymous:1994:Ed

[Ano94n] Anonymous. From the editor. Computers & Security, 13 (3):190–191, May 1994. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/016740489490068X.

Anonymous:1994:Ee

[Ano94o] Anonymous. From the editor. Computers & Security, 13 (4):276–277, ???? 1994. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404894900159.

Anonymous:1994:Ef

[Ano94p] Anonymous. From the editor. Computers & Security, 13 (5):370–371, ???? 1994. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404894900280.

Anonymous:1994:Eg

[Ano94q] Anonymous. From the editor. Computers & Security, 13 (6):456–457, ???? 1994. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404891901306.

Anonymous:1994:Eh

[Ano94r] Anonymous. From the editor. Computers & Security, 13 (7):538–539, ???? 1994. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404894900019.

Anonymous:1994:Ei

[Ano94s] Anonymous. From the editor. Computers & Security, 13 (8):620–621, ???? 1994. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/016740489490040X.

Anonymous:1994:GAa

[Ano94t] Anonymous. Guide for authors. Computers & Security, 13(1): 50, February 1994. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/0167404894900930.

Anonymous:1994:GAb

[Ano94u] Anonymous. Guide for authors. Computers & Security, 13(2): 188, April 1994. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/0167404894900671.

Anonymous:1994:GAc

[Ano94v] Anonymous. Guide for authors. Computers & Security, 13(3): 274, May 1994. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/0167404894900825.

Anonymous:1994:GAd

[Ano94w] Anonymous. Guide for authors. Computers & Security, 13(4): 368, ???? 1994. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/0167404894900272.

Anonymous:1994:GAe

[Ano94x] Anonymous. Guide for authors. Computers & Security, 13(5): 454, ???? 1994. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/0167404894900396.

Anonymous:1994:GAf

[Ano94y] Anonymous. Guide for authors. Computers & Security, 13(6): 535, ???? 1994. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/016740489190142Z.

Anonymous:1994:GAg

[Ano94z] Anonymous. Guide for authors. Computers & Security, 13(7): 618, ???? 1994. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/0167404894900140.

Anonymous:1994:GAh

[Ano94-27] Anonymous. Guide for authors. Computers & Security, 13(8): 702, ???? 1994. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/0167404894900531.

Anonymous:1994:ITCa

[Ano94-28] Anonymous. IFIP technical committee 11. Computers & Security, 13(1):47–49, February 1994. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/0167404894900922.

Anonymous:1994:ITCb

[Ano94-29] Anonymous. IFIP technical committee 11. Computers & Security, 13(2):185–187, April 1994. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404894900663.

Anonymous:1994:ITCc

[Ano94-30] Anonymous. IFIP technical committee 11. Computers & Security, 13(3):270–272, May 1994. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404894900817.

Anonymous:1994:ITCd

[Ano94-31] Anonymous. IFIP technical committee 11. Computers & Security, 13(4):365–367, ???? 1994. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/0167404894900264.

Anonymous:1994:ITCe

[Ano94-32] Anonymous. IFIP technical committee 11. Computers & Security, 13(5):451–453, ???? 1994. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/0167404894900388.

Anonymous:1994:ITCf

[Ano94-33] Anonymous. IFIP technical committee 11. Computers & Security, 13(6):532–534, ???? 1994. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/016740489190141Y.

Anonymous:1994:ITCg

[Ano94-34] Anonymous. IFIP technical committee 11. Computers & Security, 13(7):615–617, ???? 1994. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/0167404894900132.

Anonymous:1994:ITCh

[Ano94-35] Anonymous. IFIP technical committee 11. Computers & Security, 13(8):699–701, ???? 1994. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/0167404894900523.

Anonymous:1994:IBRa

[Ano94-36] Anonymous. International board of referees. Computers & Security, 13(1):51, February 1994. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404894900949.

Anonymous:1994:IBRb

[Ano94-37] Anonymous. International board of referees. Computers & Security, 13(3):229, May 1994. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404894900779.

Anonymous:1994:IBRc

[Ano94-38] Anonymous. International board of referees. Computers & Security, 13(4):315, ???? 1994. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404894900221.

Anonymous:1994:IBRd

[Ano94-39] Anonymous. International board of referees. Computers & Security, 13(5):409, ???? 1994. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404894900353.

Anonymous:1994:IBRe

[Ano94-40] Anonymous. International board of referees. Computers & Security, 13(6):493, ???? 1994. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404891901373.

Anonymous:1994:IBRf

[Ano94-41] Anonymous. International board of referees. Computers & Security, 13(7):571, ???? 1994. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404894900078.

Anonymous:1994:IBRg

[Ano94-42] Anonymous. International board of referees. Computers & Security, 13(8):659, ???? 1994. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404894900477.

Anonymous:1994:PM

[Ano94-43] Anonymous. Pages 190–274 (May 1994). Computers & Security, 13(3):??, May 1994. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:1994:PF

[Ano94-44] Anonymous. Pages 2–94 (February 1994). Computers & Security, 13(1):??, February 1994. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:1994:Pb

[Ano94-45] Anonymous. Pages 276–368 (1994). Computers & Security, 13 (4):??, ???? 1994. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:1994:Pc

[Ano94-46] Anonymous. Pages 370–454 (1994). Computers & Security, 13 (5):??, ???? 1994. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:1994:Pd

[Ano94-47] Anonymous. Pages 456–535 (1994). Computers & Security, 13 (6):??, ???? 1994. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:1994:Pe

[Ano94-48] Anonymous. Pages 538–618 (1994). Computers & Security, 13 (7):??, ???? 1994. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:1994:Pg

[Ano94-49] Anonymous. Pages 620–702 (1994). Computers & Security, 13 (8):??, ???? 1994. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:1994:PA

[Ano94-50] Anonymous. Pages 96–188 (April 1994). Computers & Security, 13(2):??, April 1994. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:1994:RPI

[Ano94-51] Anonymous. Refereed papers: International board of referees. Computers & Security, 13(2):135, April 1994. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/0167404894900620.

Anonymous:1994:SIV

[Ano94-52] Anonymous. Subject index for vol. 12. Computers & Security, 13(1):86–94, February 1994. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404894901007.

Anonymous:1995:AIV

[Ano95a] Anonymous. Author index for volume 13. Computers & Security, 14(2):124, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404895900659.

Anonymous:1995:Ca

[Ano95b] Anonymous. Calendar. Computers & Security, 14(1):40–41, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/0167404895900950.

Anonymous:1995:Cb

[Ano95c] Anonymous. Calendar. Computers & Security, 14(2):122– 123, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/0167404895900632.

Anonymous:1995:Cc

[Ano95d] Anonymous. Calendar. Computers & Security, 14(3):221– 222, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/016740489590381X.

Anonymous:1995:Cd

[Ano95e] Anonymous. Calendar. Computers & Security, 14(4):303– 304, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/0167404895900454.

Anonymous:1995:Ce

[Ano95f] Anonymous. Calendar. Computers & Security, 14(5):415– 416, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404896900371.

Anonymous:1995:Cf

[Ano95g] Anonymous. Calendar. Computers & Security, 14(6):523– 524, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/0167404895903313.

Anonymous:1995:Cg

[Ano95h] Anonymous. Calendar. Computers & Security, 14(7):617– 618, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/0167404895901361.

Anonymous:1995:Ch

[Ano95i] Anonymous. Calendar. Computers & Security, 14(8):702– 703, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404895900322.

Anonymous:1995:EB

[Ano95j] Anonymous. Editorial Board. Computers & Security, 14(1): 2, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/016740489590073X.

Anonymous:1995:Eg

[Ano95k] Anonymous. Exhibitors. Computers & Security, 14(5):429– 432, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404896900528.

Anonymous:1995:Ea

[Ano95l] Anonymous. From the editor. Computers & Security, 14(1):2– 3, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/0167404895900756.

Anonymous:1995:Ec

[Ano95m] Anonymous. From the editor. Computers & Security, 14(2):88– 89, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/0167404895900519.

Anonymous:1995:Ed

[Ano95n] Anonymous. From the editor. Computers & Security, 14 (3):186–187, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404895903615.

Anonymous:1995:Ee

[Ano95o] Anonymous. From the editor. Computers & Security, 14 (4):272–273, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404895900314.

Anonymous:1995:Ef

[Ano95p] Anonymous. From the editor. Computers & Security, 14 (5):362–363, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404896900310.

Anonymous:1995:Eh

[Ano95q] Anonymous. From the Editor. Computers & Security, 14 (6):478–479, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404895903216.

Anonymous:1995:Ei

[Ano95r] Anonymous. From the editor. Computers & Security, 14 (7):572–573, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404895901264.

Anonymous:1995:Ej

[Ano95s] Anonymous. From the editor. Computers & Security, 14 (8):658–659, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404895900268.

Anonymous:1995:ITCa

[Ano95t] Anonymous. IFIP technical committee 11. Computers & Security, 14(1):83–85, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404895900993.

Anonymous:1995:ITCb

[Ano95u] Anonymous. IFIP technical committee 11. Computers & Security, 14(2):181–183, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/0167404895900713.

Anonymous:1995:ITCc

[Ano95v] Anonymous. IFIP technical committee 11. Computers & Security, 14(3):267–269, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/0167404895903852.

Anonymous:1995:ITCd

[Ano95w] Anonymous. IFIP technical committee 11. Computers & Security, 14(4):358–360, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/0167404895900497.

Anonymous:1995:ITCe

[Ano95x] Anonymous. IFIP technical committee 11. Computers & Security, 14(5):473–475, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404896900541.

Anonymous:1995:ITCf

[Ano95y] Anonymous. IFIP technical committee 11. Computers & Security, 14(6):567–569, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/0167404895903356.

Anonymous:1995:ITCg

[Ano95z] Anonymous. IFIP technical committee 11. Computers & Security, 14(7):653–655, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/016740489590140X.

Anonymous:1995:ITCh

[Ano95-27] Anonymous. IFIP technical committee 11. Computers & Security, 14(8):739–741, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404895900365.

Anonymous:1995:IBRa

[Ano95-28] Anonymous. International board of referees. Computers & Security, 14(1):43, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404895900977.

Anonymous:1995:IBRb

[Ano95-29] Anonymous. International Board of Referees. Computers & Security, 14(2):133–134, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/0167404895900691.

Anonymous:1995:IBRc

[Ano95-30] Anonymous. International Board of Referees. Computers & Security, 14(3):223–224, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/0167404895903836.

Anonymous:1995:IBRd

[Ano95-31] Anonymous. International board of referees. Computers & Security, 14(4):305–306, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/0167404895900470.

Anonymous:1995:IBRe

[Ano95-32] Anonymous. International Board of Referees. Computers & Security, 14(5):433–434, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S016740489690053X.

Anonymous:1995:IBRf

[Ano95-33] Anonymous. International board of referees. Computers & Security, 14(6):525–526, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/016740489590333X.

Anonymous:1995:IBRg

[Ano95-34] Anonymous. International board of referees. Computers & Security, 14(7):619–620, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/0167404895901388.

Anonymous:1995:IBRh

[Ano95-35] Anonymous. International Board of Referees. Computers & Security, 14(8):705–706, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404895900349.

Anonymous:1995:LE

[Ano95-36] Anonymous. Letter to the editor. Computers & Security, 14(5):373–375, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/016740489597087Q.

Anonymous:1995:Pc

[Ano95-37] Anonymous. Pages 186–270 (1995). Computers & Security, 14 (3):??, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:1995:Pa

[Ano95-38] Anonymous. Pages 2–85 (1995). Computers & Security, 14(1): ??, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:1995:Pd

[Ano95-39] Anonymous. Pages 272–360 (1995). Computers & Security, 14 (4):??, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:1995:Pe

[Ano95-40] Anonymous. Pages 362–475 (1995). Computers & Security, 14 (5):??, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:1995:Pf

[Ano95-41] Anonymous. Pages 478–570 (1995). Computers & Security, 14 (6):??, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:1995:Pg

[Ano95-42] Anonymous. Pages 572–655 (1995). Computers & Security, 14 (7):??, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:1995:Ph

[Ano95-43] Anonymous. Pages 658–741 (1995). Computers & Security, 14 (8):??, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:1995:Pb

[Ano95-44] Anonymous. Pages 88–184 (1995). Computers & Security, 14 (2):??, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:1995:SIV

[Ano95-45] Anonymous. Subject index for volume 13. Computers & Security, 14(2):125–132, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/0167404895900675.

Anonymous:1996:ACS

[Ano96a] Anonymous. Access control systems: from host-centric to network centric computing. Computers & Security, 15(5):402, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/0167404896825971.

Anonymous:1996:AOS

[Ano96b] Anonymous. Approach to object security in distributed SOM. Computers & Security, 15(5):402, ???? 1996. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/016740489682596X.

Anonymous:1996:ASQ

[Ano96c] Anonymous. Ask stupid questions; listen for the unsaid: the EDP auditor as consultant: Alan Krull. Computers & Security, 15(5):409, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404896826071.

Anonymous:1996:ADS

[Ano96d] Anonymous. An attack detection systems for secure computer systems — design of the ADS: I. Kantzavelou and A. Patel, Department of Computer Science, University College Dublin Belfired, Dublin, Ireland. Computers & Security, 15(5):395, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/0167404896825727.

Anonymous:1996:AIV

[Ano96e] Anonymous. Author index for volume 14. Computers & Security, 15(2):137–138, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404896902837.

Anonymous:1996:BMV

[Ano96f] Anonymous. Beware of a macro virus epidemic. Computers & Security, 15(5):398, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404896825806.

Anonymous:1996:BVP

[Ano96g] Anonymous. Blaming the victim: Prof. Gene Spafford, Purdue University. Computers & Security, 15(5):418, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/0167404896826368.

Anonymous:1996:BDP

[Ano96h] Anonymous. Blind date — problems with the year 2000: Doc Farmer, S. G. Warburg. Computers & Security, 15(5):418, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/0167404896826356.

Anonymous:1996:BRE

[Ano96i] Anonymous. Book review: Efficient and probably secure key agreement: N. Alexandris, V. Chrissikopoulos and D. Peppes, University of Piraeus, Greece and M. Burmester, R.H.- University of London, Surrey, UK. Computers & Security, 15 (5):394, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/0167404896825673.

Anonymous:1996:BRI

[Ano96j] Anonymous. Book review: Internetwork access control using public key certificates: H. Park and R. Chow, Computer and Information Sciences and Engineering Department, University of Florida, Gainesville, FL. Computers & Security, 15 (5):394, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/0167404896825661.

Anonymous:1996:BRM

[Ano96k] Anonymous. Book review: MoFAC: a model for fine-grained access control: J. S. von Solms, M. S. Olivier and S. H. von Solms, Department of Computer Science, Rand Afrikaans University, Johannesburg, South Africa. Computers & Security, 15 (5):395, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/0167404896825697.

Anonymous:1996:BRS

[Ano96l] Anonymous. Book review: Security profile for interconnected open distributed systems with varying vulnerability: N. Nikitakos, Hellenic Naval Academy, Piraeus, Greece; S. Gritzalis and P. Georgiadis, University of Athens, Athens, Greece. Computers & Security, 15(5):396, ???? 1996. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/0167404896825752.

Anonymous:1996:BCP

[Ano96m] Anonymous. Business continuity planning: John Sherwood. Computers & Security, 15(5):411, ???? 1996. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/0167404896826186.

Anonymous:1996:Ca

[Ano96n] Anonymous. Calendar. Computers & Security, 15(1):50–51, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404896900589.

Anonymous:1996:Cb

[Ano96o] Anonymous. Calendar. Computers & Security, 15(2):129– 130, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404896902813.

Anonymous:1996:Cc

[Ano96p] Anonymous. Calendar. Computers & Security, 15(3):227– 228, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404896903132.

Anonymous:1996:Cd

[Ano96q] Anonymous. Calender. Computers & Security, 15(4):323– 324, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404896902217.

Anonymous:1996:Ce

[Ano96r] Anonymous. Calender. Computers & Security, 15(5):403– 404, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404896903557.

Anonymous:1996:Cf

[Ano96s] Anonymous. Calender. Computers & Security, 15(6):521– 522, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404896901091.

Anonymous:1996:Cg

[Ano96t] Anonymous. Calender. Computers & Security, 15(7):591– 592, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404896901157.

Anonymous:1996:Ch

[Ano96u] Anonymous. Calender. Computers & Security, 15(8):691– 692, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404896902400.

Anonymous:1996:CCA

[Ano96v] Anonymous. Cashless cybercrime: August Bequai. Computers & Security, 15(5):414, 417, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/0167404896826290.

Anonymous:1996:CSIb

[Ano96w] Anonymous. CIPFA survey of IT audit in UK public service organizations 1996: Vernon Poole, Aid to Industry. Computers & Security, 15(5):413, ???? 1996. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/016740489682623X.

Anonymous:1996:CLH

[Ano96x] Anonymous. Closing the loopholes in host-based authentication: Jim Reid, RTFM. Computers & Security, 15(5):414, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/0167404896826265.

Anonymous:1996:CEa

[Ano96y] Anonymous. COMPSEC ’96 exhibitors. Computers & Security, 15(5):419–420, 423–428, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404896903569.

Anonymous:1996:CEO

[Ano96z] Anonymous. Computer evidence: an overview of forensic procedures, tools and techniques: Edward wilding, network security management. Computers & Security, 15(5):417, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/0167404896826332.

Anonymous:1996:CGI

[Ano96-27] Anonymous. Corporate governance and IT: Chris Hurford, Audit Commission. Computers & Security, 15(5):406, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/0167404896826046.

Anonymous:1996:CCT

[Ano96-28] Anonymous. Countering chip theft: Philip a clearly, probe FX. Computers & Security, 15(5):410, ???? 1996. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/0167404896826125.

Anonymous:1996:CSIa

[Ano96-29] Anonymous. Crypto system initialization: Simplifying the distribution of initial keys: Carl Meyer, IBM. Computers & Security, 15(5):406, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404896826034.

Anonymous:1996:CKD

[Ano96-30] Anonymous. Cryptographic key distribution and authentication protocols for secure group communication: A. C. M. Hutchison and K. Bauknecht, Department of Computer Science, University of Zurich, Switzerland. Computers & Security, 15(5): 396, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/0167404896825740.

Anonymous:1996:DDN

[Ano96-31] Anonymous. Danger! Deadly new computer viruses want to kill your PC. Computers & Security, 15(5):400, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/0167404896825879.

Anonymous:1996:DUS

[Ano96-32] Anonymous. Dealing with a Unix system penetration: Steve Bailey, Reflex Magnetics. Computers & Security, 15(5):409– 410, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/0167404896826113.

Anonymous:1996:DRP

[Ano96-33] Anonymous. Disaster recovery: the people issues: Andrew hiles, survive! Computers & Security, 15(5):411, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/0167404896826162.

Anonymous:1996:DRK

[Ano96-34] Anonymous. Distributed registration and key distribution (DiRK): R. Oppliger, University of Berne, Institute for Computer Science, Berne, Switzerland and Andres Albanese, International Computer Science Institute, Berkeley, CA. Computers & Security, 15(5):394–395, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/0167404896825685.

Anonymous:1996:DMS

[Ano96-35] Anonymous. Does Microsoft have security all locked up? Computers & Security, 15(5):399, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/016740489682582X.

Anonymous:1996:DBI

[Ano96-36] Anonymous. Doing business on the Internet: marketing and security aspects. Computers & Security, 15(5):400, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/0167404896825892.

Anonymous:1996:EB

[Ano96-37] Anonymous. Editorial Board. Computers & Security, 15(1): 2, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404896900553.

Anonymous:1996:EPS

[Ano96-38] Anonymous. Elite panel slams US crypto policy. Computers & Security, 15(5):397, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404896825764.

Anonymous:1996:FSR

[Ano96-39] Anonymous. Factors in the selection of a risk assessment method. Computers & Security, 15(5):401, ???? 1996. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/0167404896825946.

Anonymous:1996:FTB

[Ano96-40] Anonymous. Feds turn blind eye to systems security. Computers & Security, 15(5):401, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/0167404896825910.

Anonymous:1996:Ea

[Ano96-41] Anonymous. From the editor. Computers & Security, 15(1):2– 3, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404896900565.

Anonymous:1996:Ec

[Ano96-42] Anonymous. From the editor. Computers & Security, 15(2):88– 89, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404896902801.

Anonymous:1996:Ed

[Ano96-43] Anonymous. From the editor. Computers & Security, 15 (3):182–183, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404896902102.

Anonymous:1996:Ee

[Ano96-44] Anonymous. From the editor. Computers & Security, 15 (4):268–269, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404896902138.

Anonymous:1996:Ef

[Ano96-45] Anonymous. From the editor. Computers & Security, 15 (5):354–355, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404896903521.

Anonymous:1996:Eg

[Ano96-46] Anonymous. From the editor. Computers & Security, 15 (6):456–457, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404896901054.

Anonymous:1996:Eh

[Ano96-47] Anonymous. From the editor. Computers & Security, 15 (7):550–551, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404897881096.

Anonymous:1996:Ei

[Ano96-48] Anonymous. From the editor. Computers & Security, 15 (8):652–653, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404897896976.

Anonymous:1996:FDC

[Ano96-49] Anonymous. Fundamental DES design concepts: Carl Meyer, IBM. Computers & Security, 15(5):406, ???? 1996. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/0167404896826022.

Anonymous:1996:GGA

[Ano96-50] Anonymous. GSSP (Generally-Accepted system Security Principles): a trip to Abilene: Alan Krull. Computers & Security, 15 (5):417, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/0167404896826307.

Anonymous:1996:HTS

[Ano96-51] Anonymous. An hierarchical threshold scheme with unique partial keys: H. Hassler, V. Hassler, and R. Posch, Institute for Applied Information Processing and Communications, Graz, Austria. Computers & Security, 15(5):393, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/0167404896825636.

Anonymous:1996:ITCa

[Ano96-52] Anonymous. IFIP technical committee 11. Computers & Security, 15(1):83–85, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404896900607.

Anonymous:1996:ITCb

[Ano96-53] Anonymous. IFIP technical committee 11. Computers & Security, 15(2):177–179, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404896902850.

Anonymous:1996:ITCc

[Ano96-54] Anonymous. IFIP technical committee 11. Computers & Security, 15(3):263–266, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404896903156.

Anonymous:1996:ITCd

[Ano96-55] Anonymous. IFIP technical committee 11. Computers & Security, 15(4):349–351, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404896902254.

Anonymous:1996:ITCe

[Ano96-56] Anonymous. IFIP technical committee 11. Computers & Security, 15(5):451–453, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404896903582.

Anonymous:1996:ITCf

[Ano96-57] Anonymous. IFIP technical committee 11. Computers & Security, 15(6):545–547, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404896901133.

Anonymous:1996:ITCh

[Ano96-58] Anonymous. IFIP technical committee 11. Computers & Security, 15(8):733–735, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404896902424.

Anonymous:1996:ITCg

[Ano96-59] Anonymous. IFIP technical committee 11rch in computer security. Computers & Security, 15(7):647–649, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404896901194.

Anonymous:1996:ISC

[Ano96-60] Anonymous. IFIP/SEC’96 conference proceeding. Computers & Security, 15(5):393, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404896825624.

Anonymous:1996:DMI

[Ano96-61] Anonymous. In defence of MoD IT privatization. Computers & Security, 15(5):399, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404896825855.

Anonymous:1996:ICD

[Ano96-62] Anonymous. Industry to Clinton: drop encryption policy. Computers & Security, 15(5):397–398, ???? 1996. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/0167404896825788.

Anonymous:1996:IOC

[Ano96-63] Anonymous. Information owners, custodians and users. Computers & Security, 15(5):401, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/0167404896825934.

Anonymous:1996:IWY

[Ano96-64] Anonymous. Information warfare: Your company needs you!: Tim Porter, bull information systems. Computers & Security, 15(5):414, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404896826277.

Anonymous:1996:ISRb

[Ano96-65] Anonymous. Intelligent security reporting — auditing security logs: Norman Crocker. Computers & Security, 15(5):417– 418, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/0167404896826344.

Anonymous:1996:IMA

[Ano96-66] Anonymous. Intention modelling: Approximating computer user intentions for detection and prediction of intrusions: T. Spyrou and J. Darzentas, Department of Mathematics University of the Aegean, Greece. Computers & Security, 15(5): 395, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/0167404896825703.

Anonymous:1996:IBRa

[Ano96-67] Anonymous. International board of referees. Computers & Security, 15(1):53–54, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404896900590.

Anonymous:1996:IBRb

[Ano96-68] Anonymous. International board of referees. Computers & Security, 15(2):139–140, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404896902849.

Anonymous:1996:IBRc

[Ano96-69] Anonymous. International Board of Referees. Computers & Security, 15(3):229–230, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404896903144.

Anonymous:1996:IBRd

[Ano96-70] Anonymous. International board of referees. Computers & Security, 15(4):325–326, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404896902230.

Anonymous:1996:IBRe

[Ano96-71] Anonymous. International board of referees. Computers & Security, 15(5):429–430, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404896903570.

Anonymous:1996:IBRf

[Ano96-72] Anonymous. International board of referees. Computers & Security, 15(6):523–524, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S016740489690111X.

Anonymous:1996:IBRg

[Ano96-73] Anonymous. International board of referees. Computers & Security, 15(7):593–594, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404896901170.

Anonymous:1996:IBRh

[Ano96-74] Anonymous. International board of referees. Computers & Security, 15(8):693–694, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404896902412.

Anonymous:1996:IVD

[Ano96-75] Anonymous. Internet = Virusnet?: Dr. David Aubrey-Jones, Reflex Magnetics. Computers & Security, 15(5):406, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/0167404896826058.

Anonymous:1996:ISF

[Ano96-76] Anonymous. Internet security — firewall horror stories: Dr. Stuart Broderick, Zergo. Computers & Security, 15(5):405, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/0167404896825995.

Anonymous:1996:IAR

[Ano96-77] Anonymous. Intranet as an audit resource: Doc Farmer, S. G. Warburg. Computers & Security, 15(5):413, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/0167404896826216.

Anonymous:1996:ISPb

[Ano96-78] Anonymous. Intranet security: Peter wood, first base. Computers & Security, 15(5):405, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/0167404896825983.

Anonymous:1996:ERS

[Ano96-79] Anonymous. Is encryption the road to security? Computers & Security, 15(5):401, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404896825922.

Anonymous:1996:ISPa

[Ano96-80] Anonymous. IT security and privacy education: L. Yngstrom, Department of Computer and Systems Sciences, Stockholm University and Royal Institute of Technology, Kista, Sweden. Computers & Security, 15(5):395–396, ???? 1996. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/0167404896825739.

Anonymous:1996:ISB

[Ano96-81] Anonymous. IT security is a business issue not a technical one: Rod Parkin. Computers & Security, 15(5):411, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/0167404896826174.

Anonymous:1996:ISRa

[Ano96-82] Anonymous. It’s the infrastructure, stupid: Ray Kaplan. Computers & Security, 15(5):413, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/0167404896826241.

Anonymous:1996:KIL

[Ano96-83] Anonymous. Keeping it legit. Computers & Security, 15(5): 399, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/0167404896825831.

Anonymous:1996:MSC

[Ano96-84] Anonymous. Mobile state — challenges for the 21st century: August Bequai. Computers & Security, 15(5):413, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/0167404896826228.

Anonymous:1996:NDC

[Ano96-85] Anonymous. Network disaster contingency planning: Peter wood, first base. Computers & Security, 15(5):411, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/0167404896826204.

Anonymous:1996:NIS

[Ano96-86] Anonymous. The new information security age — electronic commerce and trusted third parties: Professor Henry Beker, Zergo Holdings plc. Computers & Security, 15(5):409, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/0167404896826101.

Anonymous:1996:OCR

[Ano96-87] Anonymous. One company’s roadmap to securing it’s intranet: Bob Frank, SRI International. Computers & Security, 15(5): 414, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/0167404896826289.

Anonymous:1996:Pc

[Ano96-88] Anonymous. Pages 182–266 (1996). Computers & Security, 15 (3):??, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:1996:Pa

[Ano96-89] Anonymous. Pages 2–86 (1996). Computers & Security, 15(1): ??, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:1996:Pd

[Ano96-90] Anonymous. Pages 268–351 (1996). Computers & Security, 15 (4):??, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:1996:Pe

[Ano96-91] Anonymous. Pages 354–453 (1996). Computers & Security, 15 (5):??, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:1996:Pf

[Ano96-92] Anonymous. Pages 456–547 (1996). Computers & Security, 15 (6):??, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:1996:Pg

[Ano96-93] Anonymous. Pages 550–649 (1996). Computers & Security, 15 (7):??, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:1996:Ph

[Ano96-94] Anonymous. Pages 652–735 (1996). Computers & Security, 15 (8):??, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:1996:Pb

[Ano96-95] Anonymous. Pages 88–179 (1996). Computers & Security, 15 (2):??, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:1996:PM

[Ano96-96] Anonymous. Password master. Computers & Security, 15(5): 397, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/0167404896903545.

Anonymous:1996:PSS

[Ano96-97] Anonymous. A policy for sending secret information over communications networks. Computers & Security, 15(5):400, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/0167404896825867.

Anonymous:1996:PAH

[Ano96-98] Anonymous. Practical application of high level risk analysis and security healthchecks for telecommunications systems: Kevin Sloan, Syntegra Ltd. SISL Operations. Computers & Security, 15(5):410, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404896826137.

Anonymous:1996:PFS

[Ano96-99] Anonymous. A practical framework for secure corporate teleworking: Mark Curtis, PC security. Computers & Security, 15 (5):410, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/0167404896826149.

Anonymous:1996:PI

[Ano96-100] Anonymous. Privacy and the Internet. Computers & Security, 15(5):398, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/016740489682579X.

Anonymous:1996:RAI

[Ano96-101] Anonymous. Reviewing and assessing IT projects under development: Paul Williams, Binder Hamlyn, Anderson Worldwide. Computers & Security, 15(5):409, ???? 1996. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/0167404896826083.

Anonymous:1996:RSC

[Ano96-102] Anonymous. Right sizing without capsizing: P. J. Corum, The Corum Group. Computers & Security, 15(5):409, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/0167404896826095.

Anonymous:1996:SMD

[Ano96-103] Anonymous. SALSA: a method for developing the enterprise security architecture and strategy: John Sherwood. Computers & Security, 15(5):406, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/016740489682606X.

Anonymous:1996:SCF

[Ano96-104] Anonymous. SAP — control and fraud prevention strategies: Nigel Iyer, Network Security Management. Computers & Security, 15(5):411, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404896826198.

Anonymous:1996:SW

[Ano96-105] Anonymous. The secret war. Computers & Security, 15(5):398– 399, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/0167404896825818.

Anonymous:1996:SIA

[Ano96-106] Anonymous. Secure IP access at last with virtual TCP online. Computers & Security, 15(5):399, ???? 1996. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/0167404896825843.

Anonymous:1996:SEM

[Ano96-107] Anonymous. Securely executing multilevel transactions: S. Jajodia, K. P. Smith. B. T. Blaustein, and L. Notargiacomo, The MITRE Corporation, McLean, VA. Computers & Security, 15 (5):394, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/0167404896825648.

Anonymous:1996:SRA

[Ano96-108] Anonymous. Securing remote access to corporate systems: Colin tankard, digital pathways. Computers & Security, 15(5): 414, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/0167404896826253.

Anonymous:1996:SIVb

[Ano96-109] Anonymous. Security issues in the virtual corporation: Bob Frank, SRI International. Computers & Security, 15(5):405– 406, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/0167404896826010.

Anonymous:1996:SBS

[Ano96-110] Anonymous. Smartcards — is Britain smart enough: Alan Laird, Bull Information Systems. Computers & Security, 15(5): 417, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/0167404896826320.

Anonymous:1996:SQA

[Ano96-111] Anonymous. Software quality assurance: the underlying framework for achieving secure and reliable software systems: S. A. Frangos, Intrasoft, Athens, Greece. Computers & Security, 15 (5):394, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/016740489682565X.

Anonymous:1996:SOV

[Ano96-112] Anonymous. Some observations on vulnerabilities: Prof. Gene Spafford, Purdue University. Computers & Security, 15(5): 405, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/0167404896826009.

Anonymous:1996:SIVa

[Ano96-113] Anonymous. Subject index for volume 14. Computers & Security, 15(2):131–136, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404896902825.

Anonymous:1996:TC

[Ano96-114] Anonymous. Tales from the crypt. Computers & Security, 15 (5):400, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/0167404896825909.

Anonymous:1996:TAH

[Ano96-115] Anonymous. A taxonomy for analyzing hazards to information systems: R. Baskerville, Copenhagen Business School, Frederiksberg, Denmark. Computers & Security, 15(5):395, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/0167404896825715.

Anonymous:1996:TFE

[Ano96-116] Anonymous. Telecommunications fraud the evolving threat: Tom Mulhall, BT. Computers & Security, 15(5):410, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/0167404896826150.

Anonymous:1996:TYP

[Ano96-117] Anonymous. Testing your plan is more important than the plan itself. Computers & Security, 15(5):400, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/0167404896825880.

Anonymous:1996:TBS

[Ano96-118] Anonymous. The truth behind single sign-on: Gary Hardy, Zergo. Computers & Security, 15(5):417, ???? 1996. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/0167404896826319.

Anonymous:1996:WSS

[Ano96-119] Anonymous. Web server: security lockdown. Computers & Security, 15(5):401–402, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/0167404896825958.

Anonymous:1997:ADS

[Ano97a] Anonymous. An attack detection system for secure computer systems-outline of the solution. Computers & Security, 16(3): 207, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404897845221.

Anonymous:1997:AIV

[Ano97b] Anonymous. Author index for volume 15. Computers & Security, 16(2):172, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404897883095.

Anonymous:1997:BCA

[Ano97c] Anonymous. Between the cracks: Alan Earls, Computerworld, February 10, 1997, pp. 79–80. Computers & Security, 16(1):64– 65, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404897857999.

Anonymous:1997:Cd

[Ano97d] Anonymous. Calendar. Computers & Security, 16(4):335– 336, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404897901577.

Anonymous:1997:Ce

[Ano97e] Anonymous. Calendar. Computers & Security, 16(6):513– 514, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S016740489790273X.

Anonymous:1997:Cf

[Ano97f] Anonymous. Calendar. Computers & Security, 16(7):625– 626, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404897901620.

Anonymous:1997:Cg

[Ano97g] Anonymous. Calendar. Computers & Security, 16(8):683– 684, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404897900596.

Anonymous:1997:Ca

[Ano97h] Anonymous. Calender. Computers & Security, 16(1):67–68, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404897900341.

Anonymous:1997:Cb

[Ano97i] Anonymous. Calender. Computers & Security, 16(2):133– 134, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404897902261.

Anonymous:1997:Cc

[Ano97j] Anonymous. Calender. Computers & Security, 16(3):219– 220, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404897902674.

Anonymous:1997:CSB

[Ano97k] Anonymous. Can security baselines replace risk analysis? Computers & Security, 16(3):206–207, ???? 1997. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S016740489784518X.

Anonymous:1997:CPS

[Ano97l] Anonymous. Code of practice: a standard for information security management. Computers & Security, 16(3):206, ????

1997. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404897845166.

Anonymous:1997:CCF

[Ano97m] Anonymous. A common criteria framework for the evaluation of information technology systems security. Computers & Security, 16(3):207, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S016740489784521X.

Anonymous:1997:CSC

[Ano97n] Anonymous. A comparison of schemes for certification authorities/trusted third parties. Computers & Security, 16(3): 210, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404897845312.

Anonymous:1997:CTM

[Ano97o] Anonymous. A compile-time model for safe information flow in object-oriented databases. Computers & Security, 16(3): 206, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404897845178.

Anonymous:1997:CNK

[Ano97p] Anonymous. A comprehensive need-to-know access control system and its application for medical information systems. Computers & Security, 16(3):209, ???? 1997. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404897845294.

Anonymous:1997:CEa

[Ano97q] Anonymous. Compsec ’97 exhibitors. Computers & Security, 16(6):529–532, 534, 536, 537–540, ???? 1997. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404897902741.

Anonymous:1997:CDP

[Ano97r] Anonymous. Computer disasters: prevention and recovery. Computers & Security, 16(3):187–188, ???? 1997. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404897845099.

Anonymous:1997:CCV

[Ano97s]  Anonymous.  Constructing computer virus phylogenies.

Computers & Security, 16(3):187, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404897845087.

Anonymous:1997:CIA

[Ano97t] Anonymous. Controlling Internet access at an educational institution. Computers & Security, 16(3):210, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404897845324.

Anonymous:1997:CKG

[Ano97u] Anonymous. Cryptography: a key to growth — and crime: Nicholas Bray, Wall Street Journal Europe, January 27, 1997. Computers & Security, 16(1):62, ???? 1997. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404897857914.

Anonymous:1997:CCL

[Ano97v] Anonymous. Cyberclash over cybercash: Lisa Picarille. Computerworld, December 9, 1996, p. 12. Computers & Security, 16(1):62–63, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404897857938.

Anonymous:1997:DBT

[Ano97w] Anonymous. Data backup technology. Computers & Security, 16(3):188, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404897845105.

Anonymous:1997:DAP

[Ano97x] Anonymous. Deriving authorization from process analysis in legacy information systems. Computers & Security, 16(3):206, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404897845154.

Anonymous:1997:EB

[Ano97y] Anonymous. Editorial Board. Computers & Security, 16(1): 2, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404897900006.

Anonymous:1997:FMA

[Ano97z] Anonymous. A formal model to aid documenting and harmonizing of information security requirements. Computers & Security, 16(3):205–206, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404897845142.

Anonymous:1997:FFP

[Ano97-27] Anonymous. Fortifyingyour firewall: Peter Morrissey. Network Computing, February 15, 1997, pp. 57–71. Computers & Security, 16(1):66, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404897858051.

Anonymous:1997:Ea

[Ano97-28] Anonymous. From the editor. Computers & Security, 16(2):84– 85, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404897882879.

Anonymous:1997:Ec

[Ano97-29] Anonymous. From the editor. Computers & Security, 16 (3):178–179, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404897844987.

Anonymous:1997:Ed

[Ano97-30] Anonymous. From the editor. Computers & Security, 16 (4):264–265, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404897801885.

Anonymous:1997:Ee

[Ano97-31] Anonymous. From the editor. Computers & Security, 16 (6):452–453, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404897846391.

Anonymous:1997:Ef

[Ano97-32] Anonymous. From the editor. Computers & Security, 16 (7):570–571, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404897807924.

Anonymous:1997:Eg

[Ano97-33] Anonymous. From the editor. Computers & Security, 16 (8):648–649, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404897875773.

Anonymous:1997:FNN

[Ano97-34] Anonymous. The future of networks and network security. Computers & Security, 16(3):209, ???? 1997. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404897845270.

Anonymous:1997:GCG

[Ano97-35] Anonymous. Gotcha! Con games on the Internet: Sharon Gaudin. Computerworld, February 10, 1997, pp. 43–44. Computers & Security, 16(1):63, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S016740489785794X.

Anonymous:1997:HFT

[Ano97-36] Anonymous. Holding the fort: Thomas Parkhurst. Computerworld, February 10, 1977, pp. 73–77. Computers & Security, 16 (1):65, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404897858014.

Anonymous:1997:HDC

[Ano97-37] Anonymous. How digital certificates work: LAN times, January 6, 1997, p. 39. Computers & Security, 16(1):62, ????

1997. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404897857926.

Anonymous:1997:HTS

[Ano97-38] Anonymous. How to trust systems. Computers & Security, 16 (3):210, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404897845348.

Anonymous:1997:ITCa

[Ano97-39] Anonymous. IFIP technical committee 11. Computers & Security, 16(1):80–82, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404897900407.

Anonymous:1997:ITCb

[Ano97-40] Anonymous. IFIP technical committee 11. Computers & Security, 16(2):173–175, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404897902285.

Anonymous:1997:ITCc

[Ano97-41] Anonymous. IFIP technical committee 11. Computers & Security, 16(3):259–261, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404897902698.

Anonymous:1997:ITCd

[Ano97-42] Anonymous. IFIP technical committee 11. Computers & Security, 16(4):345–347, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404897901619.

Anonymous:1997:ITCe

[Ano97-43] Anonymous. IFIP technical committee 11. Computers & Security, 16(6):565–567, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404897902765.

Anonymous:1997:ITCf

[Ano97-44]  Anonymous.  IFip technical committee 11.  Comput-

ers & Security, 16(7):643–645, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404897901644.

Anonymous:1997:ITCg

[Ano97-45] Anonymous. IFIP technical committee 11. Computers & Security, 16(8):721–723, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404897900638.

Anonymous:1997:IIS

[Ano97-46] Anonymous. Integrating information security in the development of telematic systems. Computers & Security, 16(3): 210, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404897845336.

Anonymous:1997:IFP

[Ano97-47] Anonymous. Internal firewalls protect against inside jobs: Rivka Tadjer. Computing, December 1, 1996, pp. 28–30. Computers & Security, 16(1):64, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404897857975.

Anonymous:1997:IBRa

[Ano97-48]  Anonymous.  International Board of Referees.  Com-

puters & Security, 16(1):69–70, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404897900365.

Anonymous:1997:IBRb

[Ano97-49] Anonymous. International board of referees. Computers & Security, 16(2):135–136, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404897902273.

Anonymous:1997:IBRc

[Ano97-50] Anonymous. International Board of Referees. Computers & Security, 16(3):221–222, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404897902686.

Anonymous:1997:IBRd

[Ano97-51] Anonymous. International board of referees. Computers & Security, 16(4):337–338, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404897901590.

Anonymous:1997:IBRe

[Ano97-52] Anonymous. International board of referees. Computers & Security, 16(6):541–542, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404897902753.

Anonymous:1997:IBRf

[Ano97-53] Anonymous. International board of referees. Computers & Security, 16(7):627–628, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404897901632.

Anonymous:1997:IBRg

[Ano97-54] Anonymous. International board of referees. Computers & Security, 16(8):685–686, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404897900614.

Anonymous:1997:DYK

[Ano97-55] Anonymous. It’s 1997. do you know where your data is?: Robert Moskowitz. Network Computing, February 15, 1997, pp. 39–40. Computers & Security, 16(1):64, ???? 1997. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404897857987.

Anonymous:1997:NBC

[Ano97-56] Anonymous. A new browser could let others tap into your bank account: Dan stets. the Philadelphia inquirer, February 16, 1997. Computers & Security, 16(1):61, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404897857896.

Anonymous:1997:NOF

[Ano97-57] Anonymous. New organizational forms for information security management. Computers & Security, 16(3):210, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S016740489784535X.

Anonymous:1997:NVI

[Ano97-58] Anonymous. New vistas on info-system security. Computers & Security, 16(3):207, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404897845208.

Anonymous:1997:NIA

[Ano97-59] Anonymous. Non-intrusive authentication. Computers & Security, 16(3):209, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404897845282.

Anonymous:1997:OSS

[Ano97-60] Anonymous. Online security is Swiss cheese, a survey finds: Michael Tarsala. Investor’s Business Daily, January 20, 1997. Computers & Security, 16(1):62, ???? 1997. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404897857902.

Anonymous:1997:Pc

[Ano97-61] Anonymous. Pages 178–261 (1997). Computers & Security, 16 (3):??, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:1997:Pa

[Ano97-62] Anonymous. Pages 2–82 (1997). Computers & Security, 16(1): ??, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:1997:Pd

[Ano97-63] Anonymous. Pages 264–347 (1997). Computers & Security, 16 (4):??, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:1997:Pe

[Ano97-64] Anonymous. Pages 350–449 (1997). Computers & Security, 16 (5):??, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:1997:Pf

[Ano97-65] Anonymous. Pages 452–567 (1997). Computers & Security, 16 (6):??, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:1997:Pg

[Ano97-66] Anonymous. Pages 570–645 (1997). Computers & Security, 16 (7):??, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:1997:Ph

[Ano97-67] Anonymous. Pages 648–723 (1997). Computers & Security, 16 (8):??, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:1997:Pb

[Ano97-68] Anonymous. Pages 84–175 (1997). Computers & Security, 16 (2):??, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:1997:PAP

[Ano97-69] Anonymous. Pseudonymous audit for privacy enhanced intrusion detection. Computers & Security, 16(3):207, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404897845191.

Anonymous:1997:PKC

[Ano97-70] Anonymous. Public-key cryptosystems based on linear codes. Computers & Security, 16(3):187, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404897845063.

Anonymous:1997:RBB

[Ano97-71] Anonymous. Random bits and bytes. Computers & Security, 16(5):356–364, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404897822419.

Anonymous:1997:RDS

[Ano97-72] Anonymous. Review: A digital signature scheme based upon the theory of quadratic residues: C. C. Chang, J.K Jan and H. C. Kowng. Cryptologia, January, 1997, pp. 55– 70. Computers & Security, 16(1):66, ???? 1997. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S016740489785804X.

Anonymous:1997:RSAa

[Ano97-73] Anonymous. Review: Screening for applets: Sharon Machlis. Computerworld, February 10, 1977, pp. 51–52. Computers & Security, 16(1):65, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404897858002.

Anonymous:1997:RWS

[Ano97-74] Anonymous. Review: Web security: how much is enough?: Vance McCarthy. Datamation January 1997, pp. 112–117. Computers & Security, 16(1):65–66, ???? 1997. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404897858038.

Anonymous:1997:RSAb

[Ano97-75]  Anonymous.  Roaming security agents enabling intelli-

gent access control and network protection. Computers & Security, 16(3):209–210, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404897845300.

Anonymous:1997:SPY

[Ano97-76] Anonymous. SecureDesktop protects your PC like a vault: Christopher Smith, Network Computing, February 15, 1997, pp. 52–54. Computers & Security, 16(1):65, ???? 1997. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404897858026.

Anonymous:1997:SRS

[Ano97-77] Anonymous. Security requirements and solution in distributed electronic health records. Computers & Security, 16(3):208– 209, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404897845269.

Anonymous:1997:SWO

[Ano97-78] Anonymous. Security when outsourcing: concepts, constructs, compliance. Computers & Security, 16(3):208, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404897845233.

Anonymous:1997:SVB

[Ano97-79] Anonymous. Stop viruses before they replicate: Doug van Kirk. LAN times, December 9, 1996, p. 91. Computers & Security, 16(1):63–64, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404897857963.

Anonymous:1997:SIV

[Ano97-80] Anonymous. Subject index for volume 15. Computers & Security, 16(2):165–171, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404897883083.

Anonymous:1997:TEC

[Ano97-81]  Anonymous.  A taxonomy of electronic case schemes.

Computers & Security, 16(3):208, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404897845257.

Anonymous:1997:TKW

[Ano97-82] Anonymous. Tools keep Web surfing safe: Monica Snell. LAN times, January 6, 1997, p. 7, 10. Computers & Security, 16(1): 63, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404897857951.

Anonymous:1997:THV

[Ano97-83] Anonymous. Towards a holistic view of security and safety of enterprise information and communication technologies: adapting to a changing paradigm. Computers & Security, 16(3): 208, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404897845245.

Anonymous:1997:VKS

[Ano97-84] Anonymous. Virus killers: Stan Miastkowski. PC World, March 1997, pp. 181–190. Computers & Security, 16(1):66, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404897858063.

Anonymous:1997:WHA

[Ano97-85] Anonymous. Web hack attacks bring down servers, but businesses stay: Mitch Wavner and Kim Girard. Computerworld, January 6, 1997, p. 1, 16. Computers & Security, 16(1):61, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404897857884.

Anonymous:1997:WES

[Ano97-86] Anonymous. When encryption slows you down: Datamation, January 1997, p. 13, 14. Computers & Security, 16(1): 61, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404897857872.

Anonymous:1998:AIV

[Ano98a] Anonymous. Author index for volume 16. Computers & Security, 17(2):162, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404897820147.

Anonymous:1998:Ca

[Ano98b] Anonymous. Calendar. Computers & Security, 17(1):65–66, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404898903637.

Anonymous:1998:Cb

[Ano98c] Anonymous. Calendar. Computers & Security, 17(2):160– 161, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404898902693.

Anonymous:1998:Cc

[Ano98d] Anonymous. Calendar. Computers & Security, 17(3):229– 230, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S016740489890253X.

Anonymous:1998:Cd

[Ano98e] Anonymous. Calendar. Computers & Security, 17(4):332– 334, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404898902784.

Anonymous:1998:Ce

[Ano98f] Anonymous. Calendar. Computers & Security, 17(5):412– 414, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404898902474.

Anonymous:1998:Cf

[Ano98g] Anonymous. Calendar. Computers & Security, 17(6):522– 524, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404898900591.

Anonymous:1998:Cg

[Ano98h] Anonymous. Calendar. Computers & Security, 17(7):619– 620, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404898903078.

Anonymous:1998:Ch

[Ano98i] Anonymous. Calendar. Computers & Security, 17(8):721– 722, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404898903546.

Anonymous:1998:CIa

[Ano98j] Anonymous. Compsec 98 International. Computers & Security, 17(4):361–365, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404898902814.

Anonymous:1998:CIb

[Ano98k] Anonymous. Compsec 98 International. Computers & Security, 17(6):556–560, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404898900657.

Anonymous:1998:EB

[Ano98l] Anonymous. Editorial Board. Computers & Security, 17(1): 2, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404898903625.

Anonymous:1998:E

[Ano98m] Anonymous. Exhibitors. Computers & Security, 17(7):621– 631, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S016740489890308X.

Anonymous:1998:ITN

[Ano98n] Anonymous. IFIP TC11 news. Computers & Security, 17 (2):151–152, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S016740489890267X.

Anonymous:1998:ITCa

[Ano98o] Anonymous. IFIP technical committee 11. Computers & Security, 17(1):91–93, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404898903650.

Anonymous:1998:ITCb

[Ano98p] Anonymous. IFIP technical committee 11. Computers & Security, 17(2):185–187, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404898902735.

Anonymous:1998:ITCc

[Ano98q] Anonymous. IFIP technical committee 11. Computers & Security, 17(3):273–275, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404898902577.

Anonymous:1998:ITCd

[Ano98r] Anonymous. IFIP technical committee 11. Computers & Security, 17(4):358–360, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404898902802.

Anonymous:1998:ITCe

[Ano98s] Anonymous. IFIP technical committee 11. Computers & Security, 17(5):455–457, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404898902516.

Anonymous:1998:ITCf

[Ano98t] Anonymous. IFIP technical committee 11. Computers & Security, 17(6):553–555, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404898900633.

Anonymous:1998:ITCg

[Ano98u] Anonymous. IFIP technical committee 11. Computers & Security, 17(7):632–634, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404898903091.

Anonymous:1998:ITCh

[Ano98v] Anonymous. IFIP technical committee 11. Computers & Security, 17(8):733–735, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S016740489890356X.

Anonymous:1998:IBRa

[Ano98w] Anonymous. International board of referees. Computers & Security, 17(1):67–68, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404898903649.

Anonymous:1998:IBRb

[Ano98x] Anonymous. International board of referees. Computers & Security, 17(2):169–170, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404898902711.

Anonymous:1998:IBRc

[Ano98y] Anonymous. International board of referees. Computers & Security, 17(3):231–232, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404898902553.

Anonymous:1998:IBRd

[Ano98z] Anonymous. International board of referees. Computers & Security, 17(4):335–336, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404898902796.

Anonymous:1998:IBRe

[Ano98-27] Anonymous. International board of referees. Computers & Security, 17(5):415–416, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404898902498.

Anonymous:1998:IBRf

[Ano98-28] Anonymous. International Board of Referees. Computers & Security, 17(6):525–526, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S016740489890061X.

Anonymous:1998:IBRg

[Ano98-29] Anonymous. International Board of Referees. Computers & Security, 17(7):635–636, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404898903108.

Anonymous:1998:IBRh

[Ano98-30] Anonymous. International board of referees. Computers & Security, 17(8):723–724, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404898903558.

Anonymous:1998:Pc

[Ano98-31] Anonymous. Pages 190–275 (1998). Computers & Security, 17 (3):??, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:1998:Pa

[Ano98-32] Anonymous. Pages 2–93 (1998). Computers & Security, 17(1): ??, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:1998:Pd

[Ano98-33] Anonymous. Pages 278–366 (1998). Computers & Security, 17 (4):??, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:1998:Pe

[Ano98-34] Anonymous. Pages 368–457 (1998). Computers & Security, 17 (5):??, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:1998:Pf

[Ano98-35] Anonymous. Pages 460–560 (1998). Computers & Security, 17 (6):??, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:1998:Pg

[Ano98-36] Anonymous. Pages 562–650 (1998). Computers & Security, 17 (7):??, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:1998:Ph

[Ano98-37] Anonymous. Pages 652–735 (1998). Computers & Security, 17 (8):??, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:1998:Pb

[Ano98-38] Anonymous. Pages 96–187 (1998). Computers & Security, 17 (2):??, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:1998:SV

[Ano98-39] Anonymous. Security views. Computers & Security, 17(5):370– 380, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404898800023.

Anonymous:1998:SIV

[Ano98-40] Anonymous. Subject index for volume 16. Computers & Security, 17(2):163–168, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404897820159.

Anonymous:1999:ARA

[Ano99a] Anonymous. Abstracts of recent articles and literature. Computers & Security, 18(8):705–710, ???? 1999. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404899801345.

Anonymous:1999:AEH

[Ano99b] Anonymous. ADSL enables hacker access. Computers & Security, 18(7):598, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S016740489982032X.

Anonymous:1999:AIV

[Ano99c] Anonymous. Author index for volume 17. Computers & Security, 18(2):157, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404899800509.

Anonymous:1999:BOE

[Ano99d] Anonymous. Back orifice — an excuse to legalize hacking? Computers & Security, 18(7):597, ???? 1999. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404899820288.

Anonymous:1999:BFCa

[Ano99e] Anonymous. ‘bug’ fears cut demand for Japanese bonds. Computers & Security, 18(7):594, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404899820112.

Anonymous:1999:BFCb

[Ano99f] Anonymous. Businesses fail to comply with data protection principles. Computers & Security, 18(7):596–597, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404899820240.

Anonymous:1999:Ca

[Ano99g] Anonymous. Calendar. Computers & Security, 18(1):81–82, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404899900032.

Anonymous:1999:Cb

[Ano99h] Anonymous. Calendar. Computers & Security, 18(2):155– 156, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S016740489990055X.

Anonymous:1999:Cc

[Ano99i] Anonymous. Calendar. Computers & Security, 18(3):237– 238, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404899909501.

Anonymous:1999:Cd

[Ano99j] Anonymous. Calendar. Computers & Security, 18(4):341– 342, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S016740489990737X.

Anonymous:1999:Ce

[Ano99k] Anonymous. Calendar. Computers & Security, 18(5):438– 439, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404899906028.

Anonymous:1999:Cf

[Ano99l] Anonymous. Calendar. Computers & Security, 18(6):525– 526, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404899908660.

Anonymous:1999:Cg

[Ano99m] Anonymous. Calendar. Computers & Security, 18(7):599– 600, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404899820367.

Anonymous:1999:Ch

[Ano99n] Anonymous. Calendar. Computers & Security, 18(8):711– 712, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404899907824.

Anonymous:1999:CMR

[Ano99o] Anonymous. Canada may reap benefit of software privacy rule. Computers & Security, 18(7):593, ???? 1999. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404899820069.

Anonymous:1999:CEC

[Ano99p] Anonymous. Canada’s eavesdropping on citizens, ex-spy says. Computers & Security, 18(7):596, ???? 1999. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404899820215.

Anonymous:1999:CAC

[Ano99q] Anonymous. Clinton administration calls for network monitoring. Computers & Security, 18(7):597, ???? 1999. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404899820276.

Anonymous:1999:CEa

[Ano99r] Anonymous. Compsec ’99 exhibitors. Computers & Security, 18(7):608–616, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404899820380.

Anonymous:1999:CPA

[Ano99s] Anonymous. Compsec ’99 paper abstracts. Computers & Security, 18(7):601–607, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404899820379.

Anonymous:1999:DGP

[Ano99t] Anonymous. Don’t go proprietary, crypto expert urges. Computers & Security, 18(7):597, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S016740489982029X.

Anonymous:1999:MPR

[Ano99u] Anonymous. E-mail privacy rule. Computers & Security, 18(7): 596, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404899820239.

Anonymous:1999:EB

[Ano99v] Anonymous. Editorial Board. Computers & Security, 18(1): 2, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404899900019.

Anonymous:1999:EWP

[Ano99w] Anonymous. Enemy within poses a fraud risk. Computers & Security, 18(7):593–594, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404899820094.

Anonymous:1999:FWP

[Ano99x] Anonymous. Firms warned not to play e-mail ‘big brother’ role. Computers & Security, 18(7):595–596, ???? 1999. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404899820203.

Anonymous:1999:FPI

[Ano99y] Anonymous. France to propose Internet legislation; focus is on security. Computers & Security, 18(7):594, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404899820148.

Anonymous:1999:E

[Ano99z] Anonymous. From the editor. Computers & Security, 18 (8):644–645, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404899907812.

Anonymous:1999:GSH

[Ano99-27] Anonymous. Government sites hacked due to NATO’s bombing. Computers & Security, 18(7):598, ???? 1999. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404899820306.

Anonymous:1999:HSF

[Ano99-28] Anonymous. Hotmail ‘slow to fix flaw in security’. Computers & Security, 18(7):598, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404899820331.

Anonymous:1999:ITCa

[Ano99-29] Anonymous. IFIP technical committee 11. Computers & Security, 18(1):83–85, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404899900056.

Anonymous:1999:ITCb

[Ano99-30] Anonymous. IFIP technical committee 11. Computers & Security, 18(2):178–180, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404899900597.

Anonymous:1999:ITCc

[Ano99-31] Anonymous. IFIP technical committee 11. Computers & Security, 18(3):271–273, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404899909525.

Anonymous:1999:ITCd

[Ano99-32] Anonymous. IFIP technical committee 11. Computers & Security, 18(4):365–367, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404899907393.

Anonymous:1999:ITCe

[Ano99-33] Anonymous. IFIP technical committee 11. Computers & Security, 18(5):440–442, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S016740489990603X.

Anonymous:1999:ITCf

[Ano99-34] Anonymous. IFIP technical committee 11. Computers & Security, 18(6):545–547, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404899908684.

Anonymous:1999:ITCg

[Ano99-35] Anonymous. IFIP technical committee 11. Computers & Security, 18(7):640–642, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404899820422.

Anonymous:1999:ITCh

[Ano99-36] Anonymous. IFIP technical committee 11. Computers & Security, 18(8):734–736, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404899907848.

Anonymous:1999:IBRa

[Ano99-37] Anonymous. International board of referees. Computers & Security, 18(2):163–164, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404899900573.

Anonymous:1999:IBRb

[Ano99-38] Anonymous. International board of referees. Computers & Security, 18(3):239–240, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404899909513.

Anonymous:1999:IBRc

[Ano99-39] Anonymous. International Board of Referees. Computers & Security, 18(4):343–344, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404899907381.

Anonymous:1999:IBRd

[Ano99-40] Anonymous. International Board of Referees. Computers & Security, 18(5):443–444, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404899906041.

Anonymous:1999:IBRe

[Ano99-41] Anonymous. International board of referees. Computers & Security, 18(6):531–532, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404899908672.

Anonymous:1999:IBRf

[Ano99-42] Anonymous. International board of referees. Computers & Security, 18(8):713–714, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404899907836.

Anonymous:1999:ICE

[Ano99-43] Anonymous. IPv6 chaos to eclipse Y2K. Computers & Security, 18(7):594, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404899820136.

Anonymous:1999:MFU

[Ano99-44] Anonymous. Microsoft faces user pressure on licences. Computers & Security, 18(7):595, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404899820161.

Anonymous:1999:MHP

[Ano99-45] Anonymous. Millennium hackers plan city assault. Computers & Security, 18(7):596, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404899820227.

Anonymous:1999:MIA

[Ano99-46] Anonymous. Multi-industry alliance to beat piracy. Computers & Security, 18(7):595, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S016740489982015X.

Anonymous:1999:NTS

[Ano99-47] Anonymous. NASA technology sends porn into space. Computers & Security, 18(7):595, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404899820173.

Anonymous:1999:NNA

[Ano99-48] Anonymous. NEC’s network architecture saves data from crashes. Computers & Security, 18(7):593, ???? 1999. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404899820070.

Anonymous:1999:NPL

[Ano99-49] Anonymous. New Polish law to fight intellectual property theft. Computers & Security, 18(7):594, ???? 1999. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404899820100.

Anonymous:1999:Pc

[Ano99-50] Anonymous. Pages 182–273 (1999). Computers & Security, 18 (3):??, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:1999:Pa

[Ano99-51] Anonymous. Pages 2–85 (1999). Computers & Security, 18(1): ??, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:1999:Pd

[Ano99-52] Anonymous. Pages 276–367 (1999). Computers & Security, 18 (4):??, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:1999:Pe

[Ano99-53] Anonymous. Pages 370–453 (1999). Computers & Security, 18 (5):??, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:1999:Pf

[Ano99-54] Anonymous. Pages 456–547 (1999). Computers & Security, 18 (6):??, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:1999:Pg

[Ano99-55] Anonymous. Pages 550–642 (1999). Computers & Security, 18 (7):??, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:1999:Ph

[Ano99-56] Anonymous. Pages 644–736 (1999). Computers & Security, 18 (8):??, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:1999:Pb

[Ano99-57] Anonymous. Pages 88–180 (1999). Computers & Security, 18 (2):??, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:1999:RP

[Ano99-58] Anonymous. Refereed papers. Computers & Security, 18 (7):617–618, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404899820392.

Anonymous:1999:SFR

[Ano99-59] Anonymous. Security flaws reported in BC’s prescription drug network. Computers & Security, 18(7):594, ???? 1999. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404899820124.

Anonymous:1999:SV

[Ano99-60] Anonymous. Security view. Computers & Security, 18(8):646– 659, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404899801291.

Anonymous:1999:SIV

[Ano99-61] Anonymous. Subject index for volume 17. Computers & Security, 18(2):158–162, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404899800510.

Anonymous:1999:TSS

[Ano99-62] Anonymous. Taxman says sorry for year 2000 blunder. Computers & Security, 18(7):598, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404899820343.

Anonymous:1999:TWP

[Ano99-63] Anonymous. Telcos wake up to phreaking menace. Computers & Security, 18(7):595, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404899820185.

Anonymous:1999:TTH

[Ano99-64] Anonymous. Termination threat hangs over thin client software. Computers & Security, 18(7):597, ???? 1999. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404899820264.

Anonymous:1999:TFY

[Ano99-65] Anonymous. Top firms on Y2K danger list. Computers & Security, 18(7):595, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404899820197.

Anonymous:1999:T

[Ano99-66] Anonymous. Twinkle. Computers & Security, 18(7):593, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404899820082.

Anonymous:1999:UBC

[Ano99-67] Anonymous. US backs cyber-crime bill. Computers & Security, 18(7):598, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404899820355.

Anonymous:1999:ULU

[Ano99-68] Anonymous. Users lose under new law. Computers & Security, 18(7):598, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404899820318.

Anonymous:1999:VWA

[Ano99-69] Anonymous. Virus is wrapped and all ready for Christmas day. Computers & Security, 18(7):597, ???? 1999. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404899820252.

Antion:1992:LE

[Ant92] Karen Antion. Letter to the editor. Computers & Security, 11(7):623–624, November 1992. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/016740489290139I.

Associates:1992:ISA

[Ass92] Janus Associates. Information security administration model: a management model to help identify the best practices of the administration function within the security group. Computers & Security, 11(4):327–340, July 1992. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/016740489290172N.

Abrams:1995:SC

[AZ95] Marshall D. Abrams and Marvin V. Zelkowitz. Striving for correctness. Computers & Security, 14(8):719–738, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/0167404895000224.

Bacon:1997:BGY

[Bac97a] Michael Bacon. Bang goes your business. Computers & Security, 16(6):523, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404897846883.

Bacon:1997:EE

[Bac97b] Michael Bacon. The enemy everywhere. Computers & Security, 16(6):516, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404897846652.

Bailey:1997:IAN

[Bai97] Steve Bailey. Intelligent agents: New security issues. Computers & Security, 16(6):526, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404897847001.

Banks:1990:SP

[Ban90] Simon Banks. Security policy. Computers & Security, 9(7): 605–610, November 1990. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404890900582.

Banks:1991:UIC

[Ban91] Simon Banks. Understanding IBM’s CICS systems to improve security. Computers & Security, 10(3):205–212, May 1991. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/016740489190035C.

Banes:1995:PLN

[Ban95] David Banes. Physical layer network security: What your LAN can do for you. Computers & Security, 14(5):422, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/0167404896900437.

Baskerville:1991:RAS

[Bas91] Richard Baskerville. Risk analysis as a source of professional knowledge. Computers & Security, 10(8):749–764, December 1991. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/016740489190094T.

Borcherding:1998:ETK

[BB98] Birgit Borcherding and Malte Borcherding. Efficient and trustworthy key distribution in webs of trust. Computers & Security, 17(5):447–454, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404898000017.

Bauspiess:1992:RCH

[BD92] Fritz Bauspiess and Frank Damm. Requirements for cryptographic hash functions. Computers & Security, 11(5):427–437, September 1992. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/016740489290007E.

Backhouse:1995:MCC

[BD95] James Backhouse and Gurpreet Dhillon. Managing computer crime: a research outlook. Computers & Security, 14(7):645– 651, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/0167404895000240.

Bernaschi:1999:ENS

[BdAR99] M. Bernaschi, E. d’ Aiutolo, and P. Rughetti. Enforcing network security: a real cease study in a research organization. Computers & Security, 18(6):533–543, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404899801175.

Badenhorst:1990:CSM

[BE90] K. P. Badenhorst and Jan H. P. Eloff. Computer security methodology: Risk analysis and project definition. Computers & Security, 9(4):339–346, June 1990. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/0167404890901042.

Badenhorst:1994:TFA

[BE94] Karin P. Badenhorst and Jan H. P. Eloff. TOPM: a formal approach to the optimization of information technology risk management. Computers & Security, 13(5):411–435, ???? 1994. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/0167404894900361.

Booysen:1995:COI

[BE95] H. A. S. Booysen and J. H. P. Eloff. Classification of objects for improved access control. Computers & Security, 14(3):251– 265, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/016740489500001O.

Benton:1998:PIS

[Ben98] J. R. Benton. Physical IT security. Computers & Security, 17(5):389–391, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404898800059.

Bequai:1998:BLC

[Beq98a] August Bequai. Balancing legal concerns over crime and security in cyberspace. Computers & Security, 17(4):293–298, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404898800084.

Bequai:1998:EAC

[Beq98b] August Bequai. Employee abuses in cyberspace: Management’s legal quagmire. Computers & Security, 17(8):667–670, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404898800977.

Bequai:1998:GCC

[Beq98c] August Bequai. A guide to cyber-crime investigations. Computers & Security, 17(7):579–582, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S016740489980056X.

Bequai:1998:HTS

[Beq98d] August Bequai. High-tech security and the failings of President Clinton’s Commission on Critical Infrastructure Protection. Computers & Security, 17(1):19–21, ???? 1998. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404897802441.

Bequai:1998:SPM

[Beq98e] August Bequai. Software pirating and management’s quagmire. Computers & Security, 17(1):22–26, ???? 1998. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404897802453.

Bequai:1998:TCF

[Beq98f] August Bequai. Techno-crimes: Failings of the legal edifice. Computers & Security, 17(5):381–384, ???? 1998. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404898800035.

Bequai:1999:CCU

[Beq99a] August Bequai. Cyber-crime the US experience. Computers & Security, 18(1):16–18, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404899800030.

Bequai:1999:EIL

[Beq99b] August Bequai. Employers and the Internet: Legal woes and concerns. Computers & Security, 18(4):303–306, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404899800741.

Bequai:1999:MEC

[Beq99c] August Bequai. Monitoring employee communications in cyberspace: the pitfalls facing management. Computers & Security, 18(2):109–112, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404899800285.

Berman:1992:WCW

[BGK92] Jerry Berman, John Gilmore, and Mitchell Kapor. Wake up: the cold war is over. Computers & Security, 11(5):485–486, September 1992. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/016740489290017L.

Bergen:1992:DSF

[BH92] Helen A. Bergen and James M. Hogan. Data security in a fixed-model arithmetic coding compression algorithm. Computers & Security, 11(5):445–461, September 1992. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/016740489290011F.

Bergen:1993:CPA

[BH93] Helen A. Bergen and James M. Hogan. A chosen plaintext attack on an adaptive arithmetic coding compression algorithm. Computers & Security, 12(2):157–167, March 1993. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/016740489390099Q.

Birch:1997:WNS

[Bir97] Peter Birch. Windows NT security today and the future. Computers & Security, 16(6):521, ???? 1997. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404897846809.

Bishop:1995:ISS

[BK95] Matt Bishop and Daniel V. Klein. Improving system security via proactive password checking. Computers & Security, 14(3):233– 249, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/016740489500003Q.

Black:1992:UOS

[Bla92] George Black. Update on open systems security. Computers & Security, 11(8):699–702, December 1992. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/0167404892900889.

Black:1993:DSS

[Bla93] David K. Black. The digital signature standard: Overview and current status. Computers & Security, 12(5):437–446, August 1993. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/016740489390062A.

Blatchford:1996:IP

[Bla96]  Clive Blatchford. Internet as a pornotopia?  Com-

puters & Security, 15(3):203–208, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/0167404896897823.

Burns:1990:SSR

[BM90] John Burns and Chris J. Mitchell. A security scheme for resource sharing over a network. Computers & Security, 9(1):67– 75, February 1990. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/016740489090159Q.

Booker:1992:GCF

[BN92] Ellis Booker and Jim Nash. Great Chicago flood of ’92: IS groups stay high and dry. Computers & Security, 11(5):486, September 1992. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/016740489290018M.

Bologna:1990:EII

[Bol90] Jack Bologna. Ethical issues of the information era. Computers & Security, 9(8):689–692, December 1990. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/016740489090110F.

Bologna:1991:FEA

[Bol91] Jack Bologna. A framework for the ethical analysis of information technologies. Computers & Security, 10(4):303–307, June 1991. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/016740489190105M.

Bolton:1995:DAT

[Bol95] Andrew Bolton. Derivatives and assessing their risks. Computers & Security, 14(5):426, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/0167404895971495.

Bontchev:1996:PMV

[Bon96] Vesselin Bontchev. Possible macro virus attacks and how to prevent them. Computers & Security, 15(7):595–626, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S016740489788131X.

Bontchev:1998:MVI

[Bon98] Vesselin Bontchev. Macro virus identification problems. Computers & Security, 17(1):69–89, ???? 1998. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404897802751.

Bontchev:1999:PWV

[Bon99] Vesselin Bontchev. The problems of WordMacro virus upconversion. Computers & Security, 18(3):241–255, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404899800686.

Booker:1992:RSE

[Boo92] Ellis Booker. Retinal scanners eye-dentify inmates. Computers & Security, 11(5):491, September 1992. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/016740489290253N.

Born:1996:ELO

[Bor96] Eike Born. Enforcing legal ownership rights by an access control system. Computers & Security, 15(3):212–220, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/0167404896897847.

Bouniol:1997:PTL

[Bou97] Stephane G. Bouniol. The puzzle theorem — the less I know, the less I can disclose . . .. Computers & Security, 16(2):115– 125, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404897882909.

Bovelander:1995:SCB

[Bov95] Ernst Bovelander. Smart cards and biometrics: an overview of current technologies, threats and opportunities. Computers & Security, 14(5):418–419, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/016740489597119U.

Bowe:1992:LE

[Bow92] Rennie Bowe. Letter to the editor. Computers & Security, 11(6):529–530, October 1992. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/016740489290186U.

Boyce:1997:CEC

[Boy97] Brian Boyce. Cyber extortion — the corporate response. Computers & Security, 16(1):25–28, ???? 1997. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404897857847.

Burk:1990:VES

[BP90] Holger Bürk and Andreas Pfitzmann. Value exchange systems enabling security and unobservability. Computers & Security, 9(8):715–721, December 1990. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/0167404890901149.

Bunnell:1997:CAC

[BPH^{+}97] Julie Bunnell, John Podd, Ron Henderson, Renee Napier, and James Kennedy-Moffat. Cognitive, associative and conventional passwords: Recall and guessing rates. Computers & Security, 16(7):629–641, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404897000084.

Borrell:1996:ISV

[BR96] Joan Borrell and Josep Rifà. An implementable secure voting scheme. Computers & Security, 15(4):327–338, ???? 1996.

CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/0167404896000028.

Brunnstein:1995:VOT

[Bru95] Prof. Klaus Brunnstein. Viruses, an overestimated threat? Computers & Security, 14(5):419, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/0167404896900395.

Born:1994:DAC

[BS94] Eike Born and Helmut Stiegler. Discretionary access control by means of usage conditions. Computers & Security, 13(5):437– 450, ???? 1994. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/016740489490037X.

Buckwell:1995:SSN

[Buc95] Mark Buckwell. The spook solution — now open for business. Computers & Security, 14(5):420, 422, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/016740489597127V.

Buckwell:1996:SSN

[Buc96] Mark Buckwell. The spook solution — now open for business. Computers & Security, 15(1):17–26, ???? 1996. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/016740489687619X.

Bulmer:1999:EDU

[Bul99] Garry Bulmer. Encryption — do or die for UK business. Computers & Security, 18(4):312–316, ???? 1999. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404899800765.

Buzzard:1999:CSW

[Buz99] Keith Buzzard. Computer security — what should you spend your money on? Computers & Security, 18(4):322–334, ????

1999. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404899800789.

Candia:1990:HVK

[Can90] Tanya Candia. How VMS keeps out intruders. Computers & Security, 9(6):499–502, October 1990. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/016740489090125D.

Carrol:1990:CRB

[Car90a] John M. Carrol. Conference report: 2300 brave virus threat to attend national security conference. Computers & Security, 9(1):45–49, February 1990. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/016740489090155M.

Carroll:1990:DIY

[Car90b] John M. Carroll. “do-it-yourself” cryptography. Computers & Security, 9(7):613–619, November 1990. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/0167404890900607.

Carroll:1990:SCI

[Car90c] John M. Carroll. Security and credibility in an informationintensive society. Computers & Security, 9(6):489–498, October 1990. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/016740489090124C.

Carr:1991:EDI

[Car91] J. Carr. Electronic data interchange — security risk or not? Computers & Security, 10(1):69–72, February 1991. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/016740489190057K.

Caelli:1999:PEC

[CDR99] William J. Caelli, Edward P. Dawson, and Scott A. Rea. PKI, elliptic curve cryptography, and digital signatures. Computers & Security, 18(1):47–66, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S016740489980008X.

Caelli:1993:CAP

[CGO93] W. Caelli, I. Graham, and L. O’Connor. Cryptographic application programming interfaces (APIs). Computers & Security, 12(7):640–645, November 1993. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/016740489390079K.

Chen:1994:IBC

[CH94] Jau Liang Chen and Tzonelih Hwang. Identity-based conference key broadcast schemes with user authentication. Computers & Security, 13(1):53–57, February 1994. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/0167404894900957.

Chih-Hung:1997:MCH

[CHT97] Wang Chih-Hung and Hwang Tzonelih. Modified Chen-Hwang identity-based conference key broadcast schemes with user authentication. Computers & Security, 16(4):339–344, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404897000060.

Ciechanowicz:1997:RAR

[Cie97] Zbigniew Ciechanowicz. Risk analysis: requirements, conflicts and problems. Computers & Security, 16(3):223–232, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404897000047.

Cleary:1995:IAC

[CIRM95] John G. Cleary, Sean A. Irvine, and Ingrid Rinsma-Melchert. On the insecurity of arithmetic coding. Computers & Security, 14(2):167–180, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/016740489597050K.

Chang:1994:SDR

[CJB94] C. C. Chang, J. K. Jan, and D. J. Buehrer. A scheme to determine the relationship between two users in a hierarchy. Computers & Security, 13(3):255–261, May 1994. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/0167404894900795.

Chor:1998:GAS

[CJC98] Leong Peng Chor, Hsu Wen Jing, and Tan Peng Chong. A geometric approach for shared secrets, a refinement. Computers & Security, 17(8):725–732, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404898801028.

Coomaraswamy:1991:FOL

[CKM91] Gnanesh Coomaraswamy, Srikanta P. R. Kumar, and Michel E. Marhic. Fiber-optic LAN/WAN systems to support confidential communication. Computers & Security, 10(8):765–776, December 1991. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/016740489190095U.

Chang:1994:RPA

[CL94] Chin-Chin Chang and Wen-Yuan Liao. A remote password authentication scheme based upon ElGamal’s signature scheme. Computers & Security, 13(2):137–144, April 1994. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/0167404894900639.

Clark:1996:CCE

[CL96] Andrew J. Clark and Sapher Services Ltd. Cryptographic controls — the eternal triangle. Computers & Security, 15(7):576– 584, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404897881138.

Clark:1997:KRW

[Cla97] Andrew J. Clark. Key recovery — why, how, who? Computers & Security, 16(8):669–674, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404897875815.

Clyde:1993:DSN

[Cly93] Robert A. Clyde. DECnet security (not necessarily an oxymoron). Computers & Security, 12(2):123–127, March 1993. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/016740489390091I.

Cerullo:1992:CCP

[CM92] Michael J. Cerullo and R. Steve McDuffie. Computer contingency plans and the auditors: a survey of businesses affected by hurricane hugo. Computers & Security, 11(7):620–622, November 1992. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/016740489290138H.

Collins:1993:SYB

[CM93] Brian S. Collins and Steve Mathews. Securing your business process. Computers & Security, 12(7):629–633, November 1993. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/016740489390077I.

Coffey:1998:RMK

[CN98] T. Coffey and T. Newe. Realisation of a minimumknowledge identification and signature scheme. Computers & Security, 17(3):253–264, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404898803398.

Cogle:1995:S

[Cog95] Graham Cogle. OS/2 security. Computers & Security, 14(5): 424, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/0167404896900486.

Cohen:1991:CAT

[Coh91a] Fred Cohen. A cost analysis of typical computer viruses and defenses. Computers & Security, 10(3):239–250, May 1991. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/016740489190040K.

Cohen:1991:BPI

[Coh91b] Frederick B. Cohen. A DOS-based POset implementation. Computers & Security, 10(6):541–551, October 1991. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/016740489190080W.

Cohen:1991:NHI

[Coh91c] Fredrick B. Cohen. A note on high-integrity PC bootstrapping. Computers & Security, 10(6):535–539, October 1991. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/016740489190079S.

Cohen:1992:FDC

[Coh92a] F. B. Cohen. A formal definition of computer worms and some related results. Computers & Security, 11(7):641–652, November 1992. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/016740489290144G.

Cohen:1992:DDA

[Coh92b] Frederick B. Cohen. Defense-in-depth against computer viruses. Computers & Security, 11(6):563–579, October 1992. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/016740489290192T.

Cohen:1992:LE

[Coh92c] Frederick B. Cohen. Letter to the editor. Computers & Security, 11(2):140–142, April 1992. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/016740489290039T.

Cohen:1993:OSP

[Coh93] Frederick B. Cohen. Operating system protection through program evolution. Computers & Security, 12(6):565–584, October 1993. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/0167404893900549.

Cohen:1996:NDC

[Coh96a] Frederick B. Cohen. A note on distributed coordinated attacks. Computers & Security, 15(2):103–121, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/0167404896893230.

Cohen:1996:SWW

[Coh96b] Frederick B. Cohen. A secure World-Wide-Web daemon. Computers & Security, 15(8):707–724, ???? 1996. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404896000090.

Cohen:1997:ISA

[Coh97a] Fred Cohen. Information system attacks: a preliminary classification scheme. Computers & Security, 16(1):29–46, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404897857859.

Cohen:1997:ISD

[Coh97b] Fred Cohen. Information system defences: a preliminary classification scheme. Computers & Security, 16(2):94–114, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404897882892.

Cohen:1997:IDR

[Coh97c] Fred Cohen. Intrusion detection and response. Computers & Security, 16(6):516, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S016740489784669X.

Cohen:1998:NRD

[Coh98] Fred Cohen. A note on the role of deception in information protection. Computers & Security, 17(6):483–506, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404898800710.

Cohen:1999:SCA

[Coh99] Fred Cohen. Simulating cyber attacks, defences, and consequences. Computers & Security, 18(6):479–518, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404899801151.

Cole:1990:MSD

[Col90] Robert Cole. A model for security in distributed systems. Computers & Security, 9(4):319–330, June 1990. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/016740489090102Y.

Collinson:1994:ARAa

[Col94a] Helen Collinson. Abstracts of recent articles and literature. Computers & Security, 13(2):127–132, April 1994. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/0167404894900604.

Collinson:1994:ARAb

[Col94b] Helen Collinson. Abstracts of recent articles and literature. Computers & Security, 13(3):221–225, May 1994. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/0167404894900752.

Collinson:1994:ARAc

[Col94c] Helen Collinson. Abstracts of recent articles and literature. Computers & Security, 13(4):306–311, ???? 1994. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/0167404894900205.

Collinson:1994:ARAd

[Col94d] Helen Collinson. Abstracts of recent articles and literature. Computers & Security, 13(5):401–406, ???? 1994. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/0167404894900337.

Collinson:1994:ARAe

[Col94e] Helen Collinson. Abstracts of recent articles and literature. Computers & Security, 13(6):485–489, ???? 1994. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/016740489190135Z.

Collinson:1994:ARAf

[Col94f] Helen Collinson. Abstracts of recent articles and literature. Computers & Security, 13(7):564–569, ???? 1994. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/0167404894900051.

Collinson:1994:ARAg

[Col94g] Helen Collinson. Abstracts of recent articles and literature. Computers & Security, 13(8):651–656, ???? 1994. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/0167404894900450.

Collinson:1995:H

[Col95a] Helen Collinson. 5 H5a4BRTES7S. Computers & Security, 14 (8):701, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/0167404896817275.

Collinson:1995:AIC

[Col95b] Helen Collinson. Alarm at increase in civil service crime. Computers & Security, 14(7):611, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/0167404896816865.

Collinson:1995:AOG

[Col95c] Helen Collinson. America online girds against hacker breakins. Computers & Security, 14(6):519, ???? 1995. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/016740489599317I.

Collinson:1995:ABF

[Col95d] Helen Collinson. Another brick in the firewall. Computers & Security, 14(6):521, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/016740489599321F.

Collinson:1995:AVS

[Col95e] Helen Collinson. Anti-virus software. Computers & Security, 14(8):700–701, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/016740489681724X.

Collinson:1995:SIA

[Col95f] Helen Collinson. Are suppliers infected with an apathy virus? Computers & Security, 14(5):412, ???? 1995. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/0167404896900358.

Collinson:1995:BVF

[Col95g] Helen Collinson. Banks, vendors focus on security. Computers & Security, 14(5):409–410, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/016740489597092O.

Collinson:1995:BBN

[Col95h] Helen Collinson. Big business on the net? Not yet. Computers & Security, 14(4):301, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/016740489597080T.

Collinson:1995:BCP

[Col95i] Helen Collinson. BSA cracks down on piracy in NZ. Computers & Security, 14(1):37, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/0167404895900853.

Collinson:1995:BTP

[Col95j] Helen Collinson. BSA toughens up on pirates. Computers & Security, 14(7):612, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404896816907.

Collinson:1995:BGF

[Col95k] Helen Collinson. Build great firewalls. Computers & Security, 14(4):301–302, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/016740489597082L.

Collinson:1995:CC

[Col95l] Helen Collinson. Calling in confidence. Computers & Security, 14(2):120, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/016740489597040H.

Collinson:1995:CCT

[Col95m] Helen Collinson. Cellular carriers try new tricks to beat bandits. Computers & Security, 14(2):119, ???? 1995. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/016740489597036A.

Collinson:1995:CRS

[Col95n] Helen Collinson. China’s raids on software pirates move to interior. Computers & Security, 14(5):411, ???? 1995. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/016740489597096S.

Collinson:1995:CSC

[Col95o] Helen Collinson. Computer security centre to be set up. Computers & Security, 14(1):35, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/016740489597013Z.

Collinson:1995:CD

[Col95p] Helen Collinson. Cops of the dataway. Computers & Security, 14 (1):36, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/0167404895970164.

Collinson:1995:CPS

[Col95q] Helen Collinson. Copy protection, the sequel: coming to a PC near you. Computers & Security, 14(1):39, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/0167404895970212.

Collinson:1995:CCP

[Col95r] Helen Collinson. Costly cellular phone fraud rises with number ‘cloning’. Computers & Security, 14(2):119, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/016740489597038C.

Collinson:1995:CSE

[Col95s] Helen Collinson. Cracking a social engineer. Computers & Security, 14(8):700, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404896817238.

Collinson:1995:CCN

[Col95t] Helen Collinson. Crash course in ’net security. Computers & Security, 14(2):120, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404895970418.

Collinson:1995:CCW

[Col95u] Helen Collinson. Cryptic cross words. Computers & Security, 14 (5):413, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/016740489597103H.

Collinson:1995:CSB

[Col95v] Helen Collinson. Crypto system battles video pirates. Computers & Security, 14(1):38–39, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/0167404895900896.

Collinson:1995:CWF

[Col95w] Helen Collinson. CSIS watchdog finds possible breach of secrets in fighting high-tech crime. Computers & Security, 14(1):

37, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/0167404895900837.

Collinson:1995:CL

[Col95x] Helen Collinson. Cyberspace lawsuits. Computers & Security, 14(3):220, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/016740489597065I.

Collinson:1995:CMP

[Col95y] Helen Collinson. Cyberspace magistrates are proposed. Computers & Security, 14(8):698, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/0167404896817172.

Collinson:1995:DSU

[Col95z] Helen Collinson. Dial a star? Utahn faces felony for accessing numbers. Computers & Security, 14(1):36, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/0167404895900799.

Collinson:1995:DPL

[Col95-27] Helen Collinson. Digital privacy … take two. Computers & Security, 14(8):696–697, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/016740489681710X.

Collinson:1995:DDC

[Col95-28] Helen Collinson. Direct debiting: a case of serious fraud. Computers & Security, 14(5):410, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/016740489597093P.

Collinson:1995:DDV

[Col95-29] Helen Collinson. Document describing virus contains one. Computers & Security, 14(6):519, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/0167404895903259.

Collinson:1995:MSI

[Col95-30] Helen Collinson. E-mail security ignored. Computers & Security, 14(7):614, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404896816932.

Collinson:1995:EHS

[Col95-31] Helen Collinson. Economical, high-security payphone systems with Watermark Magnetics cards from Ascom. Computers & Security, 14(8):701, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404896817287.

Collinson:1995:EPA

[Col95-32] Helen Collinson. An efficient password authentication scheme based on a unit circle. Computers & Security, 14(3):220, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/016740489597066J.

Collinson:1995:EWS

[Col95-33] Helen Collinson. The electronic wallet: secure payment methods sought for Internet. Computers & Security, 14(1):36, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/0167404895970175.

Collinson:1995:ELB

[Col95-34] Helen Collinson. An expensive lesson for bulletin board user. Computers & Security, 14(6):520, ???? 1995. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/0167404895903275.

Collinson:1995:FPS

[Col95-35] Helen Collinson. Faced with a password solution. Computers & Security, 14(7):613, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404895901345.

Collinson:1995:FMC

[Col95-36] Helen Collinson. A fast method for cryptanalysis of substitution ciphers. Computers & Security, 14(5):414, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/016740489597105J.

Collinson:1995:FLI

[Col95-37] Helen Collinson. Feds limit info warfare role. Computers & Security, 14(6):522, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/016740489599324I.

Collinson:1995:FTI

[Col95-38] Helen Collinson. Feds target the Internet. Computers & Security, 14(3):218–219, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/016740489597061E.

Collinson:1995:FCU

[Col95-39] Helen Collinson. Firewalls come under fire. Computers & Security, 14(3):218, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/016740489597060N.

Collinson:1995:FBS

[Col95-40] Helen Collinson. Forum to bring security to ATM. Computers & Security, 14(4):300–301, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/016740489597079P.

Collinson:1995:FDM

[Col95-41] Helen Collinson. Fresh drive on mobile phone fraud. Computers & Security, 14(3):218, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/016740489597059J.

Collinson:1995:GHS

[Col95-42] Helen Collinson. Gaping holes of solicitors and accountants. Computers & Security, 14(5):410–411, ???? 1995. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/016740489597094Q.

Collinson:1995:GAF

[Col95-43] Helen Collinson. Group assails federal security proposals. Computers & Security, 14(8):698, ???? 1995. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/0167404896817184.

Collinson:1995:HAI

[Col95-44] Helen Collinson. Hacker alerts intuit to problems with its US tax filing system. Computers & Security, 14(2):117–118, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/0167404895970326.

Collinson:1995:HAT

[Col95-45] Helen Collinson. Hacker, awaiting trial, attends DefCon. Computers & Security, 14(5):409, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/016740489597091N.

Collinson:1995:HHM

[Col95-46] Helen Collinson. Hacker heaven: so many computers so few safe-guards. Computers & Security, 14(4):300, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/0167404895900438.

Collinson:1995:HTH

[Col95-47] Helen Collinson. Hacker-trackers hunt electronic cybercrooks. Computers & Security, 14(3):216–217, ???? 1995. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/016740489597058I.

Collinson:1995:HCB

[Col95-48] Helen Collinson. Hackers crack BSykB code. Computers & Security, 14(5):409, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404896900334.

Collinson:1995:HLA

[Col95-49] Helen Collinson. A hacker’s lines of attack. Computers & Security, 14(2):120, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404895970429.

Collinson:1995:HTR

[Col95-50] Helen Collinson. Hackers take revenge on the author of new book on cyberspace wars. Computers & Security, 14(1):35, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/0167404895970142.

Collinson:1995:HWS

[Col95-51] Helen Collinson. Help: working smarter lock, stock and password. Computers & Security, 14(1):39, ???? 1995. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/0167404895970223.

Collinson:1995:HN

[Col95-52] Helen Collinson. Holes in the net. Computers & Security, 14(5): 413, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/016740489597102G.

Collinson:1995:HHT

[Col95-53] Helen Collinson. Hot on the hackers trail. Computers & Security, 14(3):215, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404895903658.

Collinson:1995:HSS

[Col95-54]  Helen Collinson. How secure is security?  Comput-

ers & Security, 14(7):615–616, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/0167404896817007.

Collinson:1995:HPD

[Col95-55] Helen Collinson. How to protect distributed data. Computers & Security, 14(6):521, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/016740489599323H.

Collinson:1995:IUW

[Col95-56] Helen Collinson. IBM unveils web servers, Internet firewall. Computers & Security, 14(7):616, ???? 1995. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/0167404896817044.

Collinson:1995:ITM

[Col95-57] Helen Collinson. IMB targets market for anti-virus software. Computers & Security, 14(4):299, ???? 1995. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/0167404895900373.

Collinson:1995:RVH

[Col95-58] Helen Collinson. In Rindos v. Hardwick. Computers & Security, 14(6):520–521, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/016740489599320O.

Collinson:1995:ISS

[Col95-59] Helen Collinson. Industrial-strength security for the Internet. Computers & Security, 14(8):699, ???? 1995. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/0167404896817202.

Collinson:1995:ISB

[Col95-60] Helen Collinson. Internet security boost. Computers & Security, 14(3):217, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/016740489597057H.

Collinson:1995:ISH

[Col95-61] Helen Collinson. Internet security holes abound. Computers & Security, 14(3):216, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404895903674.

Collinson:1995:IUF

[Col95-62] Helen Collinson. Internet users fret over spies’ new role. Computers & Security, 14(6):518, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/016740489599314F.

Collinson:1995:IP

[Col95-63] Helen Collinson. Invasions of privacy. Computers & Security, 14 (7):616, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/0167404896817032.

Collinson:1995:ASA

[Col95-64] Helen Collinson. Is anything safe anymore? Computers & Security, 14(1):39, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404895900934.

Collinson:1995:SIT

[Col95-65] Helen Collinson. Is security an issue? Then a BBS may be for you. Computers & Security, 14(5):413–414, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/016740489690036X.

Collinson:1995:JSC

[Col95-66] Helen Collinson. A job for security counsel. Computers & Security, 14(5):411–412, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/016740489597097T.

Collinson:1995:JAG

[Col95-67] Helen Collinson. Joint action group fights computer crime. Computers & Security, 14(7):613, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/0167404895901329.

Collinson:1995:KBS

[Col95-68] Helen Collinson. Keeping business safe on the Internet. Computers & Security, 14(3):219, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/016740489597062F.

Collinson:1995:LDI

[Col95-69]  Helen Collinson.  Living dangerously on the Internet.

Computers & Security, 14(3):217, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/0167404895903712.

Collinson:1995:LDM

[Col95-70] Helen Collinson. Logicode’s datasafe modem provides security at 28.8 kbps. Computers & Security, 14(2):121, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/016740489597044B.

Collinson:1995:LBI

[Col95-71] Helen Collinson. London Borough innocent victim of computer fraud. Computers & Security, 14(1):36–37, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/0167404895900810.

Collinson:1995:MRC

[Col95-72] Helen Collinson. Making the right connections. Computers & Security, 14(8):700, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404896817251.

Collinson:1995:MWV

[Col95-73] Helen Collinson. Malicious word virus found on the Internet. Computers & Security, 14(6):519, ???? 1995. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/0167404895903232.

Collinson:1995:MOS

[Col95-74] Helen Collinson. MCII offers secure transaction service. Computers & Security, 14(7):614–615, ???? 1995. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/0167404896816968.

Collinson:1995:MRP

[Col95-75] Helen Collinson. Medical records, privacy on the line. Computers & Security, 14(2):118, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/0167404895970348.

Collinson:1995:MCI

[Col95-76] Helen Collinson. Microsoft criticized over Internet security claims. Computers & Security, 14(6):520, ???? 1995. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/016740489599319K.

Collinson:1995:MTW

[Col95-77] Helen Collinson. Microsoft tackles win 95 pirates. Computers & Security, 14(5):412, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/016740489597098U.

Collinson:1995:MTN

[Col95-78] Helen Collinson. Minnesota teen nabbed in computer hacker case. Computers & Security, 14(1):36, ???? 1995. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/0167404895970153.

Collinson:1995:MSW

[Col95-79] Helen Collinson. More security woes plague the Internet. Computers & Security, 14(7):613–614, ???? 1995. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/0167404896816920.

Collinson:1995:MMP

[Col95-80] Helen Collinson. Motorola, MOBIUS pact aims at halting hackers. Computers & Security, 14(2):117, ???? 1995. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/0167404895900535.

Collinson:1995:MVB

[Col95-81] Helen Collinson. Mystery ‘virus’ bugged budget. Computers & Security, 14(2):117, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404895970315.

Collinson:1995:NSR

[Col95-82] Helen Collinson. Net security reawakening. Computers & Security, 14(4):299–300, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/016740489597076M.

Collinson:1995:NSC

[Col95-83] Helen Collinson. Net surfer convicted for cache of child porn disks. Computers & Security, 14(7):612, ???? 1995. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/0167404895901302.

Collinson:1995:NSU

[Col95-84] Helen Collinson. Netscape signs up 19 users for its system of Internet security. Computers & Security, 14(2):118–119, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/0167404895900594.

Collinson:1995:NCSa

[Col95-85] Helen Collinson. New computer system causes chaos. Computers & Security, 14(2):120–121, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/016740489597043A.

Collinson:1995:NCSb

[Col95-86] Helen Collinson. New computer system causes chaos. Computers & Security, 14(3):216, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/016740489597056G.

Collinson:1995:NSP

[Col95-87] Helen Collinson. A new standard is proposed for electronic verification. Computers & Security, 14(4):298, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/016740489597073J.

Collinson:1995:NUN

[Col95-88] Helen Collinson. NT, Unix network security shapes up. Computers & Security, 14(8):697, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/0167404896817111.

Collinson:1995:OVL

[Col95-89] Helen Collinson. Old viruses licked, but new ones find fodder in Windows 95. Computers & Security, 14(8):699, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/0167404896817214.

Collinson:1995:OFA

[Col95-90] Helen Collinson. Online ‘firewalls’ aim to keep spies in cold. Computers & Security, 14(4):298, ???? 1995. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/016740489597072I.

Collinson:1995:ODC

[Col95-91] Helen Collinson. Only disconnect: can pornography in cyberspace be regulated? Computers & Security, 14(4):297, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/0167404895900330.

Collinson:1995:OI

[Col95-92] Helen Collinson. Operational insurance. Computers & Security, 14(8):700, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404896817226.

Collinson:1995:OEN

[Col95-93] Helen Collinson. Oracle extinguishes ’net security fears. Computers & Security, 14(7):615, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/016740489681697X.

Collinson:1995:PFP

[Col95-94] Helen Collinson. PeopleSoft fixes password security flaw. Computers & Security, 14(8):697–698, ???? 1995. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/0167404896817147.

Collinson:1995:PNA

[Col95-95] Helen Collinson. A piece of the ’net action. Computers & Security, 14(8):697, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404896817123.

Collinson:1995:PBCa

[Col95-96] Helen Collinson. Police break CD pirate ring in Cambridge. Computers & Security, 14(5):411, ???? 1995. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/016740489597095R.

Collinson:1995:PLW

[Col95-97] Helen Collinson. Present law will not stop the virus writers. Computers & Security, 14(8):697, ???? 1995. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/0167404896817135.

Collinson:1995:PBCb

[Col95-98] Helen Collinson. Pressure builds on chip thieves. Computers & Security, 14(7):612–613, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/0167404896816919.

Collinson:1995:PCE

[Col95-99] Helen Collinson. Privacy called endangered species. Computers & Security, 14(2):118, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404895900551.

Collinson:1995:PYM

[Col95-100] Helen Collinson. Protect your e-mail. Computers & Security, 14 (7):614, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/0167404896816944.

Collinson:1995:PYP

[Col95-101] Helen Collinson. Protecting your privacy. Computers & Security, 14(3):219–220, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/016740489597064H.

Collinson:1995:PSN

[Col95-102] Helen Collinson. Providing a safety net. Computers & Security, 14(5):413, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/016740489597101F.

Collinson:1995:PWP

[Col95-103] Helen Collinson. Pssst, what’s password again? Computers & Security, 14(5):412, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/016740489597099V.

Collinson:1995:RTP

[Col95-104] Helen Collinson. Raids target piracy suspects. Computers & Security, 14(7):612, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404896816890.

Collinson:1995:RHH

[Col95-105] Helen Collinson. A recipe for hacker heartburn. Computers & Security, 14(1):39, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404895970245.

Collinson:1995:RGS

[Col95-106] Helen Collinson. Routers get security auditing software. Computers & Security, 14(6):521, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/016740489599322G.

Collinson:1995:SSc

[Col95-107] Helen Collinson. Safe and secure. Computers & Security, 14(8): 696, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404895900306.

Collinson:1995:SSb

[Col95-108] Helen Collinson. Safe secret. Computers & Security, 14(6):519– 520, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/016740489599318J.

Collinson:1995:SDB

[Col95-109] Helen Collinson. Safeguarding databases: basic concepts revisited. Computers & Security, 14(4):302, ???? 1995. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/016740489597084N.

Collinson:1995:SM

[Col95-110] Helen Collinson. Safety on the move. Computers & Security, 14 (8):698, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/0167404896817159.

Collinson:1995:SCD

[Col95-111] Helen Collinson. SATAN and Courtney: a devil of a team. Computers & Security, 14(6):522, ???? 1995. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/016740489599325J.

Collinson:1995:SGE

[Col95-112] Helen Collinson. Satan: good or evil? Computers & Security, 14 (4):300, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/016740489597078O.

Collinson:1995:SPa

[Col95-113] Helen Collinson. SCO patches. Computers & Security, 14(1): 38, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/016740489597020B.

Collinson:1995:SPb

[Col95-114] Helen Collinson. Secret plans. Computers & Security, 14(3): 219, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/016740489597063G.

Collinson:1995:SAS

[Col95-115] Helen Collinson. Secure all stations. Computers & Security, 14 (4):301, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/016740489597081K.

Collinson:1995:SEI

[Col95-116] Helen Collinson. Secure EDI over the Internet. Computers & Security, 14(1):39, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404895970234.

Collinson:1995:SUE

[Col95-117] Helen Collinson. Secure Unix for enterprise computing. Computers & Security, 14(3):219, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/0167404895903798.

Collinson:1995:SOD

[Col95-118] Helen Collinson. Securicor offers data strongroom. Computers & Security, 14(4):299, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/016740489597074K.

Collinson:1995:SI

[Col95-119] Helen Collinson. Securing the Internet. Computers & Security, 14(3):218, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404895903771.

Collinson:1995:SAL

[Col95-120] Helen Collinson. Security across the LAN. Computers & Security, 14(1):39, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404895900918.

Collinson:1995:SA

[Col95-121] Helen Collinson. Security alert. Computers & Security, 14(3): 216, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/0167404895903690.

Collinson:1995:SDI

[Col95-122] Helen Collinson. Security and data integrity for LANS and WANs. Computers & Security, 14(7):616, ???? 1995. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/0167404896817020.

Collinson:1995:SAC

[Col95-123] Helen Collinson. Security aspects of card systems. Computers & Security, 14(2):120, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/016740489597039D.

Collinson:1995:SER

[Col95-124] Helen Collinson. Security environment reviews. Computers & Security, 14(7):616, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404896817019.

Collinson:1995:SEW

[Col95-125] Helen Collinson. Security expert warns of night-time intruders. Computers & Security, 14(7):611, ???? 1995. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/0167404896816877.

Collinson:1995:SFF

[Col95-126] Helen Collinson. Security flaw found in Netscape software. Computers & Security, 14(6):518, ???? 1995. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/016740489599315G.

Collinson:1995:SLB

[Col95-127] Helen Collinson. Security: on-line business: safer than you think? Computers & Security, 14(4):302, ???? 1995. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/016740489597085O.

Collinson:1995:SO

[Col95-128] Helen Collinson. Security overkill? Computers & Security, 14 (8):701, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/0167404896817263.

Collinson:1995:SPL

[Col95-129] Helen Collinson. Security plans lag computer crime rate. Computers & Security, 14(8):699, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/0167404896817196.

Collinson:1995:SSA

[Col95-130] Helen Collinson. Security starts with access control. Computers & Security, 14(2):118, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404895970337.

Collinson:1995:SCC

[Col95-131] Helen Collinson. Security: the complex crypto relationship. Computers & Security, 14(4):299, ???? 1995. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/016740489590039X.

Collinson:1995:SCA

[Col95-132] Helen Collinson. Sentry cuts access to naughty bits. Computers & Security, 14(7):615, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/0167404896816993.

Collinson:1995:SHC

[Col95-133] Helen Collinson. Sign here — caught in the cross fire of digital signatures. Computers & Security, 14(2):121, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/016740489597046D.

Collinson:1995:SSBa

[Col95-134] Helen Collinson. Simplified sign-on brings enhanced security features. Computers & Security, 14(8):698, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/0167404896817160.

Collinson:1995:SHS

[Col95-135] Helen Collinson. Six hackers stung by undercover cybercops: Gary Anthes. Computerworld, September 18, 1995, p.

12. Computers & Security, 14(6):518–519, ???? 1995. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/016740489599316H.

Collinson:1995:SPD

[Col95-136] Helen Collinson. Software piracy declines in Asia, elsewhere. Computers & Security, 14(2):118, ???? 1995. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/0167404895970359.

Collinson:1995:SPT

[Col95-137] Helen Collinson. Software pirates treasure win 95. Computers & Security, 14(7):614, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404896816956.

Collinson:1995:SFY

[Col95-138] Helen Collinson. Step away from that firewall with your hands up. Computers & Security, 14(4):300, ???? 1995. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/016740489597077N.

Collinson:1995:SC

[Col95-139] Helen Collinson. Stomping out cybersmut. Computers & Security, 14(4):299, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404895900411.

Collinson:1995:STS

[Col95-140] Helen Collinson. Strong theoretical stenography. Computers & Security, 14(5):414, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/016740489597104I.

Collinson:1995:SAG

[Col95-141] Helen Collinson. Summit addresses growth, security issues for Internet. Computers & Security, 14(3):218, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/0167404895903755.

Collinson:1995:SPR

[Col95-142] Helen Collinson. Swedish pirates rumbled. Computers & Security, 14(4):297, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404895900357.

Collinson:1995:TMN

[Col95-143] Helen Collinson. Taking the measure of NetWare antivirus software. Computers & Security, 14(6):522, ???? 1995. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/016740489599326K.

Collinson:1995:TE

[Col95-144] Helen Collinson. Tales from the encrypt. Computers & Security, 14(4):299, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/016740489597075L.

Collinson:1995:TDR

[Col95-145] Helen Collinson. Testing the disaster recovery plan. Computers & Security, 14(4):302, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/016740489597083M.

Collinson:1995:TYV

[Col95-146] Helen Collinson. Think you’re virus-proof?. Think again. Computers & Security, 14(6):521–522, ???? 1995. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/0167404895903291.

Collinson:1995:TBD

[Col95-147] Helen Collinson. Time bombs, drop dead devices and electronic software repossession. Computers & Security, 14(2): 121, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/016740489597045C.

Collinson:1995:TCS

[Col95-148] Helen Collinson. Top cop says fingerprint system is fraudbeater. Computers & Security, 14(7):611–612, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/0167404896816889.

Collinson:1995:UTA

[Col95-149] Helen Collinson. Underground tools aid fledgling hackers. Computers & Security, 14(7):615, ???? 1995. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/0167404896816981.

Collinson:1995:UCR

[Col95-150] Helen Collinson. Unravelling copyright rules for cyberspace. Computers & Security, 14(2):119, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/016740489597037B.

Collinson:1995:URS

[Col95-151] Helen Collinson. US rattles a sabre at Chinese pirates. Computers & Security, 14(1):37–38, ???? 1995. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/0167404895970197.

Collinson:1995:UPF

[Col95-152] Helen Collinson. US to propose federal agency to secure information highway. Computers & Security, 14(4):297–298, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/016740489597071H.

Collinson:1995:VTG

[Col95-153] Helen Collinson. Virus threat gains momentum. Computers & Security, 14(3):215–216, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/016740489597055F.

Collinson:1995:VWE

[Col95-154] Helen Collinson. Virus wars escalate. Computers & Security, 14 (8):696, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404895900284.

Collinson:1995:WDH

[Col95-155] Helen Collinson. War declared on hackers. Computers & Security, 14(3):217, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404895903739.

Collinson:1995:WCC

[Col95-156] Helen Collinson. What’s copyrightable in computer software? Computers & Security, 14(2):118, ???? 1995. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/0167404895900578.

Collinson:1995:WHS

[Col95-157] Helen Collinson. White house seeks copyright protection for text, software published online. Computers & Security, 14(5):

411, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/0167404896900346.

Collinson:1995:WPV

[Col95-158] Helen Collinson. Will privacy vanish along with wires? Computers & Security, 14(1):37, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/0167404895970186.

Collinson:1995:WL

[Col95-159] Helen Collinson. Winners and losers. Computers & Security, 14 (3):215, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/0167404895903631.

Collinson:1995:WVA

[Col95-160] Helen Collinson. Word is out on virus that attacks macros. Computers & Security, 14(5):412–413, ???? 1995. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/016740489597100O.

Collinson:1995:WAC

[Col95-161] Helen Collinson. Workstation access control. Computers & Security, 14(2):121, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404895900616.

Collinson:1995:WIS

[Col95-162] Helen Collinson. World institute for security enhancement. Computers & Security, 14(1):38, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/016740489590087X.

Collinson:1996:AAR

[Col96a] Helen Collinson. Abstract of article and recent literature. Computers & Security, 15(1):44–49, ???? 1996. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/016740489687622X.

Collinson:1996:ANS

[Col96b] Helen Collinson. Adding ‘net security. Computers & Security, 15(2):127, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404896893448.

Collinson:1996:BH

[Col96c] Helen Collinson. Battening down the hatches. Computers & Security, 15(2):127–128, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/0167404896893461.

Collinson:1996:CGT

[Col96d] Helen Collinson. Cylink, GTE target ATM encryption. Computers & Security, 15(2):124–125, ???? 1996. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/016740489689331X.

Collinson:1996:LSS

[Col96e] Helen Collinson. DE-light shadowed by security concerns. Computers & Security, 15(2):126–127, ???? 1996. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/0167404896893412.

Collinson:1996:MPG

[Col96f] Helen Collinson. E-mail privacy: a glass almost half full. Computers & Security, 15(2):126, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/0167404896893370.

Collinson:1996:ESG

[Col96g] Helen Collinson. Encryption spec gains momentum. Computers & Security, 15(2):123, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404896893266.

Collinson:1996:EEF

[Col96h] Helen Collinson. Expert: encryption flaws no secret. Computers & Security, 15(2):125, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/0167404896893357.

Collinson:1996:FGW

[Col96i] Helen Collinson. Firewall gets Web capabilities. Computers & Security, 15(2):123–124, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/0167404896893278.

Collinson:1996:FSU

[Col96j] Helen Collinson. Firewall services unveiled. Computers & Security, 15(2):124, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404896893291.

Collinson:1996:FTH

[Col96k] Helen Collinson. Firewall turns up the heat on Internet hackers. Computers & Security, 15(2):127, ???? 1996. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/016740489689345X.

Collinson:1996:HFH

[Col96l] Helen Collinson. HP finds high security ‘net niche. Computers & Security, 15(2):123, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404896893242.

Collinson:1996:ISH

[Col96m] Helen Collinson. Internet security: how much is enough? Computers & Security, 15(2):128, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/0167404896893473.

Collinson:1996:ILY

[Col96n] Helen Collinson. Intranets: living in your own private Internet. Computers & Security, 15(2):127, ???? 1996. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/0167404896893436.

Collinson:1996:JCT

[Col96o] Helen Collinson. Javascript can trace user tracks. Computers & Security, 15(2):126, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404896893394.

Collinson:1996:KCC

[Col96p] Helen Collinson. Kerberos code crack raises broader issues. Computers & Security, 15(2):123, ???? 1996. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/0167404896893254.

Collinson:1996:NGM

[Col96q] Helen Collinson. NT gets more secure. Computers & Security, 15(2):124, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404896893308.

Collinson:1996:OSP

[Col96r] Helen Collinson. Online security a pipe dream? Computers & Security, 15(2):125, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404896893345.

Collinson:1996:SS

[Col96s] Helen Collinson. The secret struggle. Computers & Security, 15 (2):127, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/0167404896893424.

Collinson:1996:SRR

[Col96t] Helen Collinson. Security risk remains for Web mail. Computers & Security, 15(2):124, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/016740489689328X.

Collinson:1996:SFC

[Col96u] Helen Collinson. Still a few chinks in Java’s armour. Computers & Security, 15(2):125, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/0167404896893333.

Collinson:1996:SCS

[Col96v] Helen Collinson. Swiss-cheese secutity. Computers & Security, 15(2):125, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404896893321.

Collinson:1996:UST

[Col96w] Helen Collinson. Uncle Sam has three-tier scheme for data security. Computers & Security, 15(2):126, ???? 1996. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/0167404896893400.

Collinson:1996:WSF

[Col96x] Helen Collinson. Webscan sets up ‘firewall’. Computers & Security, 15(2):126, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404896893382.

Collinson:1996:WSM

[Col96y] Helen Collinson. Win 95 security measures mollify users. Computers & Security, 15(2):125–126, ???? 1996. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/0167404896893369.

Collins:1999:TCPa

[Col99a] Michael Collins. Telecommunications crime — Part 1. Computers & Security, 18(7):577–586, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404899820045.

Collins:1999:TCPb

[Col99b] Michael Collins. Telecommunications crime — Part 2. Computers & Security, 18(8):683–692, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404899801321.

Corum:1995:CSP

[Cor95a] P. J. Corum. Client server — promises, problems and solutions. Computers & Security, 14(5):422, ???? 1995. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/016740489597129X.

Corum:1995:OOA

[Cor95b] P. J. Corum. Object oriented: Another silver bullet. Computers & Security, 14(5):426, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/0167404895971473.

Corcoran:1997:IWD

[Cor97] Michael J. Corcoran. Information warfare and defending the UK nation state. Computers & Security, 16(6):525, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404897846950.

Courtney:1994:LE

[Cou94] Robert H. Courtney. Letter to the editor. Computers & Security, 13(1):34–36, February 1994. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/0167404894900892.

Cohen:1998:CEM

[CPS^{+}98] Fred Cohen, Cynthia Phillips, Laura Painton Swiler, Timothy Gaylor, Patricia Leary, Fran Rupley, and Richard Isler. A cause and effect model of attacks on information systems: Some analysis based on that model, and the application of that model for CyberWarfare in CID. Computers & Security, 17(3):211– 221, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S016740489880312X.

Cumming:1993:LE

[CR93] Andrea Cumming and Steve Roberts. Letters to the editor. Computers & Security, 12(2):136–138, March 1993. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/016740489390094L.

Crichard:1997:ILQ

[Cri97] Mark Crichard. The Internet — a legal quagmire. Computers & Security, 16(6):526, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404897847013.

Crocker:1995:AAT

[Cro95a] Norman Crocker. Automated audit — tools and techniques. Computers & Security, 14(5):423–424, ???? 1995. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/016740489597138Z.

Crocker:1995:MSA

[Cro95b] Norman Crocker. MVS/SMF audit trail processing. Computers & Security, 14(5):420, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/016740489597122Q.

Crocker:1997:KCR

[Cro97] Norman Crocker. Key concerns in a review of CAACF2/ MVS. Computers & Security, 16(6):524, ???? 1997. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404897846925.

Crocker:1998:KCR

[Cro98] Norman Crocker. Key concerns in a review of CA-ACF2/ MVS. Computers & Security, 17(1):42–53, ???? 1998. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404897802490.

Caelli:1992:EHU

[CRR92] W. J. Caelli, A. W. Rhodes, and N. C. Russell. An evaluation of HP-UX (UNIX) for database protection using the European ITSEC. Computers & Security, 11(5):463–479, September 1992. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/016740489290012G.

Chang:1994:ACB

[CSW94] Chin-Chen Chang, Jao-Ji Shen, and Tzong-Chen Wu. Access control with binary keys. Computers & Security, 13(8):681– 686, ???? 1994. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/0167404894900507.

Curtis:1995:EWW

[Cur95] Mark Curtis. Evaluation of Windows 95 and Windows NT security controls and architecture. Computers & Security, 14(5): 426, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/0167404895971509.

Caminada:1998:ISI

[CvdRvZvD98] Martin Caminada, Reind van de Riet, Arjen van Zanten, and Leendert van Doom. Internet security incidents, a survey within Dutch organizations. Computers & Security, 17(5):417–433, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404898800667.

Courtney:1994:WDW

[CW94] Robert H. Courtney and Willis H. Ware. What do we mean by integrity? Computers & Security, 13(3):206–208, May 1994. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/0167404894900701.

David:1990:NV

[Dav90a] Jon David. The Novell virus. Computers & Security, 9(7): 593–599, November 1990. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/016740489090056Y.

David:1990:NG

[Dav90b] Jon David. The numbers game. Computers & Security, 9 (5):403–407, August 1990. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/016740489090070A.

David:1992:LSS

[Dav92] Jon David. LAN security standards. Computers & Security, 11(7):607–619, November 1992. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/016740489290137G.

Davies:1993:LE

[Dav93] Donald Davies. Letter to the editor. Computers & Security, 12(1):32, February 1993. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/016740489390007R.

David:1996:NFV

[Dav96] Jon David. The new face of the virus threat. Computers & Security, 15(1):13–16, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/0167404896876188.

Deane:1995:PAB

[DBHM95] Frank Deane, Kate Barrelle, Ron Henderson, and Doug Mahar. Perceived acceptability of biometric security systems. Computers & Security, 14(3):225–231, ???? 1995. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/016740489500005S.

Davida:1992:PVV

[DD92] George I. Davida and Yvo G. Desmedt. Passports and visas versus IDs. Computers & Security, 11(3):253–258, May 1992. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/0167404892902056.

Dawson:1994:BSS

[DD94] Ed Dawson and Diane Donovan. The breadth of Shamir’s secret-sharing scheme. Computers & Security, 13(1):69–78, February 1994. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/0167404894900973.

Devargas:1999:SCI

[Dev99] Mario Devargas. Survival is not compulsory: an introduction to business continuity planning. Computers & Security, 18(1):35– 46, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404899800078.

Domingo-Ferrer:1991:ASA

[DF91] Joseph Domingo-Ferrer. Algorithm-sequenced access control. Computers & Security, 10(7):639–652, November 1991. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/016740489190122T.

Domingo-Ferrer:1990:SNB

[DFHR90] Josep Domingo-Ferrer and Llorenç Huguet-Rotger. Secure network bootstrapping: an algorithm for authentic key exchange and digital signitures. Computers & Security, 9(2):145–152, April 1990. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/016740489090088B.

Dowland:1999:CCA

[DFIR99] P. S. Dowland, S. M. Furnell, H. M. Illingworth, and P. L. Reynolds. Computer crime and abuse: a survey of public attitudes and awareness. Computers & Security, 18(8):715– 726, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404899801357.

Dawson:1998:MME

[DG98] E. P. Dawson and H. M. Gustafson. A method for measuring entropy of symmetric cipher key generators. Computers & Security, 17(2):177–184, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S016740489800011X.

DiTuri:1993:SEM

[Di 93] Barbara B. Di Turi. Security for EDIFACT messages. Computers & Security, 12(5):447–455, August 1993. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/016740489390063B.

deKoning:1995:MDS

[dK95] W. Fred. de Koning. A methodology for the design of security plans. Computers & Security, 14(7):633–643, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/0167404895000208.

Denault:1994:IDA

[DKGS94] Michel Denault, Dimitris Karagiannis, Dimitris Gritzalis, and Paul Spirakis. Intrusion detection: Approach and performance issues of the SECURENET system. Computers & Security, 13(6):495–508, ???? 1994. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/0167404891901384.

Dixon:1992:RJC

[DMC92] R. Dixon, C. Marston, and P. Collier. A report on the joint CIMA and IIA computer fraud survey. Computers & Security, 11(4):307–313, July 1992. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/016740489290170V.

Doumas:1995:DNN

[DMGK95] Anastasia Doumas, Konstantinos Mavroudakis, Dimitris Gritzalis, and Sokratis Katsikas. Design of a neural network for recognition and classification of computer viruses. Computers & Security, 14(5):435–448, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/016740489500008V.

Docherty:1997:IE

[Doc97] Paul Docherty. Internet encryption. Computers & Security, 16 (6):518, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404897846731.

Donovan:1993:SPD

[Don93] Steve Donovan. Security of PCs in a distributed environment. Computers & Security, 12(1):28–31, February 1993. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/016740489390006Q.

Dosdale:1995:ESI

[Dos95] Terry Dosdale. EDI security and the Internet. Computers & Security, 14(5):426, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/016740489597145Z.

Dosdale:1997:ECW

[Dos97] Terry Dosdale. Electronic commerce — the way forward. Computers & Security, 16(6):523, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S016740489784686X.

deRu:1996:RAM

[dRE96] W. G. de Ru and J. H. P. Eloff. Risk analysis modelling with the use of fuzzy logic. Computers & Security, 15(3):239– 248, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/0167404896000089.

Docherty:1999:MAW

[DS99] Paul Docherty and Peter Simpson. Macro attacks: What next after Melissa? Computers & Security, 18(5):391–395, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404899800844.

Duncan:1995:TSC

[Dun95] Rebecca J. Duncan. There are some cracks in the cornerstone of information security. Computers & Security, 14(8):675– 680, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/016740489681707X.

deVries:1992:IT

[dV92] T. P. de Vries. The implementation of TSS. Computers & Security, 11(4):315–325, July 1992. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/016740489290171M.

Eloff:1996:SCD

[EHT96] J. H. P. Eloff, R. Holbein, and S. Teufel. Security classification for documents. Computers & Security, 15(1):55–71, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/0167404895000232.

Eloff:1993:CFR

[ELB93] J. H. P. Eloff, L. Labuschagne, and K. P. Badenhorst. A comparative framework for risk analysis methods. Computers & Security, 12(6):597–603, October 1993. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/016740489390056B.

Evans:1990:CRC

[EM90] Paul Evans and Tina Monk. Conference report: COMPSEC ’89. Computers & Security, 9(1):41–44, February 1990. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/016740489090154L.

Ekenberg:1995:CMM

[EOO95] Love Ekenberg, Subhash Oberoi, and István Orci. A cost model for managing information security hazards. Computers & Security, 14(8):707–717, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/0167404895000216.

Evans:1990:HCP

[Eva90] Paul Evans. How can professional standards be set and maintained? Computers & Security, 9(4):313–316, June 1990. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/0167404890901008.

Evans:1991:CFS

[Eva91] Paul Evans. Computer fraud — the situation, detection and training. Computers & Security, 10(4):325–327, June 1991. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/016740489190107O.

Evans:1992:EAG

[Eva92] David Evans. East Asian gang counterfeits Microsoft MS-DOS holograms. Computers & Security, 11(5):487, September 1992. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/016740489290020R.

Evans:1996:IPP

[Eva96] Paul Evans. Information protection for publishers. Computers & Security, 15(3):209–211, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/0167404896897835.

Fagan:1993:OII

[Fag93] Peter Fagan. Organizational issues in IT security. Computers & Security, 12(8):710–715, December 1993. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/0167404893900332.

Faak:1991:CVH

[Fåk91] Viiveke Fåk. Computer verification of human users’ identity: a theoretical model and some evaluation criteria. Computers & Security, 10(7):626–636, November 1991. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/0167404891901203.

Farquhar:1991:OAR

[Far91] Bill Farquhar. One approach to risk assessment. Computers & Security, 10(1):21–23, February 1991. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/016740489190051E.

Fawcett:1992:FPT

[Faw92a] Neil Fawcett. FAST pushes for tough action on Unix piracy. Computers & Security, 11(5):487–488, September 1992. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/016740489290024L.

Fawcett:1992:GPA

[Faw92b] Neil Fawcett. German pirates avoid blacklist. Computers & Security, 11(5):489, September 1992. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/016740489290033N.

Ferelli:1992:NDS

[Fer92] Mark Ferelli. New disk strategies remove security risks. Computers & Security, 11(5):490, September 1992. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/016740489290246N.

Fernandez-Gonzalez:1993:EEC

[FGAR93] J. Fernández-González, G. B. Agnew, and A. Ribagorda. Encryption and error correction codes for reliable file storage. Computers & Security, 12(5):501–510, August 1993. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/016740489390070L.

Fitch:1993:SPN

[FH93] John A. Fitch and Lance J. Hoffman. A shortest path network security model. Computers & Security, 12(2):169–189, March 1993. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/016740489390100J.

Finne:1996:ISC

[Fin96] Thomas Finne. The information security chain in a company. Computers & Security, 15(4):297–316, ???? 1996. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/0167404896889413.

Finne:1997:ISI

[Fin97a] Thomas Finne. Information security implemented in: the theory on stock market efficiency, Markowitz’s portfolio theory and porter’s value chain. Computers & Security, 16(6):469– 479, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S016740489784641X.

Finne:1997:WIS

[Fin97b] Thomas Finne. What are the information security risks in decision support systems and data warehousing? Computers & Security, 16(3):197–204, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404897845129.

Finne:1998:CFI

[Fin98a] Thomas Finne. A conceptual framework for information security management. Computers & Security, 17(4):303–307, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404898800102.

Finne:1998:TCD

[Fin98b] Thomas Finne. The three categories of decision-making and information security. Computers & Security, 17(5):397–405, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S016740489880045X.

Foley:1997:BCW

[Fol97] Simon N. Foley. Building Chinese walls in standard Unix^{T M}. Computers & Security, 16(6):551–563, ???? 1997. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404897000102.

Ford:1998:MYL

[For98a] Richard Ford. May you live in interesting times. Computers & Security, 17(7):575–578, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404899800558.

Ford:1998:WLD

[For98b] Richard Ford. What a little definition can do . . .. Computers & Security, 17(2):110–114, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404897819785.

Ford:1999:MTR

[For99a] Richard Ford. Malware: Troy revisited. Computers & Security, 18(2):105–108, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404899800273.

Ford:1999:NSM

[For99b] Richard Ford. No suprises in Melissa land. Computers & Security, 18(4):300–302, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S016740489980073X.

France:1995:DPI

[Fra95] Elizabeth France. Data protection in the information superhighway era. Computers & Security, 14(5):424, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/016740489597141V.

Frank:1996:SIV

[Fra96] Robert L. (Bob) Frank. Security issues in the virtual corporation. Computers & Security, 15(6):471–476, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404897831203.

Farmer:1995:IVE

[FW95] Doc Farmer and S. G. Warburg. Internal vs. external IT audits — or — mapping out a war zone? Computers & Security, 14(5):419–420, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/016740489597121P.

Furnell:1999:CHC

[FW99] S. M. Furnell and M. J. Warren. Computer hacking and cyber terrorism: the real threats in the new millennium? Computers & Security, 18(1):28–34, ???? 1999. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404899800066.

Galliers:1995:ATC

[Gal95] Kevin Galliers. AI techniques to counter financial fraud. Computers & Security, 14(5):426, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/0167404895971462.

Ganley:1994:DST

[Gan94] Michael J. Ganley. Digital signatures and their uses. Computers & Security, 13(5):385–391, ???? 1994. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/0167404894900310.

Gollmann:1993:ASD

[GBD93] Dieter Gollmann, Thomas Beth, and Frank Damm. Authentication services in distributed systems. Computers & Security, 12(8):753–764, December 1993. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404893900413.

Gustafson:1994:CPM

[GDNC94] H. Gustafson, E. Dawson, L. Nielsen, and W. Caelli. A computer package for measuring the strength of encryption algorithms. Computers & Security, 13(8):687–697, ???? 1994. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/0167404894900515.

Gentile:1994:SEI

[GGG^{+}94] Francesco Gentile, Luigi Giuri, Franco Guida, Emilio Montolivo, and Michele Volpe. Security evaluation in information technology standards. Computers & Security, 13(8):647–650, ???? 1994. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/0167404894900442.

Gilbert:1992:BSS

[Gil92] Françoise Gilbert. Breach of system security and theft of data: Legal aspects and preventive measures. Computers & Security, 11(6):508–517, October 1992. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/016740489290183R.

Gritzalis:1992:ZKP

[GKG92] Dimitris Gritzalis, Sokratis Katsikas, and Stefanos Gritzalis. A zero knowledge probabilistic login protocol. Computers & Security, 11(8):733–745, December 1992. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/016740489290129F.

Gritzalis:1992:DAR

[GKKT92] D. Gritzalis, S. Katsikas, J. Keklikoglou, and A. Tomaras. Determining access rights for medical information systems. Computers & Security, 11(2):149–161, April 1992. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/016740489290042P.

Gennaro:1997:TPC

[GKM^{+}97] Rosario Gennaro, Paul Karger, Stephen Matyas, Mohammad Peyravian Allen Roginsky, David Safford, Michael Willett, and Nev Zunic. Two-phase cryptographic key recovery system. Computers & Security, 16(6):481–506, ???? 1997. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404897846421.

Gliss:1990:EGH

[Gli90] Hans Gliss. East Germany — haven for software piracy. Computers & Security, 9(5):391–393, August 1990. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/0167404890900685.

Gordon:1995:I

[Gor95a] Sarah Gordon. Internet 101. Computers & Security, 14(7):599– 604, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/016740489681666X.

Gordon:1995:PVT

[Gor95b] Sarah Gordon. Publication of vulnerabilities and tools. Computers & Security, 14(5):417, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/0167404896900383.

Gordon:1995:SER

[Gor95c] Sarah Gordon. Social engineering — recognizing intruders. Computers & Security, 14(5):424, ???? 1995. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/016740489597144Y.

Gordon:1995:TEC

[Gor95d] Sarah Gordon. Technologically enabled crime: Shifting paradigms for the year 2000. Computers & Security, 14(5):391– 402, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/016740489597090W.

Gordon:1997:IFW

[Gor97] Sarah Gordon. Information flow within the globally connected environment. Computers & Security, 16(6):525, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404897846974.

Gordon:1998:BC

[Gor98] Sarah Gordon. Breaking the chain. Computers & Security, 17(7):586–588, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404899800583.

Gordon:1999:TT

[Gor99] Sarah Gordon. Telephone tales. Computers & Security, 18 (2):120–123, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404899900494.

Gritzalis:1997:BSP

[Gri97] Dimitris Gritzalis. A baseline security policy for distributed healthcare information systems. Computers & Security, 16 (8):709–719, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404897000096.

Goyal:1991:ACD

[GS91] M. L. Goyal and G. V. Singh. Access control in distributed heterogeneous database management systems. Computers & Security, 10(7):661–669, November 1991. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/016740489190124V.

Gritzalis:1991:DSM

[GTKK91] D. Gritzalis, A. Tomaras, S. Katsikas, and J. Keklikoglou. Data security in medical information systems: the Greek case. Computers & Security, 10(2):141–159, April 1991. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/0167404891900082.

Graham:1992:PSN

[GW92] I. G. Graham and S. H. Wieten. The PC as a secure network workstation. Computers & Security, 11(3):237–244, May 1992. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/0167404892902023.

Ha:1992:CFN

[Ha92] Tu Thanh Ha. Computer files need more safeguards, hospitals told. Computers & Security, 11(5):487, September 1992. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/016740489290022J.

Hancock:1995:DRA

[Han95a] Bill Hancock. Dial-up and remote access facilities for LANs. Computers & Security, 14(5):424, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/016740489597142W.

Hancock:1995:FIS

[Han95b] Bill Hancock. Future of Internet security. Computers & Security, 14(5):417, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/016740489597108M.

Hanson:1995:IAS

[Han95c] Lesley Hanson. The impact of ATM on security in the data network. Computers & Security, 14(5):422, ???? 1995. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/0167404896900425.

Hancock:1998:Ea

[Han98a] Bill Hancock. From the editor. Computers & Security, 17(1):2– 4, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404897802416.

Hancock:1998:Eb

[Han98b] Bill Hancock. From the Editor. Computers & Security, 17(2): 96–97, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404897819761.

Hancock:1998:Ec

[Han98c] Bill Hancock. From the editor. Computers & Security, 17 (3):190–191, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404898803088.

Hancock:1998:Ed

[Han98d] Bill Hancock. From the editor. Computers & Security, 17 (4):278–279, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404898800060.

Hancock:1998:Ee

[Han98e] Bill Hancock. From the editor. Computers & Security, 17 (5):368–369, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404898800011.

Hancock:1998:Ef

[Han98f] Bill Hancock. From the editor. Computers & Security, 17 (6):460–461, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404898800679.

Hancock:1998:Eg

[Han98g] Bill Hancock. From the editor. Computers & Security, 17 (7):562–563, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404899800534.

Hancock:1998:Eh

[Han98h] Bill Hancock. From the editor. Computers & Security, 17 (8):652–653, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404898800953.

Hancock:1998:SGP

[Han98i] Bill Hancock. Security goes prime-time. Computers & Security, 17(1):5–13, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404897802428.

Hancock:1998:STG

[Han98j] Bill Hancock. Security tradeshows — good or bad? Computers & Security, 17(6):462–474, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404898800680.

Hancock:1998:SVa

[Han98k] Bill Hancock. Security views. Computers & Security, 17(3):192– 200, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S016740489880309X.

Hancock:1998:SVb

[Han98l] Bill Hancock. Security views. Computers & Security, 17(4):280– 292, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404898800072.

Hancock:1998:SVc

[Han98m] Bill Hancock. Security views. Computers & Security, 17(8):654– 666, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404898800965.

Hancock:1998:WDY

[Han98n] Bill Hancock. Who do you call for, CISSP. Computers & Security, 17(2):99–109, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404897819773.

Hancock:1998:WRM

[Han98o] Bill Hancock. Windows-98, RIP and my adrenaline rush. Computers & Security, 17(7):564–574, ???? 1998. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404899800546.

Hancock:1999:CFS

[Han99a] Bill Hancock. 1999 CSI/FBI survey: Cyberattacks on the rise. Computers & Security, 18(3):188–189, ???? 1999. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404899909161.

Hancock:1999:BFM

[Han99b] Bill Hancock. 45% of big firms monitor workers. Computers & Security, 18(4):284–285, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404899907034.

Hancock:1999:IEM

[Han99c] Bill Hancock. And if that is not enough — Mainland China and Taiwan engage in cyberwar. Computers & Security, 18(6):465– 466, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404899908350.

Hancock:1999:ASC

[Han99d] Bill Hancock. Anonymizer service could compromise Kosovar residents need for privacy. Computers & Security, 18(3):191– 192, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404899909203.

Hancock:1999:ACP

[Han99e] Bill Hancock. Appeals-court panel says expor ban on encryption software is unlawful. Computers & Security, 18(4):278– 279, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404899800716.

Hancock:1999:ASP

[Han99f] Bill Hancock. Application service provider (ASP) hosting centers are a security problem. Computers & Security, 18(5):381– 382, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404899905928.

Hancock:1999:BOP

[Han99g] Bill Hancock. Back orifice prevention tool to be released. Computers & Security, 18(3):195, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404899909276.

Hancock:1999:BAY

[Han99h] Bill Hancock. Being all that you can virtually be — cyberwar education gets formal. Computers & Security, 18(4):282– 283, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404899907009.

Hancock:1999:BBB

[Han99i] Bill Hancock. Better Business Bureau gets involved in online privacy. Computers & Security, 18(2):99–100, ???? 1999.

CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404899900378.

Hancock:1999:BCG

[Han99j] Bill Hancock. British cyberpolice get new powers for good vs. evil. Computers & Security, 18(6):468–469, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404899908398.

Hancock:1999:CGI

[Han99k] Bill Hancock. China glides into Internet mainstream. Computers & Security, 18(3):186–188, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404899909148.

Hancock:1999:CIW

[Han99l] Bill Hancock. China’s internal war with the Falun Gong sect spills onto Internet. Computers & Security, 18(6):464–465, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404899908349.

Hancock:1999:CAS

[Han99m] Bill Hancock. Cisco announces security associate partnership programme. Computers & Security, 18(6):460–461, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404899908301.

Hancock:1999:CBC

[Han99n] Bill Hancock. Clinton backs cyber-plan against Milosevic — report. Computers & Security, 18(5):372, ???? 1999. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404899905849.

Hancock:1999:CPP

[Han99o] Bill Hancock. Clinton proposes privacy controls. Computers & Security, 18(4):293–294, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404899907137.

Hancock:1999:CAC

[Han99p] Bill Hancock. Conference to address combating Internet child pornography. Computers & Security, 18(6):469, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404899908404.

Hancock:1999:CPO

[Han99q] Bill Hancock. Congressional probe of online banking finds flaws. Computers & Security, 18(6):466–467, ???? 1999. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404899908374.

Hancock:1999:CDN

[Han99r] Bill Hancock. Cybersquatting on domain names gets bitten by new laws in US. Computers & Security, 18(6):469–470, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404899908416.

Hancock:1999:EAS

[Han99s] Bill Hancock. eBay auction site gets compromised. Computers & Security, 18(3):186, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404899909136.

Hancock:1999:ECC

[Han99t] Bill Hancock. Elliptical curve cryptography and standards for efficient cryptography group. Computers & Security, 18(1):4– 15, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404899800029.

Hancock:1999:EBH

[Han99u] Bill Hancock. EU bill holds ISPs liable for web content. Computers & Security, 18(4):279–280, ???? 1999. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404899906971.

Hancock:1999:EPP

[Han99v] Bill Hancock. European Parliament promotes a safer Internet. Computers & Security, 18(2):98–99, ???? 1999. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404899900330.

Hancock:1999:EUC

[Han99w] Bill Hancock. Experts: Use caution with white-hat hackers. Computers & Security, 18(4):285, ???? 1999. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404899907046.

Hancock:1999:FSA

[Han99x] Bill Hancock. Finalists selected in AES bake-off. Computers & Security, 18(6):458, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404899908283.

Hancock:1999:FBD

[Han99y] Bill Hancock. Fingerprint biometrics devices just not catching on in the real world. Computers & Security, 18(5):382– 383, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S016740489990593X.

Hancock:1999:FIM

[Han99z] Bill Hancock. First Internet e-mail corporate usage report; concludes e-mail abuse at epidemic levels. Computers & Security, 18(3):188, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S016740489990915X.

Hancock:1999:FLC

[Han99-27] Bill Hancock. France lightens up cryptographic use. Computers & Security, 18(2):100, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404899900391.

Hancock:1999:FMS

[Han99-28] Bill Hancock. ‘free-mail’ systems great targets for hackers. Computers & Security, 18(3):189, ???? 1999. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404899909173.

Hancock:1999:Ea

[Han99-29] Bill Hancock. From the editor. Computers & Security, 18(1):2– 3, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404899800017.

Hancock:1999:Eb

[Han99-30] Bill Hancock. From the editor. Computers & Security, 18(2): 88–89, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S016740489990007X.

Hancock:1999:Ec

[Han99-31] Bill Hancock. From the editor. Computers & Security, 18 (3):182–183, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404899800625.

Hancock:1999:Ed

[Han99-32] Bill Hancock. From the editor. Computers & Security, 18 (4):276–277, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404899800704.

Hancock:1999:Ee

[Han99-33] Bill Hancock. From the editor. Computers & Security, 18 (5):370–371, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404899800820.

Hancock:1999:Ef

[Han99-34] Bill Hancock. From the editor. Computers & Security, 18 (6):456–457, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404899801126.

Hancock:1999:Eg

[Han99-35] Bill Hancock. From the editor. Computers & Security, 18 (7):550–552, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404899820008.

Hancock:1999:FAY

[Han99-36] Bill Hancock. FTC accuses young investor site of privacy violations, but what’s a webmaster to do? Call Mom and Dad on each page access? Computers & Security, 18(4):280–282, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404899906983.

Hancock:1999:HAU

[Han99-37] Bill Hancock. Hackers attack US government web sites in protest of Chinese embassy bombing. Computers & Security, 18 (4):279, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S016740489990696X.

Hancock:1999:HPI

[Han99-38] Bill Hancock. Hardening of an OS by putting it out for hack attacks. Computers & Security, 18(6):463–464, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404899908337.

Hancock:1999:IMW

[Han99-39] Bill Hancock. Information managers want more and better IT security. Computers & Security, 18(5):383–385, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404899905941.

Hancock:1999:ISE

[Han99-40] Bill Hancock. Information society in the EU: Digital signatures. Computers & Security, 18(4):292–293, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404899907125.

Hancock:1999:IWN

[Han99-41] Bill Hancock. Information warfare — the next front. Computers & Security, 18(2):101–102, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404899900433.

Hancock:1999:IAG

[Han99-42] Bill Hancock. The InfraGard alliance — getting the Feds and IT managers on the same page. Computers & Security, 18(5):385– 386, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404899905953.

Hancock:1999:IIT

[Han99-43] Bill Hancock. Internal information theft in US on the rise. Computers & Security, 18(3):184, ???? 1999. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404899909124.

Hancock:1999:IUF

[Han99-44] Bill Hancock. Internet users furious over SingTel’s computer security probe. Computers & Security, 18(4):288–289, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404899907095.

Hancock:1999:ITG

[Han99-45] Bill Hancock. Internet2 testbed going online at a network near you. Computers & Security, 18(2):93–94, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404899900159.

Hancock:1999:ICK

[Han99-46] Bill Hancock. Interpol calls for key recovery. Computers & Security, 18(3):195, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404899909264.

Hancock:1999:IVO

[Han99-47] Bill Hancock. ISPs and VPNs are at odds with each other. Computers & Security, 18(5):376–380, ???? 1999. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404899905898.

Hancock:1999:VCH

[Han99-48] Bill Hancock. It’s vogue to capture hackers, not just watch. Computers & Security, 18(6):458–459, ???? 1999. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404899801138.

Hancock:1999:JCC

[Han99-49] Bill Hancock. Join Chaos Computer Club for that summer retreat you have been longing to attend . . .. Computers & Security, 18(6):466, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404899908362.

Hancock:1999:KMF

[Han99-50] Bill Hancock. Kevin Mitnick finally gets a plea — and a sentence. Computers & Security, 18(3):196–197, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S016740489990929X.

Hancock:1999:MOS

[Han99-51] Bill Hancock. Making the operating system trusted — a trend. Computers & Security, 18(6):462–463, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404899908325.

Hancock:1999:MRA

[Han99-52] Bill Hancock. Medical records accidentally posted to the web. Computers & Security, 18(2):98, ???? 1999. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404899900299.

Hancock:1999:MMA

[Han99-53] Bill Hancock. Melissa the macrovirus has an author . . .. Computers & Security, 18(3):184–186, ???? 1999. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404899800637.

Hancock:1999:MOP

[Han99-54] Bill Hancock. Microsoft to offer privacy tools. Computers & Security, 18(4):285–286, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404899907058.

Hancock:1999:MWG

[Han99-55] Bill Hancock. Microsoft’s Windows OS getting a security makeover. Computers & Security, 18(5):386, ???? 1999. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404899905965.

Hancock:1999:MCM

[Han99-56] Bill Hancock. Mobile code malicious applets cause problems for IT managers. Computers & Security, 18(5):373–374, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404899905862.

Hancock:1999:MFV

[Han99-57] Bill Hancock. More fraud: Visa International states e-commerce is major fraud source. Computers & Security, 18(3):194, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404899909252.

Hancock:1999:MOV

[Han99-58] Bill Hancock. MS Office 2000 virus infestation — already. Computers & Security, 18(2):94–95, ???? 1999. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404899900196.

Hancock:1999:MS

[Han99-59] Bill Hancock. Music security? Computers & Security, 18(5): 376, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404899905886.

Hancock:1999:NIT

[Han99-60] Bill Hancock. New IETF task force on intrusion detection. Computers & Security, 18(2):96, ???? 1999. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404899900251.

Hancock:1999:NIL

[Han99-61] Bill Hancock. New Internet legislation attempts in the US. Computers & Security, 18(4):282, ???? 1999. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404899906995.

Hancock:1999:NSP

[Han99-62] Bill Hancock. The new school paradigm: Metal detectors, roving patrols, firewalls, cyberwalls and encryption . . .. Computers & Security, 18(5):380, ???? 1999. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404899905904.

Hancock:1999:NMC

[Han99-63] Bill Hancock. Novell maybe causing privacy concerns with digitalme. Computers & Security, 18(3):190–191, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404899909197.

Hancock:1999:NPS

[Han99-64] Bill Hancock. NT passwords for sale. Computers & Security, 18 (2):92, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404899900135.

Hancock:1999:OTL

[Han99-65] Bill Hancock. One in three local authorities in the UK vulnerable to privacy cyberattacks. Computers & Security, 18(3):197– 198, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404899909306.

Hancock:1999:OFI

[Han99-66] Bill Hancock. Online fraud increasing rapidly. Computers & Security, 18(3):194, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404899909240.

Hancock:1999:OSE

[Han99-67] Bill Hancock. Outsourced security — EDS tries to go large with new unit. Computers & Security, 18(5):386–387, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404899905977.

Hancock:1999:OSC

[Han99-68] Bill Hancock. Outsourced security — the corporate information security burglar alarm hits IT. Computers & Security, 18(6):461–462, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404899908313.

Hancock:1999:PSB

[Han99-69] Bill Hancock. Parents seek blocks for web sites. Computers & Security, 18(4):291–292, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404899907113.

Hancock:1999:PII

[Han99-70] Bill Hancock. Pentium III ID cracked by German group. Computers & Security, 18(3):192–193, ???? 1999. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404899909227.

Hancock:1999:PCA

[Han99-71] Bill Hancock. Pessimism in cyberspace is alive and well. Computers & Security, 18(2):95–96, ???? 1999. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404899900238.

Hancock:1999:PPR

[Han99-72] Bill Hancock. Privacy problems on the rise. Computers & Security, 18(2):90–92, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404899900111.

Hancock:1999:RSG

[Han99-73] Bill Hancock. Risks to security growing as fast as numbers of systems. Computers & Security, 18(4):287–288, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404899907071.

Hancock:1999:SBC

[Han99-74] Bill Hancock. SAFE bill clears next hurdle in US legislature. Computers & Security, 18(3):193, ???? 1999. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404899909239.

Hancock:1999:SRA

[Han99-75] Bill Hancock. Security research alliance begins. Computers & Security, 18(2):95, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404899900214.

Hancock:1999:SV

[Han99-76] Bill Hancock. Security views. Computers & Security, 18(7):553– 564, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S016740489982001X.

Hancock:1999:SSE

[Han99-77] Bill Hancock. Statistics show European security spending will quadruple by 2003. Computers & Security, 18(6):459–460, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404899908295.

Hancock:1999:SVH

[Han99-78] Bill Hancock. Students in Virginia hacks NASA. Computers & Security, 18(2):102–103, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404899900457.

Hancock:1999:TCB

[Han99-79] Bill Hancock. Teen crackers being used by professional thieves. Computers & Security, 18(2):100–101, ???? 1999. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S016740489990041X.

Hancock:1999:TIW

[Han99-80] Bill Hancock. Theft of information in the workplace creates growing concern for corporate America. Computers & Security, 18(4):283, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404899907010.

Hancock:1999:UEC

[Han99-81] Bill Hancock. UK electronic commerce bill may require escrow. Computers & Security, 18(5):380–381, ???? 1999. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404899905916.

Hancock:1999:URB

[Han99-82] Bill Hancock. UK report blows government Internet security sky high. Computers & Security, 18(4):286–287, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S016740489990706X.

Hancock:1999:UCA

[Han99-83] Bill Hancock. A unique Canadian approach — invite the hacker to nail your networks. Computers & Security, 18(2):103–104, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404899900470.

Hancock:1999:UAB

[Han99-84] Bill Hancock. US authorities bust 45m Internet credit card billing scam. Computers & Security, 18(5):389–390, ????

1999. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404899906004.

Hancock:1999:UDD

[Han99-85] Bill Hancock. U.S. Department of Defense worries about event exposure via aggregation techniques. Computers & Security, 18 (2):97, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404899900275.

Hancock:1999:UFA

[Han99-86] Bill Hancock. US feds ask for public sector help to prosecute cybercrime. Computers & Security, 18(3):189–190, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404899909185.

Hancock:1999:UGS

[Han99-87] Bill Hancock. US Government states protections needed for companies to come forward with cyberattack reports. Computers & Security, 18(3):192, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404899909215.

Hancock:1999:UMM

[Han99-88] Bill Hancock. US moving for more privacy controls over data. Computers & Security, 18(5):372–373, ???? 1999. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404899800832.

Hancock:1999:UNL

[Han99-89] Bill Hancock. US nuclear labs find counterespionage difficult. Computers & Security, 18(4):288, ???? 1999. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404899907083.

Hancock:1999:UPG

[Han99-90] Bill Hancock. US plan for government-owned infrastructure intrusion detection system draws fire. Computers & Security, 18(6):467–468, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404899908386.

Hancock:1999:URN

[Han99-91] Bill Hancock. US releases new data privacy guidelines. Computers & Security, 18(4):284, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404899907022.

Hancock:1999:VBM

[Han99-92] Bill Hancock. VPNs build momentum. Computers & Security, 18(4):289–291, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404899907101.

Hancock:1999:WFP

[Han99-93] Bill Hancock. Want a free PC? give up your privacy, first. Computers & Security, 18(2):90, ???? 1999. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404899900093.

Hancock:1999:WYD

[Han99-94] Bill Hancock. What you don’t know about your online rights might hurt you. Computers & Security, 18(5):375–376, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404899905874.

Hancock:1999:WCT

[Han99-95] Bill Hancock. When Cisco talks about VPNs . . .. Computers & Security, 18(2):94, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404899900172.

Hancock:1999:WRR

[Han99-96] Bill Hancock. Windows-2000 requires re-building of code. Computers & Security, 18(2):99, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404899900354.

Hancock:1999:WNS

[Han99-97] Bill Hancock. Windows-NT service Pack 4 security woes. Computers & Security, 18(2):98, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404899900317.

Hancock:1999:WEE

[Han99-98] Bill Hancock. Wiretaps in Europe — EU debates a resolution. Computers & Security, 18(5):388–389, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404899905990.

Hancock:1999:WWW

[Han99-99] Bill Hancock. World wide web abusive use widespread. Computers & Security, 18(3):195–196, ???? 1999. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404899909288.

Hancock:1999:YMR

[Han99-100] Bill Hancock. You may be right, there may be law on your side, and you may not necessarily win when you sue . . .. Computers & Security, 18(5):387–388, ???? 1999. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404899905989.

Hardy:1992:CSC

[Har92] Gary Hardy. Case study: Closing ‘windows’ against intruders. Computers & Security, 11(8):711–714, December 1992. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/0167404892900926.

Hardy:1993:CAI

[Har93] Gary Hardy. Commercial accreditation of information security. Computers & Security, 12(8):716–729, December 1993. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/0167404893900343.

Hardy:1995:SNC

[Har95] Gary Hardy. Standards — the need for a common framework. Computers & Security, 14(5):426–427, ???? 1995. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/016740489597151Y.

Hardy:1997:PT

[Har97] Gary Hardy. Penetration testing. Computers & Security, 16(6): 524, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404897846913.

Hawkins:1999:SWM

[Haw99] Nick Hawkins. Security within multimedia networked conferencing. Computers & Security, 18(5):419–422, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S016740489980087X.

Hinke:1997:PDI

[HDW97] Thomas H. Hinke, Harry S. Delugach, and Randall P. Wolf. Protecting databases from inference attacks. Computers & Security, 16(8):687–708, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404897876079.

Heinlein:1993:WM

[Hei93a] Edwin B. Heinlein. I worked with Murphy. Computers & Security, 12(7):627–628, November 1993. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/016740489390076H.

Heinlein:1993:RM

[Hei93b] Edwin B. Heinlein. Recovery management. Computers & Security, 12(4):334–337, June 1993. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404893900192.

Heinlein:1993:WTC

[Hei93c] Edwin B. Heinlein. World trade center — a retrospective. Computers & Security, 12(5):435–436, August 1993. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/0167404893900619.

Heinlein:1994:NCS

[Hei94a] Edwin B. Heinlein. 17th national computer security conference — an opinion. Computers & Security, 13(6):481–483, ???? 1994. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/016740489190134Y.

Heinlein:1994:CSP

[Hei94b] Edwin B. Heinlein. Communications, security, privacy and the law. Computers & Security, 13(2):119–121, April 1994. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/0167404894900582.

Heinlein:1995:PIS

[Hei95a] Edwin B. Heinlein. Principles of information systems security. Computers & Security, 14(3):197–198, ???? 1995. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/016740489597052C.

Heinlein:1995:RSC

[Hei95b] Edwin B. Heinlein. The RSA security conference 1995. Computers & Security, 14(1):24–26, ???? 1995. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/016740489596992C.

Heinlein:1996:CSC

[Hei96a] Edwin B. Heinlein. Computer security in China. Computers & Security, 15(5):369–375, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/0167404896825594.

Heinlein:1996:MRS

[Hei96b] Edwin B. Heinlein. Medical records security. Computers & Security, 15(2):100–102, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/0167404896893229.

Heinlein:1996:YRS

[Hei96c] Edwin B. Heinlein. Year 2000 — a real IS security issue. Computers & Security, 15(6):499–500, ???? 1996. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404897831239.

Harris:1999:FC

[HH99] Brendon Harris and Ray Hunt. Fairwall certification. Computers & Security, 18(2):165–177, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404899800522.

Hickson:1995:BCP

[Hic95] Nigel Hickson. Beyond the code of practice: New challenges. Computers & Security, 14(5):426, ???? 1995. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/0167404896900516.

Hickson:1997:EPUa

[Hic97a] Nigel Hickson. Encryption policy — a UK perspective. Computers & Security, 16(6):516, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404897846640.

Hickson:1997:EPUb

[Hic97b] Nigel Hickson. Encryption policy — a UK perspective. Computers & Security, 16(7):583–589, ???? 1997. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404897807948.

Highland:1990:RBBa

[Hig90a] Harold Joseph Highland. Random bits and bytes. Computers & Security, 9(1):4–11, February 1990. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/016740489090149N.

Highland:1990:RBBb

[Hig90b] Harold Joseph Highland. Random bits and bytes. Computers & Security, 9(2):102–109, April 1990. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/0167404890900814.

Highland:1990:RBBc

[Hig90c] Harold Joseph Highland. Random bits and bytes. Computers & Security, 9(3):196–208, May 1990. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/016740489090164O.

Highland:1990:RBBd

[Hig90d] Harold Joseph Highland. Random bits and bytes. Computers & Security, 9(4):282–294, June 1990. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/016740489090095B.

Highland:1990:RBBe

[Hig90e] Harold Joseph Highland. Random bits and bytes. Computers & Security, 9(5):366–378, August 1990. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/0167404890900663.

Highland:1990:RBBf

[Hig90f] Harold Joseph Highland. Random bits and bytes. Computers & Security, 9(6):470–482, October 1990. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/016740489090122A.

Highland:1990:RBBg

[Hig90g] Harold Joseph Highland. Random bits and bytes. Computers & Security, 9(7):570–586, November 1990. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/016740489090054W.

Highland:1990:RBBh

[Hig90h] Harold Joseph Highland. Random bits and bytes. Computers & Security, 9(8):660–675, December 1990. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/0167404890901086.

Highland:1991:CUA

[Hig91a] Harold Joseph Highland. Computers under attack. Computers & Security, 10(3):190–191, May 1991. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/0167404891900318.

Highland:1991:EEP

[Hig91b] Harold Joseph Highland. Evaluating an encryption package. Computers & Security, 10(3):189–190, May 1991. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/016740489190030H.

Highland:1991:MU

[Hig91c] Harold Joseph Highland. Microcomputer utilities. Computers & Security, 10(3):191–195, May 1991. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/0167404891900329.

Highland:1991:NMV

[Hig91d] Harold Joseph Highland. A non-mathematical view of encryption. Computers & Security, 10(3):184–188, May 1991. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/016740489190029D.

Highland:1991:NVI

[Hig91e] Harold Joseph Highland. NTIS volumes of interest. Computers & Security, 10(3):195–198, May 1991. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/016740489190033A.

Highland:1991:RBBa

[Hig91f] Harold Joseph Highland. Random bits and bytes. Computers & Security, 10(1):4–16, February 1991. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/016740489190049J.

Highland:1991:RBBb

[Hig91g] Harold Joseph Highland. Random bits and bytes. Computers & Security, 10(2):92–100, April 1991. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/016740489190002U.

Highland:1991:RBBc

[Hig91h] Harold Joseph Highland. Random bits and bytes. Computers & Security, 10(5):384–397, August 1991. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/016740489190064K.

Highland:1991:RBBd

[Hig91i] Harold Joseph Highland. Random bits and bytes. Computers & Security, 10(6):486–495, October 1991. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/016740489190074N.

Highland:1991:RBBe

[Hig91j] Harold Joseph Highland. Random bits and bytes. Computers & Security, 10(7):586–600, November 1991. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/016740489190116U.

Highland:1991:RBBf

[Hig91k] Harold Joseph Highland. Random bits and bytes. Computers & Security, 10(8):690–699, December 1991. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/016740489190088U.

Highland:1992:BRP

[Hig92a] Harold Joseph Highland. Book review: Practical Unix security: Simon Garfinkel and Gene Spafford O’Reilly and Associates, Inc. (Sebastopol, CA 95472, U.S.A.) U.S. Price: $29.95 ISBN:0-939195-72-2. Computers & Security, 11(2):134–139, April 1992. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/016740489290038S.

Highland:1992:ITNf

[Hig92b] Harold Joseph Highland. IFIP TC 11 news. Computers & Security, 11(6):531–534, October 1992. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/016740489290187V.

Highland:1992:ITNh

[Hig92c] Harold Joseph Highland. IFIP TC 11 news. Computers & Security, 11(8):715–719, December 1992. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/0167404892900937.

Highland:1992:ITNa

[Hig92d] Harold Joseph Highland. IFIP TC11 news. Computers & Security, 11(1):35–38, March 1992. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/016740489290218G.

Highland:1992:ITNb

[Hig92e] Harold Joseph Highland. IFIP TC11 news. Computers & Security, 11(2):145–146, April 1992. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/016740489290040X.

Highland:1992:ITNc

[Hig92f] Harold Joseph Highland. IFIP TC11 news. Computers & Security, 11(3):247–250, May 1992. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404892902034.

Highland:1992:ITNd

[Hig92g] Harold Joseph Highland. IFIP TC11 news. Computers & Security, 11(4):341–343, July 1992. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/016740489290173O.

Highland:1992:ITNe

[Hig92h] Harold Joseph Highland. IFIP TC11 news. Computers & Security, 11(5):441–442, September 1992. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/016740489290009G.

Highland:1992:ITNg

[Hig92i] Harold Joseph Highland. IFIP TC11 news. Computers & Security, 11(7):625–628, November 1992. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/016740489290140M.

Highland:1992:RBBa

[Hig92j] Harold Joseph Highland. Random bits and bytes. Computers & Security, 11(1):4–15, March 1992. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/016740489290213B.

Highland:1992:RBBb

[Hig92k] Harold Joseph Highland. Random bits and bytes. Computers & Security, 11(2):110–120, April 1992. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/016740489290035P.

Highland:1992:RBBc

[Hig92l] Harold Joseph Highland. Random bits and bytes. Computers & Security, 11(3):200–209, May 1992. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/016740489290197Y.

Highland:1992:RBBd

[Hig92m] Harold Joseph Highland. Random bits and bytes. Computers & Security, 11(4):294–306, July 1992. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/016740489290169R.

Highland:1992:RBBe

[Hig92n] Harold Joseph Highland. Random bits and bytes. Computers & Security, 11(5):390–406, September 1992. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/0167404892900029.

Highland:1992:RBBf

[Hig92o] Harold Joseph Highland. Random bits and bytes. Computers & Security, 11(6):498–507, October 1992. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/016740489290182Q.

Highland:1992:RBBg

[Hig92p] Harold Joseph Highland. Random bits and bytes. Computers & Security, 11(7):592–601, November 1992. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/016740489290135E.

Highland:1992:RBBh

[Hig92q] Harold Joseph Highland. Random bits and bytes. Computers & Security, 11(8):682–694, December 1992. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/0167404892900867.

Highland:1992:SIN

[Hig92r] Harold Joseph Highland. The security impact of networks, telecommunications, and office automation. Computers & Security, 11(3):227–232, May 1992. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/016740489290200B.

Highland:1993:NCS

[Hig93a] Harold Joseph Highland. 16th national computer security conference. Computers & Security, 12(8):730–734, December 1993. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/0167404893900354.

Highland:1993:CR

[Hig93b] Harold Joseph Highland. Conference report. Computers & Security, 12(3):249–253, May 1993. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/016740489390111H.

Highland:1993:ITNa

[Hig93c] Harold Joseph Highland. IFIP TC 11 news. Computers & Security, 12(1):33–35, February 1993. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/016740489390008S.

Highland:1993:ITNb

[Hig93d] Harold Joseph Highland. IFIP TC 11 news. Computers & Security, 12(2):139–142, March 1993. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/016740489390095M.

Highland:1993:ITNc

[Hig93e] Harold Joseph Highland. IFIP TC 11 news. Computers & Security, 12(3):259–261, May 1993. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/016740489390113J.

Highland:1993:ITNd

[Hig93f] Harold Joseph Highland. IFIP TC 11 news. Computers & Security, 12(4):362–364, June 1993. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/016740489390022W.

Highland:1993:ITNe

[Hig93g] Harold Joseph Highland. IFIP TC 11 news. Computers & Security, 12(5):456–458, August 1993. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/016740489390064C.

Highland:1993:ITNf

[Hig93h] Harold Joseph Highland. IFIP TC 11 news. Computers & Security, 12(8):735–739, December 1993. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/0167404893900365.

Highland:1993:ISC

[Hig93i] Harold Joseph Highland. IFIP/Sec ’93 conference report. Computers & Security, 12(6):550–555, October 1993. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/016740489390050F.

Highland:1993:RBBa

[Hig93j] Harold Joseph Highland. Random bits and bytes. Computers & Security, 12(1):4–14, February 1993. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/016740489390003N.

Highland:1993:RBBb

[Hig93k] Harold Joseph Highland. Random bits and bytes. Computers & Security, 12(2):106–116, March 1993. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/016740489390089N.

Highland:1993:RBBc

[Hig93l] Harold Joseph Highland. Random bits and bytes. Computers & Security, 12(3):216–224, May 1993. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/016740489390107G.

Highland:1993:RBBd

[Hig93m] Harold Joseph Highland. Random bits and bytes. Computers & Security, 12(5):424–433, August 1993. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/016740489390060I.

Highland:1993:RBBe

[Hig93n] Harold Joseph Highland. Random bits and bytes. Computers & Security, 12(6):518–526, October 1993. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/0167404893900468.

Highland:1993:RBBf

[Hig93o] Harold Joseph Highland. Random bits and bytes. Computers & Security, 12(7):612–622, November 1993. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/016740489390074F.

Highland:1993:RBBg

[Hig93p] Harold Joseph Highland. Random bits and bytes. Computers & Security, 12(8):698–709, December 1993. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/016740489390032Z.

Highland:1993:VIS

[Hig93q] Harold Joseph Highland. A view of information security tomorrow. Computers & Security, 12(7):634–639, November 1993. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/016740489390078J.

Highland:1994:ITN

[Hig94a] Harold Joseph Highland. IFIP TC 11 news. Computers & Security, 13(2):123–125, April 1994. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/0167404894900590.

Highland:1994:RBBa

[Hig94b] Harold Joseph Highland. Random bits and bytes. Computers & Security, 13(1):4–14, February 1994. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/016740489490085X.

Highland:1994:RBBb

[Hig94c] Harold Joseph Highland. Random bits and bytes. Computers & Security, 13(2):98–106, April 1994. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/0167404894900558.

Highland:1994:RBBc

[Hig94d] Harold Joseph Highland. Random bits and bytes. Computers & Security, 13(3):192–205, May 1994. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/0167404894900698.

Highland:1994:RBBd

[Hig94e] Harold Joseph Highland. Random bits and bytes. Computers & Security, 13(4):278–286, ???? 1994. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/0167404894900167.

Highland:1994:RBBe

[Hig94f] Harold Joseph Highland. Random bits and bytes. Computers & Security, 13(5):372–378, ???? 1994. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/0167404894900299.

Highland:1994:RBBf

[Hig94g] Harold Joseph Highland. Random bits and bytes. Computers & Security, 13(6):458–465, ???? 1994. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/016740489190131V.

Highland:1994:RBBg

[Hig94h] Harold Joseph Highland. Random bits and bytes. Computers & Security, 13(7):540–546, ???? 1994. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/0167404894900027.

Highland:1994:RBBh

[Hig94i] Harold Joseph Highland. Random bits and bytes. Computers & Security, 13(8):622–627, ???? 1994. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/0167404894900418.

Highland:1995:AII

[Hig95a] Harold Joseph Highland. Artificial intelligence and intrusion detection: Current and future directions: Jeremy Frank, University of California, Davis, CA. Computers & Security, 14(1): 31, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/0167404895970108.

Highland:1995:BCM

[Hig95b] Harold Joseph Highland. Belief in correctness: Marshall Abrams, The MITRE Corporation, McLean, VA. Computers & Security, 14(1):27–28, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/016740489596995F.

Highland:1995:BRF

[Hig95c] Harold Joseph Highland. Book review: A Financial Management Approach for Selecting Optimal, Cost-effective Safeguards Upgrades for Computer- and Information-Security Risk Management: Suzanne Smith, Baranca Inc, Los Almos, NM, and Stephen Gale and William Malampy, University of Pennsylvania, Philadelphia, PA. Computers & Security, 14(1):28– 29, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/016740489597000Z.

Highland:1995:BRP

[Hig95d] Harold Joseph Highland. Book review: A Pattern Matching Model for Misuse Intrusion Detection: Sandeep Kumar and Eugene Spafford, Purdue University, West Lafayette, IN. Computers & Security, 14(1):28, ???? 1995. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/016740489596997H.

Highland:1995:BRSb

[Hig95e] Harold Joseph Highland. Book review: A software architecture to support misuse intrusion detection: Sandeep Kumar and Eugene H. Spafford, Purdue University, West Lafayette, IN, USA. Computers & Security, 14(7):607, ???? 1995. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/0167404896816713.

Highland:1995:BRAb

[Hig95f] Harold Joseph Highland. Book review: AIN’T misbehaving — a taxonomy of anti-intrusion techniques: Lawrence R. Halme, Arca Systems Inc., San Jose, CA, USA. Computers & Security, 14(7):606, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404896816695.

Highland:1995:BRO

[Hig95g] Harold Joseph Highland. Book review: An open trusted enterprise network architecture: Gray Grossman and Jeremy Epstein, Cordant, Inc., Reston, VA, USA; and Roger Schell, Novell, Inc. Provo, UT, USA. Computers & Security, 14(7):609, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/0167404896816804.

Highland:1995:BRAa

[Hig95h] Harold Joseph Highland. Book review: Architectures for C2 DOS/Windows-based Personal Computers — Securing an ‘Undesirable’ Operating System: Jeremy Epstein, Gary Grossman, Frederick Maxwell, Nobel Veirs III, Albert Donaldson, and Cornelius Hal ey; Cordant Inc, Reston, VA. Computers & Security, 14(1):30, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/016740489597005U.

Highland:1995:BRAc

[Hig95i] Harold Joseph Highland. Book review: Authorship analysis: Identifying the author of a program: Ivan Krsul and Eugene H. Spafford, Purdue University, West Lafayette, IN, USA. Computers & Security, 14(7):609–610, ???? 1995. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/016740489681683X.

Highland:1995:BRC

[Hig95j] Harold Joseph Highland. Book review: Component architectures for trusted netware: Jeremy Epstein and Gary Grossman, Cordant Inc., Reston, VA, USA; and Roger Schell, Novell, Inc. Provo, UT, USA. Computers & Security, 14(7):609, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/0167404896816828.

Highland:1995:BREb

[Hig95k] Harold Joseph Highland. Book review: ECMA’s approach for IT security evaluations: Alexander Herrigel, R3 Security Engineering AG, Switzerland; Roger French, Digital Equipment Corp, USA; and Haruki Tabuchi, Fujitsu, Ltd, Japan. Computers & Security, 14(7):608, ???? 1995. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/0167404896816774.

Highland:1995:BREc

[Hig95l] Harold Joseph Highland. Book review: Evaluting the strength of ciphers: John C. Higgins, Brigham Young University, Provo, UT, USA. Computers & Security, 14(7):608, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/0167404896816786.

Highland:1995:BRI

[Hig95m] Harold Joseph Highland. Book review: Implementing Role Based, Clark-Wilson Enforcement Rules in a Bl Online Transaction Processing System: Barbara Smith-Thomas, AT&T Bell Laboratories, Greensboro, NC and Wang Chao-Yeuh, Institute for Information Industry, Taipei, Taiwan. Computers & Security, 14(1):29, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/016740489597003S.

Highland:1995:BRMb

[Hig95n] Harold Joseph Highland. Book review: Maintaining secretary and integrity in multilevel databases: a practical approach: Sushil Jajodia, George Mason University, Fairfax, VA; Don Marks, Department of Defense, Ft. Meade; and Elisa Bertino, Universita di Milano, Milano, Italy. Computers & Security, 14(7):605–606, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/0167404896816683.

Highland:1995:BRMa

[Hig95o] Harold Joseph Highland. Book review: Modern Multilevel Security (MLS): Practical Approaches for Integration, Certification and Accreditation: Bill Neugent, Mike Burgoon,Jeanne Firev and Mundv Rudell. The MITRE Corporation. McLean, VA. Computers & Security, 14(1):31–32, ???? 1995. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/016740489597012Y.

Highland:1995:BRSc

[Hig95p] Harold Joseph Highland. Book review: Securing local area and metropolitan area network: a practical approach: Prof. Vijay Vradharajan, University of Western Sydney, Nepean, Australia. Computers & Security, 14(7):607, ???? 1995. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/0167404896816725.

Highland:1995:BRSa

[Hig95q] Harold Joseph Highland. Book review: Support for Security in Distributed Systems Using Messiahs: Steve Chapin, Kent State University, Kent, OH and Eugene Spafford, Purdue University, West Lafayette, IN. Computers & Security, 14(1):

30, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/016740489597006V.

Highland:1995:BRTa

[Hig95r] Harold Joseph Highland. Book review: Testing Intrusion Detection Systems: Design Methodologies and Results from an Early Prototype: Nicholas Puketza, Kiswanath Mukherjee, Ronald Olsson and Kui Zhang, University of California, Davis, CA. Computers & Security, 14(1):28, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/016740489596998I.

Highland:1995:BREa

[Hig95s] Harold Joseph Highland. Book review: The Electronic Intrusion Threat to National Security and Emergency Preparedness Telecommunications: an Awareness Document: Dr. Joseph Frizzell, National Communications System, Arlington, VA and Ted Phillips and Traigh Grover, Booz, Alien and Hamilton Inc, McLean, VA. Computers & Security, 14(1):30–31, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/016740489597008X.

Highland:1995:BRTb

[Hig95t] Harold Joseph Highland. Book review: Towards a Formal Verification of a Secure and Distributed System and its Applications: Cui Zhang, Rob Shaw, Mark Heckman, Gregory Benson, Myla Archer, Karl Levitt and Ronald Olsson, Dept. Computer Science, University of California, Davis, CA. Computers & Security, 14(1):29, ???? 1995. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/016740489597002R.

Highland:1995:BRUa

[Hig95u] Harold Joseph Highland. Book review: Using Application Profiles to Detect Computer Misuse: Nancy Kelem and Daniel Sterne, Trusted Information Systems, Mountain View, CA; Debra Anderson, SRI International, Menio Park, CA; and Linda Lankewicz, Spring Hill College, Mobile, AL. Computers & Security, 14(1):27, ???? 1995. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/016740489596993D.

Highland:1995:BRUb

[Hig95v] Harold Joseph Highland. Book review: Using network traffic analysis as a security tool: Peter Troxell, Curry Bartlett and Nicholas Gill, Digital Equipment Corp, Beavercreek, OH, USA. Computers & Security, 14(7):607, ???? 1995. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/0167404896816737.

Highland:1995:BRV

[Hig95w] Harold Joseph Highland. Book review: Virtual View Model to Design a Secure Object-Oriented Database: N. Boulahia-Cuppens, F. Cuppens, A. Gabillon, K. Yazdaman, ONERA/ CERT, Cedex, France. Computers & Security, 14(1):31, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/016740489597011X.

Highland:1995:BRW

[Hig95x] Harold Joseph Highland. Book review: What Color is Your Assurance?: David Wichers, Joel Sachs and Douglas Landoll, Area Systems Inc, Columbia, MD. Computers & Security, 14 (1):28, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/016740489596999J.

Highland:1995:CCC

[Hig95y] Harold Joseph Highland. Can computer crime be deterred: Sanford Sherizen, Data Security Systems, Inc, Natick, MA. Computers & Security, 14(1):28, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/016740489596996G.

Highland:1995:CPW

[Hig95z] Harold Joseph Highland. Contingency planning: What to do when bad things happen to good systems: Jay. J. Kahn and Marshall D. Abrams, The MITRE Corp., McLean, VA, USA. Computers & Security, 14(7):609, ???? 1995. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/0167404896816798.

Highland:1995:CAS

[Hig95-27] Harold Joseph Highland. The controlled application set paradigm for trusted systems: Daniel Sterne, Trusted Information Systems Inc, Glenwood, MD, USA and Glenn Benson, European Computer-Industry Research Centre, Munich, Germany. Computers & Security, 14(7):605, ???? 1995. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/0167404896816671.

Highland:1995:CFK

[Hig95-28] Harold Joseph Highland. Critical factors of key escrow encryption systems: Dorothy E. Denning, Georgetown University, Washington, DC, USA. Computers & Security, 14(7): 608, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/0167404896816750.

Highland:1995:EEA

[Hig95-29] Harold Joseph Highland. Exporting evaluation: an analysis of US and Canadian criteria for trust: Paul Olson. Computers & Security, 14(1):31, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/016740489597009Y.

Highland:1995:ITNa

[Hig95-30] Harold Joseph Highland. IFIP TC 11 news. Computers & Security, 14(1):33–34, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/0167404895900772.

Highland:1995:ITNb

[Hig95-31] Harold Joseph Highland. IFIP TC 11 news. Computers & Security, 14(5):405–406, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404896900322.

Highland:1995:ISA

[Hig95-32] Harold Joseph Highland. Internet sniffer attacks: E. Eugene Schultz, SRI International, Menlo Park, CA and Thomas Longstaff, Carnegie Mellon University, Pittsburgh, PA, USA. Computers & Security, 14(7):610, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/0167404896816841.

Highland:1995:MPE

[Hig95-33] Harold Joseph Highland. Maintaining privacy in electronic transactions: Carnegie Mellon University, Pittsburgh, PA, USA. Computers & Security, 14(7):606–607, ???? 1995. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/0167404895901280.

Highland:1995:PSA

[Hig95-34] Harold Joseph Highland. Partitioning the security analysis of complex systems: Howard Holm, National Security Agency, Fort Meade, MD. Computers & Security, 14(1):30, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/016740489597004T.

Highland:1995:PEU

[Hig95-35] Harold Joseph Highland. A perspective of evaluation in the UK versus the US: Alan Borrett, UK ITSEC Scheme, Cheltenham, Glos., UK. Computers & Security, 14(7):607–608, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/0167404896816749.

Highland:1995:RBBa

[Hig95-36] Harold Joseph Highland. Random bits and bytes. Computers & Security, 14(1):4–13, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/016740489596990K.

Highland:1995:RBBb

[Hig95-37] Harold Joseph Highland. Random bits and bytes. Computers & Security, 14(2):90–102, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/0167404895970289.

Highland:1995:RBBc

[Hig95-38] Harold Joseph Highland. Random bits and bytes. Computers & Security, 14(3):188–196, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/016740489597051B.

Highland:1995:RBBd

[Hig95-39] Harold Joseph Highland. Random bits and bytes. Computers & Security, 14(4):274–282, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/016740489597067K.

Highland:1995:RBBe

[Hig95-40] Harold Joseph Highland. Random bits and bytes. Computers & Security, 14(5):365–371, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/016740489597086P.

Highland:1995:RBBf

[Hig95-41] Harold Joseph Highland. Random bits and bytes. Computers & Security, 14(6):480–488, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/016740489599310L.

Highland:1995:RBBg

[Hig95-42] Harold Joseph Highland. Random bits and bytes. Computers & Security, 14(7):574–585, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/0167404896816646.

Highland:1995:RBBh

[Hig95-43] Harold Joseph Highland. Random bits and bytes. Computers & Security, 14(8):660–666, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/0167404896817056.

Highland:1995:SVP

[Hig95-44] Harold Joseph Highland. The Sea View Prototype: Project summary: Teresa Lunt and Peter Boucher, SRI International, Menio Park, CA. Computers & Security, 14(1):29, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/016740489597001Q.

Highland:1995:SSR

[Hig95-45] Harold Joseph Highland. Security and software reuse: George W. Rogers, Jr. and Jerry C. Crabb, The Analysis Corporation, Falls Church, VA, USA. Computers & Security, 14(7): 608, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/0167404896816762.

Highland:1995:SAP

[Hig95-46] Harold Joseph Highland. Security awareness and the persuasion of managers: Dennis Poindexter, Professionalization Directorate, TGB. Arlington, VA. Computers & Security, 14(1): 27, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/016740489596994E.

Highland:1995:SEO

[Hig95-47] Harold Joseph Highland. Social engineering; the only real test of information systems security plans: Ira S. Winkler, Science Applications International Corporation, Annapolis, MD, USA. Computers & Security, 14(7):609, ???? 1995. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/0167404896816816.

Highland:1995:TCT

[Hig95-48] Harold Joseph Highland. The table of contents for the 1st through the 17th National Conputer Security Conferences: Jack Holloran, National Computer Security Center and Darlene Affeldt, National Security Agency. Computers & Security, 14(7): 610, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/0167404896816853.

Highland:1995:TSS

[Hig95-49] Harold Joseph Highland. A taxonomy for security standards: Wayne Jansen, National Institute of Standards and Technology, Gaithersburg, MD. Computers & Security, 14(1):30, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/016740489597007W.

Highland:1995:TIL

[Hig95-50] Harold Joseph Highland. TCP/IP (lack of) security: Jasper M. Johansson, University of Maryland, College Park, MD, USA. Computers & Security, 14(7):606, ???? 1995. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/0167404896816701.

Highland:1996:ITN

[Hig96a] Dr. Harold Joseph Highland. IFIP TC 11 news. Computers & Security, 15(5):387–390, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/0167404896825612.

Highland:1996:EW

[Hig96b] Harold Joseph Highland. The enemy within. Computers & Security, 15(8):657, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404896902382.

Highland:1996:IWD

[Hig96c] Harold Joseph Highland. Internet/web directory. Computers & Security, 15(8):656–657, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404896902370.

Highland:1996:WBN

[Hig96d] Harold Joseph Highland. It’s a web browser! No, it’s a weapon! Computers & Security, 15(8):655–656, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404896902369.

Highland:1996:NVI

[Hig96e] Harold Joseph Highland. Ntis volumes of interest. Computers & Security, 15(8):658–660, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404896902394.

Highland:1996:PWI

[Hig96f] Harold Joseph Highland. Passwords, the web, and the Internet. Computers & Security, 15(8):654, ???? 1996. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404897896988.

Highland:1996:PP

[Hig96g] Harold Joseph Highland. Pornography and privacy. Computers & Security, 15(8):654–655, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404896902357.

Highland:1996:RBBa

[Hig96h] Harold Joseph Highland. Random bits and bytes. Computers & Security, 15(1):4–11, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404896900577.

Highland:1996:RBBb

[Hig96i] Harold Joseph Highland. Random bits and bytes. Computers & Security, 15(2):90–99, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/0167404896893217.

Highland:1996:RBBc

[Hig96j] Harold Joseph Highland. Random bits and bytes. Computers & Security, 15(3):184–197, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/016740489689780X.

Highland:1996:RBBd

[Hig96k] Harold Joseph Highland. Random bits and bytes. Computers & Security, 15(4):270–280, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404896902151.

Highland:1996:RBBe

[Hig96l] Harold Joseph Highland. Random bits and bytes. Computers & Security, 15(5):356–366, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404896903533.

Highland:1996:RBBf

[Hig96m] Harold Joseph Highland. Random bits and bytes. Computers & Security, 15(6):458–469, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404897831197.

Highland:1996:RBBg

[Hig96n] Harold Joseph Highland. Random bits and bytes. Computers & Security, 15(7):552–560, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404897881102.

Highland:1997:ITNa

[Hig97a] Dr. Harold Joseph Highland. IFIP TC 11 news. Computers & Security, 16(3):211–212, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404897902662.

Highland:1997:CLN

[Hig97b]  Harold Joseph Highland.  Current literature the 19th

national information systems security conference. Computers & Security, 16(1):47–59, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404897857860.

Highland:1997:DEN

[Hig97c] Harold Joseph Highland. Data encryption: a non-mathematical approach. Computers & Security, 16(5):369–386, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404897822432.

Highland:1997:HBB

[Hig97d] Harold Joseph Highland. Historical bits and bytes. Computers & Security, 16(5):387–411, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404897822444.

Highland:1997:HCVa

[Hig97e] Harold Joseph Highland. A history of computer viruses — introduction. Computers & Security, 16(5):412–415, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404897822456.

Highland:1997:HCVb

[Hig97f] Harold Joseph Highland. A history of computer viruses — the famous ‘trio’. Computers & Security, 16(5):416–429, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404897822468.

Highland:1997:HCVc

[Hig97g] Harold Joseph Highland. A history of computer viruses: Three special viruses. Computers & Security, 16(5):430–438, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S016740489782247X.

Highland:1997:ITNb

[Hig97h] Harold Joseph Highland. IFIP TC 11 news. Computers & Security, 16(4):327–328, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404897901553.

Highland:1997:ISC

[Hig97i] Harold Joseph Highland. IFIP/SEC’97 conference proceedings. Computers & Security, 16(3):205, ???? 1997. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404897845130.

Highland:1997:NS

[Hig97j] Harold Joseph Highland. News scan. Computers & Security, 16(3):180–181, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404897902601.

Highland:1997:NSN

[Hig97k] Harold Joseph Highland. News scan and notes . . .. Computers & Security, 16(2):86–87, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404897882880.

Highland:1997:NVIa

[Hig97l] Harold Joseph Highland. NTIS volumes of interest. Computers & Security, 16(2):91–93, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S016740489790225X.

Highland:1997:NVIb

[Hig97m] Harold Joseph Highland. NTIS volumes of interest. Computers & Security, 16(3):187, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404897845075.

Highland:1997:PRC

[Hig97n] Harold Joseph Highland. Procedures to reduce the computer virus threat. Computers & Security, 16(5):439–449, ????

1997. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404897822481.

Highland:1997:RBBb

[Hig97o] Harold Joseph Highland. Random bits and bytes. Computers & Security, 16(4):266–275, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404897801897.

Highland:1997:RBBa

[Hig97p] Harold Joseph Highland. Random bits and bytes and bytes. Computers & Security, 16(1):4–13, ???? 1997. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404897857823.

Highland:1997:SS

[Hig97q] Harold Joseph Highland. Salvage and security. Computers & Security, 16(2):90–91, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404897902248.

Highland:1997:SC

[Hig97r] Harold Joseph Highland. Security in computing. Computers & Security, 16(3):181, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404897902613.

Highland:1997:TW

[Hig97s] Harold Joseph Highland. The threats on the web. Computers & Security, 16(5):365–368, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404897822420.

Highland:1997:VSM

[Hig97t] Harold Joseph Highland. Virus scanners for multiple OSes. Computers & Security, 16(3):182–184, ???? 1997. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404897902637.

Highland:1997:WSD

[Hig97u] Harold Joseph Highland. Web site directory. Computers & Security, 16(3):186–187, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404897902650.

Highland:1997:WBC

[Hig97v] Harold Joseph Highland. Why I bought a Compaq. Computers & Security, 16(3):184–186, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404897902649.

Highland:1997:WNR

[Hig97w] Harold Joseph Highland. Windows NT registry. Computers & Security, 16(2):88–90, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404897902236.

Highland:1997:Y

[Hig97x] Harold Joseph Highland. Year 2000. Computers & Security, 16(3):181–182, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404897902625.

Highland:1997:YMW

[Hig97y] Harold Joseph Highland. You must be on the web. Computers & Security, 16(2):87–88, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404897902224.

Hiles:1991:CFR

[Hil91] Andrew Hiles. Computer fire risks. Computers & Security, 10(4):299–302, June 1991. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/016740489190104L.

Hiles:1995:WWM

[Hil95] Andrew Hiles. Why waste money on disaster recovery. Computers & Security, 14(5):423, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/0167404896900462.

Hinde:1998:CPM

[Hin98a] Stephen Hinde. Contingency planning for the millennium. Computers & Security, 17(4):299–302, ???? 1998. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404898800096.

Hinde:1998:CGD

[Hin98b] Stephen Hinde. Corporate governance and disasters. Computers & Security, 17(7):583–585, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404899800571.

Hinde:1998:CWO

[Hin98c] Stephen Hinde. Cyber wars and other threats. Computers & Security, 17(2):115–118, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404897819797.

Hinde:1998:HWI

[Hin98d] Stephen Hinde. Hot water, icebergs and other disasters. Computers & Security, 17(1):31–33, ???? 1998. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404897802477.

Hinde:1998:PSD

[Hin98e] Stephen Hinde. Privacy and security — the drivers for growth of e-commerce. Computers & Security, 17(6):475–478, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404898800692.

Hinde:1998:RSS

[Hin98f] Stephen Hinde. Recent security surveys. Computers & Security, 17(3):207–210, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404898803118.

Hinde:1998:SRB

[Hin98g] Stephen Hinde. Solar radiation is bad for you. Computers & Security, 17(5):385–388, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404898800047.

Hinde:1998:THW

[Hin98h] Stephen Hinde. Time, history and war: a personal voyage through Compsec ’98. Computers & Security, 17(8):671–681, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404898800989.

Hinde:1999:YC

[Hin99a] Stephen Hinde. Are you in control? Computers & Security, 18(3):199–203, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404899800649.

Hinde:1999:BTB

[Hin99b] Stephen Hinde. The bell tolls for the bell-shaped curve: or how hot magnets, species distribution and things that go bump in the air will change our view of risk. Computers & Security, 18(7):565–572, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404899820021.

Hinde:1999:EG

[Hin99c] Stephen Hinde. E for exponential growth. Computers & Security, 18(4):295–299, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404899800728.

Hinde:1999:MCS

[Hin99d] Stephen Hinde. E-mail can seriously damage your health. Computers & Security, 18(5):396–406, ???? 1999. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404899800856.

Hinde:1999:OOG

[Hin99e] Stephen Hinde. One down, one to go. Computers & Security, 18(1):19–22, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404899800042.

Hinde:1999:SSN

[Hin99f] Stephen Hinde. Step into a secure new world — Compsec ’99 report. Computers & Security, 18(8):660–669, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404899801308.

Hinde:1999:YU

[Hin99g] Stephen Hinde. Year 2000 update. Computers & Security, 18(2):113–119, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404899800297.

Hitchings:1995:DTA

[Hit95] Jean Hitchings. Deficiencies of the traditional approach to information security and the requirements for a new methodology. Computers & Security, 14(5):377–383, ???? 1995. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/016740489597088R.

Hochberg:1993:NAS

[HJS^{+}93] Judith Hochberg, Kathleen Jackson, Cathy Stallings, J. F. McClary, David DuBois, and Josephine Ford. NADIR: an automated system for detecting network intrusion and misuse. Computers & Security, 12(3):235–248, May 1993. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/016740489390110Q.

Harn:1990:CKG

[HL90] Lein Harn and Hung-Yu Lin. A cryptographic key generation scheme for multilevel data security. Computers & Security, 9(6):539–546, October 1990. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/016740489090132D.

Harn:1992:SAS

[HLY92] Lein Harn, Hung-Yu Lin, and Shoubao Yang. A software authentication system for information integrity. Computers & Security, 11(8):747–752, December 1992. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/016740489290130J.

Hui-Min:1995:CID

[HMCC95] Tsai Hui-Min and Chang Chin-Chen. A cryptographic implementation for dynamic access control in a user hierarchy. Computers & Security, 14(2):159–166, ???? 1995. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/016740489597049G.

Hoath:1997:TFG

[Hoa97] Peter Hoath. Telecommunications fraud, the gory details. Computers & Security, 16(6):521, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404897846780.

Holloway:1995:CDSa

[Hol95a] Chris Holloway. Controlling digital signature services using a smartcard. Computers & Security, 14(5):419, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/016740489597118T.

Holloway:1995:CUCa

[Hol95b] Chris Holloway. Controlling the usage of cryptographic keys. Computers & Security, 14(5):419, ???? 1995. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/016740489597116R.

Holloway:1995:CDSb

[Hol95c] Christopher J. Holloway. Controlling digital signature services using a smartcard. Computers & Security, 14(8):681– 690, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/0167404896817081.

Holloway:1995:CUCb

[Hol95d] Christopher J. Holloway. Controlling the use of cryptographic keys. Computers & Security, 14(7):587–598, ???? 1995. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/0167404896816658.

Homer:1990:SSE

[Hom90] Steve Homer. Setting standards in Europe. Computers & Security, 9(4):295–300, June 1990. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/016740489090096C.

Hoppe:1994:ACS

[Hop94] Norman Hoppé. Achieving consistent security controls throughout a multinational organization. Computers & Security, 13 (1):23–29, February 1994. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404894900876.

Hopwood:1997:CBJ

[Hop97] David Hopwood. A comparison between Java and activeX security. Computers & Security, 16(6):526, ???? 1997. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404897847025.

Horwitt:1992:STF

[Hor92] Elisabeth Horwitt. Sprint, telematic fight phone fraud. Computers & Security, 11(5):489–490, September 1992. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/016740489290244L.

Herschberg:1990:FFF

[HP90a] I. S. Herschberg and R. Paans. Friday the 13th, facts and fancies. Computers & Security, 9(2):125–130, April 1990. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/0167404890900847.

Herschberg:1990:SSS

[HP90b] I. S. Herschberg and R. Paans. Sloppy, sloppier, sloppiest. Computers & Security, 9(3):223–228, May 1990. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/016740489090167R.

Hruska:1995:NOS

[Hru95] Jan Hruska. New operating systems and old viruses. Computers & Security, 14(5):419, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/016740489597117S.

Harris:1994:DDS

[HS94a] Duncan Harris and David Sidwell. Distributed database security. Computers & Security, 13(7):547–557, ???? 1994. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/0167404894900035.

Hudoklin:1994:CSS

[HS94b] Alenka Hudoklin and Branislav Smitek. Computer systems security in Slovenia. Computers & Security, 13(1):30–33, February 1994. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/0167404894900884.

Hardjono:1995:ASA

[HS95a] Thomas Hardjono and Jennifer Seberry. Applications of smartcards for anonymous and verifiable databases. Computers & Security, 14(5):465–472, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/016740489500011V.

Hill:1995:RMC

[HS95b] Scott Hill and Martin Smith. Risk management and corporate security: a viable leadership and business solution designed to enhance corporations in the emerging marketplace. Computers & Security, 14(3):199–204, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/016740489597053D.

Hwang:1996:ACS

[HTY96] Min-Shiang Hwang, Wen-Guey Tzeng, and Wei-Pang Yang. An access control scheme based on Chinese remainder theorem and time stamp concept. Computers & Security, 15(1): 73–81, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/0167404895000283.

Hughes:1995:AAU

[Hug95a] John Hughes. Auditing and accounting: Usage and requirements. Computers & Security, 14(5):427, ???? 1995. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/016740489597152Z.

Hughes:1995:XSC

[Hug95b] John Hughes. X.400 security: Current and future usage. Computers & Security, 14(5):422, ???? 1995. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/016740489597128W.

Hughes:1999:CIO

[Hug99] John Hughes. Certificate inter-operability — white paper. Computers & Security, 18(3):221–230, ???? 1999. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404899800674.

Hardjono:1994:DAR

[HZS94] Thomas Hardjono, Yuliang Zheng, and Jennifer Seberry. Database authentication revisited. Computers & Security, 13 (7):573–580, ???? 1994. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404894900086.

IW1:1994:IIS

[ILM94] IFIP Working Group 11.5, William List, and Rob Melville. Integrity in information systems — executive summary. Computers & Security, 13(4):295–301, ???? 1994. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/0167404894900183.

Jackson:1990:CSC

[Jac90] Keith Jackson. Confessions of a security consultant. Computers & Security, 9(6):509–511, October 1990. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/016740489090127F.

Janusz:1992:PPP

[Jan92] Cheryl Janusz. Power protection plays a vital role in network security. Computers & Security, 11(5):490, September 1992. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/016740489290249Q.

Jarvis:1999:CEB

[Jar99] Neil Jarvis. E-commerce and encryption: Barriers to growth. Computers & Security, 18(5):429–431, ???? 1999. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404899800893.

Jones:1993:PCV

[JATC93] M. C. Jones, K. P. Arnett, J.-T. E. Tang, and N.-S. Chen. Perceptions of computer viruses: a cross-cultural assessment. Computers & Security, 12(2):191–197, March 1993. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/016740489390101A.

Jan:1991:DKL

[JCW91] J. K. Jan, C. C. Chang, and S. J. Wang. A dynamic key-lock-pair access control scheme. Computers & Security, 10(2):129– 139, April 1991. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/016740489190007Z.

Jing-Jang:1996:CAS

[JJ96] Hwang Jing-Jang. A conventional approach to secret balloting in computer networks. Computers & Security, 15(3):249– 262, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/0167404896000065.

Juang:1996:CFS

[JL96] Wen-Shenq Juang and Chin-Laung Lei. A collision-free secret ballot protocol for computerized general elections. Computers & Security, 15(4):339–348, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/0167404896000119.

Joachim:1997:YP

[Joa97] Margaret Joachim. The year 2000 problem. Computers & Security, 16(6):526, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404897847037.

Jones:1997:PTSa

[Jon97a] Andy Jones. Penetration testing and system audit — experience gained and lessons learned during the investigation of systems within the United Kingdom. Computers & Security, 16(6):525– 526, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404897846998.

Jones:1997:PTSb

[Jon97b] Andy Jones. Penetration testing and system audit — experience gained during the investigation of systems within the UK. Computers & Security, 16(7):595–602, ???? 1997. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404897807961.

Jansen:1996:MEM

[JvdV96] Cees J. A. Jansen and Piet van der Vlist. Message encipherment with minimal expansion and redundancy-doing better than ISO-10126. Computers & Security, 15(7):627–632, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404896000132.

Jones:1990:IMC

[JW90] Scott K. Jones and Clinton E. White. The IPM model of computer virus management. Computers & Security, 9(5):411–417, August 1990. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/0167404890900722.

Kanungo:1994:IAH

[Kan94] Shivraj Kanungo. Identity authentication in heterogeneous computing environments: a comparative study for an integrated framework. Computers & Security, 13(3):231–253, May 1994. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/0167404894900787.

Kapidzic:1998:CSA

[Kap98] Nada Kapidzic. Creating security applications based on the global certificate management system. Computers & Security, 17(6):507–515, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404898800722.

Kelman:1995:DLR

[Kel95] Alistair Kelman. Developing legal risks in multimedia. Computers & Security, 14(5):422–423, ???? 1995. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/016740489597132T.

Kemp:1997:ESOa

[Kem97a] Alasdair Kemp. Embedded systems: the other problem. Computers & Security, 16(6):527, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404897847050.

Kemp:1997:ESOb

[Kem97b] Alasdair Kemp. Embedded systems: the other problem. Computers & Security, 16(8):663–668, ???? 1997. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404897875803.

Kensey:1993:CVT

[Ken93] Michael F. Kensey. Computer viruses — towards better solutions. Computers & Security, 12(6):536–541, October 1993. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/016740489390048A.

Kim:1995:BIV

[Kim95] Hyun-Jung Kim. Biometrics, is it a viable proposition for identity authentication and access control? Computers & Security, 14(3):205–214, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/016740489597054E.

King:1992:XS

[Kin92a] Judith King. X.400 security. Computers & Security, 11(8): 707–710, December 1992. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/016740489290090E.

King:1992:CT

[Kin92b] Julia King. It’s C.Y.A. time. Computers & Security, 11(5):491, September 1992. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/016740489290254O.

Kailay:1995:RPE

[KJ95] Muninder P. Kailay and Peter Jarratt. RAMeX: a prototype expert system for computer security risk analysis and management. Computers & Security, 14(5):449–463, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/016740489500013X.

Kwok:1996:SOW

[KL96] Lam For Kwok and Dennis Longley. A security officer’s workbench. Computers & Security, 15(8):695–705, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404897897222.

Klopp:1990:MSS

[Klo90a] Charlotte Klopp. Microcomputer security systems attempt to lock up desktops. Computers & Security, 9(2):139–141, April 1990. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/0167404890900869.

Klopp:1990:MOP

[Klo90b] Charlotte Klopp. More options for physical access control. Computers & Security, 9(3):229–232, May 1990. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/016740489090168S.

Klopp:1990:SVC

[Klo90c] Charlotte Klopp. Software vendors capitalize on security needs. Computers & Security, 9(1):33–35, February 1990. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/016740489090152J.

Klopp:1990:VAI

[Klo90d] Charlotte Klopp. Vulnerability awareness improves contingency planning. Computers & Security, 9(4):309–311, June 1990. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/016740489090099F.

Koerner:1993:STV

[Koe93] Frank Koerner. System threats and vulnerabilities and the contrary principle. Computers & Security, 12(8):775–779, December 1993. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/0167404893900435.

Kovacich:1995:RCG

[Kov95] Gerald Kovacich. Recovering computer-generated evidence. Computers & Security, 14(4):289–296, ???? 1995. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/016740489597070Q.

Kovacich:1996:ENS

[Kov96] Gerald L. Kovacich. Establishing a network security programme. Computers & Security, 15(6):486–498, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404897831227.

Kovacich:1997:ICD

[Kov97a] Dr. Gerald Kovacich. ISSO career development. Computers & Security, 16(6):455–458, 461–462, 464–468, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404897846408.

Kovacich:1997:HMY

[Kov97b] Gerald Kovacich. How to market yourself as an ISSO. Computers & Security, 16(8):657–662, ???? 1997. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404897875797.

Kovacich:1997:ISS

[Kov97c] Gerald Kovacich. Information systems security metrics management. Computers & Security, 16(7):610–618, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404897807985.

Kovacich:1997:IMU

[Kov97d] Gerald Kovacich. The ISSO must understand the business and management environment. Computers & Security, 16(4):321– 326, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404897801964.

Kovacich:1997:IWI

[Kov97e] Gerald L. Kovacich. Information warfare and the information systems security professional. Computers & Security, 16(1):14– 24, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404897857835.

Kovacich:1998:EIB

[Kov98a] Gerald Kovacich. Electronic-Internet business and security. Computers & Security, 17(2):129–135, ???? 1998. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404897819827.

Kovacich:1998:EIS

[Kov98b] Gerald Kovacich. Establishing an information systems security organization (ISSO). Computers & Security, 17(7):600– 612, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404899800601.

Kovacich:1999:HFF

[Kov99a] Gerald L. Kovacich. Hackers: Freedom fighters of the 21-st century. Computers & Security, 18(7):573–576, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404899820033.

Kovacich:1999:WRC

[Kov99b] Gerald L. Kovacich. I-way robbery: Crime on the Internet. Computers & Security, 18(3):211–220, ???? 1999. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404899800662.

Kowalski:1991:ANC

[Kow91a] Stewart Kowalski. The ABCs and Es of national computer security policies. Computers & Security, 10(3):213–216, May 1991. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/016740489190036D.

Kowalski:1991:CAN

[Kow91b] Stewart Kowalski. Cybernetic analysis of national computer security. Computers & Security, 10(3):217–227, May 1991. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/016740489190037E.

Kowalski:1991:CAU

[Kow91c] Stewart Kowalski. Cybernetic analysis of the 12th U.S. National Computer Security Conference. Computers & Security, 10(3):228–235, May 1991. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/016740489190038F.

King:1999:STA

[KP99] R. A. King and Mr T. C. Phipps. Shannon, TESPAR and approximation strategies. Computers & Security, 18(5):445– 453, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404899801114.

Kruys:1991:PSD

[Kru91] J. P. Kruys. Progress in secure distributed systems. Computers & Security, 10(5):429–441, August 1991. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/016740489190069P.

Krull:1995:CNMa

[Kru95a] Alan Krull. Controls in the next millennium: Anticipating the IT-enabled future. Computers & Security, 14(5):418, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/016740489597112N.

Krull:1995:MIS

[Kru95b] Alan Krull. Marketing information systems audit. Computers & Security, 14(5):423, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404896900474.

Krull:1995:CNMb

[Kru95c] Alan R. Krull. Controls in the next millennium: Anticipating the IT-enabled future. Computers & Security, 14(6):491– 495, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/016740489599312D.

Krull:1996:GGA

[Kru96] Alan R. Krull. GSSP (Generally-Accepted System Security Principles): a trip to Abilene? Computers & Security, 15 (7):567–575, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404897881126.

Krsul:1997:AAI

[KS97] Ivan Krsul and Eugene H. Spafford. Authorship analysis: identifying the author of a program. Computers & Security, 16(3):233–257, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404897000059.

Kuhnhauser:1999:PG

[Küh99] Winfried E. Kühnhauser. Policy groups. Computers & Security, 18(4):351–363, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404899800819.

Laferriere:1990:DIS

[Laf90] Claude Laferriere. A discussion of implementation strategies for secure database management systems. Computers & Security, 9(3):235–244, May 1990. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/016740489090170X.

Laferriere:1991:PBP

[Laf91] Claude Laferriere. Predicate based polyinstantiation in multilevel secure DBMS. Computers & Security, 10(1):41–49, February 1991. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/016740489190055I.

Laird:1997:SBG

[Lai97] Alan Laird. Smartcards — is Britain getting smarter? Computers & Security, 16(6):526, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404897847049.

Lathrop:1990:PIC

[Lat90] David L. Lathrop. Perestroika and its implications for computer security in the U.S.S.R. Computers & Security, 9(8):693–696, December 1990. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/0167404890901116.

Lathrop:1992:SAW

[Lat92] David L. Lathrop. Security aspects of wireless local area networks. Computers & Security, 11(5):421–426, September 1992. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/016740489290006D.

Lathrop:1993:BRN

[Lat93] David L. Lathrop. Book review: Network security in the ’90s: Issues and solutions for managers: Thomas W. Madron John Wiley and Sons Inc, New York, 1992, 293 pages, $32.95, ISBN: 0-471-54777-8. Computers & Security, 12(2):135, March 1993. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/016740489390093K.

Laun:1992:AUA

[Lau92] Rotraut Laun. Asymmetric user authentication. Computers & Security, 11(2):173–183, April 1992. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/016740489290044R.

Lau:1998:TCS

[Lau98] Oliver Lau. The ten commandments of security. Computers & Security, 17(2):119–123, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404897819803.

Lawrence:1992:RSG

[Law92] L. G. Lawrence. RACIF security goals comment on a paper by r. paans. Computers & Security, 11(5):407–411, September 1992. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/016740489290003A.

Lawrence:1993:DSE

[Law93a] L. G. Lawrence. Digital signatures — explanation and usage. Computers & Security, 12(3):230–233, May 1993. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/016740489390109I.

Lawrence:1993:RR

[Law93b] L. G. Lawrence. The role of roles. Computers & Security, 12(1):15–21, February 1993. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/016740489390004O.

Laurichesse:1991:OIR

[LB91] Denis Laurichesse and Laurent Blain. Optimized implementation of RSA cryptosystem. Computers & Security, 10(3):263– 267, May 1991. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/016740489190042C.

Looi:1994:NST

[LC94] Mark H. Looi and William J. Caelli. A note on supplying a trusted clock via a secure device. Computers & Security, 13(7):611–613, ???? 1994. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/0167404894900124.

Laih:1994:APP

[LCC94] Chi-Sung Laih, Wen-Hong Chiou, and Chin-Chen Chang. Authentication and protection of public keys. Computers & Security, 13(7):581–585, ???? 1994. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/0167404894900094.

Labuschagne:1997:ISA

[LE97] L. Labuschagne and J. H. P. Eloff. Improved systemaccess control using complementary technologies. Computers & Security, 16(6):543–549, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404897000072.

Labuschagne:1998:URT

[LE98] L. Labuschagne and J. H. R. Eloff. The use of real-time risk analysis to enable dynamic activation of countermeasures. Computers & Security, 17(4):347–357, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404898800436.

Leach:1995:DAS

[Lea95] John Leach. Dynamic authentication for smartcards. Computers & Security, 14(5):385–389, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/016740489597089S.

Lemieux:1996:URT

[Lem96] Jacques R. Lemieux. Using RAD tools to develop secure client/server applications. Computers & Security, 15(4):289– 295, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/0167404896889401.

Lindgreen:1994:VBP

[LH94] E. E. O. Roos Lindgreen and I. S. Herschberg. On the validity of the Bell–La Padula model. Computers & Security, 13(4):317– 333, ???? 1994. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/016740489490023X.

Laferriere:1990:SAT

[LHB90] Claude Laferriere, G. Owen Higginson, and Gordon G. Bell. Security architectures for textual databases. Computers & Security, 9(7):621–630, November 1990. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/016740489090061W.

Liaw:1994:DCK

[Lia94] Horng-Twu Liaw. A dynamic cryptographic key generation and information broadcasting scheme in information systems. Computers & Security, 13(7):601–610, ???? 1994. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/0167404894900116.

Lichtenstein:1998:IRC

[Lic98] Sharman Lichtenstein. Internet risks for companies. Computers & Security, 17(2):143–150, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404897819840.

Lillywhite:1994:ISS

[Lil94] Thomas Lillywhite. IT security standards and the adoption of a benchmarking approach. Computers & Security, 13(6):467– 471, ???? 1994. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/016740489190132W.

Lindballe:1994:LE

[Lin94a] Jørgen Lindballe. Letter to the editor. Computers & Security, 13(4):303–305, ???? 1994. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404894900191.

Lindup:1994:CA

[Lin94b] Ken Lindup. The cyberpunk age. Computers & Security, 13(8):637–645, ???? 1994. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404894900434.

Lindup:1995:NMI

[Lin95] Kenneth R. Lindup. A new model for information security policies. Computers & Security, 14(8):691–695, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/0167404896817093.

Lindup:1996:RIS

[Lin96] Ken Lindup. The role of information security in corporate governance. Computers & Security, 15(6):477–485, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404897831215.

Lindup:1997:USP

[Lin97] Ken Lindup. Update on the state of play in the hacker community. Computers & Security, 16(6):521, ???? 1997. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404897846779.

Laribi:1992:PMI

[LK92] Atika Laribi and Dennis Kafura. A protection model incorporating both authorization and constraints. Computers & Security, 11(1):57–73, March 1992. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/016740489290221C.

Lin:1998:CDE

[LL98]  Chu-Hsing Lin and Tien-Chi Lee.  A confused docu-

ment encrypting scheme and its implementation. Computers & Security, 17(6):543–551, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404898800941.

Lloyd:1990:SEH

[Llo90] Ivor Lloyd. Software evaluation in high integrity systems. Computers & Security, 9(5):419–429, August 1990. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/0167404890900733.

Lo:1995:MMC

[LLO95] Raymond W. Lo, Karl N. Levitt, and Ronald A. Olsson. MCF: a malicious code filter. Computers & Security, 14(6):541–566, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/016740489500012W.

Longley:1992:ASS

[LR92] D. Longley and S. Rigby. An automatic search for security flaws in key management schemes. Computers & Security, 11(1):75– 89, March 1992. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/016740489290222D.

Labuschagne:1993:DSX

[LR93] Frans Labuschagne and Jan Roos. Data security in X.25 networks. Computers & Security, 12(5):469–475, August 1993. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/016740489390068G.

Longstaff:1993:BPA

[LS93] Thomas A. Longstaff and E. Eugene Schultz. Beyond preliminary analysis of the WANK and OILZ worms: a case study of malicious code. Computers & Security, 12(1):61–77, February 1993. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/016740489390013U.

Ladkin:1994:CPV

[LT94] Peter Ladkin and Harold Thimbleby. Comments on a paper by voas, Payne and Cohen: ‘a model for detecting the existence of software corruption in real time’. Computers & Security, 13(6):527–531, ???? 1994. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/0167404891901409.

Lunt:1992:AAT

[Lun92a] Teresa Lunt. Automated audit trail analysis for intrusion detection. Computers & Security, 11(5):492, September 1992. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/016740489290256Q.

Lunt:1992:SDS

[Lun92b] Teresa F. Lunt. Security in database systems: a research perspective. Computers & Security, 11(1):41–56, March 1992.

CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/016740489290220L.

Lunt:1993:SID

[Lun93] Teresa F. Lunt. A survey of intrusion detection techniques. Computers & Security, 12(4):405–418, June 1993. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/0167404893900295.

Liepens:1992:IDR

[LV92] G. E. Liepens and H. S. Vaccaro. Intrusion detection: Its role and validation. Computers & Security, 11(4):347–355, July 1992. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/016740489290175Q.

Longley:1994:EKG

[LV94] D. Longley and S. Vasudevan. Effect of key generators on the automatic search for flaws in key management schemes. Computers & Security, 13(4):335–347, ???? 1994. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/0167404894900248.

Lynch:1991:BRE

[Lyn91] Paul Lynch. Beyond RACF: Extending user authentication controls. Computers & Security, 10(8):711–722, December 1991. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/016740489190090Z.

Madsen:1992:GES

[Mad92a] Wayne Madsen. Government eavesdropping-should consumers pay? Computers & Security, 11(5):492, September 1992. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/016740489290258S.

Madsen:1992:GSC

[Mad92b] Wayne Madsen. Government-sponsored computer warfare and sabotage. Computers & Security, 11(3):233–236, May 1992.

CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/0167404892902012.

Madsen:1995:SAP

[Mad95] Wayne Madsen. Securing access and privacy on the Internet. Computers & Security, 14(5):418, ???? 1995. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/016740489597110V.

Maddison:1998:YEB

[Mad98] Nicky Maddison. Y2000 enters the body politic. Computers & Security, 17(3):201–206, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404898803106.

Magruder:1994:HLL

[Mag94] Scott Magruder. High-level language computer viruses — a new threat? Computers & Security, 13(3):263–269, May 1994. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/0167404894900809.

Manber:1996:SSM

[Man96] Udi Manber. A simple scheme to make passwords based on one-way functions much harder to crack. Computers & Security, 15(2):171–176, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/016740489600003X.

McClure:1997:ABM

[McC97a] Ben McClure. Area banks mount elaborate electronic defences. Computers & Security, 16(8):676, ???? 1997. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404897875864.

McClure:1997:BI

[McC97b] Ben McClure. Bulletproof IP. Computers & Security, 16(8): 679, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404897876018.

McClure:1997:CBL

[McC97c] Ben McClure. Code breakers lured for crypto challenge. Computers & Security, 16(8):680, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404897876043.

McClure:1997:CEI

[McC97d] Ben McClure. Concerns ease over Internet security. Computers & Security, 16(8):675, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404897875840.

McClure:1997:DMB

[McC97e] Ben McClure. Dutch move to block French crypto laws. Computers & Security, 16(8):678, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404897875979.

McClure:1997:EHT

[McC97f] Ben McClure. Encryption: hot topic on the Hill. Computers & Security, 16(8):678, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404897875955.

McClure:1997:ED

[McC97g] Ben McClure. Exorcising demons. Computers & Security, 16 (8):676, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404897875852.

McClure:1997:FCV

[McC97h] Ben McClure. Fighting computer viruses. Computers & Security, 16(8):676–677, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S016740489787589X.

McClure:1997:FHT

[McC97i] Ben McClure. Former high-tech thief warns of scams. Computers & Security, 16(8):675, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404897875839.

McClure:1997:HTC

[McC97j] Ben McClure. High-tech crime’s hidden face. Computers & Security, 16(8):679, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404897875992.

McClure:1997:IRS

[McC97k] Ben McClure. Internet Rx: Should I be checking downloaded Java applets for viruses or other potential security violations? Computers & Security, 16(8):677, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S016740489787592X.

McClure:1997:MAB

[McC97l] Ben McClure. Man accused of being UK’s first Internet pirate. Computers & Security, 16(8):675, ???? 1997. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404897875827.

McClure:1997:MBR

[McC97m] Ben McClure. Microsoft battles with Russian pirates. Computers & Security, 16(8):676, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404897875876.

McClure:1997:NSS

[McC97n] Ben McClure. Network security: safety is next. Computers & Security, 16(8):680–681, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404897876067.

McClure:1997:OLT

[McC97o] Ben McClure. OECD leader targets Internet security accords. Computers & Security, 16(8):679–680, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S016740489787602X.

McClure:1997:PPP

[McC97p] Ben McClure. Paring password pileup. Computers & Security, 16(8):677, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404897875918.

McClure:1997:PP

[McC97q] Ben McClure. Portable protection. Computers & Security, 16 (8):677, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404897875931.

McClure:1997:SEM

[McC97r] Ben McClure. Secure electronic-mail: return to sender. Computers & Security, 16(8):678–679, ???? 1997. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404897875980.

McClure:1997:SGE

[McC97s] Ben McClure. Security gets easier, cheaper. Computers & Security, 16(8):680, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404897876055.

McClure:1997:SOP

[McC97t] Ben McClure. Should only the paranoid get e-mail protection? Computers & Security, 16(8):677, ???? 1997. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404897875906.

McClure:1997:ST

[McC97u] Ben McClure. Stop, thief! Computers & Security, 16(8): 676, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404897875888.

McClure:1997:ANS

[McC97v] Ben McClure. That ain’t no standard spec. Computers & Security, 16(8):678, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404897875967.

McClure:1997:TT

[McC97w] Ben McClure. Trials and tribulations. Computers & Security, 16 (8):680, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404897876031.

McClure:1997:WGT

[McC97x] Ben McClure. Who goes there? Computers & Security, 16(8): 679, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404897876006.

McClure:1997:WCI

[McC97y] Ben McClure. Wireless crime’s inside story. Computers & Security, 16(8):678, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404897875943.

McClure:1998:ECC

[McC98a] Ben McClure. 2 encryption cases cast shadow on academia. Computers & Security, 17(3):227, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404898803337.

McClure:1998:AHN

[McC98b] Ben McClure. The Achilles heel of next-generation satellites. Computers & Security, 17(4):328–329, ???? 1998. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404898800308.

McClure:1998:AMS

[McC98c] Ben McClure. Adoption of S/MIME still lagging. Computers & Security, 17(4):328, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S016740489880028X.

McClure:1998:ATF

[McC98d] Ben McClure. Airports told of flaw in security system. Computers & Security, 17(2):158, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S016740489782010X.

McClure:1998:AMW

[McC98e] Ben McClure. Airtight e-mail is on the way. Computers & Security, 17(3):224, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404898803209.

McClure:1998:AVR

[McC98f] Ben McClure. Anti-virus rivals gang up over critical study. Computers & Security, 17(4):330, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404898800382.

McClure:1998:AUF

[McC98g] Ben McClure. AOL users fall prey to hackers’ dirty tricks. Computers & Security, 17(2):153, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404897819864.

McClure:1998:AC

[McC98h] Ben McClure. Authentication central. Computers & Security, 17(4):326, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404898800175.

McClure:1998:BSI

[McC98i] Ben McClure. Banking on a secure Internet. Computers & Security, 17(2):156, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404897820020.

McClure:1998:CWY

[McC98j] Ben McClure. Be careful what you say on Y2K. Computers & Security, 17(3):222, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404898803143.

McClure:1998:BYF

[McC98k]  Ben McClure. Burned by your firewall?  Comput-

ers & Security, 17(3):224–225, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404898803234.

McClure:1998:CCC

[McC98l] Ben McClure. Calling for a cyber cease-fire. Computers & Security, 17(3):228, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404898803374.

McClure:1998:CSP

[McC98m] Ben McClure. Can security pierce the electronic veil? Computers & Security, 17(2):157, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404897820068.

McClure:1998:C

[McC98n] Ben McClure. Certifiable. Computers & Security, 17(4):328, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404898800254.

McClure:1998:CSS

[McC98o] Ben McClure. Cisco strives to secure the enterprise. Computers & Security, 17(4):328, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404898800291.

McClure:1998:CYA

[McC98p] Ben McClure. Covering your assets, electronically. Computers & Security, 17(3):226, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404898803301.

McClure:1998:CH

[McC98q] Ben McClure. Cry havoc. Computers & Security, 17(3):227, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404898803350.

McClure:1998:CIW

[McC98r] Ben McClure. Cyber-IDs will soon be at your fingertips. Computers & Security, 17(2):158, ???? 1998. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404897820093.

McClure:1998:EDH

[McC98s] Ben McClure. Encryption debate heats up in Washington. Computers & Security, 17(4):329, ???? 1998. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404898800333.

McClure:1998:FSL

[McC98t] Ben McClure. Finjan SurfinGate: the lifeguard is on duty. Computers & Security, 17(3):222–223, ???? 1998. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404898803155.

McClure:1998:FBF

[McC98u] Ben McClure. The firewall behind the firewall. Computers & Security, 17(4):325, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404898800138.

McClure:1998:FEH

[McC98v] Ben McClure. Firms exposed to hackers. Computers & Security, 17(4):329–330, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404898800357.

McClure:1998:FFD

[McC98w] Ben McClure. Firms to be forced to disclose year 2000 risks. Computers & Security, 17(2):154, ???? 1998. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404897819918.

McClure:1998:FCH

[McC98x] Ben McClure. Forensic computing helps firms track hightech fraud. Computers & Security, 17(3):223–224, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404898803192.

McClure:1998:TRS

[McC98y] Ben McClure. From toy rings to sophisticated codes. Computers & Security, 17(4):330, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404898800370.

McClure:1998:GN

[McC98z] Ben McClure. Good news. Computers & Security, 17(4): 329, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S016740489880031X.

McClure:1998:GSI

[McC98-27] Ben McClure. GTE security for any ISP. Computers & Security, 17(4):327, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404898800229.

McClure:1998:IUS

[McC98-28] Ben McClure. IBM unveils secure e-business products. Computers & Security, 17(4):329, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404898800345.

McClure:1998:ICS

[McC98-29] Ben McClure. Immigration computer system less than secure. Computers & Security, 17(2):154, ???? 1998. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404897819931.

McClure:1998:IBC

[McC98-30] Ben McClure. Internet boosts cryptography. Computers & Security, 17(4):327, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404898800230.

McClure:1998:ISH

[McC98-31] Ben McClure. Internet security holes widen. Computers & Security, 17(4):328, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404898800278.

McClure:1998:ISB

[McC98-32] Ben McClure. Internet security on a budget. Computers & Security, 17(2):155–156, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404897819992.

McClure:1998:IST

[McC98-33] Ben McClure. Internet security: trouble in the stack, rik. Computers & Security, 17(4):327, ???? 1998. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404898800217.

McClure:1998:IVP

[McC98-34] Ben McClure. Internet vandals pose threat by using new attack mode called ‘smurfing’. Computers & Security, 17(2): 154, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S016740489781989X.

McClure:1998:ICI

[McC98-35] Ben McClure. IPSec for communities of interest. Computers & Security, 17(3):225–226, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404898803283.

McClure:1998:JKG

[McC98-36] Ben McClure. Java kit gets better security. Computers & Security, 17(2):158, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404897820081.

McClure:1998:KFT

[McC98-37] Ben McClure. Kicking firewall tires. Computers & Security, 17 (3):223, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404898803167.

McClure:1998:LAC

[McC98-38] Ben McClure. Labour acts to clean up digital certificate mess. Computers & Security, 17(2):157, ???? 1998. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404897820044.

McClure:1998:MHR

[McC98-39] Ben McClure. Major hacks raise hackles, spur defenders. Computers & Security, 17(3):225, ???? 1998. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404898803271.

McClure:1998:MKE

[McC98-40] Ben McClure. Managers keep an eye on security. Computers & Security, 17(3):223, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404898803180.

McClure:1998:MMS

[McC98-41] Ben McClure. MasterCard makes SET more attractive. Computers & Security, 17(2):158, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404897820111.

McClure:1998:MCS

[McC98-42] Ben McClure. Mergers cut security options. Computers & Security, 17(4):327, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404898800242.

McClure:1998:NAB

[McC98-43] Ben McClure. Network Associates to buy Dr Solomon’s for $640 million. Computers & Security, 17(4):325, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S016740489880014X.

McClure:1998:NAP

[McC98-44] Ben McClure. New antivirus packages get serious. Computers & Security, 17(2):156, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404897820019.

McClure:1998:NSC

[McC98-45] Ben McClure. No SET commitment, yet. Computers & Security, 17(2):155, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404897819955.

McClure:1998:NFT

[McC98-46] Ben McClure. NT firewalls: tough enough. Computers & Security, 17(3):227–228, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404898803362.

McClure:1998:OSL

[McC98-47] Ben McClure. Online security leak. Computers & Security, 17 (2):155, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404897819979.

McClure:1998:PMC

[McC98-48] Ben McClure. PC manager at center of $2m grocery scam. Computers & Security, 17(3):225, ???? 1998. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404898803258.

McClure:1998:PG

[McC98-49] Ben McClure. PGP grows up. Computers & Security, 17(3): 223, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404898803179.

McClure:1998:PSH

[McC98-50] Ben McClure. Plugging security holes. Computers & Security, 17(2):154–155, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404897819943.

McClure:1998:PNE

[McC98-51] Ben McClure. Powerful new encryption standard delayed by weakness. Computers & Security, 17(4):330, ???? 1998. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404898800369.

McClure:1998:PFF

[McC98-52] Ben McClure. Putting firewall in firmware. Computers & Security, 17(4):328, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404898800266.

McClure:1998:RSS

[McC98-53] Ben McClure. RCMP seizes $90 000 in software bust. Computers & Security, 17(4):330, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404898800394.

McClure:1998:RSC

[McC98-54] Ben McClure. Reno to set up center to thwart ‘cyberattackers’. Computers & Security, 17(2):158–159, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404897820123.

McClure:1998:RFF

[McC98-55] Ben McClure. Report faults FAA on year 2000 hiccup. Computers & Security, 17(2):153–154, ???? 1998. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404897819888.

McClure:1998:RPY

[McC98-56] Ben McClure. A rescue plan for your LAN. Computers & Security, 17(4):326, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404898800163.

McClure:1998:RST

[McC98-57] Ben McClure. Reuters subsidiary target of US inquiry into theft of data from Bloomberg. Computers & Security, 17(2): 157, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404897820056.

McClure:1998:RM

[McC98-58] Ben McClure. The revenue men. Computers & Security, 17(3): 226, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404898803295.

McClure:1998:RHE

[McC98-59] Ben McClure. Russian hacker enters fraud plea in Citicorp case. Computers & Security, 17(2):154, ???? 1998. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S016740489781992X.

McClure:1998:SN

[McC98-60] Ben McClure. Safety in numbers? Computers & Security, 17 (2):159, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404897820135.

McClure:1998:SC

[McC98-61] Ben McClure. Security and culture. Computers & Security, 17 (2):153, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404897819852.

McClure:1998:SAI

[McC98-62] Ben McClure. A security architecture for the Internet protocol. Computers & Security, 17(2):155, ???? 1998. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404897819980.

McClure:1998:SCH

[McC98-63] Ben McClure. Security companies hype up Java risks. Computers & Security, 17(3):226, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404898803313.

McClure:1998:SGM

[McC98-64] Ben McClure. Security giant moves to plug leaks in PC safety. Computers & Security, 17(2):153, ???? 1998. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404897819876.

McClure:1998:SS

[McC98-65] Ben McClure. Security services. Computers & Security, 17(3): 224, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404898803210.

McClure:1998:STS

[McC98-66] Ben McClure. Security tools, specs offer more protection. Computers & Security, 17(3):226–227, ???? 1998. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404898803325.

McClure:1998:SPS

[McC98-67] Ben McClure. Senate probes state department security. Computers & Security, 17(3):225, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404898803246.

McClure:1998:SDC

[McC98-68] Ben McClure. Serving up digital certificates. Computers & Security, 17(4):326–327, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404898800199.

McClure:1998:SPL

[McC98-69] Ben McClure. Signs point to looser encryption rules. Computers & Security, 17(3):225, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S016740489880326X.

McClure:1998:S

[McC98-70] Ben McClure. Smartcards. Computers & Security, 17(4):331, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404898800412.

McClure:1998:MOV

[McC98-71] Ben McClure. S/MIME and OpenPGP vie for security title. Computers & Security, 17(4):325–326, ???? 1998. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404898800151.

McClure:1998:SSA

[McC98-72] Ben McClure. Study: security audits often neglected by many. Computers & Security, 17(2):155, ???? 1998. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404897819967.

McClure:1998:TCD

[McC98-73] Ben McClure. Taking the confusion out of digital certificates. Computers & Security, 17(4):330–331, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404898800400.

McClure:1998:TN

[McC98-74] Ben McClure. Target: NT. Computers & Security, 17(2): 156, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404897820007.

McClure:1998:TMJ

[McC98-75] Ben McClure. Thwarting malicious Java attacks. Computers & Security, 17(4):326, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404898800187.

McClure:1998:UTN

[McC98-76] Ben McClure. U-turn on net trade security. Computers & Security, 17(3):222, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404898803131.

McClure:1998:ULB

[McC98-77] Ben McClure. US losing battle on control of data encryption, study says. Computers & Security, 17(2):154, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404897819906.

McClure:1998:URN

[McC98-78] Ben McClure. US rethinks net security plans after hijack plot. Computers & Security, 17(2):156–157, ???? 1998. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404897820032.

McClure:1998:VSV

[McC98-79] Ben McClure. Verisign secures VPN links. Computers & Security, 17(4):329, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404898800321.

McClure:1998:VPS

[McC98-80] Ben McClure. Virus protection software. Computers & Security, 17(3):227, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404898803349.

McClure:1998:WPB

[McC98-81] Ben McClure. Want to prevent break ins? Computers & Security, 17(4):327, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404898800205.

McClure:1998:WS

[McC98-82] Ben McClure. Weak SNMP? Computers & Security, 17(2): 157, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S016740489782007X.

McClure:1998:WNC

[McC98-83] Ben McClure. Windows NT casts a shaky Web. Computers & Security, 17(3):224, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404898803222.

McLaughlin:1992:YAM

[McL92] Robert McLaughlin. Yet another machine to break DES. Computers & Security, 11(5):492, September 1992. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/016740489290259T.

McNulty:1995:CGI

[McN95a] Lynn McNulty. Cryptography and the global information infrastructure. Computers & Security, 14(5):420, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/016740489597123R.

McNulty:1995:UAC

[McN95b] Lynn McNulty. US approaches to cryptographic key escrowing (session presentation). Computers & Security, 14(5):426, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/0167404896900504.

Mohammed:1994:DDU

[MD94] Imtiaz Mohammed and David M. Dilts. Design for dynamic user-role-based security. Computers & Security, 13(8):661–671, ???? 1994. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/0167404894900485.

Menkus:1990:CLH

[Men90a] Belden Menkus. The complexity of “long haul” data communication security. Computers & Security, 9(7):587–592, November 1990. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/016740489090055X.

Menkus:1990:HAT

[Men90b] Belden Menkus. How an “audit trail” aids in maintaining information integrity … as illustrated in retailing. Computers & Security, 9(2):111–116, April 1990. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/0167404890900825.

Menkus:1990:UDC

[Men90c] Belden Menkus. Understanding data communication security vulnerabilities. Computers & Security, 9(3):209–213, May 1990. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/016740489090165P.

Menkus:1990:WDC

[Men90d] Belden Menkus. Why data communications are insecure. Computers & Security, 9(6):483–487, October 1990. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/016740489090123B.

Menkus:1991:CFS

[Men91a] Belden Menkus. “control” is fundamental to successful information security. Computers & Security, 10(4):293–297, June 1991. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/016740489190103K.

Menkus:1991:HKA

[Men91b] Belden Menkus. “Hackers”: Know the adversary. Computers & Security, 10(5):405–409, August 1991. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/016740489190066M.

Menkus:1991:HBD

[Men91c] Belden Menkus. How to begin dealing with computer security. Computers & Security, 10(3):199–203, May 1991. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/016740489190034B.

Menkus:1992:CCS

[Men92a] Belden Menkus. Concerns in computer security. Computers & Security, 11(3):211–215, May 1992. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/016740489290198Z.

Menkus:1992:HRB

[Men92b] Belden Menkus. A high rise building fire case study. Computers & Security, 11(1):19–23, March 1992. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/016740489290215D.

Menkus:1992:ICS

[Men92c] Belden Menkus. Introduction to computer security. Computers & Security, 11(2):121–127, April 1992. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/016740489290036Q.

Menkus:1992:LGC

[Men92d] Belden Menkus. The lessons of the great Chicago flood of 1992. Computers & Security, 11(5):417–420, September 1992. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/016740489290005C.

Menkus:1992:RED

[Men92e] Belden Menkus. Re-examining the data encryption standard. Computers & Security, 11(8):703–706, December 1992. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/016740489290089A.

Menkus:1992:UES

[Men92f] Belden Menkus. Understanding EDI security issues. Computers & Security, 11(6):525–528, October 1992. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/016740489290185T.

Menkus:1994:NIB

[Men94] Belden Menkus. The new importance of “business continuity” in data processing disaster recovery planning. Computers & Security, 13(2):115–118, April 1994. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/0167404894900574.

Mercer:1990:FDR

[Mer90a] Lindsay C. J. Mercer. Fraud detection via regression analysis. Computers & Security, 9(4):331–338, June 1990. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/016740489090103Z.

Mercer:1990:TMA

[Mer90b] Lindsay C. J. Mercer. Tailor-made auditing of information systems for the detection of fraud. Computers & Security, 9(1):59– 66, February 1990. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/016740489090158P.

Messmer:1993:ARA

[Mes93] Ellen Messmer. Abstracts of recent articles and literature. Computers & Security, 12(4):365–370, June 1993. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/016740489390023X.

Meyer:1993:IKB

[Mey93] John Meyer. The 1993 IFIP Kristian Beckman Award. Computers & Security, 12(4):338–340, June 1993. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/0167404893900206.

Meyer:1995:CSI

[Mey95] Carl Meyer. Crypto system initialization: Simplifying the distribution of initial keys. Computers & Security, 14(5):419, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/016740489597120Y.

Meyer:1996:AAA

[Mey96a] Helen Meyer. Access all areas. Computers & Security, 15 (3):221–222, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S016740489690214X.

Meyer:1996:AF

[Mey96b] Helen Meyer. Antivirus firewalls. Computers & Security, 15(7): 586, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404897881175.

Meyer:1996:AHW

[Mey96c] Helen Meyer. Attack highlights Web security risks. Computers & Security, 15(6):517, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404897831331.

Meyer:1996:BRA

[Mey96d] Helen Meyer. Bank restless with app-level security. Computers & Security, 15(3):221, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404896902126.

Meyer:1996:BWM

[Mey96e] Helen Meyer. Beware the word macro virus. Computers & Security, 15(8):689–690, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404897897209.

Meyer:1996:BBP

[Mey96f] Helen Meyer. Biology battles PC viruses. Computers & Security, 15(3):222–223, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404896902229.

Meyer:1996:BAE

[Mey96g] Helen Meyer. A business approach to effective information technology risk analysis and management. Computers & Security, 15(4):321, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404896889619.

Meyer:1996:CMC

[Mey96h] Helen Meyer. Cheyenne, McAfee cure software viruses. Computers & Security, 15(7):587–588, ???? 1996. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404897881229.

Meyer:1996:CSI

[Mey96i] Helen Meyer. Commercial security on the Internet. Computers & Security, 15(4):321, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404896889607.

Meyer:1996:CER

[Mey96j] Helen Meyer. A computer emergency response team policy. Computers & Security, 15(4):320, ???? 1996. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/0167404896889565.

Meyer:1996:CHT

[Mey96k] Helen Meyer. Computer hackers turn to pager systems. Computers & Security, 15(7):585–586, ???? 1996. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404897881163.

Meyer:1996:CVM

[Mey96l] Helen Meyer. Computer viruses: myth vs reality. Computers & Security, 15(7):585, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S016740489788114X.

Meyer:1996:CDG

[Mey96m] Helen Meyer. Constructing difficult-to-guess passwords. Computers & Security, 15(4):321, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/0167404896889589.

Meyer:1996:CA

[Mey96n] Helen Meyer. Controlling access. Computers & Security, 15(7): 588, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404897881242.

Meyer:1996:CD

[Mey96o] Helen Meyer. Cope with disaster. Computers & Security, 15(6): 519, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404897831410.

Meyer:1996:CBS

[Mey96p] Helen Meyer. Corporates blasé over software theft. Computers & Security, 15(8):685, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404897897027.

Meyer:1996:CCW

[Mey96q] Helen Meyer. Cracking the complex world of security. Computers & Security, 15(3):224, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404896903077.

Meyer:1996:CYN

[Mey96r] Helen Meyer. Curing your NT security headache. Computers & Security, 15(8):689, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404897897179.

Meyer:1996:CTN

[Mey96s] Helen Meyer. Cybercop takes on network snoops. Computers & Security, 15(8):685, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404897897015.

Meyer:1996:CCC

[Mey96t] Helen Meyer. Cypherwiz cracks cybersecurity. Computers & Security, 15(3):221, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404896897859.

Meyer:1996:DSR

[Mey96u] Helen Meyer. Data security report. Computers & Security, 15 (4):320, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/0167404896889577.

Meyer:1996:DEP

[Mey96v] Helen Meyer. Decoding encryption policy. Computers & Security, 15(3):225, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404896903107.

Meyer:1996:DSF

[Mey96w] Helen Meyer. Denial of service, the future of dirty tricks. Computers & Security, 15(7):586, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404897881187.

Meyer:1996:DTC

[Mey96x] Helen Meyer. Desperate times call for desperate measures. Computers & Security, 15(3):222, ???? 1996. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404896902205.

Meyer:1996:DCC

[Mey96y] Helen Meyer. Developers can control Java access with firewall. Computers & Security, 15(8):688, ???? 1996. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404897897143.

Meyer:1996:DWA

[Mey96z] Helen Meyer. Developers worried about safety of Java and ActiveX. Computers & Security, 15(8):687, ???? 1996. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404897897106.

Meyer:1996:DSH

[Mey96-27] Helen Meyer. Digital signatures: how they work. Computers & Security, 15(3):223, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404896903041.

Meyer:1996:FLW

[Mey96-28] Helen Meyer. Fear and loathing on the world wide Web. Computers & Security, 15(4):319, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/0167404896902175.

Meyer:1996:FMS

[Mey96-29] Helen Meyer. Feeling more secure about intranets. Computers & Security, 15(6):516–517, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S016740489783132X.

Meyer:1996:FF

[Mey96-30] Helen Meyer. Firestorm over firewalls. Computers & Security, 15(8):689, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404897897180.

Meyer:1996:FFI

[Mey96-31] Helen Meyer. Firewall fights intranet threat. Computers & Security, 15(6):519, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404897831422.

Meyer:1996:F

[Mey96-32] Helen Meyer. Firewalls. Computers & Security, 15(3):225– 226, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404896903120.

Meyer:1996:FL

[Mey96-33] Helen Meyer. Firewalls light up. Computers & Security, 15(3): 222, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404896902163.

Meyer:1996:FTC

[Mey96-34] Helen Meyer. Firm turns China piracy into contract. Computers & Security, 15(6):515, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/0167404896901078.

Meyer:1996:GPN

[Mey96-35] Helen Meyer. Getting personal on the net. Computers & Security, 15(6):517, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404897831355.

Meyer:1996:GRA

[Mey96-36] Helen Meyer. Getting real about virtual private data nets. Computers & Security, 15(7):588–589, ???? 1996. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404897881254.

Meyer:1996:HC

[Mey96-37] Helen Meyer. Hacker with a cause? Computers & Security, 15 (6):518, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404897831379.

Meyer:1996:HH

[Mey96-38] Helen Meyer. Hazards of hooking up. Computers & Security, 15 (4):321, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/0167404896889620.

Meyer:1996:HSY

[Mey96-39] Helen Meyer. How safe is your LAN? Computers & Security, 15 (4):321, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/0167404896889632.

Meyer:1996:HPY

[Mey96-40] Helen Meyer. How to protect your data. Computers & Security, 15(7):586–587, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404897881199.

Meyer:1996:T

[Mey96-41] Helen Meyer. The I-team. Computers & Security, 15(6):519– 520, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404897831434.

Meyer:1996:NSW

[Mey96-42] Helen Meyer. In the ’net shall we trust. Computers & Security, 15(4):318, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404896889486.

Meyer:1996:ISB

[Mey96-43] Helen Meyer. Information security in business environments. Computers & Security, 15(4):321–322, ???? 1996. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/0167404896889644.

Meyer:1996:IS

[Mey96-44] Helen Meyer. Inner security. Computers & Security, 15(4):319– 320, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/0167404896889528.

Meyer:1996:IPR

[Mey96-45] Helen Meyer. Internet poses risk, CIBC head says. Computers & Security, 15(6):516, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404897831288.

Meyer:1996:ISH

[Mey96-46] Helen Meyer. Internet security: how much is enough? Computers & Security, 15(4):322, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/0167404896889656.

Meyer:1996:IUS

[Mey96-47] Helen Meyer. Internet users sued for software piracy. Computers & Security, 15(8):688, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404897897155.

Meyer:1996:IBI

[Mey96-48] Helen Meyer. InterScan blocks Internet viruses. Computers & Security, 15(3):224–225, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404896903089.

Meyer:1996:IMV

[Mey96-49] Helen Meyer. InterScan E-Mail Viruswall 1.5. Computers & Security, 15(8):686, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404897897064.

Meyer:1996:IFO

[Mey96-50] Helen Meyer. Intranet firewalls offer no protection against the enemy within. Computers & Security, 15(8):688–689, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404897897167.

Meyer:1996:KYN

[Mey96-51] Helen Meyer. Keep your network virus-free. Computers & Security, 15(3):224, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404896903065.

Meyer:1996:LMV

[Mey96-52] Helen Meyer. Laroux macro virus strain hits Excel. Computers & Security, 15(8):686, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404897897040.

Meyer:1996:MIA

[Mey96-53] Helen Meyer. Making Internet access safe. Computers & Security, 15(7):587, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404897881205.

Meyer:1996:MIS

[Mey96-54] Helen Meyer. Making the Internet safe for e-commerce. Computers & Security, 15(7):587, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404897881217.

Meyer:1996:MGT

[Mey96-55] Helen Meyer. Managers get third access security spec. Computers & Security, 15(6):520, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404897831446.

Meyer:1996:MST

[Mey96-56] Helen Meyer. Mobile secure telephones. Computers & Security, 15(4):320, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404896889541.

Meyer:1996:MTS

[Mey96-57] Helen Meyer. MOD tooled up on security issue. Computers & Security, 15(8):685–686, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404897897039.

Meyer:1996:MSE

[Mey96-58] Helen Meyer. More secure electronic payment. Computers & Security, 15(8):689, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404897897192.

Meyer:1996:NPP

[Mey96-59] Helen Meyer. Net porn prosecutions would be a ‘nightmare’. Computers & Security, 15(5):397, ???? 1996. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/0167404896825776.

Meyer:1996:NSS

[Mey96-60] Helen Meyer. Net server to secure smaller offices. Computers & Security, 15(4):318, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404896889462.

Meyer:1996:NUE

[Mey96-61] Helen Meyer. The new US encryption policy. Computers & Security, 15(7):585, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404897881151.

Meyer:1996:OSP

[Mey96-62] Helen Meyer. Online sales pose security concerns. Computers & Security, 15(8):687, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404897897088.

Meyer:1996:PP

[Mey96-63] Helen Meyer. Photon protection. Computers & Security, 15(8): 687, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S016740489789709X.

Meyer:1996:PRR

[Mey96-64] Helen Meyer. Police ruling reveals CMA weakness. Computers & Security, 15(8):686, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404897897052.

Meyer:1996:PIS

[Mey96-65] Helen Meyer. Privacy issues stir online passions. Computers & Security, 15(8):687–688, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S016740489789712X.

Meyer:1996:PMT

[Mey96-66] Helen Meyer. A proposed mode for triple-DES encryption. Computers & Security, 15(4):322, ???? 1996. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/0167404896889668.

Meyer:1996:PYL

[Mey96-67] Helen Meyer. Protecting your link to the net. Computers & Security, 15(4):321, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404896889590.

Meyer:1996:PW

[Mey96-68] Helen Meyer. Protection on the web. Computers & Security, 15 (4):319, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/0167404896889516.

Meyer:1996:PSU

[Mey96-69] Helen Meyer. Providing security for Unix networks. Computers & Security, 15(4):320, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/016740489688953X.

Meyer:1996:PSD

[Mey96-70] Helen Meyer. Pstt! Security designed for your eyes only. Computers & Security, 15(4):317–318, ???? 1996. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/0167404896889449.

Meyer:1996:PBY

[Mey96-71] Helen Meyer. Puffer 2.0 buys you some e-mail security via easy encryption. Computers & Security, 15(4):318–319, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/0167404896889498.

Meyer:1996:RS

[Mey96-72] Helen Meyer. Real security. Computers & Security, 15(3):223– 224, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404896903053.

Meyer:1996:SSF

[Mey96-73]  Helen Meyer.  Seattle software’s firewall keeps watch.

Computers & Security, 15(6):517, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404897831343.

Meyer:1996:SYM

[Mey96-74] Helen Meyer. Secure your Mac. Computers & Security, 15(7): 590, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404897881308.

Meyer:1996:SB

[Mey96-75] Helen Meyer. Security blanket. Computers & Security, 15(6): 518, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404897831380.

Meyer:1996:SCP

[Mey96-76] Helen Meyer. Security for cellular phones on the way. Computers & Security, 15(6):515, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404897831264.

Meyer:1996:SGU

[Mey96-77] Helen Meyer. Security gap uncovered in ActiveX components. Computers & Security, 15(6):516, ???? 1996. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404897831306.

Meyer:1996:SIL

[Mey96-78] Helen Meyer. Security issues lurk behind bank doors. Computers & Security, 15(4):319, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/0167404896889504.

Meyer:1996:SRS

[Mey96-79] Helen Meyer. Security risk or security solution? Computers & Security, 15(4):317, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404896889437.

Meyer:1996:SS

[Mey96-80] Helen Meyer. Selling security. Computers & Security, 15(6): 519, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404897831409.

Meyer:1996:SGS

[Mey96-81] Helen Meyer. Servers get single sign-on. Computers & Security, 15(4):318, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404896889450.

Meyer:1996:SSP

[Mey96-82] Helen Meyer. Setting up a security perimeter for distributed networks. Computers & Security, 15(3):223, ???? 1996. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404896902242.

Meyer:1996:SIC

[Mey96-83] Helen Meyer. Shoot-out at the IP corral. Computers & Security, 15(8):690, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404897897210.

Meyer:1996:SSI

[Mey96-84] Helen Meyer. Strategic security for IP networks. Computers & Security, 15(7):590, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404897881291.

Meyer:1996:SJR

[Mey96-85]  Helen Meyer.  Striking just the right firewall match.

Computers & Security, 15(7):589, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404897881278.

Meyer:1996:TFP

[Mey96-86] Helen Meyer. Telephone fraud poses threat to UK business. Computers & Security, 15(8):686–687, ???? 1996. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404897897076.

Meyer:1996:TTR

[Mey96-87] Helen Meyer. Teleworking: threats, risks and solutions. Computers & Security, 15(4):320, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/0167404896902199.

Meyer:1996:TBP

[Mey96-88] Helen Meyer. Terrorism bill pits safety vs. privacy. Computers & Security, 15(8):687, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404897897118.

Meyer:1996:TSY

[Mey96-89] Helen Meyer. Tips for safeguarding your digital assets. Computers & Security, 15(7):588, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404897881230.

Meyer:1996:TFH

[Mey96-90] Helen Meyer. Tough firewalls help users avoid getting burned. Computers & Security, 15(4):320, ???? 1996. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/0167404896889553.

Meyer:1996:TAS

[Mey96-91] Helen Meyer. Transarc addresses’net security concerns. Computers & Security, 15(8):688, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404897897131.

Meyer:1996:ULK

[Mey96-92] Helen Meyer. Under lock and key. Computers & Security, 15(6):517–518, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404897831367.

Meyer:1996:UAP

[Mey96-93] Helen Meyer. US assailed on plans for encryption export. Computers & Security, 15(6):515, ???? 1996. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404897831276.

Meyer:1996:UFS

[Mey96-94] Helen Meyer. Users fall short on ’net security planning. Computers & Security, 15(6):518–519, ???? 1996. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404897831392.

Meyer:1996:VSB

[Mey96-95] Helen Meyer. Viruses: sucker bait. Computers & Security, 15(7):589–590, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S016740489788128X.

Meyer:1996:WDF

[Mey96-96] Helen Meyer. Web developers face security quandary. Computers & Security, 15(4):317, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/0167404896889425.

Meyer:1996:WSK

[Mey96-97] Helen Meyer. Web security is key to HP’s ’net strategy. Computers & Security, 15(6):516, ???? 1996. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S016740489783129X.

Meyer:1996:WSS

[Mey96-98] Helen Meyer. Web server: security lockdown. Computers & Security, 15(7):589, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404897881266.

Meyer:1996:WHL

[Mey96-99] Helen Meyer. White house launches cybershield. Computers & Security, 15(4):318, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404896889474.

Meyer:1996:WSH

[Mey96-100] Helen Meyer. Will spies hold your keys. Computers & Security, 15(3):225, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404896903119.

Meyer:1996:WAU

[Mey96-101] Helen Meyer. Windows 95 antivirus utilities: the latest in PC protection. Computers & Security, 15(3):222, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404896902187.

Meyer:1996:WNG

[Mey96-102] Helen Meyer. Windows NT to gain built-in security. Computers & Security, 15(6):516, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404897831318.

Meyer:1996:YDL

[Mey96-103] Helen Meyer. Your defensive line. Computers & Security, 15(3): 225, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404896903090.

Meyer:1997:YHI

[Mey97a] Carl Meyer. 20 years of DES — how it was designed. Computers & Security, 16(6):518, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S016740489784672X.

Meyer:1997:UTO

[Mey97b] Carl Meyer. Update on triple DES operations and crypto system initialization. Computers & Security, 16(6):518, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404897846706.

Meyer:1997:A

[Mey97c] Helen Meyer. Analysis. Computers & Security, 16(3):216, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404897845518.

Meyer:1997:AVS

[Mey97d] Helen Meyer. Anti-virus software gets shot in the arm. Computers & Security, 16(2):129, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404897882983.

Meyer:1997:AD

[Mey97e] Helen Meyer. Averting disaster. Computers & Security, 16(4): 332, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404897802088.

Meyer:1997:BCN

[Mey97f] Helen Meyer. Backdoor closes, for now, on crypto plan. Computers & Security, 16(7):619, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404897807997.

Meyer:1997:BFN

[Mey97g] Helen Meyer. Banks fear net losses. Computers & Security, 16(3):216–217, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404897845531.

Meyer:1997:BMA

[Mey97h] Helen Meyer. The big macro attack. Computers & Security, 16 (4):334, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S016740489780212X.

Meyer:1997:BBB

[Mey97i] Helen Meyer. Bridging the business-to-business authentication gap. Computers & Security, 16(6):512, ???? 1997. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404897846597.

Meyer:1997:BBR

[Mey97j] Helen Meyer. Browser bugs raise security concerns. Computers & Security, 16(2):127–128, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404897882934.

Meyer:1997:BW

[Mey97k] Helen Meyer. Bugs in the web. Computers & Security, 16 (4):332–333, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S016740489780209X.

Meyer:1997:CSB

[Mey97l] Helen Meyer. Certificates still being tuned. Computers & Security, 16(4):333, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404897802118.

Meyer:1997:CKK

[Mey97m] Helen Meyer. The cherished keys to the kingdom. Computers & Security, 16(6):511, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404897846561.

Meyer:1997:CYC

[Mey97n] Helen Meyer. Cleansing your computer’s palate. Computers & Security, 16(4):333, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404897802106.

Meyer:1997:CDB

[Mey97o] Helen Meyer. Computer doctor busy fighting ‘viruses’. Computers & Security, 16(6):511, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404897846573.

Meyer:1997:CF

[Mey97p] Helen Meyer. Confessing to the FBI. Computers & Security, 16 (3):213, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404897845385.

Meyer:1997:CK

[Mey97q] Helen Meyer. Crypt keeper. Computers & Security, 16(4): 330, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404897802003.

Meyer:1997:DNI

[Mey97r] Helen Meyer. Detecting network intruders. Computers & Security, 16(7):622, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404897808103.

Meyer:1997:DYK

[Mey97s] Helen Meyer. Do you know the security code? Computers & Security, 16(2):129, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404897882995.

Meyer:1997:DSI

[Mey97t] Helen Meyer. Driving safely on the information highway. Computers & Security, 16(3):214–215, ???? 1997. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404897845452.

Meyer:1997:MAC

[Mey97u] Helen Meyer. E-mail attacks can clog systems. Computers & Security, 16(2):129, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404897883009.

Meyer:1997:SSC

[Mey97v] Helen Meyer. E-signatures for the security conscious. Computers & Security, 16(6):509–510, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404897846512.

Meyer:1997:FRT

[Mey97w] Helen Meyer. FAST running out of time. Computers & Security, 16(2):127, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404897882922.

Meyer:1997:FRS

[Mey97x] Helen Meyer. Five rules for selling security. Computers & Security, 16(3):216, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S016740489784552X.

Meyer:1997:FSS

[Mey97y] Helen Meyer. Four steps to security. Computers & Security, 16 (4):330, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404897802015.

Meyer:1997:FNC

[Mey97z] Helen Meyer. Future of net cracked, flawed. Computers & Security, 16(6):509, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404897846494.

Meyer:1997:HFE

[Mey97-27] Helen Meyer. Hacker FAQ exposes attack strategies. Computers & Security, 16(2):130, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404897883034.

Meyer:1997:HTC

[Mey97-28] Helen Meyer. Hacker tools can give IS a boost. Computers & Security, 16(6):509, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404897846500.

Meyer:1997:HBN

[Mey97-29] Helen Meyer. Hackers blackmail net businesses. Computers & Security, 16(6):508, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404897846457.

Meyer:1997:HBC

[Mey97-30] Helen Meyer. Hacker’s brother changes his grades. Computers & Security, 16(3):214, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404897845440.

Meyer:1997:HHA

[Mey97-31] Helen Meyer. Hackers hit AOL cybervirgins. Computers & Security, 16(3):213–214, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404897845403.

Meyer:1997:HC

[Mey97-32] Helen Meyer. Handle with care. Computers & Security, 16(6): 510, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404897846524.

Meyer:1997:HFT

[Mey97-33] Helen Meyer. A history of firewall technology. Computers & Security, 16(4):331, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404897802052.

Meyer:1997:HCB

[Mey97-34] Helen Meyer. House committee backs Clinton on encryption. Computers & Security, 16(6):508–509, ???? 1997. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404897846482.

Meyer:1997:ITP

[Mey97-35] Helen Meyer. Identity thieves place credit ratings at risk. Computers & Security, 16(3):213, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404897845397.

Meyer:1997:LF

[Mey97-36] Helen Meyer. In the line of fire. Computers & Security, 16(6): 512, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404897846603.

Meyer:1997:IAb

[Mey97-37] Helen Meyer. Internet armour. Computers & Security, 16 (4):331–332, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404897802064.

Meyer:1997:IML

[Mey97-38] Helen Meyer. Internet e-mail looks for security. Computers & Security, 16(3):215, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404897845476.

Meyer:1997:IES

[Mey97-39] Helen Meyer. Internet explorer security flaws. Computers & Security, 16(3):217, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404897845555.

Meyer:1997:IAa

[Mey97-40] Helen Meyer. Intruder alert. Computers & Security, 16(3): 218, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404897845580.

Meyer:1997:YWS

[Mey97-41] Helen Meyer. Is your Web site secure? Computers & Security, 16(6):510, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404897846548.

Meyer:1997:IUS

[Mey97-42] Helen Meyer. ISS unveils 4-step security program. Computers & Security, 16(7):620, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404897808036.

Meyer:1997:LUE

[Mey97-43] Helen Meyer. Labour to unlock encryption plans. Computers & Security, 16(3):214, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404897845415.

Meyer:1997:LPW

[Mey97-44] Helen Meyer. Legal peril on the web. Computers & Security, 16(4):330–331, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404897802039.

Meyer:1997:LI

[Mey97-45] Helen Meyer. Lock IT up. Computers & Security, 16(2):130– 131, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404897883046.

Meyer:1997:LOC

[Mey97-46] Helen Meyer. Low-overhead cryptography. Computers & Security, 16(7):621, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404897808073.

Meyer:1997:MES

[Mey97-47] Helen Meyer. Microsoft establishes security zones. Computers & Security, 16(3):215, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404897845464.

Meyer:1997:MBH

[Mey97-48] Helen Meyer. Military beefing up its hackers defences. Computers & Security, 16(2):128, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404897882946.

Meyer:1997:MET

[Mey97-49] Helen Meyer. MITI encryption-technology policy criticized. Computers & Security, 16(6):508, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404897846469.

Meyer:1997:MYL

[Mey97-50] Helen Meyer. Monitoring as you like it. Computers & Security, 16(7):623, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404897808140.

Meyer:1997:NAE

[Mey97-51] Helen Meyer. New age of encryption. Computers & Security, 16(4):329–330, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S016740489780199X.

Meyer:1997:NNW

[Mey97-52] Helen Meyer. A new and nasty way to flood networks with spam. Computers & Security, 16(7):622–623, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404897808115.

Meyer:1997:NEC

[Mey97-53] Helen Meyer. A new era in cryptography. Computers & Security, 16(4):329, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404897801988.

Meyer:1997:NTS

[Mey97-54] Helen Meyer. New twists on secure VPNS. Computers & Security, 16(7):619–620, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404897808012.

Meyer:1997:NPC

[Mey97-55] Helen Meyer. No peeking: can web users thwart snoops? Computers & Security, 16(4):329, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404897801976.

Meyer:1997:NGM

[Mey97-56] Helen Meyer. Novell gets $4 million in first fraud upgrade trial. Computers & Security, 16(6):507–508, ???? 1997. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404897846445.

Meyer:1997:NHW

[Mey97-57] Helen Meyer. NT hackers warn of easy password interception. Computers & Security, 16(6):507, ???? 1997. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404897846433.

Meyer:1997:PPS

[Mey97-58] Helen Meyer. Pact promises simplified security. Computers & Security, 16(6):510–511, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S016740489784655X.

Meyer:1997:PFS

[Mey97-59] Helen Meyer. Part of the foundation for secure systems: separation of duties policy. Computers & Security, 16(3):214, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404897845427.

Meyer:1997:PFG

[Mey97-60] Helen Meyer. PGP is first to gain crypto approval. Computers & Security, 16(3):215, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404897845488.

Meyer:1997:PPB

[Mey97-61] Helen Meyer. PKI provides a base for secure transactions. Computers & Security, 16(7):620–621, ???? 1997. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S016740489780805X.

Meyer:1997:PAF

[Mey97-62] Helen Meyer. Plan ahead for firewalls. Computers & Security, 16(2):129–130, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404897883010.

Meyer:1997:PC

[Mey97-63] Helen Meyer. Preparing for cyberwar. Computers & Security, 16 (2):127, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404897882910.

Meyer:1997:PBT

[Mey97-64] Helen Meyer. Privacy better than ‘pretty good’. Computers & Security, 16(7):620, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404897808048.

Meyer:1997:REM

[Mey97-65] Helen Meyer. Router encryption made easy — the hard way. Computers & Security, 16(2):128, ???? 1997. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404897882958.

Meyer:1997:SP

[Mey97-66] Helen Meyer. Save and prosper. Computers & Security, 16 (2):131–132, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404897883071.

Meyer:1997:SMS

[Mey97-67] Helen Meyer. Scrambling e-mail on a server. Computers & Security, 16(4):331, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404897802040.

Meyer:1997:SCN

[Mey97-68] Helen Meyer. Secure communications now or later? Computers & Security, 16(3):217, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404897845567.

Meyer:1997:SEM

[Mey97-69] Helen Meyer. Secure electronic-mail: return to sender. Computers & Security, 16(7):624, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404897808176.

Meyer:1997:SV

[Mey97-70] Helen Meyer. Secure VPNs. Computers & Security, 16(3):217– 218, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404897845579.

Meyer:1997:SID

[Mey97-71] Helen Meyer. Securing intranet data with SSL client certificates. Computers & Security, 16(4):334, ???? 1997. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404897802131.

Meyer:1997:SYU

[Mey97-72] Helen Meyer. Securing your users’ Web browsers. Computers & Security, 16(7):619, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404897808000.

Meyer:1997:SMW

[Mey97-73] Helen Meyer. Self-mutating word viruses create strain. Computers & Security, 16(7):621–622, ???? 1997. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404897808097.

Meyer:1997:SSD

[Mey97-74] Helen Meyer. Servers to spam: drop dead. Computers & Security, 16(7):623, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404897808139.

Meyer:1997:SCD

[Mey97-75] Helen Meyer. Shielding company data from internal threats. Computers & Security, 16(6):508, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404897846470.

Meyer:1997:SDI

[Mey97-76]  Helen Meyer. Sign up for a digital ID here!  Com-

puters & Security, 16(2):128–129, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404897882971.

Meyer:1997:STS

[Mey97-77] Helen Meyer. Singapore tops software piracy list, survey shows. Computers & Security, 16(3):214, ???? 1997. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404897845439.

Meyer:1997:SWA

[Mey97-78] Helen Meyer. Smartcards: world access. Computers & Security, 16(7):623, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404897808127.

Meyer:1997:SAE

[Mey97-79] Helen Meyer. Software AG extends mainframe security. Computers & Security, 16(6):510, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404897846536.

Meyer:1997:SCS

[Mey97-80] Helen Meyer. Software could spoof-proof domain names. Computers & Security, 16(7):621, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404897808061.

Meyer:1997:TNS

[Mey97-81] Helen Meyer. Tackling network security can be an uphill battle. Computers & Security, 16(3):217, ???? 1997. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404897845543.

Meyer:1997:TSI

[Mey97-82] Helen Meyer. Telecoms systems insecurity, exploitation, fraud. Computers & Security, 16(6):511–512, ???? 1997. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404897846585.

Meyer:1997:TMS

[Mey97-83] Helen Meyer. TIS, Microsoft security initiative. Computers & Security, 16(4):330, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404897802027.

Meyer:1997:TT

[Mey97-84] Helen Meyer. Tracking down trouble. Computers & Security, 16 (2):131, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404897883058.

Meyer:1997:TMK

[Mey97-85] Helen Meyer. Trust in my key. Computers & Security, 16(2): 128, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S016740489788296X.

Meyer:1997:UVP

[Mey97-86] Helen Meyer. Unlocking virtual private networks. Computers & Security, 16(7):624, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404897808164.

Meyer:1997:UAV

[Mey97-87] Helen Meyer. User authentication, VPN in an appliance. Computers & Security, 16(7):621, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404897808085.

Meyer:1997:VSG

[Mey97-88] Helen Meyer. Virus strategies get closer look. Computers & Security, 16(2):130, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404897883022.

Meyer:1997:VSU

[Mey97-89] Helen Meyer. VPNS: security with an uncommon touch. Computers & Security, 16(7):624, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404897808152.

Meyer:1997:WTW

[Mey97-90] Helen Meyer. Walking through the wall of fire. Computers & Security, 16(3):215, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404897845506.

Meyer:1997:WSS

[Mey97-91] Helen Meyer. Want security? See what hacker does with a cookie. Computers & Security, 16(2):131, ???? 1997. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S016740489788306X.

Meyer:1997:WTM

[Mey97-92] Helen Meyer. Watch out for the threat of e-mail viruses. Computers & Security, 16(4):332, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404897802076.

Meyer:1997:WDY

[Mey97-93] Helen Meyer. Whom do you trust? Computers & Security, 16 (3):215, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S016740489784549X.

Meyer:1997:WSC

[Mey97-94] Helen Meyer. WorldSecure server combats e-mail plagues. Computers & Security, 16(7):620, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404897808024.

Meyer:1997:E

[Mey97-95] John Meyer. From the editor. Computers & Security, 16(1):2– 3, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404897900018.

Meyer:1997:PHJ

[Mey97-96] John Meyer. Professor Harold Joseph Highland FICS, FACM. Computers & Security, 16(5):350–355, ???? 1997. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404897822407.

Meye:1998:HHB

[Mey98a] Helen Meye. How hackers break in … and how they are caught. Computers & Security, 17(7):618, ???? 1998. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404898903030.

Meyer:1998:AVB

[Mey98b] Helen Meyer. Anti-virus belongs under the rug. Computers & Security, 17(1):59–60, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404897802647.

Meyer:1998:ASD

[Mey98c] Helen Meyer. Attacks spur detection efforts. Computers & Security, 17(5):407, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404898800503.

Meyer:1998:ASV

[Mey98d] Helen Meyer. Aventail secures VPNs down to the client. Computers & Security, 17(1):58, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404897802581.

Meyer:1998:ASC

[Mey98e] Helen Meyer. Axent’s on security consulting. Computers & Security, 17(7):616, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S016740489890298X.

Meyer:1998:BLN

[Mey98f] Helen Meyer. A better lock for ’net data. Computers & Security, 17(6):519, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404898800849.

Meyer:1998:BI

[Mey98g] Helen Meyer. Bulletproof IP. Computers & Security, 17(1): 62, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404897802726.

Meyer:1998:BOS

[Mey98h] Helen Meyer. Bundled offering secures links. Computers & Security, 17(7):614, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404898902899.

Meyer:1998:CAB

[Mey98i] Helen Meyer. Call to arms to battle year 2000 problem. Computers & Security, 17(5):407, ???? 1998. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404898800497.

Meyer:1998:CGB

[Mey98j] Helen Meyer. Coast guard beefs up security after hack. Computers & Security, 17(6):519–520, ???? 1998. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404898800862.

Meyer:1998:CFG

[Mey98k] Helen Meyer. Companies found guilty face criminal prosecutions. Computers & Security, 17(6):517–518, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404898800795.

Meyer:1998:CLI

[Mey98l] Helen Meyer. Companies lap up insurer’s anti-hacker protection plan. Computers & Security, 17(5):408, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404898800540.

Meyer:1998:CAF

[Mey98m] Helen Meyer. Cryptographic accelerators face an uncertain future. Computers & Security, 17(7):616, ???? 1998. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404898902978.

Meyer:1998:CI

[Mey98n] Helen Meyer. Cryptography for the Internet. Computers & Security, 17(7):618, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404898903054.

Meyer:1998:DSS

[Mey98o] Helen Meyer. Digital signatures shouldn’t enter the encryption debate. Computers & Security, 17(1):57, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404897802520.

Meyer:1998:DYN

[Mey98p] Helen Meyer. Do you need an intranet firewall? Computers & Security, 17(1):61, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404897802696.

Meyer:1998:DNB

[Mey98q] Helen Meyer. Don’t neglect browser security. Computers & Security, 17(5):409, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404898800588.

Meyer:1998:MGS

[Mey98r] Helen Meyer. E-mail getting a scan from server. Computers & Security, 17(1):61, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404897802702.

Meyer:1998:EPT

[Mey98s] Helen Meyer. Encryption plan targets export restriction. Computers & Security, 17(6):518, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404898800801.

Meyer:1998:EUA

[Mey98t] Helen Meyer. eSafe ups ante for Java, security. Computers & Security, 17(1):57, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404897802532.

Meyer:1998:EUE

[Mey98u] Helen Meyer. Europeans unlock encryption market, thanks to US rules. Computers & Security, 17(5):407–408, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404898800527.

Meyer:1998:EBN

[Mey98v] Helen Meyer. Experts blast NT for its security failings. Computers & Security, 17(5):407, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404898800515.

Meyer:1998:ECT

[Mey98w] Helen Meyer. Experts criticise ‘too risky’ key recovery. Computers & Security, 17(5):406, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404898800461.

Meyer:1998:FSS

[Mey98x] Helen Meyer. A false sense of security. Computers & Security, 17 (6):520, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404898800898.

Meyer:1998:FEP

[Mey98y] Helen Meyer. Firewall expands purview. Computers & Security, 17(6):519, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404898800837.

Meyer:1998:FRC

[Mey98z] Helen Meyer. Firewall reality check. Computers & Security, 17 (5):410, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S016740489880062X.

Meyer:1998:FAF

[Mey98-27] Helen Meyer. Firewalls aren’t failsafe. Computers & Security, 17 (6):517, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404898800771.

Meyer:1998:HDG

[Mey98-28] Helen Meyer. Hacker defaces German pol site. Computers & Security, 17(6):517, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404898800783.

Meyer:1998:HNW

[Mey98-29] Helen Meyer. Hacker nabbed for web antics. Computers & Security, 17(1):59, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S016740489780260X.

Meyer:1998:HCS

[Mey98-30] Helen Meyer. How computer security works. Computers & Security, 17(7):618, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404898903042.

Meyer:1998:HEY

[Mey98-31] Helen Meyer. How to ensure your Internet security. Computers & Security, 17(7):614, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404898902887.

Meyer:1998:IHB

[Mey98-32] Helen Meyer. IBM hedges its bets on SET. Computers & Security, 17(6):516, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404898800734.

Meyer:1998:IUW

[Mey98-33] Helen Meyer. IBM ups Web hosting security. Computers & Security, 17(1):58, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404897802556.

Meyer:1998:IYS

[Mey98-34] Helen Meyer. Immunizing your system. Computers & Security, 17(7):616, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404898902966.

Meyer:1998:RDE

[Mey98-35] Helen Meyer. In Russia: digital eavesdropping? Computers & Security, 17(6):519, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404898800850.

Meyer:1998:ISW

[Mey98-36] Helen Meyer. Ironing out SET’s wrinkles. Computers & Security, 17(1):58, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S016740489780257X.

Meyer:1998:NID

[Mey98-37] Helen Meyer. Is network intrusion detection software being used correctly? Computers & Security, 17(6):521, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404898800904.

Meyer:1998:SSYb

[Mey98-38] Helen Meyer. Is someone stalking your network? Computers & Security, 17(7):617, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404898903005.

Meyer:1998:JAT

[Mey98-39] Helen Meyer. Java authentication token. Computers & Security, 17(1):59, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404897802623.

Meyer:1998:LLV

[Mey98-40] Helen Meyer. LAN-to-LAN VPNs: secure enough? Computers & Security, 17(6):521, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404898800916.

Meyer:1998:LTP

[Mey98-41] Helen Meyer. Land and teardrop protection. Computers & Security, 17(1):57–58, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404897802544.

Meyer:1998:LYF

[Mey98-42] Helen Meyer. Let your fingers do the logging in. Computers & Security, 17(5):411, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404898800655.

Meyer:1998:LIA

[Mey98-43] Helen Meyer. Locking down intranets from afar — for less. Computers & Security, 17(7):617, ???? 1998. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404898903017.

Meyer:1998:MCS

[Mey98-44] Helen Meyer. Managers are cutting security corners. Computers & Security, 17(5):406, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404898800473.

Meyer:1998:MDK

[Mey98-45] Helen Meyer. Managing digital keys. Computers & Security, 17 (1):62, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404897802738.

Meyer:1998:NSB

[Mey98-46] Helen Meyer. Net security best policy for insurer. Computers & Security, 17(6):520, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404898800874.

Meyer:1998:NMU

[Mey98-47] Helen Meyer. Network managers unite over bug. Computers & Security, 17(7):614, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404898902905.

Meyer:1998:NSI

[Mey98-48] Helen Meyer. Network security interest on rebound. Computers & Security, 17(1):58, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404897802568.

Meyer:1998:NND

[Mey98-49] Helen Meyer. Networks need defence against hacker attacks. Computers & Security, 17(1):60, ???? 1998. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404897802659.

Meyer:1998:NVT

[Mey98-50] Helen Meyer. A new virus threat? Computers & Security, 17 (1):59, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404897802611.

Meyer:1998:NRS

[Mey98-51] Helen Meyer. Novell remote security freebie now for sale. Computers & Security, 17(6):518, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404898800813.

Meyer:1998:LGT

[Mey98-52] Helen Meyer. On-line groups try to beef up privacy. Computers & Security, 17(5):406–407, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404898800485.

Meyer:1998:PP

[Mey98-53] Helen Meyer. Passwords to protection. Computers & Security, 17(5):410, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404898800606.

Meyer:1998:PDS

[Mey98-54] Helen Meyer. PGP disk’s security takes a bite out of crime. Computers & Security, 17(5):409, ???? 1998. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404898800576.

Meyer:1998:PV

[Mey98-55] Helen Meyer. Protection from viruses. Computers & Security, 17(7):614, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404898902875.

Meyer:1998:PLW

[Mey98-56]  Helen Meyer. Psst … wanna buy some data?  Com-

puters & Security, 17(1):61–62, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404897802714.

Meyer:1998:SAN

[Mey98-57] Helen Meyer. Secret agents need not apply. Computers & Security, 17(1):61, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404897802684.

Meyer:1998:SCC

[Mey98-58] Helen Meyer. Security consolidations consumer-friendly. Computers & Security, 17(1):57, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404897802519.

Meyer:1998:SPS

[Mey98-59] Helen Meyer. Security in platinum setting. Computers & Security, 17(7):614, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404898902917.

Meyer:1998:SIS

[Mey98-60] Helen Meyer. Security integration to simplify user access. Computers & Security, 17(6):516, ???? 1998. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404898800746.

Meyer:1998:SMC

[Mey98-61] Helen Meyer. Security, management converge. Computers & Security, 17(6):518–519, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404898800825.

Meyer:1998:SSYa

[Mey98-62] Helen Meyer. Security systems and the Y2K problem. Computers & Security, 17(7):613–614, ???? 1998. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404898902863.

Meyer:1998:STB

[Mey98-63] Helen Meyer. Security tool blocks document misuse. Computers & Security, 17(5):409–410, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S016740489880059X.

Meyer:1998:SUO

[Mey98-64] Helen Meyer. Single unit offers Internet security. Computers & Security, 17(5):408, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404898800552.

Meyer:1998:SBD

[Mey98-65] Helen Meyer. Six biometric devices point the finger at security. Computers & Security, 17(5):410–411, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404898800631.

Meyer:1998:SSD

[Mey98-66] Helen Meyer. Slam the spam door. Computers & Security, 17(7):615–616, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404898902942.

Meyer:1998:SP

[Mey98-67] Helen Meyer. SOCKS push. Computers & Security, 17(5):408– 409, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404898800564.

Meyer:1998:SCSb

[Mey98-68] Helen Meyer. Steel-cage security for Windows. Computers & Security, 17(5):411, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404898800643.

Meyer:1998:SCSa

[Mey98-69] Helen Meyer. Sun changes security framework for Java development kit. Computers & Security, 17(1):58–59, ???? 1998.

CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404897802593.

Meyer:1998:SF

[Mey98-70] Helen Meyer. Suspicious finds. Computers & Security, 17(7): 618, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404898903066.

Meyer:1998:TTF

[Mey98-71] Helen Meyer. Technology and trust: the final analysis. Computers & Security, 17(6):520, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404898800886.

Meyer:1998:SCSc

[Mey98-72] Helen Meyer. This security catches on — slowly. Computers & Security, 17(7):616, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404898902954.

Meyer:1998:TKH

[Mey98-73] Helen Meyer. Through the key holes. Computers & Security, 17 (5):410, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404898800618.

Meyer:1998:TAC

[Mey98-74] Helen Meyer. Titans aim at crypto stalemate. Computers & Security, 17(6):516–517, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404898800758.

Meyer:1998:WCY

[Mey98-75] Helen Meyer. To whom can you entrust security? Computers & Security, 17(7):614–615, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404898902929.

Meyer:1998:TGS

[Mey98-76] Helen Meyer. TriStrata: a giant step in enterprise security. Computers & Security, 17(7):617, ???? 1998. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404898903029.

Meyer:1998:UVP

[Mey98-77] Helen Meyer. Unlocking virtual private networks. Computers & Security, 17(1):62–63, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S016740489780274X.

Meyer:1998:UFR

[Mey98-78] Helen Meyer. US finally relaxes encryption policies. Computers & Security, 17(7):615, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404898902930.

Meyer:1998:VPJ

[Mey98-79] Helen Meyer. ‘virtual’ privacy just not enough. Computers & Security, 17(7):613, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S016740489890284X.

Meyer:1998:VPA

[Mey98-80] Helen Meyer. Virus protect aims high, falls short. Computers & Security, 17(1):60, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404897802660.

Meyer:1998:VPN

[Mey98-81] Helen Meyer. Virus protection for notes and exchange. Computers & Security, 17(7):613, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404898902851.

Meyer:1998:VNC

[Mey98-82] Helen Meyer. VPNs now can link to two-way pagers securely. Computers & Security, 17(5):408, ???? 1998. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404898800539.

Meyer:1998:WUA

[Mey98-83] Helen Meyer. Warning for UK authorities over data matching. Computers & Security, 17(6):517, ???? 1998. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S016740489880076X.

Meyer:1998:WWY

[Mey98-84] Helen Meyer. Watching what you watch. Computers & Security, 17(1):60–61, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404897802672.

Meyer:1998:WPD

[Mey98-85] Helen Meyer. White pages for digital certificates. Computers & Security, 17(1):59, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404897802635.

Meyer:1998:WGT

[Mey98-86] Helen Meyer. Who goes there? Computers & Security, 17 (7):616–617, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404898902991.

Meyer:1998:WPS

[Mey98-87] Helen Meyer. WSS put its stamp on e-mail security. Computers & Security, 17(6):521, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404898800928.

Meyer:1998:BBG

[Mey98-88] John Meyer. Business is booming for the good-guy hackers. Computers & Security, 17(8):719, ???? 1998. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404898903522.

Meyer:1998:DPS

[Mey98-89] John Meyer. Data protect and survive. Computers & Security, 17(8):716–717, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S016740489890342X.

Meyer:1998:CGS

[Mey98-90] John Meyer. E-commerce to get strict privacy rules. Computers & Security, 17(8):719–720, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404898903534.

Meyer:1998:FSH

[Mey98-91] John Meyer. Firewalls stand the heat. Computers & Security, 17 (8):716, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404898903418.

Meyer:1998:FFN

[Mey98-92] John Meyer. First federal net privacy law approved. Computers & Security, 17(8):719, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404898903509.

Meyer:1998:FAS

[Mey98-93] John Meyer. Fraud artists surf new turf — cyberspace. Computers & Security, 17(8):718, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404898903480.

Meyer:1998:HCG

[Mey98-94] John Meyer. Hackers can get to you by e-mail. Computers & Security, 17(8):719, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404898903510.

Meyer:1998:HGM

[Mey98-95] John Meyer. Hacks gain in malice, frequency. Computers & Security, 17(8):718, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404898903479.

Meyer:1998:IFS

[Mey98-96] John Meyer. ISS finds security ANSA. Computers & Security, 17(8):715, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404898903388.

Meyer:1998:LFH

[Mey98-97] John Meyer. Low flying hackers pose growing threat. Computers & Security, 17(8):717–718, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404898903467.

Meyer:1998:MBF

[Mey98-98] John Meyer. Microsoft backs feds’ encryption standard. Computers & Security, 17(8):716, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404898903406.

Meyer:1998:NON

[Mey98-99] John Meyer. Netscape offers new security tool. Computers & Security, 17(8):715, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404898801016.

Meyer:1998:NSU

[Mey98-100] John Meyer. Network security under attack? Buy insurance. Computers & Security, 17(8):717, ???? 1998. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404898903455.

Meyer:1998:QCS

[Mey98-101] John Meyer. Quick cracking of secret code. Computers & Security, 17(8):718–719, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404898903492.

Meyer:1998:RPH

[Mey98-102] John Meyer. Rich pickings for hackers. Computers & Security, 17(8):717, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404898903431.

Meyer:1998:SHL

[Mey98-103] John Meyer. Spate of hacking lets the cat out of the bag. Computers & Security, 17(8):715–716, ???? 1998. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S016740489890339X.

Meyer:1998:SBA

[Mey98-104] John Meyer. Symantec buys anti-virus line. Computers & Security, 17(8):717, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404898903443.

Meyer:1999:AIC

[Mey99a] Helen Meyer. 20% annual increase on card fraud. Computers & Security, 18(3):233, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404899909392.

Meyer:1999:BAB

[Mey99b] Helen Meyer. 56-bit DES algorithm broken in record time. Computers & Security, 18(2):149–150, ???? 1999. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404899800327.

Meyer:1999:AIA

[Mey99c] Helen Meyer. Active infections attack. Computers & Security, 18(1):76–77, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404899800169.

Meyer:1999:BSE

[Mey99d] Helen Meyer. Better snapshots of enterprise security. Computers & Security, 18(3):233–234, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404899909409.

Meyer:1999:CYA

[Mey99e] Helen Meyer. Can you afford not to invest in security? Computers & Security, 18(1):77, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404899800170.

Meyer:1999:CCV

[Mey99f] Helen Meyer. Combating computer viruses. Computers & Security, 18(3):235–236, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S016740489990946X.

Meyer:1999:CLA

[Mey99g] Helen Meyer. Companies launch anti-virus education programme. Computers & Security, 18(1):76, ???? 1999. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404899800133.

Meyer:1999:CPF

[Mey99h] Helen Meyer. Compaq points fingerprints at security hole. Computers & Security, 18(1):76, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404899800157.

Meyer:1999:CPS

[Mey99i] Helen Meyer. Crypto paper set for January launch. Computers & Security, 18(1):76, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404899800145.

Meyer:1999:CS

[Mey99j] Helen Meyer. Cybercop scanner. Computers & Security, 18(3): 234, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404899909410.

Meyer:1999:DOS

[Mey99k] Helen Meyer. Doors open to simpler VPNs. Computers & Security, 18(1):78–79, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404899800224.

Meyer:1999:EOR

[Mey99l] Helen Meyer. European organizations at risk from virus infection. Computers & Security, 18(3):232, ???? 1999. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404899909355.

Meyer:1999:FSS

[Mey99m] Helen Meyer. Fingerprint sensors solve security problems. Computers & Security, 18(2):149, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404899800303.

Meyer:1999:FFS

[Mey99n] Helen Meyer. Flak flies over secrecy on virus. Computers & Security, 18(1):75, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404899800108.

Meyer:1999:FFE

[Mey99o] Helen Meyer. Free file encryption offered. Computers & Security, 18(3):231, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404899909318.

Meyer:1999:HGJ

[Mey99p] Helen Meyer. Hacker gets jail time for ISP break-ins. Computers & Security, 18(2):152, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404899800406.

Meyer:1999:HTP

[Mey99q] Helen Meyer. Hacker trial postponed. Computers & Security, 18 (1):77, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404899800182.

Meyer:1999:HCR

[Mey99r] Helen Meyer. Harmonized copyright rules across Europe. Computers & Security, 18(3):232–233, ???? 1999. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404899909379.

Meyer:1999:HSH

[Mey99s] Helen Meyer. HTML security hole in Excel can invade PCs. Computers & Security, 18(2):150, ???? 1999. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404899800352.

Meyer:1999:IAP

[Mey99t] Helen Meyer. International agreement pees encryption at 64. Computers & Security, 18(1):77–78, ???? 1999. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404899800194.

Meyer:1999:IDS

[Mey99u] Helen Meyer. Interoperability drives security alliances. Computers & Security, 18(1):78, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404899800200.

Meyer:1999:KYE

[Mey99v] Helen Meyer. Know your enemy. Computers & Security, 18(1): 80, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404899800261.

Meyer:1999:LYD

[Mey99w] Helen Meyer. Lock up your data. Computers & Security, 18(2): 153, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404899800455.

Meyer:1999:MIW

[Mey99x] Helen Meyer. Making IPSec work for you. Computers & Security, 18(2):154, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404899800492.

Meyer:1999:MMH

[Mey99y] Helen Meyer. Making the most of hacker fears. Computers & Security, 18(1):79, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404899800236.

Meyer:1999:MNS

[Mey99z] Helen Meyer. Mapping a network security strategy. Computers & Security, 18(3):236, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404899909483.

Meyer:1999:MWJ

[Mey99-27] Helen Meyer. MCI WorldCom joins managed security fray. Computers & Security, 18(3):231–232, ???? 1999. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404899909343.

Meyer:1999:MBT

[Mey99-28] Helen Meyer. More bark than bite. Computers & Security, 18 (3):236, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404899909495.

Meyer:1999:MCC

[Mey99-29] Helen Meyer. Motorola CipherNET certifies security. Computers & Security, 18(2):153, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404899800467.

Meyer:1999:NHA

[Mey99-30] Helen Meyer. The need for host authentication. Computers & Security, 18(3):235, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404899909458.

Meyer:1999:NHT

[Mey99-31] Helen Meyer. Netbus hacker tool presents a greater threat to Windows shops than BackOrifice. Computers & Security, 18(2): 150, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404899800339.

Meyer:1999:NSA

[Mey99-32] Helen Meyer. Network security for all. Computers & Security, 18(2):154, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404899800480.

Meyer:1999:PCP

[Mey99-33] Helen Meyer. Passwords could be past tense by 2002. Computers & Security, 18(1):75–76, ???? 1999. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404899800121.

Meyer:1999:PPC

[Mey99-34] Helen Meyer. PGP 6.0: Pro crypto suite gets personal. Computers & Security, 18(2):151–152, ???? 1999. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S016740489980039X.

Meyer:1999:PHM

[Mey99-35] Helen Meyer. PKI heads for mainstream. Computers & Security, 18(2):151, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404899800364.

Meyer:1999:PGB

[Mey99-36] Helen Meyer. Privacy group blasts new cryptography laws. Computers & Security, 18(2):151, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404899800388.

Meyer:1999:PCV

[Mey99-37] Helen Meyer. Protection from Chernobyl virus. Computers & Security, 18(3):231, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404899909331.

Meyer:1999:PBV

[Mey99-38] Helen Meyer. Providers beef up VPNs with security. Computers & Security, 18(3):233, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404899909380.

Meyer:1999:RGU

[Mey99-39] Helen Meyer. RSA goes down under to dodge export laws. Computers & Security, 18(2):152, ???? 1999. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S016740489980042X.

Meyer:1999:SWN

[Mey99-40] Helen Meyer. Securing windows NT server. Computers & Security, 18(3):236, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404899909471.

Meyer:1999:SES

[Mey99-41] Helen Meyer. Security elite SWAT at viruses. Computers & Security, 18(2):151, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404899800376.

Meyer:1999:STFa

[Mey99-42] Helen Meyer. Security takes front seat at Pentium preview. Computers & Security, 18(3):232, ???? 1999. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404899909367.

Meyer:1999:STFb

[Mey99-43] Helen Meyer. Security teams failing to shine. Computers & Security, 18(3):234, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404899909434.

Meyer:1999:SBJ

[Mey99-44] Helen Meyer. Server-based Java security products help guard your enterprise flank. Computers & Security, 18(2):153–154, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404899800479.

Meyer:1999:SSS

[Mey99-45] Helen Meyer. Sidewinder Security Server 4.0. Computers & Security, 18(2):153, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404899800443.

Meyer:1999:SEI

[Mey99-46] Helen Meyer. Strong encryption for internal use only? Computers & Security, 18(3):234, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404899909422.

Meyer:1999:SS

[Mey99-47] Helen Meyer. Stronghold of security. Computers & Security, 18 (1):79, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404899800248.

Meyer:1999:TS

[Mey99-48] Helen Meyer. Tapping away at security. Computers & Security, 18(1):78, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404899800212.

Meyer:1999:TTB

[Mey99-49] Helen Meyer. Think twice before becoming a hacker attacker. Computers & Security, 18(2):150, ???? 1999. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404899800340.

Meyer:1999:CIT

[Mey99-50] Helen Meyer. To catch an Internet thief. Computers & Security, 18(3):234–235, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404899909446.

Meyer:1999:UWC

[Mey99-51] Helen Meyer. Users warned on CE security flaws. Computers & Security, 18(1):75, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S016740489980011X.

Meyer:1999:VPC

[Mey99-52] Helen Meyer. Vendors plan chip-level security systems. Computers & Security, 18(2):149, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404899800315.

Meyer:1999:VRC

[Mey99-53] Helen Meyer. Vendors to roll out centralized security tools, services. Computers & Security, 18(2):152–153, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404899800431.

Meyer:1999:WDP

[Mey99-54] Helen Meyer. Web data-protection offerings readied. Computers & Security, 18(3):231, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S016740489990932X.

Meyer:1999:WWA

[Mey99-55] Helen Meyer. Will we all be secure in the new world? Computers & Security, 18(1):79–80, ???? 1999. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S016740489980025X.

Meyer:1999:YMM

[Mey99-56] Helen Meyer. Y2K may mask hacker attacks. Computers & Security, 18(2):152, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404899800418.

Miyoshi:1992:STB

[MG92] Dennis Miyoshi and Mary Green. Security and technology: a better mousetrap. Computers & Security, 11(5):491, September 1992. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/016740489290252M.

Muftic:1992:CGS

[MH92] Sead Muftic and Edina Hatunic. CISS: Generalized security libraries. Computers & Security, 11(7):653–659, November 1992. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/016740489290145H.

Mulhall:1997:PFF

[MH97] Tom Mulhall and Peter Hoath. PBX fraud, the final solution — detection, deterrent, countermeasures. Computers & Security, 16(6):521, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404897846792.

Milsom:1995:NS

[Mil95] Chris Milsom. Novell security. Computers & Security, 14(5): 424, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/0167404895971406.

Miller:1996:EWS

[Mil96] Nigel Miller. Establishing web sites — legal issues and risks. Computers & Security, 15(3):198–202, ???? 1996. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/0167404896897811.

Mitchell:1995:VAS

[Mit95] John Mitchell. Value added system development auditing. Computers & Security, 14(5):419, ???? 1995. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/0167404896900401.

Moulton:1996:ECR

[MM96] Rolf T. Moulton and Mark E. Moulton. Electronic communications risk management: a checklist for business managers. Computers & Security, 15(5):377–386, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/0167404896825600.

Michels:1996:GBO

[MNP96] Markus Michels, David Naccache, and Holger Petersen. GOST 34.10 — a brief overview of Russia’s DSA. Computers & Security, 15(8):725–732, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404896000168.

Moeller:1992:UBC

[Moe92] Mike Moeller. US banks can buy cover against computer crime. Computers & Security, 11(5):488–489, September 1992. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/016740489290029Q.

Mohrenschlager:1995:HCS

[Möh95] Manfred Möhrenschlager. Hacking: To criminalize or not? — suggestions for the legislature. Computers & Security, 14 (2):103–112, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/016740489597029A.

Molini:1997:EBD

[Mol97] James E. Molini. Electronic borders: Defining and protecting national networks. Computers & Security, 16(3):189–196, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404897845117.

Moreau:1996:PFP

[Mor96] Thierry Moreau. A probabilistic flaw in PGP design? Computers & Security, 15(1):39–43, ???? 1996. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/0167404896876218.

Moreau:1999:ELF

[Mor99] Thierry Moreau. The emergence of a legal framework for electronic transactions. Computers & Security, 18(5):423–428, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404899800881.

Moseley:1993:PKM

[Mos93] David Moseley. Public key management for X.25 network security. Computers & Security, 12(2):128–133, March 1993. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/016740489390092J.

Moulton:1998:RCC

[Mou98] Mark Moulton. Reducing charges of e-comm harassment. Computers & Security, 17(2):137–142, ???? 1998. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404897819839.

Matyas:1998:RDM

[MPRZ98] Mike Matyas, Mohammad Peyravian, Allen Roginsky, and Nev Zunic. Reversible data mixing procedure for efficient publickey encryption. Computers & Security, 17(3):265–272, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404898803404.

McKosky:1990:FIC

[MS90] Robert A. McKosky and Sajjan G. Shiva. A file integrity checking system to detect and recover from program modification attacks in multi-user computer systems. Computers & Security, 9(5):431–446, August 1990. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/0167404890900744.

Muftic:1990:TPA

[Muf90] Sead Muftic. Transaction protection by “antennas”. Computers & Security, 9(3):245–255, May 1990. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/016740489090171O.

Mulhall:1995:DTF

[Mul95] Tom Mulhall. Dial thru’ fraud the ramifications for a business. Computers & Security, 14(5):422, ???? 1995. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/016740489597131S.

Mulhall:1997:CRF

[Mul97a] Tom Mulhall. Computer-related fraud and its evolution within telephony. Computers & Security, 16(6):521, ???? 1997. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404897846767.

Mulhall:1997:WAHa

[Mul97b] Tom Mulhall. Where have all the hackers gone? A study in motivation, deterrence, and crime displacement: Part 1 — introduction and methodology. Computers & Security, 16(4):277– 284, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404897801903.

Mulhall:1997:WAHb

[Mul97c] Tom Mulhall. Where have all the hackers gone?: Part 2 — frequency of computer related abuse: the surveys, and public perception of the problem. Computers & Security, 16(4):285– 290, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404897801915.

Mulhall:1997:WAHc

[Mul97d] Tom Mulhall. Where have all the hackers gone?: Part 3 — motivation and deterrence. Computers & Security, 16(4):291– 297, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404897801927.

Mulhall:1997:WAHd

[Mul97e] Tom Mulhall. Where have all the hackers gone?: Part 4 — legislation. Computers & Security, 16(4):298–303, ????

1997. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404897801939.

Mulhall:1997:WAHe

[Mul97f] Tom Mulhall. Where have all the hackers gone?: Part 5 — conclusions. Computers & Security, 16(4):304–306, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404897801940.

Mulhall:1997:WAHf

[Mul97g] Tom Mulhall. Where have all the hackers gone?: Part 6 — appendices and bibliography. Computers & Security, 16(4):307– 315, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S016740489790153X.

Murray:1991:LE

[Mur91] William Hugh Murray. Letter to the editor. Computers & Security, 10(8):723–724, December 1991. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/016740489190091Q.

Mitchell:1991:MFC

[MV91] Chris J. Mitchell and Vijay Varadharajan. Modified forms of cipher block chaining. Computers & Security, 10(1):37–40, February 1991. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/016740489190054H.

Mostert:1994:MIC

[MvS94] D. N. J. Mostert and S. H. von Solms. A methodology to include computer security, safety and resilience requirements as part of the user requirement. Computers & Security, 13(4):349– 364, ???? 1994. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/0167404894900256.

Nacht:1997:SMF

[Nac97] Marius Nacht. The spectrum of modern firewalls. Computers & Security, 16(6):516, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404897846664.

Nacht:1998:SMF

[Nac98] Marius Nacht. The spectrum of modern firewalls. Computers & Security, 17(1):54–56, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404897802507.

Nelms:1999:IMR

[Nel99] Chris Nelms. Internet e-mail risks and concerns. Computers & Security, 18(5):409–418, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404899800868.

Newman:1995:CRS

[New95] Peter Newman. CD-ROM security issues. Computers & Security, 14(5):422, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404896900449.

Nichol:1999:APA

[Nic99a] Sandy Nichol. Anti-piracy alliance under fire. Computers & Security, 18(6):522–523, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S016740489990857X.

Nichol:1999:ATM

[Nic99b] Sandy Nichol. Audit trails might have fingered Los Alamos insiders. Computers & Security, 18(6):524, ???? 1999. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404899908623.

Nichol:1999:AHK

[Nic99c] Sandy Nichol. Aussie hackers as key escrow alternative. Computers & Security, 18(4):338, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404899907241.

Nichol:1999:SB

[Nic99d] Sandy Nichol. Be a standard bearer. Computers & Security, 18 (5):434, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404899800984.

Nichol:1999:BBW

[Nic99e] Sandy Nichol. Big brother is watching you. Computers & Security, 18(6):523, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404899908581.

Nichol:1999:BUC

[Nic99f] Sandy Nichol. BSA uncovers counterfeit of itself. Computers & Security, 18(5):435, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404899801023.

Nichol:1999:BCI

[Nic99g] Sandy Nichol. Bug circumvents IE5 passwords. Computers & Security, 18(6):521, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404899908519.

Nichol:1999:BCN

[Nic99h] Sandy Nichol. Bug cracks NT open. Computers & Security, 18 (6):524, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404899908647.

Nichol:1999:CUS

[Nic99i] Sandy Nichol. Canada urged to step up computer security. Computers & Security, 18(5):432, ???? 1999. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S016740489980090X.

Nichol:1999:CIP

[Nic99j] Sandy Nichol. Card issuers push security plan. Computers & Security, 18(6):520, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404899908465.

Nichol:1999:CSY

[Nic99k] Sandy Nichol. Chem sector Y2K ready? Computers & Security, 18(4):338, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404899907277.

Nichol:1999:CBI

[Nic99l] Sandy Nichol. City braces itself for cyber attacks. Computers & Security, 18(6):523, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404899908593.

Nichol:1999:CYC

[Nic99m] Sandy Nichol. Concerns over Y2K cut funds for security. Computers & Security, 18(5):432, ???? 1999. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404899800911.

Nichol:1999:DPD

[Nic99n] Sandy Nichol. Data protection dispute exposes core differences. Computers & Security, 18(6):519, ???? 1999. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S016740489990843X.

Nichol:1999:DHL

[Nic99o] Sandy Nichol. Deployment hurdles: look before you leap. Computers & Security, 18(4):337, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404899907228.

Nichol:1999:DCF

[Nic99p] Sandy Nichol. Developing countries ‘face Y2K trade problems’. Computers & Security, 18(6):519–520, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404899801163.

Nichol:1999:DTA

[Nic99q] Sandy Nichol. Duchy is taking up arms for cyber wars. Computers & Security, 18(5):437, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404899801084.

Nichol:1999:CBFb

[Nic99r] Sandy Nichol. E-commerce bill faces the axe. Computers & Security, 18(6):520–521, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404899908477.

Nichol:1999:CBFa

[Nic99s] Sandy Nichol. E-commerce bill is flawed and lacking in direction, say MPs. Computers & Security, 18(5):433, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404899800935.

Nichol:1999:EAS

[Nic99t] Sandy Nichol. Electronic attack: Serbs bombard Internet with mail. Computers & Security, 18(4):335, ???? 1999. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404899800790.

Nichol:1999:ERP

[Nic99u] Sandy Nichol. Encryption row, part 2. Computers & Security, 18(5):435–436, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404899801035.

Nichol:1999:EKU

[Nic99v] Sandy Nichol. Escrow key to US espionage. Computers & Security, 18(5):436, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404899801047.

Nichol:1999:EUDb

[Nic99w] Sandy Nichol. EU-US data privacy dispute to continue. Computers & Security, 18(6):523, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404899908611.

Nichol:1999:EUDa

[Nic99x] Sandy Nichol. Europe and US disagree on privacy. Computers & Security, 18(5):436, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404899801060.

Nichol:1999:FST

[Nic99y] Sandy Nichol. FBI sting targets computer hackers. Computers & Security, 18(5):433, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404899800947.

Nichol:1999:YEOb

[Nic99z] Sandy Nichol. For your ears only: inexpensive telephone encryption arrives. Computers & Security, 18(6):521, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404899908507.

Nichol:1999:YEOa

[Nic99-27] Sandy Nichol. For your eyes only. Computers & Security, 18(5):432–433, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404899800923.

Nichol:1999:GWS

[Nic99-28] Sandy Nichol. Get wise to software viruses. Computers & Security, 18(4):337–338, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S016740489990723X.

Nichol:1999:GSH

[Nic99-29] Sandy Nichol. Government site hacked. Computers & Security, 18(4):339, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404899907320.

Nichol:1999:HOD

[Nic99-30] Sandy Nichol. Hackers open door to Windows. Computers & Security, 18(6):522, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404899908556.

Nichol:1999:HND

[Nic99-31] Sandy Nichol. Hacking in the name of democracy in China. Computers & Security, 18(6):521, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404899908490.

Nichol:1999:IS

[Nic99-32] Sandy Nichol. Internet security. Computers & Security, 18 (4):339–340, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404899907344.

Nichol:1999:ISM

[Nic99-33] Sandy Nichol. Internet systems ‘must be easily accessible by police’. Computers & Security, 18(6):521, ???? 1999. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404899908489.

Nichol:1999:IIH

[Nic99-34] Sandy Nichol. IT industry hit hard by software piracy. Computers & Security, 18(5):436, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404899801059.

Nichol:1999:JSB

[Nic99-35] Sandy Nichol. Java security bug found, patched. Computers & Security, 18(5):434, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404899800960.

Nichol:1999:LC

[Nic99-36] Sandy Nichol. Laying down cyberlaw. Computers & Security, 18 (6):520, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404899908441.

Nichol:1999:LPM

[Nic99-37] Sandy Nichol. Legislation protects mail from employers’ snooping. Computers & Security, 18(5):435, ???? 1999. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404899801011.

Nichol:1999:LTC

[Nic99-38] Sandy Nichol. Legislative threat to e-commerce. Computers & Security, 18(6):522, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404899908544.

Nichol:1999:MRC

[Nic99-39] Sandy Nichol. Many Russian computers still aren’t Y2K ready. Computers & Security, 18(6):524, ???? 1999. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404899908635.

Nichol:1999:MWO

[Nic99-40] Sandy Nichol. MCI Worldcom offers outsourced global VPNs. Computers & Security, 18(6):522, ???? 1999. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404899908532.

Nichol:1999:MMC

[Nic99-41] Sandy Nichol. Melissa mutates into copycat virus via RTF. Computers & Security, 18(5):437, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404899801102.

Nichol:1999:MVW

[Nic99-42] Sandy Nichol. Melissa virus wakes up IT world. Computers & Security, 18(4):336, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404899907162.

Nichol:1999:MM

[Nic99-43] Sandy Nichol. Melissa’s mayhem. Computers & Security, 18(4):335–336, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404899907150.

Nichol:1999:NSS

[Nic99-44] Sandy Nichol. Net security scare for Microsoft. Computers & Security, 18(5):435, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S016740489980100X.

Nichol:1999:NSB

[Nic99-45] Sandy Nichol. Net surfers: Beware of hole. Computers & Security, 18(4):339, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404899907319.

Nichol:1999:NVW

[Nic99-46] Sandy Nichol. Novell virtual wallet raises security issues. Computers & Security, 18(4):338, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404899907265.

Nichol:1999:NWW

[Nic99-47] Sandy Nichol. Novell wages war on Israeli software piracy. Computers & Security, 18(4):339, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404899907307.

Nichol:1999:PVC

[Nic99-48] Sandy Nichol. Password vulnerability cited as cracker hacks Ebay site. Computers & Security, 18(4):338, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404899907289.

Nichol:1999:PSV

[Nic99-49] Sandy Nichol. PDAs spread virus epidemics. Computers & Security, 18(6):523, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S016740489990860X.

Nichol:1999:PVH

[Nic99-50] Sandy Nichol. Powerful virus hits computers. Computers & Security, 18(6):519, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404899908428.

Nichol:1999:PCP

[Nic99-51] Sandy Nichol. Privacy concerns over Pentium III launch. Computers & Security, 18(4):336, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404899907198.

Nichol:1999:RNS

[Nic99-52] Sandy Nichol. Report: NASA system vulnerable because of lax security. Computers & Security, 18(6):524, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404899908659.

Nichol:1999:RSU

[Nic99-53] Sandy Nichol. Report says US has backdoor to notes. Computers & Security, 18(6):520, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404899908453.

Nichol:1999:RMH

[Nic99-54] Sandy Nichol. Russia may hold off Y2K tests. Computers & Security, 18(5):434, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404899800972.

Nichol:1999:SWI

[Nic99-55] Sandy Nichol. SEC warns on Internet trading. Computers & Security, 18(5):433, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404899800959.

Nichol:1999:SY

[Nic99-56] Sandy Nichol. Secure yourself. Computers & Security, 18 (4):336–337, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404899907204.

Nichol:1999:SO

[Nic99-57] Sandy Nichol. Security from the outside. Computers & Security, 18(4):337, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404899907216.

Nichol:1999:SEB

[Nic99-58] Sandy Nichol. Security is in the eye of the beholder. Computers & Security, 18(4):336, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404899907174.

Nichol:1999:SMG

[Nic99-59] Sandy Nichol. Software maker gets OK to ship abroad. Computers & Security, 18(4):339, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404899907332.

Nichol:1999:SPG

[Nic99-60] Sandy Nichol. Software piracy grows in Canada, dips in US. Computers & Security, 18(6):521–522, ???? 1999. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404899908520.

Nichol:1999:TC

[Nic99-61] Sandy Nichol. Tales of the cryptochip. Computers & Security, 18(6):522, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404899908568.

Nichol:1999:TCS

[Nic99-62] Sandy Nichol. Threat of cyber sabotage increases. Computers & Security, 18(5):434–435, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404899800996.

Nichol:1999:TFF

[Nic99-63] Sandy Nichol. Top firms fail virus test. Computers & Security, 18(4):338, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404899907253.

Nichol:1999:UCT

[Nic99-64] Sandy Nichol. UK companies trust in non-existent security. Computers & Security, 18(5):436–437, ???? 1999. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404899801072.

Nichol:1999:VTS

[Nic99-65] Sandy Nichol. Veto threat shadows Y2K bill. Computers & Security, 18(5):437, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404899801096.

Nichol:1999:VUV

[Nic99-66] Sandy Nichol. VLANs usurped by virtual private networks. Computers & Security, 18(4):340, ???? 1999. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404899907356.

Nichol:1999:VCS

[Nic99-67] Sandy Nichol. VPN client software needs to be made simpler. Computers & Security, 18(4):338–339, ???? 1999. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404899907290.

Nichol:1999:VFA

[Nic99-68]  Sandy Nichol.  VPN false alarms can be diminished.

Computers & Security, 18(4):340, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404899907368.

Nichol:1999:WES

[Nic99-69] Sandy Nichol. Weigh in on encryption standard. Computers & Security, 18(4):336, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404899907186.

Nichol:1999:WDP

[Nic99-70] Sandy Nichol. What to do to prevent attack. Computers & Security, 18(4):335, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404899907149.

Norman:1992:PSI

[Nor92a] Geoff Norman. PC security issues. Computers & Security, 11(5):412–416, September 1992. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/016740489290004B.

North:1992:CC

[Nor92b] Sandra North. To catch a “crimoid”. Computers & Security, 11(5):490, September 1992. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/016740489290247O.

Norman:1993:DRA

[Nor93] Geoff Norman. Disaster recovery after downsizing. Computers & Security, 12(3):225–229, May 1993. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/016740489390108H.

Nosworthy:1999:YCP

[Nos99] Julie D. Nosworthy. Y2K contingency planning: Taking BCM into the 21-st century. Computers & Security, 18(8):693–704, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404899801333.

Nurmi:1991:SBE

[NSS91] Hannu Nurmi, Arto Salomaa, and Lila Santean. Secret ballot elections in computer networks. Computers & Security, 10 (6):553–560, October 1991. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/016740489190081N.

Olson:1990:CAC

[OA90] Ingrid M. Olson and Marshall D. Abrams. Computer access control policy choices. Computers & Security, 9(8):699–714, December 1990. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/0167404890901138.

Obaidat:1993:MIC

[Oba93] M. S. Obaidat. A methodology for improving computer access security. Computers & Security, 12(7):657–662, November 1993. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/016740489390083H.

Oliphant:1995:ITA

[Oli95a] Alan Oliphant. Internet tools for auditors. Computers & Security, 14(5):424, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404895971392.

Oliver:1995:PAA

[Oli95b] Claude Oliver. Privacy, anonymity and accountability. Computers & Security, 14(6):489–490, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/016740489599311C.

Olnes:1993:ESM

[Øln93] Jon Ølnes. EDIFACT security made simple-the EDIMED approach. Computers & Security, 12(8):765–774, December 1993. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/0167404893900424.

Olnes:1994:DSP

[Øln94] Jon Ølnes. Development of security policies. Computers & Security, 13(8):628–636, ???? 1994. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/0167404894900426.

OShea:1991:OSI

[O’S91] G. F. G. O’Shea. Operating system integrity. Computers & Security, 10(5):443–465, August 1991. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/016740489190070T.

OShea:1995:RAR

[O’S95] Greg O’Shea. Redundant access rights. Computers & Security, 14(4):323–348, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/016740489500006T.

Osborne:1995:SIC

[Osb95] Keith Osborne. Security issues in contemporary banking. Computers & Security, 14(5):426, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/0167404896900498.

Osborne:1997:AAS

[Osb97a] Keith Osborne. Advanced auditing: Selecting and testing the right controls. Computers & Security, 16(6):518, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404897846718.

Osborne:1997:AIS

[Osb97b] Keith Osborne. Auditing the IT security function. Computers & Security, 16(6):524, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404897846937.

Osborne:1998:AIS

[Osb98] Keith Osborne. Auditing the IT security function. Computers & Security, 17(1):34–41, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404897802489.

Olivier:1992:BSD

[OvS92] M. S. Olivier and S. H. von Solms. Building a secure database using self-protecting objects. Computers & Security, 11(3):259– 271, May 1992. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/0167404892902067.

Ozier:1998:GPO

[Ozi98] Will Ozier. GSSP preface/overview. Computers & Security, 17(1):14–18, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S016740489780243X.

Pfitzmann:1993:MES

[PA93]  Andreas Pfitzmann and Ralf Aßmann.  More efficient

software implementations of (generalized) DES. Computers & Security, 12(5):477–500, August 1993. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/016740489390069H.

Poole:1995:CDO

[PA95] Vernon Poole and Dean Adams. Current developments in open systems security. Computers & Security, 14(5):420, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/016740489597125T.

Paans:1991:MES

[Paa91] Ronald Paans. With MVS/ESA security labels towards B1. Computers & Security, 10(4):309–324, June 1991. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/016740489190106N.

Parker:1990:AAC

[Par90] T. A. Parker. Application access control standards for distributed systems. Computers & Security, 9(6):519–528, October 1990. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/016740489090130L.

Parker:1995:RBA

[Par95a] Tom Parker. Role based access control. Computers & Security, 14(5):418, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/016740489597115Q.

Parkin:1995:ISI

[Par95b] Rod Parkin. IT security — an implementation strategy. Computers & Security, 14(5):423, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/016740489597133U.

Parker:1997:IFA

[Par97a] Donn B. Parker. Infosec folk art and beyond. Computers & Security, 16(6):515, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404897846615.

Parker:1997:SVIa

[Par97b] Donn B. Parker. The strategic value of information in business. Computers & Security, 16(6):525, ???? 1997. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404897846986.

Parker:1997:SVIb

[Par97c] Donn B. Parker. The strategic values of information security in business. Computers & Security, 16(7):572–582, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404897807936.

Parker:1997:WCA

[Par97d] Tom Parker. What are certification authorities and what are they liable for? Computers & Security, 16(6):516, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404897846688.

Parkin:1997:IIS

[Par97e] Roderick Parkin. The importance of IT security. Computers & Security, 16(6):523, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404897846895.

Preneel:1990:CFC

[PBGV90] Bart Preneel, Antoon Bosselaers, René Govaerts, and Joos Vandewalle. Cryptanalysis of a fast cryptographic checksum algorithm. Computers & Security, 9(3):257–262, May 1990. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/016740489090172P.

Peyravian:1999:SSK

[PC99] Mohammad Peyravian and Don Coppersmith. A structured symmetric-key block cipher. Computers & Security, 18(2):134– 147, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404899900536.

Paans:1991:ASS

[PdL91] Ronald Paans and Hans de Lange. Auditing the SNA/SNI environment. Computers & Security, 10(3):251–261, May 1991. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/016740489190041B.

Peach:1991:DRU

[Pea91] S. Peach. Disaster recovery: an unnecessary cost burden or an essential feature of any DP installation? Computers & Security, 10(6):565–568, October 1991. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/016740489190084Q.

Pepper:1997:AYI

[Pep97] Bill Pepper. Addressing your information security skills shortages. Computers & Security, 16(6):522–523, ???? 1997. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404897846846.

Pfleeger:1991:FSR

[Pfl91] Shari Lawrence Pfleeger. A framework for security requirements. Computers & Security, 10(6):515–523, October 1991. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/016740489190076P.

Picciotto:1994:ELP

[PG94] Jeffrey Picciotto and Richard D. Graubart. Extended labeling policies for enhanced application support. Computers & Security, 13(7):587–599, ???? 1994. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/0167404894900108.

Paans:1990:ACM

[PH90] R. Paans and I. S. Herschberg. Auditing the change management process. Computers & Security, 9(2):161–174, April 1990. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/016740489090090G.

Post:1991:AVS

[PK91] Gerald V. Post and Karen-Ann Kievit. Accessibility vs. security: a look at the demand for computer security. Computers & Security, 10(4):331–344, June 1991. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/016740489190109Q.

Post:1998:UEA

[PK98] Gerald Post and Albert Kagan. The use and effectiveness of anti-virus software. Computers & Security, 17(7):589–599, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404899800595.

Plant:1993:GMB

[Pla93] Martin Plant. Getting management buy-in to IT security. Computers & Security, 12(7):623–626, November 1993. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/016740489390075G.

Peyravian:1999:GUB

[PMRZ99] Mohammad Peyravian, Stephen M. Matyas, Allen Roginsky, and Nev Zunic. Generating user-based cryptographic keys and random numbers. Computers & Security, 18(7):619–626, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404899820409.

Pool:1997:CAG

[Poo97] Vernon Pool. COBIT — audit guidance on effective implementation. Computers & Security, 16(6):518, 521, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404897846755.

Porter:1996:IWY

[Por96] Tim Porter. Information warfare — your company needs you! Computers & Security, 15(7):561–566, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404897881114.

Pounder:1997:FST

[Pou97] Chris Pounder. First steps towards a European Union policy on the securing of electronic communications. Computers & Security, 16(7):590–594, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S016740489780795X.

Pounder:1998:ARU

[Pou98a] Chris Pounder. The annual report of the UK data protection registrar. Computers & Security, 17(6):479–482, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404898800709.

Pounder:1998:ECT

[Pou98b] Chris Pounder. European Commission takes action to secure the Internet. Computers & Security, 17(5):392–396, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404898800448.

Pounder:1998:FDF

[Pou98c] Chris Pounder. Further developments in the field of encryption and digital signatures. Computers & Security, 17(4):308– 312, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404898800114.

Pounder:1998:HNL

[Pou98d] Chris Pounder. Homeworking: No longer an easy option? Computers & Security, 17(1):27–30, ???? 1998. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404897802465.

Pounder:1998:SND

[Pou98e] Chris Pounder. Security and the new data protection law. Computers & Security, 17(2):124–128, ???? 1998. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404897819815.

Pounder:1999:IPA

[Pou99a] Chris Pounder. An initial plan of action for compliance with the UK’s data protection act 1998. Computers & Security, 18(1): 23–27, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404899800054.

Pounder:1999:RVB

[Pou99b] Chris Pounder. The revised version of BS7799 — so what’s new? Computers & Security, 18(4):307–311, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404899800753.

Press:1991:STI

[Pre91a] Jim Press. Secure transfer of identity and privilege attributes in an open systems environment. Computers & Security, 10 (2):117–127, April 1991. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/016740489190006Y.

Preston:1991:CCV

[Pre91b] Charles Preston. Creating a corporate virus plan. Computers & Security, 10(8):701–710, December 1991. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/016740489190089V.

Press:1996:OOC

[Pre96] Jim Press. Object oriented cryptographic facility design: Export considerations. Computers & Security, 15(6):507–514, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404897831252.

Preneel:1999:SAC

[Pre99] Bart Preneel. State-of-the-art ciphers for commercial applications. Computers & Security, 18(1):67–74, ???? 1999. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404899800091.

Peyravian:1998:PHV

[PRK98] Mohammad Peyravian, Allen Roginsky, and Ajay Kshemkalyani. On probabilities of hash value matches. Computers & Security, 17(2):171–176, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404897820160.

Peyravian:1999:HBE

[PRZ99] Mohammad Peyravian, Allen Roginsky, and Nev Zunic. Hashbased encryption system. Computers & Security, 18(4):345– 350, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404899800807.

Pujals:1993:SP

[Puj93] Joseph M. Pujals. Security and privacy. Computers & Security, 12(1):22–27, February 1993. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/016740489390005P.

Pottas:1991:CMR

[PvS91] D. Pottas and S. H. von Solms. A computerized management reporting aid for a mainframe security system. Computers & Security, 10(7):653–660, November 1991. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/016740489190123U.

Ramaswamy:1990:KMA

[Ram90] Raju Ramaswamy. A key management algorithm for secure communication in open systems interconnection architecture. Computers & Security, 9(1):77–84, February 1990. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/016740489090160U.

Rangan:1992:ATT

[Ran92] P. Venkat Rangan. An axiomatic theory of trust in secure communication protocols. Computers & Security, 11(2):163– 172, April 1992. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/016740489290043Q.

Rangan:1993:ADP

[Ran93] P. Venkat Rangan. The authenticated datagram protocol: a high performance, subtransport level, secure communication protocol. Computers & Security, 12(3):305–314, May 1993. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/016740489390119P.

Ratnasingham:1997:ESR

[Rat97] Pauline Ratnasingham. EDI security — re-evaluation of controls and its implications on the organizations. Computers & Security, 16(8):650–656, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404897875785.

Ratnasingham:1998:ESI

[Rat98] Pauline Ratnasingham. EDI security: the influences of trust on EDI risks. Computers & Security, 17(4):313–324, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404898800126.

Ratnasingham:1999:ITR

[Rat99a] Pauline Ratnasingham. Implicit trust in the risk assessment process of EDI. Computers & Security, 18(4):317–321, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404899800777.

Ratnasingham:1999:RLT

[Rat99b] Pauline Ratnasingham. Risks in low trust among trading partners in electronic commerce. Computers & Security, 18(7):587– 592, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404899820057.

Reitenspiess:1993:OSS

[Rei93] Manfred Reitenspiess. Open system security standards. Computers & Security, 12(4):341–361, June 1993. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/016740489390021V.

Reid:1995:OSSa

[Rei95a] Jim Reid. Open systems security: Traps and pitfalls. Computers & Security, 14(5):420, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/016740489597124S.

Reid:1995:OSSb

[Rei95b] Jim Reid. Open systems security: Traps and pitfalls. Computers & Security, 14(6):496–517, ???? 1995. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/016740489599313E.

Reid:1996:PHH

[Rei96] Jim Reid. Plugging the holes in host-based authentication. Computers & Security, 15(8):661–671, ???? 1996. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S016740489789699X.

Rosenthal:1991:BRP

[RH91] Paul Rosenthal and Barry Himel. Business resumption planning: Exercising your emergency response teams. Computers & Security, 10(6):497–514, October 1991. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/016740489190075O.

Richards:1999:NBI

[Ric99] Kevin Richards. Network based intrusion detection: a review of technologies. Computers & Security, 18(8):671–682, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S016740489980131X.

Rihaczek:1991:HIE

[Rih91] Karl Rihaczek. The harmonized ITSEC evaluation criteria. Computers & Security, 10(2):101–110, April 1991. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/016740489190003V.

Rihaczek:1994:DIL

[Rih94] Karl Rihaczek. Data interchange and legal security — signature surrogates. Computers & Security, 13(4):287–293, ???? 1994. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/0167404894900175.

Robertson:1991:EMS

[Rob91] Bernard Robertson. Electronic mail — is is safe to use? Computers & Security, 10(1):17–19, February 1991. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/016740489190050N.

Robinson:1992:CSE

[Rob92] John Robinson. Computer security evaluation: Developments in the European ITSEC programme. Computers & Security, 11(6):518–524, October 1992. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/016740489290184S.

Roberts:1993:LE

[Rob93] Steve Roberts. Letter to the editor. Computers & Security, 12(3):255–257, May 1993. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/016740489390112I.

Robb:1999:DFT

[Rob99] David Robb. Developing firewall technology: Hardwall — white paper. Computers & Security, 18(6):471–478, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S016740489980114X.

Rotenberg:1992:FEP

[Rot92] Marc Rotenberg. FBI endangering privacy for all. Computers & Security, 11(5):489, September 1992. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/016740489290243K.

Rushinek:1991:DEF

[RR91] Avi Rushinek and Sara Rushinek. Data encryption features for computer hardware and software profitability: I/O ports, expansion slots, algorithms, cyphers and security. Computers & Security, 10(4):345–357, June 1991. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/016740489190110Y.

Russell:1993:PVA

[Rus93] Selwyn Russell. Paradigms for verification of authorization at source of electronic documents in an integrated environment. Computers & Security, 12(6):542–549, October 1993. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/016740489390049B.

Russell:1994:ARP

[Rus94] Selwyn Russell. Audit-by-receiver paradigms for verification of authorization at source of electronic documents. Computers & Security, 13(1):59–67, February 1994. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/0167404894900965.

Sandhu:1992:LBE

[San92] Ravi S. Sandhu. Lattice-based enforcement of Chinese Walls. Computers & Security, 11(8):753–763, December 1992. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/016740489290131A.

Schell:1990:BRT

[SB90] Roger R. Schell and David Bailey. Book review: Taschenlexicon Sicherheit der Informationstechnik: by Dr. Hartmut Pohl with assistance from Ludger Hutte and Ursula Weiler, DATAKONTEXT-VERLAG Gmbh, Postfach 400253, D-5000 Koln 40, F.R.G., ISBN 3-89209-000-9, 1989. Computers & Security, 9(5):459–460, August 1990. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/0167404890900777.

Schramm:1991:CCU

[Sch91] Christof Schramm. Computing centers undermine MVS security. Computers & Security, 10(4):359–370, June 1991. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/016740489190111P.

Schramm:1993:ASM

[Sch93] Christof Schramm. Added-on security for MVS: weaknesses and measures. Computers & Security, 12(4):379–388, June 1993. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/0167404893900273.

Schultz:1995:NPF

[Sch95a] Eugene Schultz. A new perspective on firewall security. Computers & Security, 14(5):417, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/016740489597107L.

Schultz:1995:WBG

[Sch95b] Eugene Schultz. Why the bad guys are winning. Computers & Security, 14(5):427, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404895971532.

Schultz:1997:STP

[Sch97a] Eugene Schultz. Securing third party connections. Computers & Security, 16(6):524–525, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404897846949.

Schultz:1997:SWN

[Sch97b] Eugene Schultz. Securing Windows NT in today’s network environment. Computers & Security, 16(6):521–522, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404897846810.

Schwartau:1998:TBS

[Sch98] Winn Schwartau. Time-based security explained: Provable security models and formulas for the practitioner and vendor. Computers & Security, 17(8):693–714, ???? 1998. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404898801004.

Schultz:1999:WNS

[Sch99a] E. Eugene Schultz. Windows NT security: Kudos, concerns, and prescriptions. Computers & Security, 18(3):204–210, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404899800650.

Schwartau:1999:SDS

[Sch99b] Winn Schwartau. Surviving denial of service. Computers & Security, 18(2):124–133, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404899900512.

Soh:1995:SOI

[SD95] B. C. Soh and T. S. Dillon. Setting optimal intrusion-detection thresholds. Computers & Security, 14(7):621–631, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/0167404895000178.

Soh:1997:SIP

[SD97] B. C. Soh and T. S. Dillon. System intrusion processes: a simulation model. Computers & Security, 16(1):71–79, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404897000011.

Spector:1994:PSN

[SG94] Yishay Spector and Jacob Ginzberg. Pass-sentence — a new approach to computer code. Computers & Security, 13(2):145– 160, April 1994. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/0167404894900647.

Schumacher:1998:FFN

[SG98] H. J. (Jerry) Schumacher and Sumit Ghosh. A fundamental framework for network security towards enabling security on demand in an ATM network. Computers & Security, 17(6):527– 542, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S016740489880093X.

Spinellis:1999:TTP

[SGI^{+}99] Diomidis Spinellis, Stefanos Gritzalis, John Iliadis, Dimitris Gritzalis, and Sokratis Katsikas. Trusted third party services for deploying secure telemedical applications over the WWW. Computers & Security, 18(7):627–639, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404899820410.

Shain:1994:SIE

[Sha94] Mr Michael Shain. Security issues with enterprise multimedia. Computers & Security, 13(1):15–22, February 1994. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/0167404894900868.

Sherizen:1990:CCR

[She90] Sanford Sherizen. Criminological concepts and research findings relevant for improving computer crime control. Computers & Security, 9(3):215–222, May 1990. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/016740489090166Q.

Sherizen:1991:EUI

[She91] Sanford Sherizen. European unification ’92 impacts on information security. Computers & Security, 10(7):601–610, November 1991. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/016740489190117V.

Sherman:1992:BF

[She92a] Robin L. Sherman. Biometrics futures. Computers & Security, 11(2):128–133, April 1992. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/016740489290037R.

Sherman:1992:DSS

[She92b] Robin L. Sherman. Distributed systems security. Computers & Security, 11(1):24–28, March 1992. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/016740489290216E.

Shepherd:1995:HSS

[She95a] Simon J. Shepherd. A high speed software implementation of the data encryption standard. Computers & Security, 14 (4):349–357, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/016740489500007U.

Sherwood:1995:BCP

[She95b] John Sherwood. Business continuity planning. Computers & Security, 14(5):423, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/016740489597137Y.

Sherwood:1995:UFP

[She95c] John Sherwood. Using firewalls to provide secure Internet access. Computers & Security, 14(5):417, ???? 1995. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/016740489597106K.

Sherwood:1996:SMD

[She96] John Sherwood. SALSA: a method for developing the enterprise security architecture and strategy. Computers & Security, 15(6):501–506, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404897831240.

Shehu:1997:US

[She97a] Suleiman Shehu. Unix security. Computers & Security, 16(6): 522, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404897846822.

Sherwood:1997:MSOa

[She97b] John Sherwood. Managing security for outsourcing contracts. Computers & Security, 16(6):523, ???? 1997. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404897846858.

Sherwood:1997:MSOb

[She97c] John Sherwood. Managing security for outsourcing contracts. Computers & Security, 16(7):603–609, ???? 1997. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404897807973.

Shao:1994:DAC

[SHW94] Bao-Min Shao, Jing-Jang Hwang, and PeCheng Wang. Distributed assignment of cryptographic keys for access control in a hierarchy. Computers & Security, 13(1):79–84, February 1994. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/0167404894900981.

Sandhu:1991:IPM

[SJ91] Ravi Sandhu and Sushil Jajodia. Integrity principles and mechanisms in database management systems. Computers & Security, 10(5):413–427, August 1991. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/016740489190068O.

Sandhu:1992:EPS

[SJ92] Ravi S. Sandhu and Sushil Jajodia. Eliminating polyinstantiation securely. Computers & Security, 11(6):547–562, October 1992. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/016740489290191S.

Shiuh-Jeng:1996:SCB

[SJJF96] Wang Shiuh-Jeng and Chang Jin-Fu. Smart card based secure password authentication scheme. Computers & Security, 15(3):231–237, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404896000053.

Spruit:1996:ISD

[SL96] Marcel E. M. Spruit and Maarten Looijen. IT security in Dutch practice. Computers & Security, 15(2):157–170, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/0167404896000016.

Smith:1992:TVG

[Smi92] Tony Smith. Typeface vendors gang up on pirates. Computers & Security, 11(5):488, September 1992. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/016740489290027O.

Smith:1995:FCI

[Smi95a] Martin Smith. Facing the challenge of IT security. Computers & Security, 14(5):423, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/016740489597134V.

Smith:1995:AIO

[Smi95b] Paul Smith. Achieving inter-operable security services in open systems products. Computers & Security, 14(5):420, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/016740489597126U.

Smith:1996:AIS

[Smi96] Paul Smith. Achieving interoperable security services in open systems products. Computers & Security, 15(1):27–37, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/0167404896876206.

Smulders:1990:TIT

[Smu90] Peter Smulders. The threat of information theft by reception of electromagnetic radiation from RS-232 cables. Computers & Security, 9(1):53–58, February 1990. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/016740489090157O.

Solomon:1992:VAS

[Sol92] Alan Solomon. The virus authors strike back. Computers & Security, 11(7):602–606, November 1992. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/016740489290136F.

Sommer:1992:FCF

[Som92] Peter Sommer. Finding a crime to fit the punishment. Computers & Security, 11(5):491–492, September 1992. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/016740489290255P.

Sommer:1994:IEA

[Som94] Peter Sommer. Industrial espionage: Analysing the risk. Computers & Security, 13(7):558–563, ???? 1994. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/0167404894900043.

Spafford:1992:OPW

[Spa92] Eugene H. Spafford. OPUS: Preventing weak password choices. Computers & Security, 11(3):273–278, May 1992. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/0167404892902078.

Spafford:1997:BV

[Spa97a] Eugene Spafford. Blaming the victim. Computers & Security, 16 (6):527, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404897847062.

Spafford:1997:SOV

[Spa97b] Eugene Spafford. Some observations on vulnerabilities. Computers & Security, 16(6):515, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404897846627.

Schwemmlein:1998:RMR

[SPP98] J. Schwemmlein, K. C. Posch, and R. Posch. RNS-modulo reduction upon a restricted base value set and its applicability to RSA cryptography. Computers & Security, 17(7):637– 650, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404899800613.

Sandhu:1993:DCB

[SS93] Ravi S. Sandhu and Gurpreet S. Suri. A distributed capability-based architecture for the transform model. Computers & Security, 12(3):285–303, May 1993. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/016740489390118O.

Smith:1995:BCP

[SS95] Martin Smith and John Sherwood. Business continuity planning. Computers & Security, 14(1):14–23, ???? 1995. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/016740489596991B.

Shepherd:1990:CSS

[SSP90] S. J. Shepherd, P. W. Sanders, and A. Patel. A comprehensive security system — the concepts, agents and protocols. Computers & Security, 9(7):631–643, November 1990. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/016740489090062X.

Stahl:1993:ISW

[Sta93] Stanley H. Stahl. Information security in workstation environments. Computers & Security, 12(2):117–122, March 1993. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/016740489390090R.

Steinke:1991:DAA

[Ste91] Gerhard Steinke. Design aspects of access control in a knowledge base system. Computers & Security, 10(7):612–625, November 1991. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/016740489190119X.

Stoll:1995:NDP

[Sto95] Frank Stoll. The need for decentralization and privacy in mobile communications networks. Computers & Security, 14(6):527– 539, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/016740489500014Y.

Strous:1994:SEC

[Str94] Leon Strous. Security evaluation criteria. Computers & Security, 13(5):379–384, ???? 1994. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404894900302.

Slack:1993:IOO

[SU93] James M. Slack and Elizabeth A. Unger. Integrity in objectoriented database systems. Computers & Security, 12(4):389– 404, June 1993. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/0167404893900284.

Svigals:1994:SSA

[Svi94] Jerome Svigals. Smartcards — a security assessment. Computers & Security, 13(2):107–114, April 1994. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/0167404894900566.

Spafford:1993:SFC

[SW93] Eugene H. Spafford and Stephen A. Weeber. Software forensics: Can we track code to its authors? Computers & Security, 12(6):585–595, October 1993. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/016740489390055A.

Symonds:1994:SDC

[Sym94] Ian M. Symonds. Security in distributed and client/server systems — a management view. Computers & Security, 13(6):473– 480, ???? 1994. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/016740489190133X.

Tabbey:1990:CCP

[Tab90] Kirk W. Tabbey. Computer crime: “Preparing the computer specific search warrant”. Computers & Security, 9(2):117–123, April 1990. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/0167404890900836.

Tanner:1995:OES

[Tan95] Ray Tanner. Outsourcing — ensuring security compliance. Computers & Security, 14(5):418, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/016740489597111M.

Taylor:1992:VPY

[Tay92] Wendy Taylor. Virus-proof your PC. Computers & Security, 11(5):490, September 1992. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/016740489290248P.

Taylor:1995:CRA

[Tay95] Frank Taylor. The changing rises associated with computer systems as reflected by disaster recovery experience. Computers & Security, 14(5):423, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/016740489597136X.

Tzong-Chen:1996:API

[TCHS96] Wu Tzong-Chen and Sung Hung-Sung. Authenticating passwords over an insecure channel. Computers & Security, 15 (5):431–439, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404896000041.

Tzong-Chen:1995:GAS

[TCWH95] Wu Tzong-Chen and He Wei-Hua. A geometric approach for sharing secrets. Computers & Security, 14(2):135–145, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/016740489597047E.

Thimbleby:1991:CVE

[Thi91]  Harold Thimbleby. Can viruses ever be useful?  Com-

puters & Security, 10(2):111–114, April 1991. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/016740489190004W.

Thuraisingham:1991:MSI

[Thu91] Bhavani Thuraisingham. Multilevel security issues in distributed database management systems II. Computers & Security, 10(8):727–747, December 1991. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/016740489190093S.

Thuraisingham:1994:SIF

[Thu94] Bhavani Thuraisingham. Security issues for federated database systems. Computers & Security, 13(6):509–525, ???? 1994. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/0167404891901395.

Tanaka:1996:CSI

[TMO96] Kiyoto Tanaka, Hiroyuki Matsumoto, and Ikuro Oyaizu. A confidentiality system for ISDN inter-PC high-speed file transfer. Computers & Security, 15(2):141–155, ???? 1996. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/0167404896000077.

Tanaka:1991:KDS

[TO91] Kazue Tanaka and Eiji Okamoto. Key distribution system for mail systems using ID-related information directory. Computers & Security, 10(1):25–33, February 1991. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/016740489190052F.

Tettero:1997:ISE

[TOFS97] O. Tettero, D. J. Out, H. M. Franken, and J. Schot. Information security embedded in the design of telematics systems. Computers & Security, 16(2):145–164, ???? 1997. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404897000035.

Thuraisingham:1992:MSI

[TR92] Bhavani Thuraisingham and Harvey H. Rubinovitz. Multilevel security issues in distributed database management systems-III. Computers & Security, 11(7):661–674, November 1992. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/016740489290146I.

Varadharajan:1991:MSD

[VB91] Vijay Varadharajan and Stewart Black. Multilevel security in a distributed object-oriented system. Computers & Security, 10(1):51–68, February 1991. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/016740489190056J.

Varadharajan:1996:ESP

[VC96] Vijay Varadharajan and Claudio Calvelli. Extending the schematic protection model-II. revocation. Computers & Security, 15(6):525–536, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404896000107.

vanderMerwe:1998:SSC

[vdME98] Gert van der Merwe and Jan H. R. Eloff. Software source code, visual risk analysis: an example. Computers & Security, 17(3):233–252, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404898803386.

Venter:1998:DPI

[VE98] H. S. Venter and J. H. P. Eloff. Data packet intercepting on the Internet: How and why? A closer look at existing data packet-intercepting tools. Computers & Security, 17(8):683– 692, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404898800990.

Venema:1995:PS

[Ven95] Wietse Venema. Project Satan. Computers & Security, 14 (5):417–418, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/016740489597109N.

Vazquez-Gomez:1994:MS

[VG94] José Vázquez-Gómez. Multidomain security. Computers & Security, 13(2):161–184, April 1994. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404894900655.

Volpano:1997:SFT

[VI97] Dennis Volpano and Cynthia Irvine. Secure flow typing. Computers & Security, 16(2):137–144, ???? 1997. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404897000023.

Vince:1995:IPI

[Vin95] Judith Vince. An international perspective on IT security. Computers & Security, 14(5):423, ???? 1995. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/016740489597135W.

Voas:1993:MDE

[VPC93] Jeffrey M. Voas, Jeffrey E. Payne, and Frederick B. Cohen. A model for detecting the existence of software corruption in real time. Computers & Security, 12(3):275–283, May 1993. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/016740489390117N.

vonSolms:1996:ISM

[vS96] Rossouw von Solms. Information security management: the second generation. Computers & Security, 15(4):281–288, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/0167404896889395.

vonSolms:1998:ECS

[vS98] S. H. von Solms. Electronic commerce with secure intelligent trade agents. Computers & Security, 17(5):435–446, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404898000121.

vonSolms:1998:CSB

[vSG98] S. H. von Solms and J. H. S. Geldenhuys. Collecting security baggage on the Internet. Computers & Security, 17(4):337– 345, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404898800424.

vonSolms:1999:MML

[vSG99] S. H. von Solms and J. H. S. Geldenhuys. Managing multi level security in a military intranet environment. Computers & Security, 18(3):257–270, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404899800698.

vonSolms:1992:BSP

[vSN92] Sebastiaan von Solms and David Naccache. On blind signatures and perfect crimes. Computers & Security, 11(6):581–583, October 1992. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/016740489290193U.

Vaughn:1993:SSI

[VSU93] R. B. Vaughn, H. Saiedian, and E. A. Unger. A survey of security issues in office computation and the application of secure computing models to office systems. Computers & Security, 12(1):79–97, February 1993. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/016740489390014V.

vonSolms:1994:MCS

[vSvdM94] S. H. von Solms and Isak van der Merwe. The management of computer security profiles using a role-oriented approach. Computers & Security, 13(8):673–680, ???? 1994. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/0167404894900493.

Warren:1992:CSF

[War92] Peter Warren. Crime sans frontiers or the end of liberty. Computers & Security, 11(5):491, September 1992. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/016740489290251L.

Watts:1995:SCS

[Wat95] Clive Watts. Secure commercial servers. Computers & Security, 14(5):426, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404895971484.

Wood:1993:HEO

[WB93] Charles Cresson Wood and William W. Banks. Human error: an overlooked but significant information security problem. Computers & Security, 12(1):51–60, February 1993. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/016740489390012T.

Webster:1995:SPU

[Web95] Geoffrey Webster. Software piracy — an update. Computers & Security, 14(5):418, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/016740489597113O.

Wu:1992:SEL

[WFZ92] Jie Wu, Eduardo B. Fernandez, and Ruiguang Zhang. Some extensions to the lattice model for computer security. Computers & Security, 11(4):357–369, July 1992. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/016740489290176R.

Whiting:1992:UVV

[Whi92] Todd Whiting. Understanding VAX/VMS security. Computers & Security, 11(8):695–698, December 1992. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/0167404892900878.

Wilson:1997:CSC

[Wil97] Heather Wilson. Creating a security culture. Computers & Security, 16(6):523, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404897846871.

Winters:1990:SSD

[Win90] Peter Winters. Secure systems design — an evolving national strategy. Computers & Security, 9(5):379–389, August 1990. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/0167404890900674.

Wiseman:1990:CCD

[Wis90] Simon Wiseman. Control of confidentiality in databases. Computers & Security, 9(6):529–537, October 1990. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/016740489090131C.

Wiseman:1995:CSS

[Wis95] Simon R. Wiseman. Classification services in the SWORD secure DBMS. Computers & Security, 14(4):307–321, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/016740489500004R.

Withers:1992:PSS

[Wit92] Lucia Withers. Putting a stop to software piracy. Computers & Security, 11(5):490–491, September 1992. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/016740489290250U.

Wolfe:1995:CSF

[Wol95] Henry B. Wolfe. Computer security: For fun and profit. Computers & Security, 14(2):113–115, ???? 1995. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/016740489597030E.

Wood:1990:SCC

[Woo90a] Charles Creson Wood. Spectacular computer crimes — what they are and how they cost American business half a billion dollars a year!: Buck Bloombecker, Don Jones-Irwin, Homewood, IL, 1990. Computers & Security, 9(6):557, October 1990. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/016740489090134F.

Wood:1990:FMF

[Woo90b] Charles Cresson Wood. Fifteen major forces driving the civilian information security market. Computers & Security, 9(8): 677–686, December 1990. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404890901097.

Wood:1990:HMI

[Woo90c] Charles Cresson Wood. How many information security staff people should you have? Computers & Security, 9(5):395– 402, August 1990. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/0167404890900696.

Wood:1990:PSI

[Woo90d] Charles Cresson Wood. Principles of secure information systems design. Computers & Security, 9(1):13–24, February 1990. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/016740489090150R.

Wood:1991:BCS

[Woo91a] Charles Cresson Wood. Burning computer security, privacy, and freedom issues. Computers & Security, 10(6):524–532, October 1991. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/016740489190077Q.

Wood:1991:UIS

[Woo91b] Charles Cresson Wood. Using information security to achieve competitive advantage. Computers & Security, 10(5):399–404, August 1991. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/016740489190065L.

Wood:1993:DTI

[Woo93a] Charles Cresson Wood. Duress terminations and information security. Computers & Security, 12(6):527–535, October 1993. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/0167404893900479.

Wood:1993:PSI

[Woo93b]  Charles Cresson Wood.  Principles of secure informa-

tion systems design with groupware examples. Computers & Security, 12(7):663–678, November 1993. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/016740489390084I.

Wood:1994:FWS

[Woo94] Charles Cresson Wood. Fifty ways to secure dial-up connections. Computers & Security, 13(3):209–215, May 1994. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/016740489490071X.

Wood:1995:SIS

[Woo95a] Charles Cresson Wood. Shifting information systems security responsibility from user organizations to vendor/publisher organizations. Computers & Security, 14(4):283–284, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/016740489597068L.

Wood:1995:WNS

[Woo95b] Charles Cresson Wood. Wireless network security. Computers & Security, 14(5):422, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404895971303.

Wood:1995:WIPa

[Woo95c] Charles Cresson Wood. Writing infosec policies. Computers & Security, 14(5):418, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/016740489597114P.

Wood:1995:WIPb

[Woo95d] Charles Cresson Wood. Writing infosec policies. Computers & Security, 14(8):667–674, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/0167404896817068.

Wood:1995:SIN

[Woo95e] Peter Wood. Security implications of network infrastructure migration. Computers & Security, 14(5):424, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/016740489597143X.

Wood:1997:ECI

[Woo97a] Charles Cresson Wood. Essential controls for Internet electronic commerce. Computers & Security, 16(6):525, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404897846962.

Wood:1997:MVI

[Woo97b] Charles Cresson Wood. A management view of Internet electronic commerce security. Computers & Security, 16(4):316– 320, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404897801952.

Wood:1997:PAD

[Woo97c] Charles Cresson Wood. Policies alone do not constitute a sufficient awareness effort. Computers & Security, 16(6):522, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404897846834.

Wood:1997:IMS

[Woo97d] Peter Wood. Internet e-mail security. Computers & Security, 16 (6):515, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404897846639.

Wood:1997:NSS

[Woo97e] Peter Wood. NT server security, audit and control. Computers & Security, 16(6):523–524, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404897846901.

White:1996:CSM

[WP96] Gregory White and Vdo Pooch. Cooperating security managers: Distributed intrusion detection systems. Computers & Security, 15(5):441–450, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/0167404896000120.

Wood:1995:IIS

[WS95] Charles Cresson Wood and Karen Snow. ISO 9000 and information security. Computers & Security, 14(4):287–288, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/016740489597069M.

Yesberg:1996:QAV

[YA96] John D. Yesberg and Mark S. Anderson. Quantitative authentication and vouching. Computers & Security, 15(7):633– 645, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404896000144.

Yen:1996:CRM

[Yen96] Sung-Ming Yen. Cryptanalysis and repair of the multi-verifier signature with verifier specification. Computers & Security, 15(6):537–544, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404896000156.

Yang:1999:PAS

[YS99] Wen-Her Yang and Shiuh-Pyng Shieh. Password authentication schemes with smart cards. Computers & Security, 18(8):727– 733, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404899801369.

Zajac:1990:WWT

[Zaj90a] Bernard P. Zajac. The 1990s-what will they hold? Computers & Security, 9(6):503–507, October 1990. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/016740489090126E.

Zajac:1990:CVC

[Zaj90b] Bernard P. Zajac. Computer viruses: Can they be prevented? Computers & Security, 9(1):25–31, February 1990. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/016740489090151I.

Zajac:1990:ICS

[Zaj90c] Bernard P. Zajac. Interview with Clifford stoll. Computers & Security, 9(7):601–603, November 1990. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/016740489090057Z.

Zajac:1990:POS

[Zaj90d] Bernard P. Zajac. People: the “other” side of computer security. Computers & Security, 9(4):301–303, June 1990. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/016740489090097D.

Zajac:1992:CVR

[Zaj92a] Bernard P. Zajac. Computer viral risks-how bad is the threat? Computers & Security, 11(1):29–34, March 1992. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/016740489290217F.

Zajac:1992:CEA

[Zaj92b] Bernard P. Zajac. Cost-effectiveness of anti-viral software. Computers & Security, 11(3):217–226, May 1992. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/0167404892901992.

Zajac:1992:LE

[Zaj92c] Bernard P. Zajac. Letter to the editor. Computers & Security, 11(5):439–440, September 1992. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/016740489290008F.

Zajac:1994:BRA

[Zaj94a] Bernard P. Zajac. Book review: Applied cryptography: Protocols, algorithms, and source code in C: Bruce Schneier, John Wiley and Sons, 1994, 618 pages, $44.95, ISBN: 0-471-59756-2. Computers & Security, 13(3):217–218, May 1994. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/0167404894900728.

Zajac:1994:BRC

[Zaj94b] Bernard P. Zajac. Book review: Commonsense computer security — Your practical guide to information protection — 2nd edition: Martin Smith, McGraw-Hill, 1993, 280 pages, £24.95, $24.95 (USA), ISBN: 0-07-707805-5. Computers & Security, 13 (3):218, May 1994. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/0167404894900736.

Zajac:1994:BRI

[Zaj94c] Bernard P. Zajac. Book review: Information security policies made easy: Charles Cresson Wood, Baseline Software, Sausalito, CA 94966-1219, USA; 1994, 262 pages, $495, ISBN: 1-881585-01-8. Computers & Security, 13(3):218–219, May 1994. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/0167404894900744.

Zajac:1994:CEA

[Zaj94d] Bernard P. Zajac. Cost effectiveness of anti-viral scanners. Computers & Security, 13(5):392–400, ???? 1994. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/0167404894900329.

Zviran:1990:CPK

[ZH90] Moshe Zviran and William J. Haga. Cognitive passwords: the key to easy access control. Computers & Security, 9(8):723–736, December 1990. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/016740489090115A.

Zviran:1990:SDS

[ZHM90] Moshe Zviran, James C. Hoge, and Val A. Micucci. SPAN — a DSS for security plan analysis. Computers & Security, 9(2):153– 160, April 1990. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/016740489090089C.